A Complete Bibliography of Publications in Annals of Applied Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.27

Title word cross-reference

2^{K} [BDR16]. 3 [LHF^{+}20, SDHZ14, WNZK14, WSK^{+}21]. ^{2} [CDM18]. ^{(}R) [GV14]. ϵ [CCD22]. γ [CHH^{+}14]. k [BJ09]. L − 1 [WZL12]. G [MRW09]. p [SW10, TDS^{+}14, ZTLW20]. Q [OW11]. t [FD11]. × [GQ10, JEAS09]. W [Goe11, Goe14].

-allele [BJ09]. -D [LHF^{+}20]. -distributions [FD11]. -SELC [MRW09]. -space [OW11]. -SUP [CHH^{+}14]. -tapering [CCD22]. -value [SW10]. -values [TDS^{+}14, ZTLW20].

107th [LRM17]. 175 [SW17]. 1876-2015 [BCR^{+}19]. 19 [BKVW22, CdVM^{+}22, GRS23, KAGK^{+}23, Sun22, ZSLH23, ZHYS23]. 19th [DCHP21].

30-day [YKHS21]. 3D [SSH^{+}11].

55 [PS15].

A/H1N1 [PPB^{+}14]. Aaron [CW13b, CRZ13, Gil13, Moh13a, RP13, Sch13, Whi13]. ability [RGSB^{+}18]. abortion [SML^{+}21]. abrupt [WKLvD16]. absence [WS10a, WS10b]. abundance [BAH22, CGW^{+}10, KMKB16, WMKG19, WHNW15]. abundances [MWW20]. abuse [GMM08, YSR22]. Academy [FW21]. accelerated [DHM^{+}17, HSFP11, KAvdW^{+}23, WZLP20, ZW19]. accelerometer [DMN20]. access [BBB^{+}18, SP20]. accessibility [Hat14, Hav14, NSS14a, NSS14b, Pad14, Ser11, SC16, Wal14]. accident [Lie19]. account [DGCT10]. Accounting [BBL22, CBvdHvdH08, GSC^{+}20, LR21, LR20b, LZ07, MKN22, RLHD21, SPsLC16, WTJ10]. Accuracy [KBB^{+}11]. accurate [LZZL18, SBSH18]. achievement [GPRR16]. across [BCR^{+}19, FBM09, GG19, SH18, SBSH18]. activation [AMR18, WDL22]. Active [ZTCS20, vdHWC^{+}12]. activities [SBD23]. activity [CD20, CD18, JSF^{+}22, PW12, RGT13, WBKJ22]. actor [NS17, RVW20]. actor-event [RVW20]. acyclic [DGL13, RKM^{+}23]. ad [CA22]. adaptation [FBH23, GRL^{+}13, HPB23]. Adapting [SPF20, CT07]. Adaptive [BvdBS^{+}15, CMJZ22, Fuk19, WGL^{+}18b, XCCL20, BR08, BBDP11, DSC^{+}23, DD16, FFW09, KDL^{+}17, KHZK23, LNW08a, LNW08b, Lee18, MB08, Mur08, Qiu08, RJ11, RTB^{+}21, Ros12, SWHO11, SIS^{+}20, TJW10, Tib08, TvdL08, WZF^{+}13, ZZD11, ZYFF19, ZZTL22]. Adaptive-weight [WGL^{+}18b]. adaptively [LT11, YY11, ZST14]. added [ML11, RGSB^{+}18]. Additive [CPG^{+}21, ZLR19, CGM10, CMJ09, FFJJ14, HR22, Kan20, KKLS15, KKLS16, SH18, SPI^{+}23]. address [TACH21]. Addressing [SHC12]. adherence [HWHWA11]. adiposity [Hua18]. adjust [YSL08]. Adjusted [VVSK18, CAS20, FND09, HHH10a, SG16, STD13, SZO12, TCZ16, ZHJZ15, dCdCAGM16]. Adjusting [HTP14, MTZZ21]. Adjustment [ZCRC18, BYZ18, BE23, BvdH09, CdVM^{+}22, LS18, LCMJ11, PD20, RHL^{+}22, WYT^{+}20]. adjustments [Fre08, Lin13]. administrative [PT12]. admixture [CJMF18]. adoption [WD10]. adults [GGFG^{+}18]. advantage [FLS16]. adversarial [KW23]. adverse [HCD^{+}21, LTL19]. adversity [Hua18]. advertisements [FHI18]. aerosol [WMA^{+}14]. aerosols [BLTG15]. affect [YHX13]. affiliation [dCP10]. Afghanistan [RHHH13]. AFLP [LL09]. African [CSZK14, MGM^{+}14]. after [YSR22, YKLK23]. against [PLM^{+}16]. Age [WK10, QWC17, SMW^{+}22, vdKvEW17, KÓ14a]. Age- [WK10, SMW^{+}22]. age-specific [QWC17, vdKvEW17]. AGEMAP [SZO12]. agents [FZSI^{+}08]. ages [GWZ19]. aggregate [CAS20, CGCN22, SFPS^{+}21, SZ20]. Aggregated [CKM21, LM10a, LM10b, QBC13]. aggregates [RTB22]. aggregating [LY16]. aggregation [CHOK14, GBMRR20, JGF08, PLCX23, SJM^{+}14, WGL^{+}18a]. Agnostic [Lin13]. agreement [HS14]. AIC [CSL^{+}08]. aided [ZBT^{+}20]. aids [AM16, Bro09, HL08, JLS^{+}17, SW17, ZW19]. Air [WBK^{+}19, BC23, BJS^{+}22, BGC20, KDL^{+}17, Lee18, LYH^{+}16, MBL^{+}17, OSL^{+}14, WMT^{+}21, WCW^{+}22, XQ23]. airway [CHS^{+}16]. alcohol [ZLZB18]. alcoholism [SSL^{+}10]. Aldrich [PBS^{+}23]. algorithm [AY12, CHH^{+}14, FLRZ23, JL19, LZLW14, LWSP17, NCHJ13, NZ12, PLCX23, QGM^{+}14, SMZ16, WMGB23, WWM^{+}14, ZJLC08, ZMB23]. algorithmic [YWL^{+}12]. algorithms [BH11, MRW09, PC19, WL08b, ZZH08]. aligned [SYZ11]. alignment [CDB11, Jam07, MFB^{+}13, RS14]. all-or-none [ZHFN23]. Allele [CJM^{+}17, BJ09, CT07, RSI16, WS10c]. Allele-specific [CJM^{+}17, RSI16]. allocation [FPC20a, LMGJ15, ZWZ19, ZLR20]. allowing [AM16]. along [SRC15, YZS^{+}13]. ALPHA [Sta23]. alterations [SPPR08]. alternative [FGMP16, FD11, RAKS14, RAKS15, WOH23]. alternatives [AK12]. Alzheimer [LZK^{+}15, QM23, RGPC19, STMC17, WZ18]. ambient [PD20, SLZS08, WMT^{+}21]. ambiguous [Gho10]. ambulance [ZM16]. ambulances [WWMH13]. ambulatory [LCSZ15]. America [GF19]. American [BBL22, Gil17, JRHM22, LMB18, McE09, MSS09, RB10a, RB11, Sav16]. American-style [RB10a, RB11]. among [EHKW12, FD20, LBK^{+}23, PHCM^{+}10, RHC23, ZNB^{+}21]. analogies [SHGA10]. Analyses [ZCD^{+}20, Hua19, HYL23, TCS^{+}23, WS14, XLDO13, YMP11, ZJBS21, ZTH19]. Analysing [BMLG21, FSPWWE18]. Analysis [Ané08, CZM10, GMLB^{+}14, KZS23, LCG09, LAS16, MM15, ML13, MG22, PHWM11, PHCM^{+}10, Pur11, AMGG13, AL16, AY12, AICV11, AV15, AT15, APW^{+}09, BYZ18, BZS19, BMM^{+}16, Ben08, Ber11, Big13, BBE^{+}18, Bir08, BCJ15, BJS^{+}22, BWBS14, BDC^{+}11, BHIK09, BGH^{+}09, CM09, CCH^{+}21, CTM14, COM22, CLW20, CL13, CLZ16, CLFC23, CW20, CSS11, CVF10, CR11, DBF^{+}16, DL11a, DBTP21, DH11, DCCP09, DTZP13, DJ11, DKLL19, DK12, DTL^{+}23, DKS18, ELD09, EHM18, ENH^{+}18, FPL10, FM17, FSM^{+}19, Feu08a, Feu08b, Feu13, FO11, Fie10a, Fie10b, FH19, FSM17, FH14, Fuc08, GDJR20, GZB^{+}11, Gil17, GCC^{+}11, GDTP23, GTW13, GY23, GBNS22, GV14, HBHM13, HU11, HS22, HY14, HW08, Hol11, HZG^{+}22a, HPF13, HVL14, HZG22b, HST19, Ing08, IWG13, JMJ^{+}21, JLB^{+}14, JEK^{+}22, JLS23, JSH^{+}22]. analysis [Joh09, JLGJL12, JLRK23, JFM11, Kaf11a, KNWJ14, KH23, Kap11, KOJ^{+}14, KDH^{+}19, KOB^{+}20, KH13, KBG21, LYRR22, LHH10, LTL19, LRMM15, LL10, LG18, LWLX19, LKB21, Lia19, LSY^{+}22, LZCW21, LRHF12, LLR15, LHMN13, LPH22, LSM15, LCSZ15, MDP21, MLM13, MM11, MW11b, ML14, MCCW09, MBK^{+}21, MBH^{+}11, MV08, MDR10, NQdB^{+}07, NECS17, NS20, NDRF17, NL11, PK18, PK19, PHM^{+}23, PZ19, PAS23, PS15, PL08, PLCX23, QW08, RMP17, RZC^{+}18, REG^{+}11, RCF^{+}13, RN14, RBF^{+}20, RG21, RS10, Ros18, Ros09, Rou11, RLHD21, Rub08, RHC23, SFC11, SGLB10, SML^{+}11, SB20, SMZ21, SDL^{+}11, SRA^{+}15, SMR11, SDT08, SDP22, SG16, SZL16, SS15a, Sme11, SL19, ST14, STJ^{+}07, SGNM22, SMC^{+}20, SML^{+}21, Sti08, SKKS14, SD10, TDS^{+}14, Thi11, Tin11, TFB14, TLH14, URZF21, VBK19, VGH14, WA11, WTCW10, WBB13]. analysis [WJF^{+}15, WHLN15, WZ17, WME17, WOC18, WACY20, WLM^{+}21, WWL22, WLML23, WL08a, WS14, WYKH07, WI07, WL10, XKS15, YLH17, YGAT20, YN14, YL11, YLL12, ZST16, ZRA^{+}21, ZJLC08, ZASM12, ZD13, ZW15, ZMA^{+}19, ZWW13, ZS22, ZSG11, ZSFS22, ZZD22, ZGS^{+}14, vDDS^{+}09, vdBR10]. analytic [ZS09]. analyze [HKP23, MHB^{+}09, MGMB19]. analyzer [WHLN15]. analyzers [MBL^{+}17]. Analyzing [APC23, GTZ^{+}21, PT12, AWL13, ACG13, JLL09, ZMA^{+}20]. anatomical [GCC^{+}11]. ancestry [LLR10]. ancillary [ZZ08]. Angeles [XS11, KB10]. anger [QYP09]. angle [AYJ^{+}09]. Anglo [Zan15]. Anglo-Saxon [Zan15]. angular [OW11]. animal [APC23, HHA15, HGS23, JPS21, Mas22, NDRF17, RHHH18, SWH22, WJT^{+}21]. anisotropic [YZAD13]. ankle [HRP10]. annotations [Wen16]. annual [Kan20, KÓ14a, MLP^{+}19]. anomaly [HWPH10, HAL21, JLRK23, MH19, TEF22]. ANOVA [VH14, ZHB09]. answering [JLLK20]. Antarctic [LHF^{+}20, WKR21]. antidepressants [CMPR22]. antitrust [MVV13]. AOAS [Ano18]. apnea [HAFFH21, JEAS09]. app [MHG18]. Application [BE23, ENH^{+}18, FHI18, GMMW17, HKP^{+}19, HEHM23, JB21, JL19, LZTB16, MBD11, RAY14, SHC12, VVSK18, YLS14, ZRCC21, AD22, AT10, AS23, AS17, ARK^{+}18, BTJ^{+}14, BFM12, BPS22, BHC^{+}20, BBB^{+}18, BLM09, BBL22, Big13, BKGJ14, BvdH19, BL11, BHW15, BDR16, BAH22, Bro09, BdHZ08, CGM17, CCdCW18, CSGD16, CL13, CSS18, CLM22, Chi12, CCH19, CLLR20, ÇL12, CA18, CDN12, CWH20, CDB11, DMA19, DBF^{+}16, DAL^{+}23, DL11b, DFN08, DVF13, ENF14, FPLM18, FS13a, Fin13, FWK^{+}13, FS14, FH19, FSM17, FND09, FSJW11, FHSJ14, FLHA15, GPR^{+}22, Gau11, Gho10, GBMRR20, GH22, GTW13, GMM08, GKZS12, GMB15, GCL^{+}15, GV14, HSH12, HZG^{+}22a, HL08, HIH^{+}21, Hun12, JL10, JRHM22, JL09, JFRS17, JDP^{+}13, JLS^{+}17, JPTO17, JD18, JEK^{+}22, JL11, JCK22, JTLE22, JLRK23, JWZBC19, KGGQ15, KKMS16]. application [KDS20, KK12, KBH^{+}11, KX12, KKLS15, KKLS16, KG11, KH13, KHZK23, KB10, KM17, LSAR12, LBA11, LM10a, LM10b, LRDD22, LL10, LHPW13, LWLW15, LZ21, LAS16, LBD18b, LZCW21, LT12, LYH^{+}16, LW17, LQNM19, LLM20, LLKP18, LSM15, LCSZ15, LCZ^{+}17, LRS12, LRS15, MAZM13, MV12, MLM13, MGTZ21, MVW^{+}23, MGRG^{+}23, MMGC22, MR15, MNR14, MSH21, MBH^{+}11, NCHJ13, NvdBCR23, PHJ22, PPB11, PZB^{+}10, PRRW11, PDM19, PPLK18, PHT15, PL08, QM23, RJP16, SML^{+}11, SJH11, SH08, SWHO11, SZ12, SG17, SM10, SFGLR15, SW17, ST11, SC16, DPT22, STD13, SW10, SZO12, TJDE17, TDS^{+}14, TACH21, TAC^{+}16, TFB^{+}20, TM22, TB22, TEF22, VGH14, WD10, WLL17, WZ17, WZ18, WYT^{+}20, WACY20, WDSJ23, WMGB23, WGL^{+}18a, WMA^{+}14, WYKH07, WHAW21, XQ23, XKG^{+}19, XBS23, YMP11, YWQG23, ZLR19, ZZL11, ZCM^{+}11, ZY12, ZSH13, ZOZ17]. application [ZW19, ZSLH23, ZSCL23, ZMB23, ZLDR17, ZGS^{+}14, dCdCAGM16]. applications [AH16, AL16, AK12, BKS21, BWBS14, BH11, BFF^{+}09, CA23, CDM18, CSS11, DH11, DPHL10, DSH^{+}13, DKZ09, EHM18, FH09, FH14, Fuk19, Goe11, GM08, HS09, HGRS17, HE14, JJRZ21, JG23, KB23, LPKP22, LCB16, MZI18, MMBL20, MDR10, PC20, PM08, PPM14, PG14, PDS13, QTL^{+}22, QW08, RKM^{+}23, RKLT19, RS09, RGSB^{+}18, SDP22, SZ20, TWZ15, WCL23, WLA^{+}21, WH11, WZF18, XCCL20, YWLL22, ZASM12, ZD13, ZKS15, ZW15, ZYXS16, ZYC^{+}17, ZMO22, ZMLS22, ZW18, ZHM^{+}19]. Applied [Cox07, BDC^{+}11, DK18, HJS22, MSJ14, ZSG11]. Applying [CSZK14, GDG^{+}16]. apportionment [CN07]. Appraisal [PL11]. approach [AL16, AS10b, AZC^{+}17, AICV11, AFS07, BSDG22, BZC^{+}19, BK21, BM11, BZN18, CPvV^{+}11, CT07, CAV^{+}19, DH10, Dup17, EKW20, ERM15, GMNP^{+}21, GBST19, GH12, Gho10, GEC13, GG19, HMT12, HZG^{+}22a, JLA16, JLL09, KKMS16, KMKB16, LLR10, LHF^{+}20, LSS^{+}20, LSL^{+}15, LYH^{+}16, LMB18, LKTJ^{+}15, LLL10, MRV10, MN15, MC17, ME18, MLCW13, MVP11, MNR14, MSH21, NBZ11, OMM^{+}14, PKP16, PKGG23, PGL^{+}19, RLH^{+}15, RN14, RWK17, RDH^{+}20, RHHH13, Ryu22, Sad14, SBJR09, SvdLMP14, SYZ15, SCV^{+}10, SCTV11, TL11, TCS^{+}23, VBK19, WYW^{+}23, WSH^{+}14, YL13, YE14, YFHE20, YSH22, YH13, ZFB14, ZLR19, ZYFF19, ZS09, ZDL10, ZLD12]. Approaches [LM10a, LM10b, BSLL10, BvdH09, FLHA15, JGF08, LZTB16, TFB14]. appropriate [LFMM23]. Approximate [JSH^{+}22, Joh09, RMS^{+}19, BK21, JGVM18, KBMF^{+}23, RODC19]. Approximating [CDN12]. approximation [DS14, FWK^{+}13, HGS23, ISR12, SJH11, SCA13, TDBM23]. approximations [GGMG23, PK11, TM22]. Arabia [LG20]. arbitrary [OE12]. arc [JFM11]. archaeological [Ben08, Bir08, Ing08, Kad08]. archeological [Bir08, Feu08a, Feu08b, Feu13, Fuc08, HW08, MV08, Sti08]. architecture [NKAY10]. Arctic [CA18, DRB21]. area [ALC09, BL19, BK21, BvdB22, IHJ16, KB10, MRSA19, MWP^{+}15, MNR14, dCdCAGM16]. areal [BBL22, BHW15]. areas [BNMG23, YZAD13]. Argo [PKGG23, YSH22]. arising [SS15a]. arm [WLM^{+}21]. arm-based [WLM^{+}21]. array [CDP^{+}17, VH14, YWL^{+}12]. arrays [JLL09, OE12, SPPR08]. arrival [MMWH11]. arrivals [SRJ07]. art [CGM17, YSG16]. arterial [GBNS22]. artery [BMM^{+}16]. articles [TCZ16]. artifacts [JWH22, YRY17]. Asia [LGK18]. aspects [KPA^{+}10, NQdB^{+}07]. assassination [STJ^{+}07]. assays [SS15a]. assemblages [FBM09]. assess [MKKN21, RSI16, DPT22]. Assessing [AT15, AM16, Bro09, ÇL12, CSB^{+}15, HS13, HCYH20, HZF22, HSVF09, LRM17, MSS09, QGFL08, RWK17, SS10b, WLA^{+}21, ZCGC21, DL09, DLM14, GGFG^{+}18, LBND13, WHC^{+}22, YTHY18]. Assessment [ASX13, KKL11, BL19, FMB^{+}12, GG19, HMP22, HYL23, HF20, JYB16, KLCM20, LCZ^{+}17, ME18, RCBB19, RGSB^{+}18, SPPR08, SKBL23, WZ18, YGAT20, ZCM^{+}11]. asset [FPC20a, MGM^{+}14]. Assignment [HBP17]. assimilation [GKP^{+}16]. assisted [LLR15]. associated [AM07, KH23]. association [AXEC18, BDL^{+}16, Bro09, CGT^{+}14, CGN22, CFRW19, DVA^{+}19, FWGS11, GEC13, GS11, HY14, HVL14, JHMC16, LvdVvWvdW13, LGL^{+}12, LZLW14, LWLW15, LG18, LZ21, LBL20, MSH21, RN14, SM20a, SSZT19, SS15a, ST11, TJDE17, WWL22, WS14, WIC^{+}10, YD23, ZLS^{+}17, ZZL11, ZCRC18, ZMA^{+}20, ZSS23, Zho17b, ZS17]. associations [KHDV20, KHDV22, KBG21, MBYWX19, RSI16, SCK19, VML^{+}21, WILW22, WGL^{+}18b, ZLJW23]. assumption [HZG22b]. assumptions [GM09]. asthma [QBC13, SRZ^{+}15, ZRCC21]. astronomical [RMS^{+}19, TMvD^{+}17]. astrophysical [WBA^{+}14, WKLvD16]. astrostatistics [LRS09]. Asymmetric [GH22, JSX16, CHH^{+}14, CCD22, DB15, ZBC16]. asymptotic [FJK10]. Atherosclerosis [OSL^{+}14, LMKC12]. atmospheric [BPS22, FGS08]. atopic [ZRCC21]. attempt [MGTZ21]. attendance [SP13]. attractive [Mas22]. attributable [LR20a]. attribute [KK12]. attributes [CCS18, MSG^{+}20, NS17]. attribution [Kip22, KN20, LCYZ23]. attrition [SRH16]. auction [GH12]. auctions [JY10, PM08, SRJ07, dCP10]. audience [CVF10]. Audit [Sta23]. audited [Sta23]. audits [Sta08a, Sta08b]. augmentation [HGRS17, LYY13, WWMH13]. Australia [WTB16]. Australian [BCR^{+}19, KGGQ15]. authentication [GQ11]. authenticity [MDR10]. authorship [Kip22, RY11]. autism [LLR15]. autocorrelation [Ané08]. autocovariances [LLKP18]. Automated [BYZ18, MBH^{+}11, ZCG^{+}09]. Automatic [HHC17, WBA^{+}14]. autonomous [FTE^{+}21, HK23a, PPB11]. autopsies [KLCM20]. autoregression [CMAC^{+}23, HKP23]. autoregressive [CPP^{+}14, CMZ19, HR22, NBZ11, PHWM11, XCCL20]. auxiliary [FDR16, Lia19, SHW18]. average [AS17, NMD19, SWPN09, SGNM22, WGL^{+}18a, ZHFN23]. averages [Bro08]. averaging [ALC09, AS23, Big13, ZLD12]. Award [HSFP11]. Awards [FW21]. aware [KKL23]. axes [HRP10].

B [AN14, BBE^{+}18, LSZL22]. B-cell [BBE^{+}18]. B-scaling [LSZL22]. B-spline [AN14]. B2B [WD10]. back [FSM17]. Background [CM09, FSG16, LS18, SC16]. backtesting [Dav17, HK17a, Kra17, NZ17a, NZ17b, Sch17, Zho17a]. Backward [CW10]. bacteria [FYB^{+}15, JGVM18]. bacterial [DSH^{+}13, PMQW14, Ryu22]. BAGEL [LNR^{+}22]. bagging [LPKP22]. balance [BDR16, ZPR14]. balanced [KZ16]. Balancing [RHC23, FHI18, KBMF^{+}23]. ballots [Sta23]. BAN [GBNS22]. bandits [CDF^{+}20]. Bangladesh [KLH^{+}16]. bank [LMMS21]. banking [Dav17, HK17a, Kra17, NZ17a, NZ17b, Sch17, Zho17a]. bankruptcy [PHM^{+}23]. bariatric [TACH21]. BARISTA [SRJ07]. Barrier [GKP^{+}16]. BART [BKGJ14, CGM10, SMC^{+}20]. base [WZ16]. baseball [HS10, JSW09, ZTH19]. based [AS23, BBM20, BFM12, BMGN21, BBL22, BG09, BZC^{+}19, BZN18, CGCA21, CQ09, CSC^{+}12, CWWW17, CDF^{+}18, CN07, COC23, DGM^{+}08, DGL13, DHM^{+}17, EOB21, ESO19, FFM^{+}21, FAL^{+}10, FZSI^{+}08, GM16, Gho10, GGFG^{+}18, HMP22, HCS18, HCKFZ21, HMM09, JGF08, JFRS17, JPTO17, JLLK20, KPA^{+}10, KBFM12, KHZK23, LHF^{+}20, LBK^{+}23, LC12, LES12, LRI21, LSL^{+}15, LZCW21, LYH^{+}16, LX18, MGM^{+}14, MDR10, PM08, RFWE22, RLH^{+}15, RBB11, Ryu22, SHF^{+}16, SM15, SHM20, SYZ15, SS15a, SS10b, SSH^{+}11, TTH21, Tal13, TWA18, TACH21, TWZ15, VBK19, VHS13, WHLN15, WLM^{+}21, WCL23, WL08a, WZL12, WWM^{+}14, WZF18, YLH07, YRY17, ZPGO21, ZPMA10, ZWZ19, ZZL11, ZW15, ZSP19, ZMA^{+}19, ZMO22, ZHO22, ZW12, ZSG14, ZSMJ19, ZSMJ20, ZZH08]. Baseline [BGC20]. bases [BSNP16]. basis [BR08, LNW08a, LNW08b, LLR15, MB08, Mur08, Qiu08, Tib08, TvdL08]. basketball [FMBG15, SB20, SFDMM22, XZC17]. Batch [ZGJ^{+}22]. Batch-sequential [ZGJ^{+}22]. batting [Bro08]. BayCount [XZX18]. Bayes [Bro08, CC19, CT07, FW21, KP15, Lia19, MNR14, MG22, Mur10, MNB^{+}12, TCW21, WOH23, ZWW13]. Bayesball [JSW09]. Bayesian [KHDV22, AMGG13, ALC09, ASX13, ARC07, AN14, AZC^{+}17, ACS^{+}23, AMB^{+}20, AMR16, BLTG15, BZS19, BTJ^{+}14, BSLL10, BSDG22, BHB^{+}21, BM22, BBL22, BML^{+}20, BM11, BE23, BMT13, BDL^{+}16, BGK^{+}15, BB11, CDF^{+}20, CdVM^{+}22, CTM14, CFLP15, COM22, CFMR18, CGT^{+}14, COD22, CGCA21, CSGD16, CSC^{+}12, CLFC23, CCJ^{+}09, CGM10, CCS18, CGN22, CWS15, COC23, CAV^{+}19, CLGK22, CDB11, DH10, DBTP21, DAL^{+}23, DTZP13, Efr12, EFDS20, FPL10, FM17, FK22, FFR^{+}08, FPC20a, FGA09, FND09, FRBT13, GMNP^{+}21, GPR^{+}22, Gau11, GKM23, GRS23, Gil17, GL18, GDTP23, GS11, GL08, GQ11, HGM15, HSSF21, HWPH10, HS13, HS14, HJS22, HISV15, HGRS17, HBW17, HCW11, HZY^{+}15, HWHWA11, HR22, HKP23, HST19, DFGY23, JRHM22, JLDQ10, JGVM18, JLA16, JCS07, JSW09, JEAS09, JEK^{+}22, JYB16, JSH^{+}22, JLRK23]. Bayesian [JG23, Kad08, KNWJ14, KDH^{+}19, KAvdW^{+}23, KKLS15, KKLS16, KG11, KDH^{+}13, KHDV20, KHBV20, KHZK23, KLCM20, KPC^{+}19, LBND13, LZK^{+}15, LMGJ15, LTL19, LRMM15, LWLW15, LZW^{+}15, LNC^{+}19, LWLX19, LNR^{+}22, LSY^{+}22, LYY13, LZTB16, LWZ19, LSM15, LKTJ^{+}15, LCSZ15, LZ11, LW18, MFB^{+}13, MN15, MBGDS11, MC17, MLM13, MRM12, MB22, MCCW09, MH19, MVP11, MDWH21, MVV13, MBD11, MKM23, NWJ20, OBHL22, PMQW14, PHJ22, PNB22, PS15, RMP17, RFWE22, RCLWW10, RLH^{+}15, RF07, RBF^{+}20, RS14, RWK17, RK22, RGPC19, RODC19, Sad14, Sad18, SHAB22, SPI^{+}23, SP13, Sav16, SGC23, SWH22, STA18, SMW^{+}22, Sco09, SRH16, SPS20, SCL^{+}13, SM20b, SML^{+}21, DPT22, SCV^{+}10, SCTV11, SJ11, SRL10, TTH21, TMvD^{+}17, TJDE17, TL11, TTB22, TAC^{+}16, TSS10, TFB^{+}20, TWHP15, VCC22, WFHZ23, WBB13, WDL22, WFH^{+}22, WLML23, WCH^{+}23]. Bayesian [WYW^{+}23, WGL^{+}18a, WP12, WS14, Wen16, WWMH13, WBKJ22, WIC^{+}10, WHNW15, WHAW21, XZX18, XCS11, YGAT20, YWLL22, YGLH08, YJD21, YOZC23, YSL08, YMP11, YY11, YLL12, ZFB14, ZLR19, ZWZ19, ZZL11, ZY12, ZCS13, ZGV^{+}16, ZYC^{+}17, ZYFF19, ZMA^{+}19, ZSLH23, ZSCL23, ZKY14, ZCG^{+}09, ZLD12, ZS17, ZHM^{+}19, ZZXL23]. BayICE [TTH21]. be [BJ09, Efr08]. behavior [JAZ15, MMGC22, PGL^{+}19, SGLB10, ZS09]. behavioral [FS13a, KHBV20, YSR22]. behaviour [DSCS19, DLS^{+}17]. being [ENH^{+}18]. beliefs [SJM^{+}14]. benefit [YGAT20]. benefit-risk [YGAT20]. best [BK21, LMB18, MRSA19]. Beta [GV14, AS23, FHSJ14, HZL^{+}15, MWW20]. beta-binomial [MWW20]. BETS [ZJBS21]. better [LRMM15, PL11]. between [Bro09, Fin13, GEC13, GM15, KH13, LvdVvWvdW13, LZ21, LBL20, RSI16, RWK17, TMvD^{+}17, WGL^{+}18b, YSG16, ZSCL23, vdKvEW17]. beverages [GQ11]. BFLCRM [LZK^{+}15]. Bi [OP09, YWLL22]. bi-clustering [YWLL22]. Bi-cross-validation [OP09]. bias [AN14, AT15, HSFP11, KP15, NZRC13, TT09, WLA^{+}21, YSL08, ZCGC21, ZJBS21, ZST14, ZB11]. bias-corrected [KP15]. bias-reduction [AN14]. Biased [SRZ^{+}15, MGTZ21]. BIC [FND09]. biclustering [CMR15, MRG21]. bid [SRJ07]. Bidimensional [LPH22]. bifurcated [JSR16]. big [Men18, WEWX21, RFWE22]. bike [BCJ15]. bilateral [BHP10]. bimicrobial [Ryu22]. binary [BLM09, DSCS19, EFJ07, FLRZ23, FLHA15, HZG^{+}22a, LHH10, LBBM21, PGL^{+}11, ST11, WD10, XLDO13, YLL12]. binary-mark [DSCS19]. binding [ZWS08]. binned [VC14]. binomial [MWW20, MMBL20, RHL^{+}22, RJP16, WHNW15]. bio [DSB19]. bio-probes [DSB19]. Biobank [QTL^{+}22]. bioinformatics [BC09]. Biological [KPA^{+}10, BH11, DBTP21, JCS07, OM12, OKGM14, SHGA10, SCTV11, ZNB^{+}21]. biology [BC09, KK12, YWLL22, ZMO22]. Biomarker [TMY17, WZ18, BWS19, Bro09, CW20, MDP21, MZI18, SVYP11, Sch23, XZW21, ZZTL22, ZBLC17]. biomarkers [DBG21, DGCT10, FAL^{+}10, FKSBS19, GWZ19, HST19, SF11, ZNSL14]. Biomass [SGCT17]. biophysical [ZGM23]. biophysics [Kou08]. biosensor [ZYFF19]. Bird [SKKS14, CA18, DMGJ20]. birth [CGCA21, CWS15, DSH^{+}13, WMGB23, ZSG^{+}13]. birth-death [CWS15, WMGB23]. births [CGCA21]. bisexual [GGFG^{+}18]. Bivariate [YOZC23, BGH10, CTB17]. black [CFRW19, LBBM21]. blind [MBR09, WG23]. blinking [JWH22]. bloc [GQ10]. Block [ZZL11, LBA11, LRDD22, PC20, WWL22, XZC17]. Block-based [ZZL11]. blockmodel [NMW^{+}21, XFS10]. blockmodels [AWL13]. blocks [CLZ16]. blogosphere [LBA11]. blood [LT12, LCSZ15, SFGLR15]. blue [DLS^{+}17, YBL^{+}17]. Board [Ano16a, Ano16b, Ano16c, Ano23a]. body [CMPR22, GPRZ17]. bonds [LMMS21]. bone [AL16]. Bonferroni [GGQY07]. book [ZSP19]. boom [PPM14]. boosting [HIH^{+}21, KB10, MHK22, ZZH08]. Bootstrap [LHPW13, LY16, AS23, CQ09, CWE18, Efr12, LSAR12, Owe07]. bootstrap-based [AS23, CQ09]. Bootstrapping [OE12]. born [vdHWC^{+}12]. borne [LSS^{+}12]. Borrowing [RHR12]. both [LR20a]. Bottom [HLK18]. Bottom-up [HLK18]. boundary [CZM10, WSH^{+}14]. bounded [GV14]. Bounding [SGNM22]. box [CFRW19]. brain [BMM^{+}16, DAL^{+}23, ESO19, FO11, GEC13, GHO23, GBNS22, KKL23, LS18, LA22, LSL^{+}15, MAM17, MLX23, PDM19, RKLT19, SDT08, SG16, SL20, SC23, WNZK14, WZD19, WG23, WYW^{+}23, WRSS15, WBKJ22, YZAD13, ZYC^{+}17]. branching [OIHH09, XKG^{+}19]. breaches [SXZ23]. break [MBGDS11]. break-points [MBGDS11]. breast [BDC^{+}11, CCH^{+}21, GBMRR20, LHPW13, PZB^{+}10, QDN^{+}21, ZY12, ZHJZ15]. breast-cancer [BDC^{+}11]. breastfeeding [HS13, SvdLMP14]. Breiman [Bic10, Büh10, Cut10, FY10, Jor10, Ols10, Sto10]. bridge [CHS^{+}16]. Bridging [WLM^{+}21, EOB21]. Britain [BvdH19, KBG21]. broiler [RLH^{+}15]. Brown [TAC^{+}16]. Brownian [BX09, Cop09, Feu09, Gen09, GFS09, Kos09, Kos13, Rém09, SR09b, SR09c]. budding [OIHH09]. Building [HS10, LRMM15]. buildings [KZ16]. built [PBSVS23, WESVS23]. bulk [TTH21, ZLDR18]. bullet [HHC17, STJ^{+}07]. bundle [GTZ^{+}21]. burden [WGL^{+}18b]. bursty [WWCZ22]. business [AT15]. bust [PPM14]. bycatch [CFW17].

C [GQ10, HZY^{+}15, WSU^{+}19]. calcium [MHB^{+}09, MVP11, PSW18]. Calculations [GRL^{+}13]. Calibrated [PSD13, EJD19, GL18, HK23a]. Calibrating [GBH^{+}15, LHF^{+}20, SHAB22]. Calibration [BC23, AHZ23, BML^{+}20, BHP10, CHOK14, CHAP16, CKK^{+}22, HRFS19, LYBA22, RODC19, SKZ14, WYT^{+}20]. California [BMAF^{+}23, BWBS14, CSS11, QBC13, SYZ15, UH20, XS11]. call [ANFM09, GCL^{+}15, MMWH11, SH08]. called [ZTH19]. calls [ZW12]. camera [ARK^{+}18]. Campylobacter [RLH^{+}15]. can [BJ09]. Canada [GV14]. cancer [BDC^{+}11, CCH^{+}21, DTZP13, GBMRR20, HZG^{+}22a, JL11, KDS20, LGL^{+}18, LTL19, LHPW13, LCG09, LWSP17, LQNM19, LPH22, LCMJ11, MCCW09, NMD19, PZB^{+}10, PHT15, QDN^{+}21, Ros12, SSD15, SHAB22, TTB22, TP11, WNX^{+}23, YWB^{+}23, ZY12, ZOZ17, ZCD^{+}20, ZHJZ15]. cancers [GPBT22, MLX23]. cannabis [FS13a]. Canonical [KH13, SML^{+}11, WJF^{+}15]. canopy [BFM12]. capacity [MHH17]. Cape [JDP^{+}13]. Capture [KMKB16, ARK^{+}18, DPR^{+}20, FS13a, FHSJ14, MC17, MBDL14, MLKQ22, WMKG19]. Capture-recapture [KMKB16, DPR^{+}20, MC17, MLKQ22, WMKG19]. Capturing [SKAL19]. carbon [MAE^{+}08, RCBB19]. carcinoma [FCC15]. cardiac [GDG^{+}16]. cardiovascular [MDP21, SP20, SGCW07, ZS09]. cards [ML14]. care [FGMP16, FND09, GG19, HSD^{+}22, YKHS21, ZCD^{+}20]. caries [JYB16]. Carlo [KRC23, RB10a, RB11, Wan11, WLK18, ZW08]. carryover [ÇL12]. Carter [DPR11]. Cascadia [GSD^{+}18]. Case [MLX23, ANFM09, CCS18, CFRW19, FW21, HYL23, JGC^{+}18, JTLE22, KDS20, Lie19, LRM17, QGFL08, SWLS14, SC16, YL13, YR21, ZCRC18, ZMA^{+}20, ZSS23]. case-background [SC16]. case-cohort [HYL23, QGFL08]. case-control [JTLE22, KDS20, YL13, ZCRC18, ZMA^{+}20, ZSS23]. case/control [SWLS14]. casualties [Gil17]. cat [AS10a]. catalogs [RMS^{+}19]. catchment [GKP^{+}16]. categorial [GT10]. Categorical [FDR16, CW20, CAL^{+}23, GPRZ17, KBG21]. categories [BNW08]. category [JGF08]. cattle [RKM^{+}23, TFB^{+}20]. Causal [BHC^{+}20, BKVW22, KSD11, NKAY10, SML^{+}21, TB22, WBK^{+}19, ZRA^{+}21, AS17, AM16, BSDG22, BMLG21, BGK^{+}15, CA22, DLKM20, GEF22, HS13, KDH^{+}19, LYRR22, LBK^{+}23, LMM15, LL19, ML14, NMD19, PD20, Rub08, STA18, STG21, SHW18, SL20, YLS14, YFM19, ZB11]. cause [KLCM20, PHM^{+}23, SMW^{+}22]. cause-specific [SMW^{+}22]. CCS [JMY^{+}14]. CD8 [WYKH07]. cDNA [BM08]. cell [BC09, BBM20, BBE^{+}18, CDF^{+}20, CSC^{+}12, FPL10, FIM^{+}21, FWK^{+}13, FGA09, FGS^{+}10, Hun12, HGB21, HCRB23, LWZ19, LZ11, MKKN21, MGSD19, NvdBCR23, OIHH09, PBS^{+}23, VKG12, WYKH07, WL22, XKG^{+}19, ZLDR18]. cell-based [CSC^{+}12]. cell-type [HGB21]. cells [CZM10, RVW20, SSH^{+}11]. cellular [FMA18]. censored [ENH^{+}18, Gau11, HMM09, HCP^{+}17, JLDQ10, KAvdW^{+}23, SH11, YWQG23, ZW19]. censoring [CMJZ22, Gho10, MKN22, SKAL19, TMY17]. census [JGF08, KFB11]. center [ANFM09, SH08, ZZ08]. centered [DTL^{+}23]. central [LMMS21, MBD11]. centric [LC12]. century [DCHP21]. cervix [FCC15]. cesarean [YKLK23]. chain [FSPWWE18, Gho10, KRC23]. chains [HS09]. Challenge [BGH^{+}09, FCPL19, KBB^{+}11]. challenges [Bir08, CM09]. challenging [FRL18]. Change [CLZ16, KÓ14a, SZ12, AK12, CKHP15, CSZK14, EKW20, FJK10, GDJR20, GZB^{+}11, HWK21, LLR09, LCYZ23, SMZ16, TMY17, TCW21, TEF22, WZHC12, ZGM23]. Change-point [SZ12, AK12, EKW20, FJK10, GDJR20, HWK21, TMY17, TCW21, TEF22]. change-points [LLR09, SMZ16]. changepoint [SBD23]. changepoints [FM17]. changes [CCH19, FZCV22, MKKN21, MKS^{+}14, WKLvD16]. Changing [BSLL10, ZCGC21, ZHO22, ZST14]. channel [CSC^{+}12]. characteristics [FPC20b, GGFG^{+}18, SS20]. Characterization [ZMB23, BPS22, KK12, QW08]. characterize [MBL^{+}17, MLX23]. Characterizing [FMBG15, LSL^{+}15, PGL^{+}19]. charters [TFG12]. charts [CQ09, WZL12, XQ23]. cheating [CLM22]. ChemCam [BML^{+}20]. Chemical [STJ^{+}07, MAB^{+}14, MDWH21]. chemosensitivity [BC09]. Chi [SPF20]. Chi-square [SPF20]. Chief [Ano18]. child [AN14, MWP^{+}15, SMW^{+}22]. childhood [FGMP16, LT12, ZRCC21]. children [HS13, SRZ^{+}15, WHC^{+}22, YSR22]. Chile [DZ23, ZPR14]. China [SFGLR15]. ChIP [SJGM13]. ChIP-Seq [SJGM13]. Chlorophyll [GGPM16]. Chlorophyll-a [GGPM16]. choice [FGA09, RAB20, WTJ10]. cholera [BHC^{+}20, KLH^{+}16]. chromatography [KPA^{+}10, KOJ^{+}14]. chromosomal [SPPR08, ZMO22]. chronic [ÇL12, ZS18]. cigarette [WSGH12]. circadian [HKP^{+}19]. circular [APC23, JB21, SG17, YR21]. circular-circular [JB21]. Cis [ZW07]. Cis-regulatory [ZW07]. citation [Cra16, JJ16a, JJ16b, KP16, KT16, RRS16, Sil16, WR16]. citizen [BC23]. City [BDR16, GRS16]. claims [REFT18, SFB16, SZ20, WNX^{+}23]. class [FSM17, GL08, JDP^{+}13, JPTO17, KBG21, REG^{+}11, XLS^{+}19]. class-specific [GL08]. classes [WSK^{+}21]. classical [FD11]. Classification [CD12, TON20, Wit11, BTJ^{+}14, BT11, BFF^{+}09, Chi12, CKM21, Deb09, FAL^{+}10, GBST19, GPZ^{+}22, JLRK23, KW23, KDH^{+}13, KB23, Loh09, MBD11, RKLT19, SLBL23, TJW10, TWZ15, YH13]. classified [MKM23, VH14]. classifier [RHL^{+}22]. classifiers [LRMM15, SSD15, SDL^{+}11]. classifying [BZC^{+}19]. classroom [LSM15]. Clauset [CW13b, CRZ13, Gil13, Moh13a, RP13, Sch13, Whi13]. clickstream [AS23]. Climate [BCR^{+}19, KN20, BLTV14, CS13, CHOK14, DFN08, FC20, HISV15, HKP23, LCYZ23, LLL10, MLP^{+}19, SFC11, SJH11, ZCGC21]. climates [PMMS16]. Clinical [SW17, HL08, HZG22b, LSY^{+}22, LCMJ11, PBS^{+}23, QM23, SHC12, XLDO13, ZHFN23, ZZ08]. clique [DVA^{+}19]. clonal [MMY^{+}16, OSB15]. Clonality [TLF^{+}19]. clones [CDP^{+}17]. close [PBSVS23]. closed [COC23]. Clostridium [CD17]. closure [FSPWWE18]. cloud [AYJ^{+}09]. cluster [FCC15, GBMRR20, JFRS17, MKN22, WDSJ23, ZYC^{+}17]. cluster-randomized [JFRS17, WDSJ23]. cluster-structured [ZYC^{+}17]. clustered [AL16, BHC^{+}20, CLTZ22, DK18, FLHA15, GMMW17, HST19, MLP^{+}19, MKN22, SvdLMP14]. Clustering [ACS^{+}23, GGPM16, JPK21, KK13, MGM^{+}14, RFB17, TSY22, ZS18, BJS^{+}22, CHH^{+}14, ESO19, FSPWWE18, FRBT13, GPZ^{+}22, HZL^{+}15, JMJ^{+}21, KDL^{+}17, KBFM12, LNC^{+}19, LRI21, LWZ19, MM22, Moh13b, MBD11, PM08, RD14, SWM13, SR23, VHS13, WBC15, WFS19, WYL^{+}22, Wit11, WJL16, YWLL22, YRY17, Zan15, ZPMA10, ZSLH23, ZCH^{+}16]. clusters [MAB^{+}14, NCHJ13]. CMH [SPF20]. Co [BJS^{+}22, NS17, SR23, JDP^{+}13, PDM19, YWL^{+}12, ZLZB18]. Co-clustering [BJS^{+}22, SR23]. co-dependence [ZLZB18]. Co-evolution [NS17, PDM19]. co-occurrence [JDP^{+}13]. co-training [YWL^{+}12]. coalescent [CP20]. coarse [WBC15]. coarsening [HJS22]. Coauthorship [Cra16, JJ16a, JJ16b, KP16, KT16, RRS16, Sil16, WR16]. cocaine [YLG15]. code [GTW13]. codes [SPH17, WLML23]. coding [LW17]. coefficient [CLR16, CLLR20, GPRZ17, HHH10b, JWL^{+}19, LWLW15, Ser11, SW17, YZS^{+}13]. coefficients [CJMF18, GMLB^{+}14, RRSM18, WLA^{+}21]. cognition [GEC13]. cognitive [HS13, SEX21]. Coherence [ESO19]. Coherence-based [ESO19]. coherent [CGC12]. cohort [HYL23, QGFL08]. cohorts [KDL^{+}17]. coli [TFB^{+}20]. collaboration [ME18]. collected [WBKJ22]. collective [JJRZ21, SWH22]. College [YFM19, CTM14]. colon [LCG09]. colorectal [KDS20, LQNM19, ZCD^{+}20]. column [Ste07]. Combating [TSS10]. combination [LY16, WZ18]. combinations [KH23, MRW09]. combinatorial [GDTP23]. combined [CW20, Efr08, GPRR16, SVYP11, YY11]. Combining [GQ10, MHC15, QGM^{+}14, RFWE22, YMP11, FW21, HLK18, LT11, PNB22, TACH21, YOZC23, ZTLW20]. commensurate [MHC15, WLM^{+}21]. Common [WZD19, BBE^{+}18, GG19, HS13, LCB16]. communicable [WOK^{+}16]. communities [BZC^{+}19, DVA^{+}19, WWM^{+}14]. Community [BBL22, JRHM22, McE09, Sav16, BZ16, GBC21, JFRS17, JLLK20, KHLB22, PC20]. community-based [JFRS17]. community-level [GBC21]. comorbidity [ZLZB18]. compactly [KBH^{+}11]. comparative [Ané08, BCJ15, HYS19, JMY^{+}14]. comparator [DHL18]. Compared [FGMP16]. Comparing [SPH17, FS14, LGL^{+}18, SWHO11]. Comparison [KSD11, ALC09, CAV^{+}19, LBL20, MSSS^{+}10, Thi11, VGH14, YH20, YSG16]. comparisons [CAS20, Ros18]. compartment [FGA09]. compartmental [HCS18]. compensating [PG13]. Competing [TVJM13, AL16, MB22, PG13]. competition [BGH^{+}09, GM15]. competitions [FW21]. competitor [ZHB09]. complementary [Zan15]. complements [RAB20]. Complete [HRFS19, BHR11, KMKB16]. completion [SHR^{+}22]. Complex [AMR18, DZ23, ACS^{+}23, BSNP16, DW21, HVL14, ISR12, KKR13, MHB^{+}09, MWP^{+}15, MTZZ21, RLSF12, SHSZ19, WZHC12, WGL^{+}18b]. Complex-valued [AMR18]. compliance [BSDG22, HZF22, WDSJ23, ZHFN23]. complications [LHH19, YKLK23]. Component [LAS16, AXEC18, AY12, AICV11, APW^{+}09, DCCP09, EHM18, KSW^{+}21, LBL20, PZ19, RG21, SG16, YLH17, ZST16, ZW15, Zho17b]. components [BWS19, JEK^{+}22, LHH10, RC23b, Sch23, TM22, WT08, ZGS^{+}14]. Composite [AD22, BJ12, CHS^{+}16, FYB^{+}15, Hua19, WYT^{+}20]. Compositional [SL19, BAH22, HHN^{+}20, MKKN21, SZL16, SXC^{+}20, THN^{+}19]. compositions [MM22, RBF^{+}20]. compound [SZ20]. compounds [MAB^{+}14]. comprehensive [KOJ^{+}14]. Compression [FC20]. compressor [NMD19]. computation [JGVM18, PDS13, RODC19, TSS10]. computational [JWZBC19, KK12, MBR09, YWLL22]. Computationally [PHJ22, DFN08, ZMA^{+}20]. computer [GTW13, GBH^{+}15, GB16, KBH^{+}11, LKB21, LBBM21, MH19, MMGC22, PTH22]. computer-use [MMGC22]. computing [LPT^{+}11, ZS18]. concentration [DLZ10, EKW20, GPZ^{+}22, JLA16, REG^{+}11, SCDD18, ZCG^{+}09]. concentrations [OSL^{+}14, PYP^{+}09, PK18, PK19, SLZS08]. Concise [JMY^{+}14]. concurrency [AH16]. condition [TEF22]. Conditional [WYL^{+}22, AL16, CDN12, FW21, JSX16, LX18, NDRF17, RTB22, TSG17, XLS^{+}19, YL11]. conditionally [PHWM11]. conditioned [HS09]. conditions [ÇL12, MRM12, ZS18]. Conex [GHO23]. Conex-Connect [GHO23]. confidence [FS13b, HHH10b]. Confident [DLL^{+}18]. confidential [BBB^{+}18]. configurations [DSCS19, MPT12]. conflict [CSS18, MMBL20, WH11]. confocal [BHP10]. conformal [JPK21, KB23]. conformation [ZMO22]. conformational [WLK18]. confounder [HHH10b]. confounding [CMPR22, FH19, HBHM13, HZG^{+}22a, Mar08, MTZZ21, PD20, SP20]. congestion [BMLG21]. Connect [GHO23]. connectivity [BM11, CA23, ESO19, FO11, GHO23, KKL23, LA22, MVP11, MAM17, PCJW15, WG23, ZYC^{+}17, ZS22]. connectome [RGPC19]. connectomics [RKLT19]. consensus [BK20, EHKW12]. conservation [BL19, STA18, ZGJ^{+}22]. Conservative [Sta08a]. considerations [DLKM20]. consistent [SSZT19, ZZH08]. constant [LMW10, ZYFF19]. constants [BE23]. constrained [KS17, LvdVvWvdW13]. constraint [QGM^{+}14]. constraints [DJ11, GHK^{+}13, MBR09, WHLN15]. constructing [RMS^{+}19]. Construction [SDL^{+}11, LHPW13, ZSR14]. constructs [LSM15]. consumer [RAB20]. consumption [WSGH12]. contact [PHLH11, PHLH12, YGLH08, vdKvEW17]. contagion [HJS22, OV17]. contained [EKW20]. content [BM22, KZ16, SHR^{+}22, VML^{+}21]. Contents [Ano16d, Ano16e, Ano16f, Ano23b]. contest [DMA19]. Context [GGG^{+}12, SM10, Tin11, WBK^{+}19]. contextualization [ZPBW^{+}18]. continent [BCR^{+}19]. contingency [JGF08]. continual [LY16, LYY13]. Continuous [HHA15, FCC15, FS13a, FFM^{+}21, FHI18, HGS23, HS09, JLL09, JPS21, KK13, LK13, NS17, XZC17, ZMB23]. Continuous-time [HHA15, HGS23, HS09, JLL09, JPS21, XZC17]. continuum [GG19]. contour [DRB21]. contours [SRC15]. contrast [SXC^{+}20]. Contrastive [JTLE22]. contributed [LYRR22]. contributions [CN07]. Control [GGQY07, XQ23, AS10a, BMAF^{+}23, BG09, CQ09, CAS20, GMM08, JTLE22, KDS20, Mar08, NPM12, RCF^{+}13, SWLS14, WZL12, Wen16, YL13, ZZD11, ZCRC18, ZMA^{+}20, ZSS23]. Controlled [CGI08, KS19, WLM^{+}21]. controller [SIS^{+}20]. controls [KDH^{+}19]. convective [DT23]. conversion [LZK^{+}15]. Convex [BST15, BvdBS^{+}15]. convolution [Lee18]. cooling [GKS17]. cooperative [CGC12]. cooperative-Lasso [CGC12]. Coordinate [BH11, WL08b, FHHT07]. coordinated [AWL13]. Copula [DL11b, CGN22, Hof07, SZ20]. copula-linked [SZ20]. copulas [SFB16]. copy [CGT^{+}14, CJM^{+}17, LvdVvWvdW13, NZ12, SZ12, ZLOS10]. coronavirus [ZJBS21]. corpora [JMY^{+}14]. corporate [LMMS21, YTHY18]. corpus [SDP22]. corrected [KP15]. Correcting [HSFP11, Sch15]. Correction [Gel13, JSR16, KKLS16, KHDV22, Kos13, LM10b, LRS15, NS20, PK19, WS10a, LLS^{+}22, TT09, YRY17, ZCGC21]. correlated [BL07a, BL07b, BZN18, DPHL10, HS22, MZA22, NKAY10, RFB17, TEF22]. correlates [KH23]. Correlation [DK12, AR18, CSB^{+}15, Dup17, HGG13, KBH^{+}11, KY07, KH13, LZ07, SML^{+}11, WJF^{+}15, ZNB^{+}21]. correlations [JSX16, MHH17, WGL^{+}18b, XDO10]. Corrigendum [RAKS15, Sch23]. Cosmic [CM09, FDKP13, SCL^{+}13]. cosmological [BGH^{+}09]. cosmology [KBH^{+}11]. cost [FND09, KB10, MBR09, WNX^{+}23]. cost-adjusted [FND09]. cost-effective [FND09]. cost-sensitive [KB10]. costs [HBHM13]. count [BvdH19, CWS15, GPR^{+}22, SGC23, SS10a, VVSK18, ZASM12, ZSCL23]. counterfactual [KN20]. counting [Bir08, CP20, ELD09]. countries [LR20a, LR20b, Sha16, WFH^{+}22]. country [KAGK^{+}23]. country-specific [KAGK^{+}23]. counts [BZC^{+}19, BAH22, CWS15, LLKP18, NvdBCR23, SRC15, WZ16, XZX18, XCCL20, XBS23]. County [XS11, ZHJZ15]. coupled [CDM18, MP11, MMBL20]. Coupling [ZW07, KOB^{+}20]. couplings [KRC23]. course [CCJ^{+}09, FRBT13, QYP09, SHM15, TCW21, WL08a, WYKH07, ZD13]. Court [RY11]. Covariance [AV15, SJH11, AT10, BX09, BBE^{+}18, Cop09, CLK^{+}13, DKZ09, Feu09, Gen09, GFS09, JS08, Kos09, Kos13, LRZ08, MSH21, NH19, PDM19, REG^{+}11, Rém09, SR09b, SR09c, ZLDR17, ZW18]. covariance-enhanced [MSH21]. covariances [LXC11, PMMS16]. Covariate [FHI18, KDL^{+}17, SDH18, BvdH09, BDR16, CMPR22, GBST19, HHH10a, KHDV20, KHDV22, LCG09, PQR21, RHL^{+}22, REG^{+}11, SG16, SKAL19, STD13, WZ18, YSR21, ZLS^{+}17, ZHJZ15, ZB11, dCdCAGM16]. Covariate-adaptive [KDL^{+}17]. covariate-adjusted [HHH10a, SG16, STD13, ZHJZ15, dCdCAGM16]. covariate-dependent [REG^{+}11]. covariate-driven [CMPR22]. covariate-modulated [ZLS^{+}17]. covariates [CCH^{+}21, DZ23, FFR^{+}08, HWHWA11, HCRB23, JPTO17, LK13, LL10, RDL23, SKAL19, SW17, YLS14, ZNSL14, ZCRC18, ZZ18, ZW19, vdHWC^{+}12]. coverage [BFM12, DW21, PHJ22]. COVID [ZJBS21, BKVW22, CdVM^{+}22, GRS23, Sun22, ZSLH23, ZHYS23, KAGK^{+}23]. COVID-19 [ZJBS21, BKVW22, CdVM^{+}22, GRS23, Sun22, ZSLH23, ZHYS23, KAGK^{+}23]. Cox [GM16, GGMG23, JG23, LZK^{+}15, PZSW23, QGFL08, QDN^{+}21, SG17, WLG17]. credible [EHKW12]. Crime [FCPL19, BJ19, Moh13b, SG17, XCCL20, ZX22]. criminal [LRDD22]. criteria [RLH^{+}15]. criterion [PCJW15, SHW18]. Critical [WCW^{+}22]. criticism [Kip22, WSH^{+}14]. critique [Lin13]. crop [NCHJ13]. cross [CGM17, MGTZ21, MHH17, MKM23, OP09, RGSB^{+}18, SWHO11, TT09, TWHP15, TEF22, VH14, WWL22]. cross-classified [MKM23, VH14]. cross-correlated [TEF22]. cross-correlations [MHH17]. cross-sectional [MGTZ21]. cross-sections [CGM17]. cross-study [TWHP15]. cross-trait [WWL22]. cross-validation [RGSB^{+}18, SWHO11, TT09]. crossed [BLM^{+}23]. crowdsourced [CLW20]. cryo [CHH^{+}14]. cryo-electron [CHH^{+}14]. cryptocurrency [GH22]. cube [HBW17]. cumulative [CQ09]. cure [KXC09, PA23]. Curiosity [BML^{+}20]. current [MGTZ21, SPH17]. curse [XCS11]. Curve [Jam07, HSH12, VBK19, dCdCAGM16]. curve-based [VBK19]. curves [Big13, CMZ19, CD20, GKZS12, GKS17, MDWH21, TMvD^{+}17, WKR21, ZCG^{+}09]. Customized [PHT15]. cycle [BMLG21, FPL10, PL08]. cycles [MLP^{+}19, PPM14]. cyclic [KPC^{+}19]. cycling [WCH^{+}23]. cyclone [MVW^{+}23, WFHZ23]. cylinders [MSJ14]. cytometry [FBH23, HCRB23, OIHH09, PLCX23]. CytOpT [FBH23].

D [LHF^{+}20, SDHZ14, WNZK14, WSK^{+}21]. Daily [KKR13, BCR^{+}19, CMAC^{+}23, REG^{+}11, XBS23, ZGLH13]. dairy [RKM^{+}23]. dangers [ZJBS21]. dark [GFW^{+}09]. Data [DSC^{+}23, RCP^{+}16, SCRS^{+}20, ZZTL22, AL16, AS10b, ACG13, AT10, AS23, APC23, AV15, AT15, Ané08, ACS^{+}23, AMR16, AK12, ABB^{+}12, BTA20, BTJ^{+}14, BBM20, BBB^{+}18, BNW08, BLM09, BR08, Big13, BPSC14, BC23, BMT13, BK20, BJS^{+}22, BHW15, BGC20, BAH22, BB11, COM22, CMAC^{+}23, CJMF18, CLW20, COD22, CHOK14, CHAP16, CKK^{+}22, CSC^{+}12, CL13, CWWW17, CD20, CLM22, CLTZ22, CDP^{+}17, CW20, CAS20, CCH19, CGCN22, CLR16, CDN12, CT07, CH14, CVF10, CWS15, CAV^{+}19, CBvdHvdH08, CGFT15, DAAM22, DH10, DBF^{+}16, DPR^{+}20, DBTP21, DH11, DGL13, DTZP13, DB22, DL11b, DW21, DSH^{+}13, DJ11, DKLL19, DHM^{+}17, DVF13, EKW20, ELD09, EOB21, EFJ07, ENH^{+}18, FM17, FLRZ23, FK22, FH09, Fie10a, Fie10b, FMB^{+}12, FWK^{+}13, FH13, FS14, FH14, FDR16, FMA18, FBH23, FRBT13]. data [FRL08, Fuk19, FLHA15, GGPM16, GMMW17, GPR^{+}22, Gau11, GGMG23, GFW^{+}09, GDG^{+}16, GLB^{+}17, GKP^{+}16, GBMRR20, GH22, GM08, GQ10, GPRR16, GPZ^{+}22, GY23, GHO23, GS13, GCL^{+}15, GV14, HG10, HSSF21, HS22, Hof15, HGRS17, HMM09, HBW17, HZG^{+}22a, HVL14, HPB23, HSVF09, HCP^{+}17, IGA22, JGF08, JSR16, JL10, JLDQ10, JCS07, JWL^{+}19, JLS^{+}17, JEK^{+}22, JLS23, JSH^{+}22, JLGJL12, JS08, Kaf12, KXC09, KFB11, KOJ^{+}14, KMKB16, KY07, KK13, KBFM12, KHDV20, KHBV20, KHDV22, KM17, KB23, KZS23, LSAR12, LNW08a, LNW08b, LHH10, LLR09, LGL^{+}12, LG18, LLZ19, LS22, Lie13, LRI21, Lin13, LSL^{+}15, LBD18b, LRHF12, LYY13, LZTB16, LYH^{+}16, LWFW16, LGK18, LWZ19, LQNM19, LLM20, LCRM21, LSZL22, LHMN13, LZ13, LCB16, LCSZ15, LCZ^{+}17, LN12, LZ11, MKKN21, MDP21, MAE^{+}08, MSSS^{+}10, MV14, MLP^{+}19, MGMB19, MMM^{+}16]. data [MC17, ML13, ML11, MJ16, MBDL14, MON20, MGTZ21, MVW^{+}23, MB08, Men18, MWP^{+}15, MLKQ22, MAB^{+}14, MVP11, MKN22, Moh13b, MMBL20, MG22, MBH^{+}11, MKM23, MNB^{+}12, MDR10, MHC15, Mur08, NCHJ13, NvdBCR23, NZRC13, NWJ20, OGP^{+}18, OSB15, OE12, PHWM11, PG13, PSD13, PHJ22, PNB22, PGL^{+}11, PAS23, PSW18, PDM19, PT12, PCJW15, PDS13, PHLH11, Pur11, QHPD19, Qiu08, QBC13, RMP17, RFWE22, RHL^{+}22, RKM^{+}23, RZC^{+}18, RSI16, RB10b, RN14, RFB17, RBF^{+}20, RHR12, RG21, RAKS14, RAKS15, RHHH13, Ryu22, SSD15, SGLB10, STM17, SGL^{+}08, SJH11, SFDMM22, SP13, SGC23, SRA^{+}15, SVYP11, SvdLMP14, SMW^{+}22, SJGM13, SS10a, SDP22, SWPN09, SHM20, Sha16, SWM13, SLBL23, SPsLC16, SIL^{+}11, SZL16, SSL^{+}10, SG17, SHC12, SM10, SS20, SBSH18, SX09, SW17, SR23, Ste09]. data [SBS14, SSH^{+}11, SH11, SZO12, SKZ14, SHH22, SXZ23, TTH21, TMN18, TON20, TACH21, TCW21, Tib08, THN^{+}19, TSS10, TvdL08, TEF22, VFH16, VVSK18, VC14, VBK19, VH14, VKG12, WD10, WR12, WNZK14, WSM^{+}16, WLL17, WZ17, WME17, WOC18, WSU^{+}19, WFS19, WEWX21, WYL^{+}22, WNX^{+}23, WMGB23, WCL23, WYW^{+}23, WMT^{+}21, WS10a, WS10b, WJT^{+}21, WGL^{+}18a, WCD23, WL08a, WP12, WS10c, WWMH13, WH11, WFC^{+}22, WKG^{+}15, Wit11, WESVS23, WOK^{+}16, WMKG19, WI07, WL10, WGL^{+}18b, WZF18, WL22, XZW21, YLLS21, YSH22, YN14, YLG15, YLC^{+}17, YL11, YRY17, YSL08, YMP11, YWQG23, YBL^{+}17, YLL12, ZST16, ZRA^{+}21, ZWS08, ZK10, ZZL11, ZCM^{+}11, ZY12, ZASM12, ZD13, ZSH13, ZW15, ZGV^{+}16, ZOZ17, ZYFF19, ZMA^{+}19, ZRCC21, ZMLS22, ZSLH23, ZSCL23, ZWW13, ZCD^{+}20, ZS09, ZDL10, ZSG11, ZHJZ15, ZSMJ19, ZSMJ20, ZCP14, ZLDR18, ZZD22, ZZXL23, DMN20, vdBN09]. Data [RFWE22]. Data-adaptive [DSC^{+}23, ZZTL22]. data-Insights [SFDMM22]. data-with [JLS^{+}17]. database [VFMD17]. databases [CD17]. dataset [ZSP19]. datasets [BBE^{+}18, Ger09, PGW18, YKLK23]. date [DGCT10]. Dating [TFG12, FLHA15]. day [YKHS21]. death [CWS15, DSH^{+}13, KLCM20, WMGB23]. deaths [BKVW22, HWK21, SP20]. debiased [PKGG23]. debit [ML14]. deciding [NCHJ13]. decision [FZSI^{+}08, JWZBC19, KHZK23, MM08, MVV13, RCBB19, SB20, SC14, YH13, ZZ22]. decision-making [FZSI^{+}08, SC14]. decision-theoretic [YH13]. decisions [FGA09, RY11]. declassified [GGCM20]. decoding [VRN^{+}11]. decomposition [ML23, WFS19, XZX18, ZSH13]. decompositions [MHB^{+}09, VH14]. Deconvolution [WZ16, BHR11, MMY^{+}16, TTH21]. deduplicating [HPB23]. deep [KW23, MTZZ21]. deeply [ZSP19]. default [ASX13, GJPS08, YTHY18]. defects [LBD^{+}18a]. defensive [FMBG15]. deformation [SDHZ14]. degeneration [GWZ19]. Degradation [ZSG11, ZSG14]. Degradation-based [ZSG14]. degree [CTM14, ZKS15]. delayed [DHL18, LYY13]. delays [TMvD^{+}17]. delinquency [KHZK23]. delivered [STG21]. delivery [YLS14, YKLK23]. delta [ZGJ^{+}22]. demand [CPG^{+}21, ZM16]. Democrats [GC08, Gel13]. demographic [MZ12, WLA^{+}21]. Dempster [ELD09]. Dengue [WDSJ23, JGC^{+}18]. densely [WOK^{+}16]. densities [BvdH19, MJ16, MBK^{+}21]. Density [JL10, CR13, CD20, CDN12, LM10a, LM10b, SGCT17, WKR21, Yua09]. density-dependent [SGCT17]. dental [JYB16]. Department [BDR16]. dependence [CCdCW18, CMZ19, CA18, Dup17, DT19, GH22, HHHV17, JJRZ21, KÓ14a, LZP16, PHCM^{+}10, RRSM18, SCDD18, SPsLC16, XLS^{+}19, YLG15, ZLZB18, ZHO22, ZW18]. dependencies [CA23]. dependent [AMR16, DMGJ20, GSC^{+}20, REG^{+}11, SGCT17, SW17, Tal15, TMPF12, WZLP20]. depends [ZB11]. depression [LNR^{+}22]. depth [BD11, HS14, LW17]. Deriving [BC09]. descent [BH11, LYBA22, SWLS14, WL08b]. Describing [EFJ07]. Design [DHL18, BFM12, BBL22, CDF^{+}20, CMJZ22, DLKM20, FFM^{+}21, HHK^{+}16, HYL23, LMM15, LMMS21, LSY^{+}22, MM08, Ros12, Rub08, SVYP11, SM10, SC16, WDSJ23, YJD21, ZGJ^{+}22]. design-based [BFM12, FFM^{+}21]. Designing [LFMM23]. designs [BM08, BDR16, BR16, DPHL10, DZ23, JL09, KMMS13, RD14, SRZ^{+}15, WDSJ23, YKLK23, ZBLC17]. detect [CSZK14, HST19, NZ12, PLM^{+}16, YLH07]. Detecting [FZCV22, Ger09, JCK22, MVW^{+}23, MNB^{+}12, RY11, Sad14, SYZ11, WKLvD16, CGFT15, HJ18, LT11, LZLW14, RS09, RDH^{+}20, YL13, ZWW13, ZLDR17]. Detection [BZ16, CLM22, DH10, HHH10a, LLR09, WSH^{+}14, AMR18, CDM18, EKW20, GZB^{+}11, HWPH10, HAL21, HZY^{+}15, JSF^{+}22, JLRK23, KGGQ15, KOJ^{+}14, LCYZ23, MBDL14, MBR09, MKS^{+}14, MH19, MZI18, PC20, SVYP11, SJGM13, SCL^{+}13, SMZ16, SBD23, SHH22, TEF22, WZHC12, WWL22, ZX22]. determinants [GPRZ17]. determination [SSD15]. detrended [Dup17]. developed [Sha16]. developing [PDM19]. Development [GG19, LSL^{+}15, LT12, WGL^{+}18a]. device [ZMLS22]. devices [FS14]. diabetes [ZZL11]. diabetic [CFH^{+}14]. diagnose [KRC23]. Diagnosis [YKHS21, FSM17, GWZ19, SEX21, WLML23, dCdCAGM16]. Diagnosis-group-specific [YKHS21]. diagram [DH11]. diagrams [KSD11, PG14]. diameter [SGCT17]. diarization [FSJW11]. dictionary [YD23]. Did [KÓ14a, RHC23]. died [Far22]. dietary [ZCM^{+}11]. difference [CSL^{+}08, WZHC12]. differences [JL11, Lie19, LRM17, MVW^{+}23, RLHD21, SG16, SPH17]. different [BvdB22, ZSG14]. differential [BL11, BAH22, CFH^{+}14, HST19, JND12, LT11, LMW10, LZCW21, PLM^{+}16, RSH12, Ros09, RJP16, RHHH18, TCS^{+}23, WZ18, ZYC^{+}17, ZS22]. differentially [BZN18, CGFT15]. differentiate [WSK^{+}21]. differentiation [WYKH07]. difficile [CD17]. diffractograms [DGM^{+}08]. diffuse [BBE^{+}18]. diffusion [CT18, DKZ09, GM15, HMT12, KL16, Laz16, OW11, SDT08, Sch16, WLP^{+}16, WLPP16, YZAD13, YZS^{+}13, ZGS^{+}14]. dimension [CHOK14, SFDMM22]. dimension-reduced [CHOK14]. dimensional [AWL13, ACG13, BW18, BHW15, CLR16, HHLC16, KOJ^{+}14, LLR09, LZW^{+}15, LAS16, LWFW16, MAB^{+}14, MDR10, PGW18, PCJW15, RGSB^{+}18, SWPN09, SM13, WL10, XBS23, ZCS13, ZYC^{+}17, ZLDR17, ZW18, ZGS^{+}14]. dimensionality [Fuk19, TJW10]. Dimensions [YSR22]. dioxide [MK21, ZGLH13]. diploid [SBS14]. diplomatic [GGCM20]. dipoles [SJA^{+}13]. Direct [HCS18, JFM11]. directed [Cha17, DGL13, RKM^{+}23, WYW^{+}23]. direction [JLGJL12, KL16, Laz16, SDT08, Sch16, WLP^{+}16, WLPP16]. directional [HRP10, Mas22]. Dirichlet [CL13, DMGJ20, FLP23, LDV^{+}10, MM22, PAS23, SX09, TMN18]. Dirichlet-multinomial [CL13, PAS23]. Dirichlet-tree [MM22, TMN18]. disability [DL11b, EFJ07, MV14]. disaggregate [CPG^{+}21]. disaggregation [BBL22]. disagree [FRL18]. disambiguations [AC12]. disclosure [CFLP15, SS10b]. DISCO [RS10]. discontinuities [vdBR10]. discontinuity [DZ23, LMM15, LMMS21]. discoveries [GGQY07]. Discovering [CLZ09, PQR21, ZPBW^{+}18, LLR10]. discovery [BYZ18, BG09, BZN18, GPR^{+}22, JCJ20, LRI21, LWSP17, Mur10, Sch08, SDT08, SWHO11, TWZ15, TP11, Wen16, YD23, ZLS^{+}17, ZW07]. discrepancy [BML^{+}20, HHHV17, Sta08b]. discrete [AS10b, BZC^{+}19, GCL^{+}15, HHA15, HGS23, KK13, YFHE20]. discrete-space [HHA15, HGS23]. discretely [HCS18]. discretization [HGS23]. discretized [COD22]. discriminability [KPC^{+}19]. discriminant [MDR10, WL10]. discriminants [ABNG14]. discriminating [Kip22]. discrimination [PG13, SCDG17, WK10, XDM15]. discriminative [BCJ15, DSC^{+}23]. discussant [Rub18]. Discussion [Ben08, Ber11, BR08, BX09, Bir08, Cop09, CR11, Cra16, CRZ13, DL11a, Dav17, Feu09, Fuc08, Gen09, Gil13, GFS09, HU11, Hat14, Hav14, HW08, Hol11, HK17a, Ing08, Kad08, KL16, Kap11, KP16, KT16, Kos09, Kos13, Kra17, Laz16, MM11, MB08, Moh13a, MV08, Mur08, NL11, Qiu08, RRS16, RP13, Rém09, Rou11, SMR11, Sch17, Sch13, Sch16, Sme11, Sti08, TF11, Tib08, TvdL08, WA11, Wal14, WR16, Whi13, Zho17a, New09, Sil16, ZPBW^{+}18, Tin11]. disease [FCGA^{+}13, GWZ19, GM16, GMMW17, HCS18, JMJ^{+}21, JGC^{+}18, MDP21, Mar08, MH14, RTB^{+}21, Ros16, RGPC19, STMC17, SP20, WLL17, WZ18, YLH17, YLH07, ZJBS21]. diseases [AH16, ENF14, HVL14, LSS^{+}12, WOK^{+}16]. Disentangling [YTHY18]. disequilibrium [ZSS23, ZW12]. disjunction [PG14]. disorders [FMB^{+}12, FLP23]. disparities [GRS16, KM17, Tal15]. disparity [Tal13]. dispersion [LYH^{+}16, RJP16]. dissecting [LC10]. Distance [PM08, BX09, CZM10, CA18, Cop09, Feu09, Gen09, GFS09, Kos09, Kos13, Rém09, SR09b, SR09c, XDM15, YBL^{+}17, ZMO22]. Distance-based [PM08]. distances [DH18]. distancing [ZHYS23]. distinct [LN12, NQdB^{+}07, TCS^{+}23]. Distributed [Tad15, BWS19, Sch23, WMT^{+}21, WHC^{+}22, ZS18]. Distribution [CQ09, BHR11, CZM10, CVF10, FJK10, Gau11, GJPS08, Gil17, HBW17, HWF15, JLA16, Joh09, Mac20a, Mac20b, QHPD19, SGCT17, TCS^{+}23, WTCW10, ZW08]. Distribution-free [CQ09]. Distributional [SHSZ19, KKLS15, KKLS16, MVW^{+}23]. Distributions [AM07, CHJCK18, FD11, FZZW17, Goe11, HK23a, MSSS^{+}10, PBSVS23, SZ20, VDP08, VC14, WJT^{+}21, WBA^{+}14, WMA^{+}14, ZKS15, dCP10]. district [PAS23]. divergence [Tal13]. Diverse [KY07, WP12]. diversity [AMR16]. divisions [FGS^{+}10]. DNA [DSH^{+}13, FSM^{+}19, FGS^{+}10, HZL^{+}15, HSSF21, LvdVvWvdW13, NZ12, SZ12, TWH13, Wei07, ZLOS10, ZSMJ20]. Do [ML14, PHM^{+}23, FDH10, YHX13]. document [CB10, SDL^{+}11]. documents [GGCM20]. Does [MM15, SM20a, XLS^{+}19, LMB18]. Dollo [KN17]. domain [BvdB22, FBH23, RHZ^{+}15, RSD22]. domain-specific [RSD22]. domains [CLZ16, FBM09, GPR^{+}22, WSU^{+}19]. domestic [YSL08]. door [Feu13, SHW18]. dose [BBDP11, LY16, LYY13, MDWH21, PTGN12, TTB22, URZF21, ZHYS23, ZMB23]. dose-finding [BBDP11, URZF21]. dose-response [PTGN12, TTB22, ZHYS23]. doses [CCD22]. dosing [CH14]. double [RCLWW10, RGSB^{+}18]. Doubly [JLS^{+}17, JLL09, MSG^{+}20, SBD23, JLDQ10, QM23, SH11, ZK10]. doubly-interval-censored [JLDQ10]. Doubly-online [SBD23]. doubly-robust [QM23]. down [HLK18]. downscaler [BGH10]. Downscaling [MSSS^{+}10, TETJ17]. DREAM3 [RKM^{+}23]. drift [BGC20]. drink [BvdH19]. drink-driving [BvdH19]. driven [BGM17, CMPR22, GGFG^{+}18]. driver [LWSP17]. drivers [RCP^{+}16, SCRS^{+}20, WCH^{+}23]. driving [BvdH19, HK23a, JAZ15, ZASM12]. drop [MKN22]. drop-out [MKN22]. dropout [LMM15, QM23, ZY12]. drug [HWHWA11, LNR^{+}22, LY16, LCRM21, PRRW11, SML^{+}11, SWHO11, TTB22, TWZ15, WGL^{+}18a, YLG15, ZCG^{+}09]. drug-combination [LY16]. drugs [GM15, LBK^{+}23, YY11]. DTI [YZAD13]. dual [MSJ14]. duality [DH11]. due [SVYP11, vdBR10]. duplicates [Sad14]. duration [MGTZ21, PHCM^{+}10]. during [DHL18, DGCT10, KAGK^{+}23, SBD23, ZSLH23]. dust [BLTG15, MAL^{+}22]. dyadic [KZS23, SS20, ZY12]. Dynamic [MBL^{+}17, RDL23, SJM^{+}14, SHF^{+}16, SJA^{+}13, SCW^{+}23, WLL17, YLH17, BPS22, BSNP16, DBF^{+}16, DMVT23, DD16, FLRZ23, FPC20a, GTW13, HSH12, HWHWA11, JPS21, LMW10, LPT^{+}11, LW18, QHPD19, RS09, SB20, SKS12, Sin11, SAV^{+}14, TWA18, TF11, VIF13, WBB13, XQ23, XFS10, ZZXL23]. Dynamical [Chi12, KSH^{+}13, MLCW13, PPB11]. dynamically [BM11]. Dynamics [GF19, BZS19, BBG^{+}12, CMZ19, CFH^{+}14, CGCN22, DH18, DKLL19, DGCT10, HE15, HL08, DFGY23, LSS^{+}20, LCSZ15, MDP21, OM12, RS09, SKS10, SH11, TB22, BHW15]. dysbiosis [MWW20].

E-loyalty [JY10]. each [Efr09]. Early [GWZ19, ZBT^{+}20, FGMP16, HZG22b, JPTO17, ZJBS21, YFM19]. Earth [Gne12]. earthquake [BWBS14, CSS11, FSG16, RK22]. earthquakes [KO14b]. East [CSZK14, vdHWC^{+}12]. ecclesiastical [JLB^{+}14]. ECG [Big13, ZS09]. ecological [EFDS20, GQ10, Thi11, WHNW15, YHE20]. ecology [CWH20, WS10a, WS10b]. economic [GC08, Gel13, KZ16]. economies [HKP23]. edge [DRB21, MH19]. edit [PSD13]. Editor [Ano18, Fio08, HTL13, Mac20a, Mac09, SST^{+}08, SH15]. Editor-in-Chief [Ano18]. Editorial [Ano16a, Ano16b, Ano16c, Ano23a, Fie07, Fie08, Pad13, Pad14, Ste11]. education [FGMP16, FH19, KBG21, STD13, BDR16]. educational [BMH16, CLM22, JL09, WBB13]. EEG [GHO23, JPTO17, KF10, LSS^{+}20, QW08]. Effect [SvdLMP14, BMLG21, BKVW22, CHOK14, CLLR20, CMPR22, DPHL10, DT23, ENH^{+}18, GEF22, HSFP11, HBHM13, HMP22, HS13, HZF22, IR13, KKMS16, KDS20, KHBV20, LMM15, LNR^{+}22, LYBA22, LKTJ^{+}15, MSG^{+}20, Mur10, NMD19, Ros18, RHC23, SH18, SP20, SGNM22, WRNR14, WTB16, YLS14, YHE20, ZHFN23, ZB11]. Effective [WYT^{+}20, BM11, CCD22, FND09, LSM15]. effectiveness [MB22, ZPR14]. effects [AMB^{+}20, AS17, AM16, BLM^{+}23, BSDG22, BMAF^{+}23, BL19, CFW17, CFLP15, CWWW17, CAS20, ÇL12, CB22, DLKM20, DLL^{+}18, FD20, FPC20b, FSPWWE18, GMMW17, GZB^{+}11, GRL^{+}13, HRP10, HHH10a, HWHWA11, Hua18, Hua19, HYS19, IWG13, JRHM22, JCK22, LBK^{+}23, LC10, LA22, LSY^{+}22, LX18, LCMJ11, LN12, MIP22, ML11, MG22, MKM23, NMD19, NECS17, OGP^{+}18, PD20, PC20, PZSW23, RBF^{+}20, STG21, SHW18, SKAL19, SL20, SJHJD20, SML^{+}21, DPT22, STD13, SHH22, TB22, URZF21, VFMD17, WTCW10, WYH^{+}14, WZ18, WH11, WSH^{+}14, YL13, YLS14, YFM19, ZLR19, ZHB09, ZGLH13, ZNSL14, ZSFS22, NS20]. efficacy [DLL^{+}18, FHI18, YGLH08]. Efficiency [ZLR20, LCYZ23, SRZ^{+}15, WHLN15, ZBLC17]. Efficient [BHB^{+}21, BWT^{+}20, KBH^{+}11, LRS12, LRS15, MBR09, MP11, PDS13, SWHO11, vdKvEW17, KMMS13, LWSP17, PHJ22, QM23, SVYP11, TSS10, ZMA^{+}20, ZZTL22]. egocentric [PHLH11]. egocentrically [KM17]. elastic [CKM21, FKSBS19]. election [GM08, JLS23, Men18, Sta08a, Sta08b]. election/poll [JLS23]. elections [RLH^{+}13, DPT22]. electoral [KSP16]. electrical [ERM15]. Electricity [Dup17, CPG^{+}21, CMZ19, Lie13, Lie19]. electroencephalographic [WOC18]. electron [CHH^{+}14, QHPD19, SSH^{+}11, SHR^{+}22]. electronic [CD17, GSC^{+}20, HCP^{+}17, JG23, SPS20, TACH21, WD10, WSM^{+}16, WLML23, ZCD^{+}20]. electrostatic [FPLM18]. elegans [HZY^{+}15]. element [PK18, PK19]. elementary [KP15, KS17]. Elevated [CN07]. Elicitability [Dav17, HK17a, Kra17, NZ17a, NZ17b, Sch17, Zho17a]. elimination [MRW09]. elliptical [SRC15]. elusive [BvdH09]. email [MKM23]. embedded [MHG18, ZZTL22]. embedding [ZMO22]. embryonic [HZY^{+}15]. emergency [MMWH11]. emerging [BWT^{+}20]. EMG [SIS^{+}20]. emission [EKW20, KDH^{+}19]. emphasizing [OMM^{+}14]. Empirical [FM17, KP15, Lia19, ME18, Sch08, XDO10, AS23, Bro08, CT07, FFR^{+}08, MRSA19, Mur10, MNB^{+}12, RRSM18, SIL^{+}11, TCW21, VC14, ZWW13]. employees [BBB^{+}18]. Employer [BHW15]. employment [ENH^{+}18, HTP14, WK10]. emulating [BJ12]. emulation [GB16, GSD^{+}18, PACB21]. emulators [KBH^{+}11]. Encoding [VRN^{+}11]. end [HSD^{+}22, HK23a, RAKS14, RAKS15]. end-of-life [HSD^{+}22]. end-to-end [HK23a]. endangered [CFW17]. endogenous [CMPR22, HBP17]. endorsement [SPH17]. endosperm [LC10]. endpoint [HS09]. endpoint-conditioned [HS09]. endpoints [HZG22b, QGFL08, STMC17]. energy [FDKP13, GFW^{+}09, HBW17, HLK18, SCL^{+}13, WKLvD16, ZW08]. enforcement [MVV13]. engagement [SGNM22]. engineered [PTGN12]. engineering [LSAR12]. English [TFG12]. enhanced [MSH21]. enough [KM16]. enriches [HGS23]. enrichment [ABB^{+}12, NQdB^{+}07]. ensemble [FMA18, SM15]. ensembles [CS13, FP08, JCCG18, JL11, NV18, SCW^{+}23]. Ensembling [TWZ15]. entering [DH10]. entity [CSS18]. entry [TFB14]. environment [AMGG13, JPS21, PBSVS23, SPI^{+}23, Sin11, TF11, WZHC12, WS14, WESVS23]. environmental [AMB^{+}20, BZS19, CDN12, FFM^{+}21, SCRS^{+}20, WZS19, WHC^{+}22, ZMC^{+}21]. environments [YHX13, ZSG14]. enzymatic [DK12]. epidemic [CPvV^{+}11, DPR^{+}20, TDBM23, vdBN09]. epidemics [BNMG23, TFB^{+}20]. epidemiological [FK22]. epidemiology [KDL^{+}17, RD14]. epigenetic [CLZ16, RSI16, ZMC^{+}21]. epigenetics [Hua18]. epigenomic [BZ16]. epilepsy [WJL16]. epistasis [ZZL11]. epoetin [CH14]. eQTL [KX12, MBYWX19, MSH21, SW10]. equal [HZG22b]. equation [CFH^{+}14, FLHA15, GFW^{+}09, LMW10, RHHH18, VIF13, ZYC^{+}17]. equations [BL11, LYY^{+}23, RSH12, SBJR09, WNX^{+}23]. equilibrium [ZGM23]. equity [Ser11]. Equivalence [FS14, FH13, SH18]. erosion [BK21].

Error [SR09a, BSNP16, CSC^{+}12, CA18, GMB15, KGGQ15, LLS^{+}22, NPM12, PZSW23, Sch15, SW17, SKZ14, TT09, WBK^{+}19, ZCM^{+}11]. Error-free [SR09a]. error-in-variables [KGGQ15]. error-prone [GMB15, SW17]. errors [CKHP15, DPHL10, GSC^{+}20, SJH11, VFMD17]. errors-in-variables [CKHP15]. erythrocyte [VGH14]. escape [APC23]. Escherichia [TFB^{+}20]. esophageal [LTL19]. ESPREE [IHJ16]. essential [CC19]. establishment [PT12]. estimands [SL20]. estimate [CCD22, JGVM18, PPB^{+}14, RFB17, SMW^{+}22, STG21, vdBR10, vdBN09]. estimated [BLTV14, WLA^{+}21]. estimates [ASY09, BBL22, BJ09, CT07, DCHP21, IHJ16, McE09, STA18, TMvD^{+}17, vdHWC^{+}12]. Estimating [AMB^{+}20, AS17, AR18, BNMG23, BMAF^{+}23, BBE^{+}18, CAS20, CB22, CW13a, CW13b, CSL^{+}08, CRZ13, DT23, ELD09, FYB^{+}15, Gil13, Gil17, GEC13, GCL^{+}15, HH21, IR13, KAGK^{+}23, KSAX10, LGL^{+}18, LBK^{+}23, LL11, LR20a, LCB16, MRMB15, MB22, Moh13a, MKM23, NMD19, PG13, PD20, PHLH11, PHLH12, RP13, Sch13, SLZS08, SL20, SC16, Sun22, TWH13, WFH^{+}22, WCW^{+}22, WJT^{+}21, WCD23, Whi13, WMKG19, WL22, YHE20, YLC^{+}17, ZGLH13, ZKS15, ZYFF19, ZHFN23, BHP10, BvdH09, CdVM^{+}22, FLHA15, GGFG^{+}18, Lee18, LYY^{+}23, LFMM23, SBJR09, SP20, TWA18, WNX^{+}23, WMT^{+}21, WHNW15, YGLH08, YSL08, YFM19, ZZ18]. Estimation [CFH^{+}14, CMPR22, DS14, FIM^{+}21, GKZS12, IHJ16, LMW10, MON20, PZSW23, RKM^{+}23, RHR12, SS20, YLS14, ALC09, AN14, AZC^{+}17, AYJ^{+}09, BLTG15, BFM12, BK21, BvdB22, BGC20, CGI08, CFLP15, CW10, CGCA21, CK14, CJM^{+}17, CSS18, CDF^{+}18, CGCN22, CD18, DPR^{+}20, DLZL16, DW21, DFN08, FFJJ14, Far22, Fin13, FBW^{+}17, Gau11, Gho10, Goe11, HRP10, HHLC16, Hof07, HCKFZ21, HGB21, HLK18, JL10, JLA16, JCJ20, JLS^{+}17, KL16, KKMS16, KFB11, KMKB16, KB10, KP15, LK13, Laz16, LRZ08, LCYZ23, LCMJ11, MAZM13, MBYWX19, MRSA19, MSG^{+}20, MZ12, MGRG^{+}23, MWP^{+}15, Moh13b, MNR14, Mur10, OKGM14, PHJ22, PNB22, PDM19, PLM^{+}16, PL08, RLSF12, RBB11, Sad18, Sav16, SGC23, SCRS^{+}20, SK22, Sch16, SKS10, SBSH18, SM13, SIS^{+}20, TMY17, TSG17, TLF^{+}19, VFH16, WDSJ23, WWMH13, WBA^{+}14]. estimation [WLP^{+}16, WLPP16, WMA^{+}14, YLLS21, Yua09, YJZ09, ZLOS10, ZZTL22, Zho17b, ZST14, ZB11, vdKvEW17]. estimator [BY13, FO11, SSZT19, SRL10, ZSP19]. estimators [FSG16, FKSBS19]. Estrogen [QDN^{+}21]. eSwatini [DPR^{+}20]. ETAS [CDM18, KO14b]. ETAS-I [CDM18]. ethnic [LMKC12, OSL^{+}14]. Euclidean [DKZ09, WCW15]. Europe [Far22, KÓ14a]. European [CCdCW18]. Europeans [FDH10]. Eurovision [DMA19]. evaluate [PC19, ZLR19]. Evaluating [AK12, CH14, DMVT23, HBHM13, LMM15, LMMS21, WSM^{+}16, HS13, HZG22b, JSW09, RCF^{+}13, SM20b, ZNSL14]. Evaluation [GKS17, BSLL10, BLM09, COC23, HK17b, HK23b, IR13, KHLB22, MLM13, RLH^{+}15, SDL^{+}11, SFPS^{+}21, ZBLC17]. event [AS10b, Bro09, COD22, ÇL12, FZSI^{+}08, GMMW17, GSC^{+}20, GMB15, HCD^{+}21, KMMS13, KN20, MDP21, OBHL22, RVW20, RC23b, SHM20, SPsLC16, SG17, SH11, TFB14, WLL17, WACY20, WZLP20, WWCZ22, YLC^{+}17, ZHO22]. event-history [FZSI^{+}08]. event-related [KMMS13, SPsLC16]. events [CW10, CW13a, CW13b, CRZ13, FCPL19, GGCM20, Gil13, HCD^{+}21, LTL19, Moh13a, RP13, RCP^{+}16, SM20a, Sch13, SCW^{+}23, Whi13, WTB16]. Evidence [DT19, KDS20, LMM15, CPvV^{+}11, COC23, PK18, PC19, PPB^{+}14, SM20b, ML14, PK19]. evolution [ABB^{+}12, CHJCK18, HS09, LL09, NS17, PBS^{+}23, PDM19, RSH12, vDDS^{+}09]. evolutionary [EFDS20, LNC^{+}19]. evolving [PMMS16, WYH^{+}14, XKS15]. Exact [FJK10, MG22, Ros12, STD13]. examine [SFGLR15]. examiner [HS14]. Examining [Tal15]. example [CWE18, ZCD^{+}20]. Examples [DH11, FRL18]. exceedance [FS13b]. excess [KAGK^{+}23]. excessive [WCL23]. exchange [DMA19]. exchanges [KZS23]. exciting [CT18, CD18, PW12]. exclusion [GM16]. excursion [FTE^{+}21]. exhibiting [ZMC^{+}21]. Exit [GQ10]. exome [CJM^{+}17]. exon [RJP16]. exoplanet [HCKFZ21, JSF^{+}22]. expand [RHC23]. expansion [RHC23]. expectancy [LR21, Sha16]. expected [HHH10a]. expensive [BJ12]. experience [KÓ14a]. experienced [YSG16]. experiment [AS17, CDF^{+}20, CK14, CB22, EKO22, GBH^{+}15, JMY^{+}14, JCK22, SM10, WYKH07]. Experimental [JL09, HEHM23, Lin13]. experiments [BM08, BY13, BM11, CWWW17, CA22, FT18, Fre08, GH12, JTLE22, KBH^{+}11, LBHB11, Lia19, LFMM23, OGP^{+}18, OIHH09, PACB21, SHM15, ZSR14]. expert [MAB^{+}14, SJM^{+}14]. expertise [JLLK20]. experts [FSPWWE18, GM08]. explainable [BY13]. explained [LHMN13]. explanatory [GT10]. explicit [CR13]. Exploiting [CDF^{+}18, GPRR16, MLM13, SC14, BHP10]. exploration [FFW09, PGW18]. exploratory [AICV11, BDC^{+}11, LSM15]. Exploring [WLK18, ZGLH13, NCHJ13]. exponential [ABB^{+}12, GRL^{+}13, Sch08]. exposure [BZS19, BPSC14, DLS^{+}17, GTZ^{+}21, LCZ^{+}17, NMD19, PD20, PZSW23, RGF^{+}20, RLHD21, SBJR09, SLZS08, WBK^{+}19]. exposures [FD20, KDL^{+}17, MBL^{+}17, WLG17, ZMC^{+}21]. expressed [CGFT15, FPL10]. expression [ABNG14, BTJ^{+}14, BBE^{+}18, CGT^{+}14, CCJ^{+}09, CZM10, FRBT13, HZY^{+}15, HVL14, HST19, JND12, KY07, KG11, KBFM12, LvdVvWvdW13, LT11, LZCW21, MKKN21, ML13, MGSD19, MCCW09, PLM^{+}16, QDN^{+}21, RSI16, Ros09, SZO12, SKZ14, TMPF12, WFS19, WL08a, WZ16, WL22, YLH17, YLLS21, YRY17, ZWS08, ZLD12]. Extended [FH19, LRDD22, IHJ16]. Extending [BSNP16, Hof07, HIH^{+}21, STA18]. extension [RS10, WNX^{+}23]. external [CAS20, FFR^{+}08, XLDO13]. Extracting [PSW18]. extraction [WWM^{+}14, WKG^{+}15]. extrapolation [RHR12]. extremal [GHO23, RC23b, ZHO22]. Extreme [BHR11, RCF^{+}13, REFT18, BYZ18, BD11, CCdCW18, CA23, DT23, ESF14, KN20, KPDO23, LBND13, MSSS^{+}10, RAY14, RS12, RCP^{+}16, TAC^{+}16, TETJ17, WTJ10, WD10, WTB16]. Extremes [ADE15, BD22, BdHZ08, CTB17, CDN12, DT19, DET23, MSSS^{+}10, RTB22, SYZ15]. extrinsic [FWK^{+}13].

Facebook [ZPBW^{+}18]. factor [COM22, DBTP21, FH14, GDTP23, HSH12, HSVF09, KG11, KLCM20, LCYZ23, LRHF12, LSM15, LSS^{+}12, LHH19, LW18, MBL^{+}17, ML13, MCCW09, RB10b, RLHD21, WP12, WOH23, ZWS08]. factorial [BM08, BDR16]. factorisation [PTH22]. factorization [LPH22, MM15, OP09, ZTCS20]. Factors [SRZ^{+}15, BR16, HJ18, KDS20, Mar08, MON20, SSD^{+}19, YSR22]. failure [CW10, GMMW17, HSFP11, KAvdW^{+}23, SFPS^{+}21, YHX13, YN14, ZW19]. failures [ERM15]. falling [KS17]. False [SDT08, TP11, AS10a, BYZ18, BG09, GPR^{+}22, GGQY07, JCJ20, Mur10, Sch08, Wen16, ZLS^{+}17]. falsification [KSP16]. families [Sch08, YL13]. family [GRL^{+}13, Goe11, KZS23, RJP16, SS20]. far [YKLK23]. Fast [BPS22, CJMF18, CHOK14, DLZL16, QHPD19, LHF^{+}20, LZLW14]. fatality [CdVM^{+}22]. fate [FGA09]. fault [CDM18]. FDR [NPM12, ZZD11]. feasibility [GGFG^{+}18]. feasible [BWT^{+}20]. Feature [AS10a, BTA20, CLR16, CLLR20, SzCT10, WKG^{+}15, BH11, COC23, DSC^{+}23, LMGJ15, LRI21, MMY^{+}16, WZF18, ZWZ19]. features [CBZG17, CAL^{+}23, EHKW12, MLX23, RSI16, WSK^{+}21, WT08, ZSCL23]. federal [BBB^{+}18]. feedback [MDP21]. feedlot [TFB^{+}20]. female [CGCA21]. fertility [LR20b]. Fiber [KL16, Laz16, Sch16, WLP^{+}16, WLPP16]. Fibre [HKT12, DSCS19]. Fibre-generated [HKT12]. fidelity [FGS^{+}10, KSH^{+}13, PACB21]. fiducial [WOH23]. field [BZC^{+}19, BRG08, Bro08, GKZS12, LSL^{+}15, WL08a, YHX13, YN14, ZPGO21]. field-based [LSL^{+}15]. fielding [JSW09]. fields [CDB11, DLZ10, FPLM18, FCGA^{+}13, FTE^{+}21, GREG15, HKT12, JYB16, MAE^{+}08, RF07, ZYXS16]. Fienberg [Ano18, Rub18]. filing [Kan20]. filter [DS14, KS19, MKS^{+}14, TM22]. filtering [BGC20, SM10, SJA^{+}13]. filters [DB15]. financial [AM16, HSVF09, ZW18]. find [Ben08, Bir08, Feu08a, Feu08b, Feu13, Fuc08, HW08, Ing08, Kad08, MV08, Sti08]. Finding [EHKW12, SWPN09, BBDP11, LY16, LYY13, URZF21]. fine [PD20, WBC15, XBS23]. fine-grained [WBC15]. Fingerprint [Kad18, DL09, DLM14]. fingerprinting [DSH^{+}13, LCYZ23]. fingerprints [KSP16]. Finite [FH13, HS09, SMZ21]. Finite-sample [FH13]. Fire [AABC^{+}19, LYRR22]. fire-contributed [LYRR22]. firms [AM16, PS12]. first [YLG15, dCP10]. first-price [dCP10]. Fisher [ZZH08]. Fisher-consistent [ZZH08]. fission [FPL10]. fit [BYZ18, HE15]. fitness [GRL^{+}13]. Fitting [DSH^{+}13, Mac20a, Mac20b, TDBM23, ZSP19, GCC^{+}11, ISR12, SBJR09]. Flexible [HCP^{+}17, STD13, BPS22, GBST19, HZG^{+}22a, JDP^{+}13, KDS20, LLKP18, Ros09, SMW^{+}22, SS10a]. floats [PKGG23]. flood [JSH^{+}22, RC23b]. Floristic [JDP^{+}13]. flow [Chi12, FBH23, HCRB23, JSR16, OIHH09, PS15, PLCX23]. flowering [LC10]. flows [PHCM^{+}10, PPLK18]. Flu [GV14]. fluorescence [DAL^{+}23, HFH10]. fluorescent [MVP11]. fluoride [GPZ^{+}22]. fluorophores [PGL^{+}19]. flux [WBA^{+}14]. fMRI [AMR18, AK12, BY13, KMMS13, PCJW15, VRN^{+}11, WZF^{+}13, WDL22, WYW^{+}23, ZGV^{+}16]. Focused [CWH20, PCJW15]. Focusing [HK17b]. folded [Mac20a, Mac20b]. folding [WLK18]. follow [GMM08]. follow-up [GMM08]. followed [ZGM23]. food [DBG21, GQ11, MDR10]. footwear [SM20b]. for-profit [ZPR14]. force [Bic10]. forcings [BLTV14]. forecast [AHZ23, CSS11, EJD19, HK17b, HK23b]. Forecasting [MMWH11, OHC^{+}17, SH08, ANFM09, BRG08, CPG^{+}21, CPP^{+}14, DRB21, FCPL19, HSH12, HE14, IGA22, JGC^{+}18, LG20, LR20a, LR21, Lie13, PACB21, RK22, SM15, SHSZ19, Sha16, TSY22, ZBG14]. forecasts [BC23, BWBS14, GRS23, LM10a, LM10b, PHM^{+}23, WFHZ23]. Forensic [BC09, COC23, Fie07, GM09, PK18, PK19, PC19, SDL^{+}11, SM20b, STJ^{+}07, WOH23]. forest [BZC^{+}19, GPZ^{+}22]. forestry [JWZBC19]. Forests [SML^{+}21, IKBL08, JLL^{+}19, SHSZ19]. form [KFB11]. formation [LBBM21]. formulation [Mar08, MGMB19]. Förster [HBW17]. forward [BG09, RJ11]. Forward-Lasso [RJ11]. founding [GM09]. fraction [LR20a]. frailty [ENF14, ZHJZ15]. framework [BNW08, CCD22, CDN12, DLM14, GGFG^{+}18, HHHV17, HSVF09, LC10, LG18, LGK18, NKAY10, QM23, RF07, SHAB22, SMW^{+}22, SCL^{+}13, SJHJD20, SXZ23, WFHZ23, WJT^{+}21, WCW15, WI07, YGLH08, ZNSL14, ZOZ17, Zho17b, ZLDR18, ZCH^{+}16]. France [BJS^{+}22]. Fréchet [Big13, PDM19]. free [CQ09, OGP^{+}18, SR09a, YLLS21]. Freedman [Lin13]. freeway [CCH19]. French [Gau11, LBA11, MGRG^{+}23]. frequencies [BvdB22, WS10c]. frequency [ASY09, CFMR18, CT07, DT23, GM16, GKM23, GL18, GS13, JSH^{+}22, Kip22, Ste09, SXZ23]. frequency-based [GM16]. frequency-calibrated [GL18]. frequency-severity [SXZ23]. friend [Rub18]. frisk [GRS16]. front [AZM11]. fuel [AMR16]. Fukushima [Lie19]. Full [KKMS16, VCC22]. Function [LLM20, ZMC^{+}21, CHS^{+}16, FS14, Goe14, JJRZ21, MAZM13, PD20, PGL^{+}11, RMP17, RY11, SRZ^{+}15, SL20, TVJM13, WZF^{+}13]. Function-on-function [ZMC^{+}21]. Function-on-scalar [LLM20]. Functional [AXEC18, FFJJ14, FZZW17, GSD^{+}18, HSH12, LA22, LRHF12, RD14, ZS09, dCdCAGM16, BNW08, BM11, BCJ15, BJS^{+}22, CA23, CMZ19, Chi12, CCH19, DCCP09, DL11b, FO11, FS14, GDJR20, Ger09, GLB^{+}17, GCC^{+}11, GKS17, JEK^{+}22, LZK^{+}15, LWLW15, Lie13, MMGC22, MBH^{+}11, MHC15, PSL^{+}16, RN14, Sha16, SG16, SS15a, SIS^{+}20, SXC^{+}20, Sun22, TSY22, TON20, VGH14, WNZK14, WJF^{+}15, WHLN15, WDL22, WYL^{+}22, WFC^{+}22, WBKJ22, YLH17, YSH22, YHE20, YLL12, ZW15, ZSLH23, ZSG11, ZZD22]. functional-data [YSH22]. functions [BJ12, BHR11, HHHV17, JDP^{+}13, KBH^{+}11, LCB16, REG^{+}11, SLZS08, SCA13]. Fund [MSS09]. fundamental [BE23]. Fused [HYS19]. fuselage [WYT^{+}20]. fusion [FDR16, LZTB16, LSZL22, LCZ^{+}17, SCRS^{+}20, WMT^{+}21]. future [CW13a, CW13b, CRZ13, Gil13, Moh13a, RP13, Sch13, TETJ17, Whi13].

GaGa [Ros09]. gage [FRL08]. gain [CCD22]. game [GNCS22, HGM15, HS10]. games [BLM^{+}23]. Gamma [VDP08, BWS19, Sch23, XBS23, HGRS17]. Gap [LSAR12]. GARCH [WCD23]. gas [EKW20, KOJ^{+}14, NMD19]. gasoline [CN07]. gastrointestinal [SvdLMP14]. Gaul [JLB^{+}14]. Gaussian [BJ12, BHB^{+}21, BMAF^{+}23, CKHP15, CC19, CMJZ22, DBF^{+}16, DL11b, FPLM18, FD20, FTE^{+}21, FJK10, GGMG23, GB16, HHHV17, HJ18, HCKFZ21, JGVM18, JYB16, JGC^{+}18, JLGJL12, JSF^{+}22, LS18, MRW09, MGMB19, MAL^{+}22, PKGG23, PHJ22, RKM^{+}23, RG21, SG17, SCA13, TJDE17, THN^{+}19, YL11, ZCGC21, ZBT^{+}20]. gay [GGFG^{+}18]. GDP [TM22]. gels [MPT12]. gender [DFGY23]. genders [LR20a]. Gene [JHMC16, LdGK^{+}17, LC12, TLH14, AMGG13, BBE^{+}18, BKGJ14, CGT^{+}14, CDP^{+}17, FD11, FWK^{+}13, FRBT13, HZY^{+}15, HVL14, JND12, JGVM18, Kaf12, KY07, KG11, LvdVvWvdW13, LT11, LZLW14, LZCW21, LN12, LRS12, LRS15, MBYWX19, ML13, MGSD19, MLCW13, MCCW09, NQdB^{+}07, RSI16, SZO12, SKZ14, TDBM23, TMPF12, WJF^{+}15, WHLN15, WFS19, WILW22, WL08a, WP12, WS14, WL22, YLLS21, YRY17, ZWS08, ZKY14, ZLD12]. Gene-centric [LC12]. gene-environment [AMGG13]. gene-gene [LRS15]. Gene-level [TLH14, WILW22]. Gene-proximity [JHMC16]. gene-set [NQdB^{+}07, WHLN15]. gene-specific [LN12]. gene-trait [TLH14]. genealogies [TDBM23]. general [AS17, AM07, BHB^{+}21, BvdH19, CLK^{+}12, DLZL16, LC10, LG18, Mar08, MNR14, SBS14, WZS19]. generalization [HEHM23, NECS17, NS20]. Generalized [FWGS11, LW17, WD10, WZLP20, WOH23, YD23, CFH^{+}14, CLLR20, CVF10, DLM14, DMVT23, FO11, Goe11, KGGQ15, LYY^{+}23, MHC15, NMW^{+}21, PGL^{+}11, XCCL20]. generated [BL11, HKT12]. generating [HGG13]. generation [Kaf12, SZ12, ZWW13, ZSMJ19]. generative [CBZG17]. generators [JCCG18]. genes [CCJ^{+}09, CGFT15, ET07, FPL10, GM09, JEAS09, LNC^{+}19, NPM12, PLM^{+}16, SCTV11, WP12, ZKY14, ZLDR17]. genetic [AXEC18, CFRW19, FWGS11, GHK^{+}13, GRL^{+}13, HVL14, LLR10, LLR15, LBL20, MRW09, NKAY10, PDS13, QGM^{+}14, RN14, SS15a, SPF20, SBS14, WS14, WOK^{+}16, WSH^{+}14, XCS11, ZCRC18, ZLZB18, ZMA^{+}20, ZLJW23]. genetical [YL11]. genetically [CLK^{+}13]. genetics [CSGD16, GM09, HHK^{+}16, NKAY10, PG14, SGLB10, WLA^{+}21]. Genome [Hua19, BDL^{+}16, GS11, HY14, JLL09, JHMC16, LZLW14, LWLW15, LX18, SSZT19, WZF18, ZLS^{+}17, Zho17b, ZS17]. Genome-wide [Hua19, BDL^{+}16, GS11, HY14, JHMC16, LZLW14, LWLW15, LX18, SSZT19, ZLS^{+}17, Zho17b, ZS17]. genomic [BBB^{+}10, CCH^{+}21, ST14, TDS^{+}14, WMGB23, WP12, Wen16, ZD13, ZMA^{+}19]. genomics [BKS21, LL10, PZB^{+}10, SGCW07, YL11, ZYXS16, ZOZ17]. genotype [SPI^{+}23, WS10c, WGL^{+}18b, ZW12]. geo [CA22]. geographic [CJMF18]. geographies [WR12]. geometric [SK22, SW10]. Geostatistical [Fin13, RTB^{+}21]. geostrophic [ZBG14]. geothermal [BB11]. geriatric [LHH19]. germ [VKG12]. Germany [KKLS15, KKLS16]. gestation [SML^{+}21]. get [FDH10]. Ghats [MBD11]. Gibbs [CMR15]. Gibbs-plaid [CMR15]. GIS [GPZ^{+}22]. GIS-informed [GPZ^{+}22]. given [CDN12]. Glasgow [LL16]. glass [PK18, PK19, PC19]. Global [AN14, CS13, CGCA21, AZC^{+}17, BL11, FC20, JCCG18, JS08, KAGK^{+}23, LdGK^{+}17, PKGG23, RDH^{+}20, SPH17, Ste07]. global-local [LdGK^{+}17, RDH^{+}20]. glucose [CFH^{+}14]. glucose-insulin [CFH^{+}14]. GMM [CDM18]. gold [KDH^{+}13, WZ18]. Good [PS12]. Goodness [HE15, BYZ18]. goodness-of-fit [BYZ18]. Google [GV14]. government [BBB^{+}18, Sun22]. gPCA [Fuk19]. GPS [CD20]. grade [DZ23]. gradient [CLK^{+}12, HIH^{+}21, KB10, LYBA22]. gradients [QBC13]. grained [WBC15]. grants [LMM15]. Graph [KKL23, PTH22, DH18, JND12, LLR10, LL10, Pur11, ZPBW^{+}18]. Graph-aware [KKL23]. graph-structured [JND12, LL10]. Graphical [MGSD19, BTJ^{+}14, BDL^{+}16, CA23, CGN22, DL11b, FD11, GCL^{+}15, LNR^{+}22, LZP16, NKAY10, PCJW15, SM13, SCV^{+}10, TJDE17, VVSK18, WMGB23, YLLS21, YL11, ZOZ17, ZRCC21]. graphs [CC19, DGL13, HHLC16, KOB^{+}20, MRV10, RKM^{+}23, XDO10]. Gravitational [KBB^{+}11]. gravitationally [TMvD^{+}17]. great [PS12, GKP^{+}16]. GREAT08 [BGH^{+}09]. GREAT10 [KBB^{+}11]. green [KZ16]. grid [SM15, dCP10]. gridded [MSSS^{+}10]. Ground [DT19, MK21, RCP^{+}16]. Ground-level [DT19, MK21]. Group [SW17, KX12, LWLW15, PHWM11, RLHD21, Sha16, YKHS21, ZHM^{+}19]. grouped [JL10, SP13]. grouping [LZW^{+}15]. groups [CGC12, MZ12, RGT13]. growing [ZPMA10]. growth [ACS^{+}23, BZS19, HHHV17, HJ18, PMQW14, PPB11, PPM14, VCC22, ZW15, ZCG^{+}09]. guaranteed [DSC^{+}23]. Guerry [DJ11]. guessing [HGM15]. guided [KX12, SzCT10]. guidelines [LQNM19]. gun [BMAF^{+}23]. gunfire [HJS22]. gut [SXC^{+}20, VML^{+}21]. GWAS [BDL^{+}16, CLLR20, SWLS14].

H1N1 [PPB^{+}14]. H7 [TFB^{+}20]. Haar [WNZK14]. habitat [SRC15]. hand [Sta23]. hand-audited [Sta23]. Handbook [BGH^{+}09, KBB^{+}11]. Handling [CAL^{+}23, Ger09, PKP16, SRH16]. handwriting [COC23]. haplotype [SX09]. hard [MZ12]. hard-to-reach [MZ12]. Hardy [ZSS23]. harmful [VFMD17]. harmonic [HAFFH21]. harvest [Mei10]. hate [SM20a]. Hawkes [FSG16, HJS22, MMBL20, WWCZ22]. hazard [LCZ^{+}17, MR15, OBHL22, TVJM13, XS11]. hazards [FSM^{+}19, PZSW23, SW17, WLG17, WK10, ZHJZ15]. HDP [FSJW11]. HDP-HMM [FSJW11]. Health [GMB15, AMB^{+}20, ÇL12, FND09, GSC^{+}20, HCP^{+}17, JFRS17, JCK22, JG23, KGGQ15, Lee18, NH19, PD20, PHJ22, PBSVS23, RGF^{+}20, SBJR09, SPS20, SBD23, Tal13, Tal15, TACH21, WSM^{+}16, WLML23, WHC^{+}22, ZMB23]. healthcare [Hat14, Hav14, NSS14a, NSS14b, Pad14, SPH17, SXZ23, Wal14]. healthy [HSFP11]. heaping [CWS15]. heart [KH13]. heat [BM22, PKGG23, SRCK16]. heavy [PS12, VDP08]. heavy-tailed [PS12, VDP08]. height [AYJ^{+}09, RAY14]. heights [TETJ17]. hematopoiesis [XKG^{+}19]. hematopoietic [FGA09, PBS^{+}23]. hemodynamic [WZF^{+}13]. hemoglobin [SPS20]. hemolysis [VGH14]. hereditary [GPBT22]. heritability [SSZT19]. Hermite [HCKFZ21]. heterogeneities [YHX13]. Heterogeneity [FS13a, FDH10, HHN^{+}20, HYS19, IR13, LMGJ15, MBDL14, MBK^{+}21, SKAL19, WHAW21, XZX18, ZPR14]. Heterogeneous [BSDG22, AV15, BBM20, BHR11, FH19, FMA18, GCL^{+}15, HZL^{+}15, JCK22, KH23, LG18, LCRM21, LW18, SML^{+}21, WYL^{+}22, WS14, WL22, ZLR19]. heterosexual [AH16]. heteroskedastic [JGC^{+}18, ZGJ^{+}22]. Hi [WSU^{+}19]. Hi-C [WSU^{+}19]. Hidden [RGT13, SPPR08, SSL^{+}10, AM07, BvdH09, CLFC23, CGFT15, DLS^{+}17, FGA09, FCGA^{+}13, HAFFH21, JAZ15, JLL09, LDV^{+}10, LNC^{+}19, PGL^{+}19, SKAL19, SM13, VCC22, WYH^{+}14, WCH^{+}23, WL08a, ZW07, DMN20]. Hierarchical [BM22, CB10, DL09, DTZP13, FBM09, Gau11, LPKP22, LHH19, SRC15, VH14, WKR21, YH20, Ané08, BST15, BB11, CGT^{+}14, CGCA21, COC23, DMGJ20, DB22, EFDS20, FLRZ23, FCC15, FLP23, HGM15, HST19, JSW09, JWL^{+}19, JYB16, KNWJ14, LSS^{+}20, LSY^{+}22, LWZ19, LSS^{+}12, LN12, LZ11, MLP^{+}19, MRM12, MNR14, PNB22, RS12, SJM^{+}14, SP13, SGC23, SG16, SX09, SM20b, DPT22, TTH21, TL11, VBK19, WFHZ23, WYKH07, YOZC23, ZSP19, ZHM^{+}19, ZCH^{+}16]. High [ASY09, BDR16, OBHL22, PACB21, SGLB10, YFM19, AWL13, ACG13, BC09, BW18, BHW15, DBTP21, FDKP13, FCPL19, GSD^{+}18, GS13, HHLC16, HH21, HKP23, LLR09, LBHB11, LZW^{+}15, LS22, LWFW16, LPT^{+}11, MAB^{+}14, MDR10, OW11, PGW18, PCJW15, RGSB^{+}18, SFDMM22, SPPR08, SWPN09, SS15a, SCL^{+}13, SM13, Ste09, VML^{+}21, WKLvD16, WL10, XBS23, ZCS13, ZYC^{+}17, ZLDR17, ZW18, ZGS^{+}14, ZPR14]. high-dimensional [AWL13, ACG13, BW18, BHW15, HHLC16, LLR09, LZW^{+}15, LWFW16, MAB^{+}14, PGW18, PCJW15, RGSB^{+}18, SM13, WL10, XBS23, ZCS13, ZYC^{+}17, ZLDR17, ZW18, ZGS^{+}14]. high-energy [WKLvD16]. High-fidelity [PACB21]. high-frequency [Ste09]. high-income [HKP23]. high-order [VML^{+}21]. High-resolution [OBHL22, FCPL19, HH21, SFDMM22]. High-throughput [SGLB10, BC09, DBTP21, LBHB11, LS22, SPPR08, SS15a]. Higher [Kip22, WSH^{+}14]. highly [VFH16]. histogram [ZMLS22]. histologic [CCH^{+}21]. histone [CLZ16]. historical [BZC^{+}19, CW13a, CW13b, CRZ13, Gil13, JSR16, Moh13a, NZRC13, RP13, Sch13, Whi13, ZCD^{+}20]. history [FZSI^{+}08, LNC^{+}19, SH11]. HIV [AXEC18, ARC07, BNMG23, Bro09, CPvV^{+}11, DPR^{+}20, DMVT23, DGCT10, HL08, HWHWA11, JLS^{+}17, JD18, KH23, KM17, LBK^{+}23, LNR^{+}22, LMW10, PRRW11, SLBL23, STD13, SH11, YWQG23, ZW19]. HIV/AIDS [Bro09, JLS^{+}17, ZW19]. HMM [FSJW11]. hmmSeq [CGFT15]. hockey [TVJM13]. holding [MHH17]. hole [LBBM21]. home [STG21]. home-delivered [STG21]. homeless [GBC21, KB10]. homelessness [GF19]. homes [BLM09]. homicide [Sad14]. homocysteine [Ros18]. homogeneous [WCH^{+}23]. homology [BMM^{+}16, MLX23]. horizontal [JGVM18]. hormone [YSR21]. horseshoe [DAL^{+}23, NV18]. Horseshoes [DGH08]. Hospital [KBMF^{+}23, CD17]. hospitalization [QBC13]. hospitals [CD17]. hotspots [GPR^{+}22, HH21, RDH^{+}20]. hourly [DLZ10]. House [YR21, NBZ11, PPM14, DPT22]. Household [BHW15, ML14, PHLH11, STA18, SCK19]. households [Gau11, MGM^{+}14]. housing [RFB17, VIF13]. human [CD20, CAV^{+}19, DM18, HRP10, JMY^{+}14, LZ21, LSL^{+}15, LPKP22, Sad18, SL20, VML^{+}21, WJL16, YZAD13, ZSLH23]. hunting [XDM15]. Huntington [GM16, GMMW17]. hurdle [CFW17, PW12]. hurricane [LLKP18, PACB21, RF07, SKKS14, XKS15]. hybrid [MAE^{+}08, SSH^{+}11, ZLZB18]. hydrodynamics [GBH^{+}15]. hypergraphs [JWZBC19]. hyperparameter [PLM^{+}16]. hypertensive [FLP23]. hypervariable [PLM^{+}16]. hypotheses [Hua19, JEAS09]. Hypothesis [BW18, GLB^{+}17, ST14, BFF^{+}09, Efr08, PG14, SZO12].

I-55 [PS15]. I/II [YY11]. IA [GFW^{+}09]. iBAG [ZMA^{+}19]. Ice [CKK^{+}22, KÓ14a, CHAP16, DRB21, LHF^{+}20, TVJM13]. ideal [TM22]. Identification [KH23, SHW18, AABC^{+}19, CC19, FKSBS19, HCW11, JD18, KHDV20, KHDV22, LMS10, LES12, LGL^{+}12, MG22, Ryu22, SF11, SAV^{+}14, WYT^{+}20, WWL22, WOH23, ZMC^{+}21, ZWZ19, ZZ22, ZBT^{+}20]. identifications [GPRZ17]. identify [GM16, LKTJ^{+}15, NQdB^{+}07, SCRS^{+}20, SSD^{+}19, WP12]. Identifying [CCH19, FD20, HAFFH21, HZL^{+}15, MZA22, RVW20, YFM19, BZC^{+}19, FPL10, JPTO17, LBD^{+}18a, PZB^{+}10, VFMD17, WSK^{+}21, WYW^{+}23, WWM^{+}14, YZAD13]. identity [ARK^{+}18, BvdH19, SWLS14]. identity-by-descent [SWLS14]. ignition [UH20]. II [Fie10b, YY11]. illustrated [RHHH13]. illustration [CCJ^{+}09]. image [BGH^{+}09, CBZG17, DLM14, KW23, LZW^{+}15, LGK18, MBH^{+}11, WNZK14, ZJLC08]. imagery [AYJ^{+}09, MVW^{+}23]. images [CCH^{+}21, CHH^{+}14, LS18, LBD^{+}18a, LWLX19, MLX23, RMS^{+}19, VRN^{+}11, YWL^{+}12]. imaging [BM11, BHP10, CSGD16, DAL^{+}23, DKZ09, GDG^{+}16, MZI18, MVP11, MBK^{+}21, OW11, PSW18, PHT15, WBKJ22, WKG^{+}15, ZGS^{+}14]. imbalance [GBST19, YSR21]. imbalanced [MKM23, WLML23]. immune [DHG19, KH23]. immunization [DW21]. Impact [DZ23, KHLB22, BGK^{+}15, HJ18, HKP23, Lee18, STA18, Sun22, WCW^{+}22]. impacts [FGMP16]. impending [ZBT^{+}20]. imperfect [BSDG22, MBDL14]. Implications [HS13, RCBB19, YHE20]. importance [CP20, LKB21, LN12, RTB^{+}21]. important [Bic10]. impressions [HHC17]. imprinting [YL13]. improve [HK23b, HEHM23, HGSJ23, JD18, SRZ^{+}15, SCRS^{+}20, WHLN15]. Improved [BD22, PMMS16, RJ11, AMR18, ZBG14]. improves [PLM^{+}16]. Improving [BHP10, BDR16, CHAP16, CT07, JSF^{+}22, Loh09, ZBLC17, ZW12, KY07, LHH19]. Imputation [LQNM19, NSMM23, TDS^{+}14, AT10, Gho10, GSC^{+}20, GG19, GPZ^{+}22, KFB11, LYY^{+}23, PSD13, PKP16, QM23, STG21, SPS20, SHC12, TACH21, WR12, ZLOS10, ZZD22]. imputation-based [Gho10]. impute [WS10c]. In-game [GNCS22]. In-season [Bro08]. incentives [Cha17]. incidence [JGC^{+}18]. income [HKP23, KKLS15, KKLS16]. Incompatibility [McE09]. incompatible [CDF^{+}18]. incomplete [BHR11, CWWW17, DPR^{+}20, GKS17, LWFW16, YN14, ZCD^{+}20]. incorporate [ZHM^{+}19]. Incorporating [HHN^{+}20, SCTV11, XLDO13, XLS^{+}19, ZBG14, HWHWA11, MBK^{+}21, Wen16]. increase [ML14]. increases [SWLS14]. increasing [DT19]. incremental [CA22, RODC19, ZBLC17]. indel [LL09]. independent [AY12, Efr09, JL19, NPM12, SG16, ZST16]. index [CMPR22, GPRZ17, OIHH09, RFB17, RWK17, Tal13, Tal15, YHE20]. indexed [GKZS12]. Indian [Gil17, MSS09]. indicator [ZHM^{+}19]. indicators [MRSA19, MNR14]. indices [SHAB22]. indirect [CAS20, MGRG^{+}23]. individual [CCH^{+}21, CJMF18, EFJ07, GQ10, LHMN13, MC17, MBDL14, SLBL23, WZD19, WFS19, WHAW21]. individual-level [EFJ07, GQ10]. individuality [DL09, DLM14]. Individualized [SKBL23, DMVT23, SHAB22]. individuals [FWGS11, SWLS14]. inducing [AMB^{+}20]. industrial [Gho10]. inequality [Gau11, KKLS15, KKLS16, LL16]. inexperienced [YSG16]. infant [BZS19, SXC^{+}20]. infants [SvdLMP14, YLS14]. infection [CdVM^{+}22, CGCN22, CPvV^{+}11, DGCT10, KH23, LMW10, SvdLMP14]. infections [CD17]. infectious [ENF14, HCS18, MH14, RTB^{+}21, YLH07, YGLH08]. Inference [BBG^{+}12, BGM17, CGCN22, FH09, GFW^{+}09, HR22, KK12, KM17, Mey08, SDHZ14, ZRCC21, BPS22, BHC^{+}20, BHB^{+}21, BD22, BCR^{+}19, BBB^{+}10, BKVW22, CGT^{+}14, CJMF18, CR13, CWE18, CA22, CCJ^{+}09, CMR18, CD17, DLL^{+}18, DB22, DHG19, Efr08, Efr12, ESO19, FGA09, FRL18, FGS^{+}10, HHN^{+}20, Hat14, Hav14, HLY^{+}21, HS13, HCS18, HBW17, HFH10, JLDQ10, JJRZ21, JFRS17, JSH^{+}22, JWZBC19, JG23, KSD11, KRC23, KF10, KY07, LHPW13, LL19, LQNM19, LL09, LZ11, MSJ14, MON20, NKAY10, NZRC13, NSS14a, NSS14b, OM12, Pad14, PZSW23, RFWE22, RCLWW10, RMS^{+}19, RHZ^{+}15, Rub08, STMC17, Sch08, STG21, SPF20, SCV^{+}10, SJ11, TJDE17, TACH21, TAC^{+}16, TFB^{+}20, TB22, VCC22, Wal14, WTCW10, WDSJ23, WBKJ22, WIC^{+}10, WBK^{+}19, XKG^{+}19, ZPMA10, ZK10, ZYC^{+}17, ZNB^{+}21, ZHYS23, ZZXL23, ZCH^{+}16]. inferences [GM09]. inferred [CLK^{+}13]. Inferring [BK20, BHR11, BGK^{+}15, DBG21, LSM15, RSD22, SHM15, WJF^{+}15, ZBC16, HGM15, HSSF21, LNR^{+}22, MVP11, SBS14, XZX18]. infinite [LHH19]. inflated [BvdH19, CVF10, CBvdHvdH08, LZCW21, LMKC12, MGSD19, NvdBCR23, RBF^{+}20, WL22, XCCL20, ZCM^{+}11]. inflation [FLHA15, JB21, KZS23]. Inflection [GBC21]. influence [AMR16, DCHP21, MM15]. Influencing [RLH^{+}13, SRZ^{+}15]. influential [CLZ09]. influenza [BWT^{+}20, OHC^{+}17, PHLH12, PPB^{+}14, TSS10]. Information [VML^{+}21, CT18, CCS18, FDR16, HE14, HLK18, JL19, KZ16, LCG09, Lia19, PCJW15, RKM^{+}23, RC23a, RRSM18, SHGA10, SzCT10, SCTV11, XLDO13, YSR21, ZW12, ZBG14]. informative [BK21, GJPS08, MKN22, PHJ22, RB10b, TFB14, WMA^{+}14]. informed [CGN22, GPZ^{+}22, PQR21]. infra [SCDG17]. infra-marginality [SCDG17]. infrastructure [HAL21]. infrequent [SH11]. inhibitory [JLA16, ZCG^{+}09]. Inhomogeneous [DAAM22, FSG16, SH08]. Initiative [GMB15]. initiatives [STA18]. inject [LBK^{+}23]. INLA [ISR12]. innovation [GM15]. inputs [JLRK23]. Insights [SFDMM22]. instantaneous [WSGH12]. institutionalization [GMM08]. instrument [KM16, SVYP11]. Instrumental [ENH^{+}18, HYL23, JCK22, KKMS16, SGNM22, YLS14, YKLK23]. instrumented [HAL21]. instruments [ENH^{+}18, Fin13, KM16, ZSG^{+}13]. insulin [CFH^{+}14]. insurance [JFRS17, JCK22, PHJ22, REFT18, SFB16, SZ20]. intake [DBG21]. Integer [KSP16, Haz15, LLKP18, ZSG^{+}13]. integer-valued [Haz15, LLKP18]. integrated [CKHP15, ISR12, JWL^{+}19, LHMN13, WKR21, WILW22]. Integrating [RTB^{+}21, WESVS23, BPSC14, ZSMJ20]. integration [BTA20, HBP17, JCS07, RG21, WMGB23, ZOZ17]. Integrative [KBFM12, PGW18, RGF^{+}20, WACY20, XZW21, HS22, MCCW09, PZB^{+}10, SWM13, ZMA^{+}19]. intellectual [Bic10]. intelligence [LZ21]. intelligent [SP19]. Intensity [SK22, DT23, FRL08]. intensive [DFN08]. inter [HS14]. inter- [HS14]. interaction [AMGG13, Fin13, LC12, LWLX19, LRS12, LRS15, SPI^{+}23, ZSCL23, dGC07]. interactions [BST15, EFDS20, FD20, LZLW14, ML13, PQR21, WJF^{+}15, WS14, WWCZ22]. interest [HK17b]. interference [AS17, AM16, BHC^{+}20, LBK^{+}23, SDHZ14]. intergenerational [KZS23, MZA22]. interlocking [FT18]. intermediate [SCW^{+}23]. intermittent [SRA^{+}15, SS15b]. international [AR18, HMT12, WH11]. Internet [PHCM^{+}10, SJ11]. Interpolating [MAE^{+}08]. Interpolation [GS13, PKGG23, Ste09]. Interpretable [LRMM15, BT11, PT12, SKBL23]. Interpretation [dGC07, CWE18, DSC^{+}23, MMGR21, SW10]. Interpreting [SGL^{+}08, FBH23]. intersection [NPM12]. interstate [PS15]. interstellar [MAL^{+}22]. interval [HCP^{+}17, JLDQ10, KAvdW^{+}23, SML^{+}21, SH11]. interval-censored [HCP^{+}17, SH11]. intervals [SYZ11]. Intervention [vdBR10, HYS19, JL09, PPLK18, SGNM22]. interventional [CC19, LZ11, SHM15]. interventions [LBK^{+}23, SP13, Sun22]. intra [HS14]. intra-examiner [HS14]. intracellular [MN15, MHB^{+}09]. intrinsic [FWK^{+}13, SFDMM22]. Introducing [DJ11, New09]. Introduction [Fie10a, Fie10b, LRS09, Sil16]. Invariant [vdHWC^{+}12, Tal13]. inverse [LPT^{+}11, MKN22, NWJ20, SWH22, TACH21, THN^{+}19, YE14, ZCS13, ZKS15]. inverse-probability [TACH21]. investigate [RCP^{+}16]. Investigating [HMT12, SG16, WCH^{+}23, WG23]. ion [MHB^{+}09]. ionosphere [GKS17]. ionospheric [FPLM18, GKZS12]. iQRAT [WILW22]. Ireland [PL11]. Irish [NMW^{+}21]. irregular [RAY14, ZRA^{+}21]. irregularly [Ger09, LCB16, LCSZ15, MLKQ22]. isoform [LZZL18]. Isolation [ZSR14]. isomorphic [MBH^{+}11]. isotonic [LRS12, LRS15, QGM^{+}14]. isotropy [FDKP13]. issues [Lyo08]. Italia [ZST16]. Italian [DPR11, DPT22]. Italy [FS13a]. item [AZM11, WBB13]. iTRAQ [MAZM13].

Jensen [YHE20]. Jesus [Fie08]. JFK [STJ^{+}07]. JIVE [LHMN13]. job [MLM13, PG13]. Joining [CDF^{+}18, ZBC16]. Joinpoint [MBGDS11]. Joint [BLM^{+}23, CCH^{+}21, FCC15, FHSJ14, HS22, HVL14, Hua18, LZZL18, LHMN13, NH19, OKGM14, SGC23, SH11, YLG15, ZW19, CGCN22, LTL19, MBYWX19, MSH21, MDWH21, MKM23, NKAY10, PC20, WJT^{+}21, WP12, YWQG23, Far22]. jointly [AS10b, LR20a, vdBN09]. Jolly [WHAW21]. journalism [RHHH13]. JTPA [HEHM23]. Jump [KGGQ15, Kan20]. junction [GZB^{+}11]. juvenile [BSLL10].

Kernel [RZC^{+}18, WHC^{+}22, DGH08, FCPL19, LC12, SML^{+}11, ZK10, ZM16]. Kernel-penalized [RZC^{+}18]. key [DPR^{+}20, ZB11]. killings [KN13]. kinetic [CGCN22]. kinetics [OMM^{+}14]. Kingdom [CHJCK18]. Kingman [CP20]. knockoff [KS19]. Knockoffs [RC23a]. knot [JWZBC19]. knowledge [FDH10]. Kullback [CSL^{+}08].

label [OGP^{+}18]. label-free [OGP^{+}18]. labeled [MVW^{+}23, SLBL23]. labeling [CWWW17, MAZM13]. labeling-based [CWWW17]. labels [WBC15]. laboratory [DHM^{+}17]. lack [HS13]. lag [PSL^{+}16, WMT^{+}21, WHC^{+}22]. Lagged [KRC23]. lakes [SCRS^{+}20]. Lambert [Goe11, Goe14]. land [HHC17, MK21]. land-use [MK21]. landscape [HHK^{+}16, MBD11, ZW08]. landscapes [ZPGO21]. landslide [OBHL22]. LAPD [KHLB22]. Laplace [GGMG23, ISR12]. Large [JJRZ21, QTL^{+}22, ZNB^{+}21, ACG13, AR18, BHB^{+}21, BBE^{+}18, CGW^{+}10, CW13a, CW13b, CDN12, CRZ13, DBF^{+}16, DTL^{+}23, DVF13, FBM09, GGMG23, Gil13, GBH^{+}15, GS11, JL09, JMY^{+}14, JG23, LRZ08, Men18, Moh13a, MKM23, PYP^{+}09, PGW18, PDS13, RP13, SJH11, Sch13, SWPN09, SPsLC16, SKAL19, VFMD17, VHS13, Whi13, YKLK23, YBL^{+}17, ZPMA10, ZZ18, ZSP19, ZRCC21, ZS17]. Large-scale [JJRZ21, QTL^{+}22, ZNB^{+}21, DTL^{+}23, GS11, JL09, JG23, MKM23, PYP^{+}09, PDS13, SPsLC16, YBL^{+}17, ZS17]. largest [Joh09]. Lasso [MRG21, GKM23, KX12, MV12, WNRZ11, WL08b, FFW09, CGC12, LRZ08, LWLW15, RJ11, Tin11]. lassoed [WT08]. last [Ber11, CR11, DL11a, HU11, Hol11, Kap11, MM11, MW11b, NL11, Rou11, SMR11, Sme11, Tin11, WA11]. Late [DCHP21, ZSG^{+}13]. latency [PZSW23]. Latent [BLM09, DMA19, GZB^{+}11, HHK^{+}16, HTM^{+}13, JPTO17, KZS23, MZ12, STM17, XBS23, BZC^{+}19, BPSC14, BSNP16, COM22, CGW^{+}10, CSB^{+}15, DBG21, FSM17, GEC13, HSVF09, HST19, JTLE22, KXC09, MMY^{+}16, MRV10, ML13, MON20, MGM^{+}14, MCCW09, MR15, NMW^{+}21, PKGG23, RB10b, SP19, SBJR09, Sch15, SRH16, SZO12, TB22, VVSK18, XLS^{+}19]. later [DZ23]. Lateral [KN17]. lattices [LBD^{+}18a]. Law [Men18, Gil17, MH14, VC14]. layered [RSH12]. lead [CN07]. lead-based [CN07]. leaded [CN07]. leading [CCdCW18, RRSM18]. League [ZTH19, JSW09, GNCS22]. Learning [DGL13, FTE^{+}21, GHO23, MLCW13, MAM17, PC19, WBC15, BSDG22, CMJ09, FP08, HK23a, KW23, LW18, MTZZ21, NvdBCR23, PHM^{+}23, SP19, SWH22, SC23, TWA18, WMGB23, XZW21, XDO10, YWB^{+}23, SC14]. learns [Sta23]. least [ACG13, DMVT23]. Lecar [DS14]. Lee [DPR11]. left [GC08, Gel13, HMM09, HCP^{+}17, YWQG23]. left-censored [YWQG23]. legal [RY11]. legislative [MM15]. legislator [LRM17]. Leibler [CSL^{+}08]. Length [MGTZ21]. Length-biased [MGTZ21]. lensed [TMvD^{+}17]. lenses [CFMR18]. Lensing [KBB^{+}11, BGH^{+}09]. Leo [Bic10, Büh10, Cut10, Fel10, FY10, Fri10, Jor10, Ols10, SBW^{+}09, Sto10, Yu10]. lesbian [GGFG^{+}18]. lesions [GMLB^{+}14, PSL^{+}16]. Letter [Fio08, HTL13, Mac20a, Mac09, SST^{+}08, SH15]. leukemia [YLH17]. Level [BR16, CKHP15, DT19, EFJ07, GBC21, GQ10, JL09, LBD^{+}18a, MRW09, MK21, OMM^{+}14, PHJ22, RHC23, RCP^{+}16, SLBL23, TLH14, WILW22, WI07, WZ16, WJL16, ZGLH13, ZHJZ15]. Level-screening [BR16]. levels [BvdB22, BR16, CAL^{+}23, PD20, Ros18, VGH14, WZ16, YLH17]. Leveraging [HEHM23, SWLS14, ZSS23]. liability [CSB^{+}15]. Lie [HHN^{+}20]. lies [CWS15]. life [Bic10, DZ23, HSD^{+}22, HMM09, LR21, OMM^{+}14, RDL23, Sha16, ZSG14]. lifetime [AL16, HMM09, ZSG11]. ligands [ABB^{+}12]. light [DAL^{+}23, TMvD^{+}17]. lightning [GGMG23]. Likelihood [HFH10, NZRC13, Ryu22, AYJ^{+}09, BvdH19, BJ09, CBZG17, FYB^{+}15, FIM^{+}21, FDH10, HCS18, Hof07, KMKB16, LK13, LES12, SvdLMP14, SIL^{+}11, SKS10, SBSH18, TFB14, YL13]. Likelihood-based [Ryu22, HCS18, LES12]. limit [SVYP11]. limitation [SS10b]. limits [CQ09, ELD09, HHH10b]. line [JL10]. lineage [FIM^{+}21, XKG^{+}19]. linear [AD22, CFLP15, CWH20, DAAM22, FWK^{+}13, FPC20a, HHH10b, LvdVvWvdW13, LZK^{+}15, LYY^{+}23, LS22, PSD13, PGL^{+}11, PDS13, PSL^{+}16, SCTV11, TDBM23, TSG17, UH20, WEWX21, WS10c, XCCL20, ZCS13]. lines [BC09]. linguistic [GGG^{+}12]. linguistics [SDP22]. link [JDP^{+}13, PGL^{+}11, PTH22]. linkage [Sad18, STG21, TL11, XLS^{+}19, ZW12, ZX22]. linked [LLZ19, LPH22, PT12, SZ20, WNX^{+}23]. Linking [CHS^{+}16, PBSVS23, SXC^{+}20, WSGH12, CLZ16]. links [YLC^{+}17]. Liquid [KPA^{+}10]. liquidity [ASY09]. list [BK20]. lists [CDF^{+}18, NPM12]. Little [KÓ14a]. liver [GDJR20]. loads [GKP^{+}16, LBND13]. Local [GRL^{+}13, SML^{+}11, YZAD13, DGL13, DGH08, GPR^{+}22, JCJ20, LdGK^{+}17, PD20, PKGG23, PL08, RDH^{+}20, SWLS14, SJGM13, SGNM22, ZLS^{+}17]. localization [AFS07, DSC^{+}23, DGM^{+}08, GMNP^{+}21, LLR09]. localized [RFB17]. Locally [DD16, Lee18]. location [CPP^{+}14, HFH10, ZBT^{+}20]. location-mixture [CPP^{+}14]. locations [FS13b, HGSJ23]. locus [FYB^{+}15, WG23]. Log [GGMG23, SXC^{+}20, AD22, CFLP15, SG17, TSG17, XCCL20, XBS23]. Log-contrast [SXC^{+}20]. log-gamma [XBS23]. Log-Gaussian [GGMG23]. log-linear [CFLP15, TSG17, XCCL20]. Logistic [HPF13, FW21, GJPS08, LHH10, MGMB19, MDGM21, NDRF17, RLH^{+}13, SLBL23]. logistic-Gaussian [MGMB19]. logit [PGL^{+}11]. Loglinear [DM18, DVA^{+}19, KKL11]. lognormal [RCLWW10, SRC15]. logs [RHHH13]. London [BMLG21]. long [BLTV14, COD22, CLFC23, HL08, KFB11, RDL23, ZASM12, ZGM23]. long-memory [BLTV14, ZGM23]. long-term [HL08]. Longitudinal [ZGS^{+}14, AS10b, AV15, BLM09, CLR16, CH14, CWS15, FAL^{+}10, GDG^{+}16, Hof15, HL08, HYS19, JEK^{+}22, KK13, LT12, LZ13, ML23, MKN22, PGL^{+}11, QM23, RDL23, SBJR09, SvdLMP14, Sco09, SIL^{+}11, SHC12, SH11, TMY17, WYL^{+}22, WNX^{+}23, WH11, YLH17, YLG15, YWQG23, ZRA^{+}21, ZY12, ZASM12, ZZXL23, BHW15, MV14]. longitudinally [LNR^{+}22]. look [YSG16]. loss [DLZL16]. losses [ZZH08]. lost [Fie08]. lots [STJ^{+}07]. low [FH09, FSM17, HH21, PD20, TAC^{+}16, WG23]. low-rank [FH09, HH21, WG23]. loyalty [JY10]. Luce [CTM14]. lucky [PS12]. luminosity [HLY^{+}21]. lung [CPP^{+}14, CHS^{+}16, HZG^{+}22a, JL11, MLX23, ZLR20]. lurking [SH18]. LWR [PS15]. Lymphangiogenesis [FCC15]. lymphoma [BBE^{+}18].

Machine [Deb09, BSDG22, LC12, PA23, PHM^{+}23, WHC^{+}22]. Macroeconomic [IWG13, ZCP14]. magnetic [BM11, GKZS12, WBKJ22, ZJLC08]. magnetization [PSL^{+}16]. magnetoencephalography [LPT^{+}11, SJA^{+}13, YE14, YFHE20]. magnetosphere [KSH^{+}13]. magnitude [ZHO22]. Mahalanobis [LW17]. main [FD20]. major [JSW09, LLKP18, ZTH19]. make [NZRC13]. making [FZSI^{+}08, RCBB19, SC14, ZZ22]. malaria [KKMS16]. Malawi [JD18]. MALDI [HCW11]. MALDI-TOF [HCW11]. mallards [WHAW21]. Mallows [CAV^{+}19]. malware [SAV^{+}14]. mammals [FGS^{+}10, LZTB16]. management [HE14, MHG18, SH08]. Manifold [SDP22, VBK19]. manifolds [JFM11, LAS16, SKKS14]. manufacturing [SH18]. Many [FZSI^{+}08, MGRG^{+}23, Bic10, BMT13, BR16, CAL^{+}23, ENH^{+}18, Far22, JEAS09, STA18]. many-household [STA18]. Many-to-One [MGRG^{+}23]. map [Deb09]. Mapping [MAL^{+}22, AFS07, BL11, FFM^{+}21, FCGA^{+}13, KX12, LCZ^{+}17, MHH17, MSH21, OMM^{+}14, OBHL22, ZZL11]. maps [RTB^{+}21, SGL^{+}08, SDT08, WDL22]. Marginal [ZASM12, CMPR22, PGL^{+}11, SYZ15]. marginality [SCDG17]. Marginally [HK23a]. marine [CFW17, LZTB16]. mark [CA18, DSCS19, LK13, LWLX19]. mark-recapture [CA18]. Marked [CT18, CR13, CDB11, XKS15]. marker [CZM10, HYL23, KXC09, LL09]. markers [GM16, HPF13, MPT12]. market [ASY09, Fin13, GH22, HSVF09]. marketing [FLRZ23, MKM23]. markets [CCdCW18, FFJJ14]. Markov [AM07, BLM09, BZC^{+}19, BSNP16, CFMR18, CLFC23, CGFT15, DLS^{+}17, FCGA^{+}13, FSPWWE18, GREG15, HAFFH21, HHN^{+}20, HS09, HGRS17, JAZ15, JYB16, JLL09, JPS21, KRC23, LDV^{+}10, LNC^{+}19, LSL^{+}15, MMGR21, PGL^{+}19, RGT13, SB20, SPPR08, SRCK16, SSL^{+}10, SM13, TB22, VCC22, WYH^{+}14, WCH^{+}23, WL08a, WWCZ22, YGAT20, ZPGO21, ZW07, DMN20]. Markov-modulated [WWCZ22]. Markov-switching [CFMR18, MMGR21, SRCK16]. marks [KPDO23]. marrow [AL16]. Mars [BML^{+}20]. masking [ZDL10]. mass [CMPR22, FZSI^{+}08, GPRZ17, HCW11, KPA^{+}10, KOJ^{+}14, LMS10, LES12, LLM20, MAZM13, OGP^{+}18, PHT15, Ryu22, WKG^{+}15]. massive [GB16, LSAR12, TSS10]. massively [BLM^{+}23]. master [PZB^{+}10]. mastery [SP19, SEX21]. matched [KZ16]. Matching [MPT12, ZPR14, CAS20, CDB11, HHC17, HGSJ23, JWZBC19, KKMS16, KM16, PPLK18, SDH18, YKLK23, ZMB23]. matching-adjusted [CAS20]. material [DH10]. materials [CLEB14, DHM^{+}17]. maternal [AZC^{+}17, FLP23, LL16, YL13, YKLK23]. Mathematics [Gne12]. matrices [AR18, DKZ09, FH09, HGG13, LRZ08, PDM19, ZLDR17, ZW18]. Matrix [SHR^{+}22, BBE^{+}18, JPTO17, LCYZ23, LS22, LPH22, MM15, OP09, PTH22, Pur11, ZTCS20, ZCH^{+}16]. matter [DBF^{+}16, PYP^{+}09, PD20, XLS^{+}19, XBS23, YZS^{+}13]. max [RS12, SYZ15]. max-stable [RS12, SYZ15]. maxima [SJGM13, ZHO22]. maximal [RRSM18]. Maximin [KMMS13]. maximin-efficient [KMMS13]. Maximizing [KZ16]. Maximum [AYJ^{+}09, BJ09, CBZG17, LK13, SKS10, TFB14, CMAC^{+}23, FIM^{+}21, FDH10, KKR13, SvdLMP14, SBSH18]. maximum-likelihood [FIM^{+}21]. MCAR [ZHB09]. MCMC [WMGB23]. MDL [LLL10]. me [Fel10]. meals [STG21]. Mean [LXC11, FJK10, GGQY07, NH19, PG13, SH18]. means [Big13, LXC11]. measles [DW21]. measure [Sta08b, ZS09]. measured [HCRB23, Mar08]. Measurement [Hat14, Hav14, LLS^{+}22, NSS14a, NSS14b, Pad14, Wal14, CSC^{+}12, CA18, FND09, PZSW23, Sch15, SKZ14, WTJ10, ZCM^{+}11]. measurements [AS10b, JSR16, OIHH09, SR09a, ZMLS22]. measures [ASY09, BD22, BDC^{+}11, GMMW17, Gau11, NH19, RRSM18, WLL17, YOZC23, ZCD^{+}20]. Measuring [CD20, HSD^{+}22, Kan20, LBHB11, LSS^{+}12, Fin13, SVYP11]. meat [RLH^{+}15]. mechanism [SHC12]. mechanistic [BHIK09, GRS23, ZW19]. media [NMW^{+}21, ZKS15]. median [HHLC16]. Mediation [KBG21, Hua18, Hua19, HYL23, KDH^{+}19, SL19, ZRA^{+}21]. mediators [KDH^{+}19]. Medicaid [RHC23]. medical [JB21, MMWH11, MRM12, MLX23, SML^{+}21, WNX^{+}23, ZCD^{+}20]. Medicare [HZG^{+}22a, WNX^{+}23]. medication [HWHWA11, VFMD17]. medicine [SHAB22, ZD13]. medieval [TFG12]. medoids [GPZ^{+}22]. MEG [KF10, THSL12]. Melbourne [ZM16]. melding [ARC07]. membership [FLP^{+}15, RSD22, SP13, WME17, XFS10, MV14]. memory [Ano18, BLTV14, COD22, Hun12, RDL23, SJHJD20, WSGH12, WYKH07, ZGM23]. men [SGCW07, vdKvEW17]. Mendelian [HIH^{+}21]. Merovingian [JLB^{+}14]. MESA [LMKC12]. mesoscale [DT23, FRL08]. Meta [YLL12, BBE^{+}18, DTL^{+}23, FPL10, HST19, KNWJ14, LSY^{+}22, ST14, TDS^{+}14, URZF21, WTCW10, WLM^{+}21, WS14]. meta-analyses [WS14]. Meta-analysis [YLL12, BBE^{+}18, DTL^{+}23, FPL10, HST19, KNWJ14, LSY^{+}22, ST14, TDS^{+}14, URZF21, WTCW10, WLM^{+}21]. meta-patterns [HST19]. metabolic [TJDE17, dCdCAGM16]. metabolomics [MON20]. Metagenomic [Pur11]. metamaterials [YJD21]. meteorological [RCP^{+}16]. method [BDC^{+}11, CTB17, CLK^{+}12, CLZ09, EJD19, Fuk19, HGG13, HCKFZ21, JCJ20, KOJ^{+}14, LBND13, LC12, LES12, LY16, LYY13, LSZL22, MRMB15, Mar08, Mur10, Sha16, SF11, SBS14, TJW10, THSL12, WME17, WG23, XZX18, ZK10, ZMA^{+}20]. methodologies [Bro08]. Methodology [ANFM09]. Methods [ZASM12, BBB^{+}10, BMT13, CW20, CSS11, CFRW19, DGH08, DS14, FYB^{+}15, FCPL19, HWPH10, HPB23, HF20, Kaf12, KDH^{+}19, LSAR12, LBL20, MP11, NQdB^{+}07, NZRC13, PL11, PMMS16, RGF^{+}20, RTB22, SFPS^{+}21, SDH18, SS10b, SL20, SHR^{+}22, TDS^{+}14, Thi11, TWHP15, WZL12, WS14, XCS11, YWL^{+}12, YWLL22, ZCG^{+}09]. methylation [FGS^{+}10, HZL^{+}15, KBFM12, MZA22]. metric [TCZ16]. mHealth [KHBV20]. mice [SGCW07]. micro [DLKM20]. micro-randomized [DLKM20]. microarray [BM08, BNW08, FH09, JL11, KBFM12, LHPW13, LN12, SSD15, SKZ14, WYKH07, WI07]. microarrays [Efr09, GZB^{+}11]. microbial [CGN22, DVA^{+}19, MWW20]. microbiological [RLH^{+}15]. microbiome [BAH22, CL13, Fuk19, JEK^{+}22, JLRK23, KHDV20, KHDV22, ML23, MM22, RZC^{+}18, RBF^{+}20, SZL16, SSD^{+}19, SL19, SXC^{+}20, TMN18, WZ17, WCL23, ZSCL23]. microbiota [PAS23, VML^{+}21]. microdata [SS10b]. microrheology [LLS^{+}22]. microRNA [SCV^{+}10]. microscopic [BHP10, QHPD19]. microscopy [CHH^{+}14, DAL^{+}23, HFH10]. Microsimulation [RODC19]. microstructure [ASY09]. Microwave [CM09]. Middle [vdHWC^{+}12]. might [HJ18]. migrants [Far22]. migration [AR18, DMGJ20, MC17, SKKS14]. migratory [WHAW21]. Milan [ZST16]. milestones [SR09a]. minimum [CCD22, JLA16, KKR13, TT09, ZCG^{+}09]. Mining [GGCM20, RCP^{+}16]. Minnesota [LCG09]. Mira [HLY^{+}21]. misaligned [DPR^{+}20]. misalignment [BGH10]. misclassification [BGM17, SS10b]. misclassification-based [SS10b]. misidentification [MBDL14]. miss [HJ18]. Missing [ML11, AT10, CGCA21, JD18, LGL^{+}12, LQNM19, LCRM21, MSG^{+}20, MON20, OGP^{+}18, PZ19, SRA^{+}15, SPS20, SHC12, TACH21, WCD23, YLS14]. missingness [RB10b]. misspecification [SD10]. Misspecified [HE15]. misuse [CCS18]. Mitigating [SP20]. mitigation [HJS22]. Mixed [DT23, FLHA15, MV14, BTA20, BL19, CFLP15, CFMR18, CWWW17, CLLR20, CB22, CWH20, DLM14, DLS^{+}17, FK22, GMMW17, GKM23, GEC13, HRP10, HWHWA11, JAZ15, JRHM22, MIP22, MAM17, MBH^{+}11, MNB^{+}12, PKP16, PGL^{+}11, PDS13, RBF^{+}20, RSD22, SFGLR15, WME17, WMGB23, WH11, XFS10, ZNB^{+}21]. mixed-effect [CLLR20]. mixed-effects [CWWW17, HWHWA11, MIP22]. Mixed-frequency [DT23, CFMR18]. mixed-mode [CB22, PKP16]. mixed-type [ZNB^{+}21]. mixing [AH16]. Mixture [DMN20, AD22, AV15, AT15, ACS^{+}23, BCJ15, CPP^{+}14, CGCA21, CWS15, DL09, DAL^{+}23, EFJ07, FSPWWE18, GM08, HJ18, HWF15, HBP17, HCRB23, DFGY23, JRHM22, JCJ20, LDV^{+}10, LMS10, LS18, MDGM21, Mur10, MHC15, RODC19, SMZ21, SRH16, SX09, WHNW15, ZKY14, ZBT^{+}20]. mixture-of-experts [FSPWWE18]. mixtures [AMB^{+}20, MM22, MBL^{+}17, RGF^{+}20, SRC15, VDP08, WCW^{+}22, WHC^{+}22, WMA^{+}14, Yua09]. mle [FJK10]. mobile [AFS07, MHG18, SGNM22, ZST16]. mobility [BKVW22, DM18, KBG21, ZST16, ZSLH23]. modalities [FMB^{+}12]. mode [CB22, PKP16]. Model [AS23, FLP23, KBFM12, LRI21, RHHH13, SH18, SD10, SRL10, VHS13, YLLS21, YRY17, ZMO22, ALC09, AN14, ACS^{+}23, AMR16, ABB^{+}12, BLM09, BL19, BRG08, BCA18, BML^{+}20, BL07a, BL07b, BE23, BSNP16, BBDP11, BCJ15, BDL^{+}16, CFW17, CFMR18, CGT^{+}14, CC19, CPP^{+}14, CHOK14, CHAP16, CKK^{+}22, CGCA21, CMR15, CSGD16, CSC^{+}12, CWWW17, CFH^{+}14, COC23, CAL^{+}23, CBvdHvdH08, CGFT15, CWH20, CSB^{+}15, DPR11, DBG21, DLM14, DLS^{+}17, DAL^{+}23, DS14, DM18, EFDS20, ENF14, EOB21, FFJJ14, FPLM18, FGA09, FC20, FLHA15, FZSI^{+}08, GPRZ17, GMMW17, GDG^{+}16, GM08, HRP10, HAFFH21, HGM15, HSFP11, HHLC16, HH21, HSSF21, HWK21, HS14, HJS22, HISV15, HWHWA11, HRFS19, HS10, HBP17, HST19, JAZ15, JSW09, JLB^{+}14, JYB16, JPS21, JWZBC19, JLLK20, KNWJ14, KMMS13, KXC09, KAvdW^{+}23, KSH^{+}13, KHDV20, KHBV20, KHDV22]. model [KPC^{+}19, LZK^{+}15, LMGJ15, Lee18, LHF^{+}20, LRMM15, LMS10, LC12, LWLX19, LNR^{+}22, LMW10, LBBM21, LT12, LWZ19, LYBA22, MKKN21, ML23, MAE^{+}08, MMY^{+}16, MBYWX19, MBGDS11, MLP^{+}19, MSJ14, ME18, MLCW13, MR15, MMBL20, MG22, MKM23, MDR10, NDRF17, OIHH09, OHC^{+}17, PA23, PNB22, PGL^{+}11, PGL^{+}19, PC20, PZSW23, PT12, PDS13, PS15, PSL^{+}16, QGFL08, QDN^{+}21, RSI16, RB10b, RS12, Ros09, RAB20, RHHH18, RODC19, SJH11, SGC23, SM15, SGCT17, SCRS^{+}20, SS10a, Ser11, SRCK16, SYZ15, SZ12, SRJ07, SHC12, SKS12, SX09, SM20b, DPT22, SS15b, TTH21, TMN18, TTB22, TCW21, WFHZ23, WHLN15, WLG17, WYT^{+}20, WFH^{+}22, WCH^{+}23, WNX^{+}23, WMGB23, WL08a, WZLP20, WH11, WIC^{+}10, Wit11, WMKG19, WYKH07, WHAW21, XZC17, XCCL20, YOZC23, YL11, YZS^{+}13, ZPMA10, ZWZ19, ZCM^{+}11, ZCS13, ZGV^{+}16, ZSP19, ZW19, ZKY14, ZLD12, ZHJZ15]. model [ZBT^{+}20]. Model-based [KBFM12, LRI21, VHS13, YRY17, ZMO22, LC12, MDR10, WHLN15, ZPMA10]. Model-robust [SRL10]. Modeling [AH16, AZM11, BWS19, CKHP15, Cha17, CGW^{+}10, CMZ19, CD18, CPvV^{+}11, DVA^{+}19, DLZ10, DGCT10, DET23, FDH10, GM15, HG10, HL08, HCRB23, LvdVvWvdW13, Lie13, LL09, MDP21, MWW20, Mas22, Moh13b, NMW^{+}21, QBC13, RAY14, RD14, Sch23, SCK19, TMPF12, TSG17, UH20, VIF13, WRNR14, XKS15, ZLZB18, ZGM23, ZHO22, ZHJZ15, AMR18, AS10b, BJ19, BLM^{+}23, BM22, BM11, BD11, BB11, CN07, DRB21, DL11b, DFN08, DHM^{+}17, Fie10a, Fie10b, Fin13, FD11, FZCV22, FHSJ14, FRL08, GGMG23, GT10, GH22, HGRS17, HWF15, DFGY23, IGA22, JWH22, JPTO17, JTLE22, KKL23, KN20, KG11, KLH^{+}16, KPDO23, Kou08, LZZL18, LSS^{+}20, LA22, LMKC12, LYH^{+}16, LGK18, LCRM21, LN12, LZ11, MHG18, ML11, MRM12, MCCW09, MDWH21, NBZ11, NH19, OSL^{+}14, PYP^{+}09, PMQW14, PK18, PK19, PTGN12, PPB11, PPM14]. modeling [QYP09, QWC17, RF07, RGPC19, RSD22, SMZ21, SJM^{+}14, SP13, SRC15, SGCW07, SFB16, SZ20, SCL^{+}13, SSH^{+}11, SCV^{+}10, SH11, SJ11, WTJ10, WP12, WKR21, WFC^{+}22, WWCZ22, XS11, YLG15, YZS^{+}13, ZLR19, ZY12, ZGJ^{+}22, ZSCL23, ZS09, ZSG11, ZW18]. Modelled [TM22, GKP^{+}16]. Modelling [DSCS19, DSB19, MC17, RTB22, SJHJD20, WTB16, COM22, CT18, CA18, CLGK22, CD17, DMA19, DMGJ20, ESF14, EKO22, GPRR16, GSD^{+}18, JGVM18, RS09, REFT18, SCDD18, WSU^{+}19, YWQG23]. models [ASX13, AD22, AT10, AS23, AV15, AM07, BJ12, BTJ^{+}14, BPS22, BLTV14, BHB^{+}21, BPSC14, BL11, BK20, BvdB22, BHW15, BWBS14, BHIK09, BGK^{+}15, BJ09, CGM17, CTM14, CFLP15, CS13, COD22, CR13, CB10, CBZG17, CA23, CLFC23, CGCN22, CLR16, CLLR20, CGN22, CD18, CB22, CSS11, CWS15, CMJ09, CWH20, DL09, DBF^{+}16, DPHL10, DB22, DL11b, DVA^{+}19, DW21, EFJ07, FCC15, FBM09, FZZW17, FPC20a, FH13, FSG16, FMA18, GJPS08, GRL^{+}13, GKM23, GEC13, GMB15, GB16, GY23, GCL^{+}15, GM15, GQ11, HHA15, HHK^{+}16, HHN^{+}20, HHHV17, HSH12, HJ18, HGS23, HCS18, HGRS17, HTP14, HCW11, HIH^{+}21, HR22, Hun12, HCP^{+}17, ISR12, JRHM22, JWL^{+}19, JHMC16, JSF^{+}22, JS08, Kan20, KKL11, KN17, KG11, KM17, KZS23, KO14b, KLCM20, LK13, LBA11, LRDD22, LDV^{+}10, LWLW15, LS18, LNC^{+}19]. models [LLZ19, LKB21, LYY^{+}23, LS22, LBD18b, LZP16, LQNM19, LSS^{+}12, LHH19, LN12, MV14, MIP22, MMM^{+}16, ML13, MBDL14, MGSD19, MMGR21, MGTZ21, MGM^{+}14, MB22, MLKQ22, MH14, MZI18, MBH^{+}11, MDGM21, NKAY10, PQR21, PHJ22, PCJW15, PW12, RGT13, RB10a, RB11, RBF^{+}20, RTB^{+}21, RLSF12, SFC11, STM17, SBJR09, SGL^{+}08, SPI^{+}23, SPPR08, SEX21, SWHO11, SSL^{+}10, SRH16, SFGLR15, SBSH18, SL20, SW17, SM13, SCTV11, SKZ14, TJDE17, TDBM23, TVJM13, TWZ15, TB22, VIF13, VFMD17, VCC22, VVSK18, VRN^{+}11, WZHC12, WBB13, WZF^{+}13, WYH^{+}14, WSM^{+}16, WME17, WCL23, WS10a, WS10b, WHC^{+}22, WK10, WHNW15, XLS^{+}19, XBS23, YD23, YHE20, YSL08, YR21, YBL^{+}17, ZPGO21, Zan15, ZYC^{+}17, ZOZ17, ZRCC21, ZNB^{+}21, ZW07, ZSFS22, dCP10, DMN20, vdBR10]. moderate [KPDO23]. moderator [NECS17, NS20]. modern [CHAP16, DH11, Kaf11a]. modes [FRL18]. modifications [CLZ16]. modified [ZRCC21]. Modifying [SPF20]. modulated [WWCZ22, ZLS^{+}17]. modules [WJF^{+}15, ZW07]. Molecular [Wen16, CDB11, DKLL19, HS09]. molecule [DK12, GMNP^{+}21]. moment [CB22, NWJ20, ZSP19]. moment-based [ZSP19]. moments [Jam07]. MONEYBaRL [SC14]. Monitoring [HCD^{+}21, BWT^{+}20, CMPR22, GDJR20, GKP^{+}16, IGA22, Ste09, SBD23, TEF22, WHNW15, XQ23, ZKS15, ZST14]. monkeys [MMGC22]. monotone [WKR21]. Monotonic [FPC20b]. monotonicity [QGM^{+}14, ZZD11]. monoxide [MAE^{+}08]. Monte [KRC23, RB10a, RB11, Wan11, WLK18, ZW08]. morphology [SSH^{+}11]. Morris [DS14]. Mortality [Sha16, AN14, AZC^{+}17, DPR11, FH14, GRS23, IWG13, KAGK^{+}23, LR20a, LR21, MWP^{+}15, NMD19, QWC17, SMW^{+}22, WRNR14, WBK^{+}19, XBS23, ZSH13]. mortgage [ASX13]. most [LR20b, VFMD17]. mother [ZMA^{+}20]. mother-offspring [ZMA^{+}20]. motifs [FSM^{+}19]. motion [BLTG15, CPP^{+}14, FHSJ14, SBW^{+}09]. motivated [CdVM^{+}22]. move [GC08, Gel13]. movement [CA18, HHA15, HGS23, JPS21, Mas22, NDRF17, RHHH18, SWH22, SHF^{+}16]. movements [HSVF09, MN15]. MPE [Gne12]. MRI [KL16, Laz16, Sch16, WLPP16, WLP^{+}16]. MSIQ [LZZL18]. much [PBSVS23]. Multi [AWL13, HWF15, LBD18b, ZZ08, AYJ^{+}09, BR08, CPvV^{+}11, CMJ09, FPLM18, JGF08, KNWJ14, KK12, KX12, KSH^{+}13, LM10a, LM10b, LNW08a, LNW08b, LL11, LMKC12, McE09, MB08, Moh13b, Mur08, NDRF17, PAS23, Qiu08, SX09, Tib08, TvdL08, WHLN15, WFS19, WMKG19, WJL16, YFM19, ZGV^{+}16, ZHM^{+}19, OSL^{+}14]. multi-angle [AYJ^{+}09]. multi-attribute [KK12]. Multi-center [ZZ08]. multi-district [PAS23]. multi-ethnic [LMKC12, OSL^{+}14]. multi-fidelity [KSH^{+}13]. multi-functional [WHLN15]. multi-individual [WFS19]. multi-level [WJL16]. multi-neuronal [LL11]. multi-omics [ZHM^{+}19]. multi-parameter [CPvV^{+}11]. multi-period [WMKG19]. multi-population [SX09]. multi-resolution [FPLM18]. multi-response [KX12]. Multi-rubric [LBD18b]. multi-scale [BR08, LNW08a, LNW08b, MB08, Mur08, Qiu08, Tib08, TvdL08]. multi-site [YFM19]. multi-source [Moh13b]. Multi-species [HWF15]. multi-state [NDRF17, WMKG19]. multi-step [LM10a, LM10b]. multi-subject [ZGV^{+}16]. multi-tissue [WFS19]. multi-type [KNWJ14]. multi-view [CMJ09]. Multi-way [AWL13, JGF08]. multi-year [McE09]. multiarmed [CDF^{+}20]. Multicategory [WL10, ZZH08]. MultiChannel [GHO23]. multidimensional [CCD22, DMA19, DGH08, YH20]. Multilayer [KS19]. Multilevel [BvdB22, CGM17, DCCP09, SFB16, SCL^{+}13, DSCS19, GPRR16, PPLK18, Sha16, WIC^{+}10]. Multilinear [Hof15]. multimodal [APC23]. multimodality [AABC^{+}19, XZW21]. multinomial [BSNP16, CL13, GL08, HTP14, MM22, MIP22, PAS23, Tad15, TMN18]. multinomials [BW18]. multiomic [GBMRR20]. multiomics [JEK^{+}22]. multiperiod [YTHY18]. multiplatform [ZMA^{+}19]. multiplayer [BLM^{+}23]. Multiple [LZP16, SJGM13, SZO12, WR12, Ber11, BPSC14, CDP^{+}17, CW20, CCH19, CGCN22, CCS18, CSZK14, CR11, DBG21, DL11a, EKO22, EOB21, FM17, FFR^{+}08, FHSJ14, GWZ19, GMLB^{+}14, GSC^{+}20, GGQY07, GG19, HU11, HS22, Hol11, HISV15, HZG^{+}22a, HLK18, JL09, JCJ20, JJRZ21, Kap11, KKL11, KS19, KDH^{+}19, LT11, LZZL18, LL19, LYY^{+}23, LBL20, LHMN13, MHB^{+}09, MLM13, MM11, MW11b, MKS^{+}14, NH19, NL11, OKGM14, PMQW14, PNB22, PC20, PSL^{+}16, PLCX23, RTB^{+}21, Ros16, RY11, Rou11, RDH^{+}20, SJH11, SP13, SMR11, STMC17, Sco09, STG21, SWM13, SPsLC16, SPS20, SHC12, Sme11, SMZ16, TJDE17, TSY22, TACH21, Tin11, VVSK18, WA11, WLL17, WLG17, WCW^{+}22, WJT^{+}21, WP12, WESVS23, WMKG19, YOZC23, ZW07, ZS17, ZGS^{+}14]. multiple-level [JL09]. multiple-model [SHC12]. multiple-neuron [VVSK18]. multiple-response [RDH^{+}20]. multiple-trial [EKO22]. multiresolution [CP20, Sav16]. multiresponse [LZ13]. Multiscale [EKO22, WZF^{+}13]. multistage [BK20]. Multistate [MLKQ22, MDP21]. multistrain [TFB^{+}20]. multistudy [DBTP21, LPKP22, GDTP23]. multitarget [YJD21]. multitask [BMAF^{+}23]. multitissue [MSH21]. MultiType [XBS23]. Multivariate [BHW15, GPRR16, GQ11, JSF^{+}22, LLKP18, MHH17, QWC17, RSD22, AS10b, AV15, BZC^{+}19, BJS^{+}22, CLM22, CDN12, CSB^{+}15, DLS^{+}17, DGL13, DJ11, DTL^{+}23, EFJ07, FPC20a, HWK21, HSVF09, JRHM22, JLDQ10, JEK^{+}22, Joh09, Kaf11a, KN20, KOB^{+}20, KK13, KZS23, LS18, MBL^{+}17, MLP^{+}19, MB22, PTGN12, PZB^{+}10, QTL^{+}22, RF07, SFC11, SJH11, SXZ23, TJW10, THN^{+}19, WME17, XBS23, ZWS08, ZHB09, ZCM^{+}11, ZNB^{+}21, ZMLS22, ZSCL23, ZZXL23]. Multiview [GBMRR20, BTA20, MM15, STM17]. muscle [DSCS19]. musical [MMGR21, YH20]. mutated [LWSP17]. mutation [DTZP13, FYB^{+}15, FSM^{+}19, OSB15, TP11, ZSMJ19]. mutations [MNB^{+}12]. my [Bic10, Sto10]. myelination [PDM19].

names [Zan15]. nanomaterials [PTGN12]. nanoparticle [LKTJ^{+}15]. nanoparticles [KDH^{+}13]. nanoscale [Kou08]. NanoString [JWL^{+}19]. narwhals [SJHJD20]. NASH [MDGM21]. national [DH10, GNCS22]. natural [BLTV14, NMD19, SH11, VRN^{+}11, ZSR14]. navigational [DCHP21]. NBA [GEF22]. nCounter [JWL^{+}19]. NCVS [YSL08]. Near [YKLK23]. Near-far [YKLK23]. nearest [BBM20, DBF^{+}16, KFB11]. nearest-neighbor [BBM20]. needlets [FDKP13]. negative [RHL^{+}22, WDSJ23]. neighbor [BBM20, DBF^{+}16, KFB11, ZBC16]. neighborhood [HPB23, MAM17]. Neolithic [BBG^{+}12]. Nested [KB23, BL11, DLKM20, DKLL19, GKM23, GG19, ISR12, LRZ08, LMS10, MKM23, RD14, WYH^{+}14, ZSP19]. net [FKSBS19]. Netherlands [CPvV^{+}11, vdHWC^{+}12]. Network [FFW09, Haz15, LLR15, OM12, RKLT19, Sin11, TF11, WSU^{+}19, WZF18, ZS22, AS17, BZ16, BBE^{+}18, DH18, DTL^{+}23, Fie10a, Fie10b, GLB^{+}17, JLB^{+}14, JLRK23, KM17, LdGK^{+}17, LBK^{+}23, LLR09, LHPW13, LLZ19, MRMB15, MBYWX19, ME18, MHK22, NKAY10, PPLK18, RVW20, STM17, SM10, SKS10, SHM15, SCV^{+}10, SKBL23, UH20, VML^{+}21, WLM^{+}21, WL08a, WH11, XZW21, XFS10, YLC^{+}17, ZKS15]. Network-based [WZF18, LBK^{+}23, WL08a]. network-linked [LLZ19]. networks [AH16, ADE15, BKS21, BC23, Cha17, CB10, CGN22, Cra16, DMA19, DAAM22, DLZL16, DD16, FD11, GY23, GPBT22, GBNS22, HG10, HWPH10, JND12, JSZZ10, JY10, JCS07, JJ16a, JJ16b, KP16, KK12, KKL23, KW23, KSAX10, KT16, LRDD22, LW18, MM15, MH19, MP11, MAM17, MVV13, NS17, OKGM14, OV17, PTH22, PC20, PHLH11, PHLH12, RKM^{+}23, RRS16, RDL23, SHF^{+}16, SK22, SDP22, SG16, SHGA10, Sil16, SS20, SC23, SJ11, TJDE17, VHS13, WYH^{+}14, WR16, WZD19, WYW^{+}23, WZS19, WP12, WWM^{+}14, WL22, XZC17, ZPMA10, ZKS15, ZST14]. neural [BC23, GTZ^{+}21, GPBT22, KKL11, KW23, LW17, RMP17, RS09, SKBL23, WCW15]. neurobehavioral [SXC^{+}20]. neurochemical [LPKP22]. neuroimages [LZ21]. neuroimaging [FMB^{+}12, GM16, GLB^{+}17, KNWJ14, LAS16, PC20, RHZ^{+}15, RG21, SIL^{+}11, YLL12]. neurological [FMB^{+}12, GWZ19]. neuron [VVSK18]. neuronal [CA23, DS14, DKS18, LL11, MVP11]. neurons [LL11, MP11, PSW18]. neuroscience [Kaf11b]. next [Feu13, Kaf12, SZ12, ZWW13, ZSMJ19]. next-generation [Kaf12, SZ12]. nicotine [ZLZB18]. NICU [YLS14]. NIJ [FCPL19]. Niño [WTB16]. nitrogen [MK21, SCRS^{+}20, ZGLH13]. Node [Mei10, Cha17]. noise [ASY09, FWK^{+}13, TDBM23, ZLJW23]. noisy [BHR11, CLW20, LBD^{+}18a, PK11]. nomination [FLP^{+}15]. Non [DKZ09, FPLM18, JYB16, NMD19, PHJ22, RG21, WCH^{+}23, ZCGC21]. Non-Euclidean [DKZ09]. non-Gaussian [FPLM18, JYB16, PHJ22, RG21, ZCGC21]. non-homogeneous [WCH^{+}23]. non-overlap [NMD19]. noncompliance [HZF22]. nonconfounding [ZCRC18]. nonconvex [BH11]. nondiscovery [AS10a]. none [ZHFN23]. nonexchangeable [MHC15]. nonhomogeneous [CLFC23, HGRS17, SZ12]. nonignorable [HTP14, OGP^{+}18, YLS14, ZY12]. noninferiority [NZRC13]. Nonlinear [BPSC14, GCC^{+}11, ST11, CFH^{+}14, CFRW19, DFN08, GZB^{+}11, HRP10, HE15, HWHWA11, Hun12, IWG13, KMMS13, LMW10, LCMJ11, MLCW13, PPB11, WYL^{+}22, YHE20, ZW19, ZGM23]. nonnegative [OP09, WFS19]. Nonparametric [CDF^{+}20, DHG19, FBW^{+}17, KG11, LKB21, Lie19, LW18, MSJ14, OW11, QW08, Sco09, WTCW10, WJL16, ZK10, APC23, AMR16, BBM20, BM11, CTM14, CFLP15, CLGK22, FSG16, FDH10, GMNP^{+}21, HSSF21, HCW11, DFGY23, JRHM22, LWLW15, LSZL22, MC17, QHPD19, RS10, RWK17, SP13, Sav16, SSD^{+}19, TWHP15, VRN^{+}11, WZL12, Yua09, YLL12, ZGV^{+}16, ZKY14]. nonparametrics [HS13]. nonprobability [WLA^{+}21]. nonrandom [MON20]. nonrectangular [Gau11]. nonresponse [AT15, HTP14, PT12]. Nonseparable [DBF^{+}16]. nonspatial [LCG09]. nonstandard [KO14b]. Nonstationary [JS08, KO14b, RCBB19, SCDD18, BM11, EKO22, FC20, GS13, HKP^{+}19, SRC15, SKS12, ZHO22]. nontransitive [CAV^{+}19]. nonuniform [JWZBC19]. Noordin [RVW20]. normal [Mac20a, Mac20b]. normalized [CSL^{+}08]. normalizing [JWL^{+}19]. norms [YHE20]. North [LMB18, TETJ17]. Norway [KN13]. not-for-profit [ZPR14]. note [LM10b, WS10a]. notes [Lin13]. novel [CCD22, EKW20, LWSP17, LSZL22, SBS14, TCS^{+}23]. novo [LWSP17]. nuclear [CLEB14, Lie19]. null [Hua19, Joh09, Sch08]. number [CGT^{+}14, CJM^{+}17, Gil17, GGQY07, HGM15, KN13, LvdVvWvdW13, LL11, MBGDS11, NCHJ13, NZ12, SZ12, SKAL19, ZLOS10, ZZ18]. numerical [PSD13]. numerically [BJ09]. nursing [BLM09].

O157 [TFB^{+}20]. obesity [CLLR20]. Object [MVV13, KDH^{+}13]. Object-oriented [MVV13]. Objective [CC19, Rub08]. objects [ZMLS22]. Oblique [JLL^{+}19]. observational [BHC^{+}20, CH14, GBST19, PL11, Ros12, Ros16, Ros18, RG23, VFH16, YSR22, YKLK23, ZLD12, ZSG^{+}13, ZPR14]. observations [BCA18, BHR11, LSM15, PZ19]. observed [BMT13, CDN12, DS14, HCS18, PM08, RAY14, XKG^{+}19]. observer [CA18]. obstacle [AC12]. occupational [HTP14, LCZ^{+}17]. occurrence [JDP^{+}13, TWH13]. occurrences [FSG16, RK22, SS15b, WWCZ22, XKS15]. ocean [BM22, DSB19, FTE^{+}21, PKGG23]. odds [SC16]. off [FK22]. offenders [BSLL10]. offspring [ZMA^{+}20]. often [LMB18]. Ohio [HWK21]. older [GGFG^{+}18]. omic [RGSB^{+}18]. omics [AS10a, LPH22, SWM13, ZHM^{+}19]. omission [HHH10b]. omnibus [SHH22]. oncogene [YLH17]. oncology [LS18, MHC15, YY11]. oncomarkers [BZ16]. One [MGRG^{+}23, BvdH19, FZSI^{+}08]. one-inflated [BvdH19]. online [BLM^{+}23, CPP^{+}14, FFJJ14, GH12, JY10, LBD18b, MHG18, PM08, SM20a, SRJ07, SBD23, ZPMA10]. only [FH13, MG22, TON20, WS10a, WS10b]. onset [GM16, HZY^{+}15]. Open [Lyo08, MMM^{+}16]. operating [YHX13]. operations [ZCGC21]. opinion [MAB^{+}14]. opioid [HWK21, SKBL23]. opposing [GEF22]. Optimal [AC12, BM08, DPHL10, FPC20a, FBH23, GH12, MBR09, PPLK18, SM10, SIS^{+}20, FLS16, JLS^{+}17, TWA18, ZZ18]. Optimization [MRW09, BvdBS^{+}15, FHHT07, LXC11, YJD21]. optimizing [LCYZ23]. options [RB10a, RB11]. Order [Hun12, GTZ^{+}21, GY23, OE12, ST14, VML^{+}21]. ordered [BYZ18, BK20, EKO22, HTP14]. ordering [RKM^{+}23]. Ordinal [MMGC22, HSVF09, LMMS21, LBD18b]. ordinary [DMVT23, ZYC^{+}17]. Oregon [JCK22]. organic [RCBB19]. organizing [SGL^{+}08]. orientations [HKT12]. oriented [MSJ14, MVV13]. origin [LC10]. origins [CW10, UH20]. Orthogonal [AICV11, WEWX21]. Oscar [HSFP11]. Oscillation [WTB16]. other [Efr09, GJPS08, GS11]. outbreak [Sun22]. outbreaks [DHL18, KLH^{+}16]. outcome [FBW^{+}17, GEC13, HZG^{+}22a, MMGC22, NH19, SCDG17, SXC^{+}20, ZNSL14]. outcomes [DZ23, FCC15, GSC^{+}20, GMB15, HS22, HS13, HTP14, HEHM23, KSD11, KAvdW^{+}23, KHDV20, KH13, LT12, MLM13, PBSVS23, Ros16, SBJR09, SHW18, TLH14, XLDO13, ZMB23, ZSG^{+}13, KHDV22]. outdoor [DHM^{+}17]. outlet [LFMM23]. outlier [Deb09]. outliers [CLM22]. Outline [ZTH19]. outlying [Ger09]. output [GB16, SFC11, SM15]. outputs [BCA18]. ovarian [Ros12]. overcome [XCS11]. overdispersion [SPF20]. overdose [HWK21]. overdoses [LCRM21]. overlap [NMD19]. Overlapping [LBA11, RVW20, ZHM^{+}19]. Oxfordshire [CD17]. ozone [BL11, DLZ10, DT19, LYH^{+}16, REG^{+}11, RCF^{+}13, RCP^{+}16, Ste07, WRNR14, YMP11]. ozone-related [WRNR14].

page [AZM11]. pain [FSM17]. paint [CN07]. paintings [YSG16]. pair [CAV^{+}19, WZ16, ZMA^{+}20]. paired [CA22, LT12, RAKS14, RAKS15, Thi11]. paired-end [RAKS14, RAKS15]. pairing [ZPR14]. pairs [ZSMJ19]. Pairwise [VGH14, CKM21, JL11]. paleoclimate [CHAP16, EHKW12, GREG15, THN^{+}19]. paleoclimatic [Tin11]. palindromes [TWH13]. PALM [DTL^{+}23, JWH22]. pan [LPH22, JMJ^{+}21]. pan-cancer [LPH22]. Pan-disease [JMJ^{+}21]. pan-omics [LPH22]. pandemic [BWT^{+}20, CdVM^{+}22, KAGK^{+}23, PPB^{+}14, ZJBS21]. panel [CFMR18, DSH^{+}13, HKP23, IWG13, SRH16, SCK19]. panels [DET23]. paper [New09]. papers [Fie10a, Fie10b, LRS09]. paradises [Men18]. paradox [Men18]. paradoxes [Men18]. Parallel [GB16]. Parameter [DFN08, KSH^{+}13, LYBA22, CPvV^{+}11, DLZL16, Gho10, RLSF12, WHLN15]. parameters [LMW10, MNR14, RJP16, Sun22]. parametric [Efr12, HS14, RS09, TFB14]. parent [LC10]. parent-of-origin [LC10]. Pareto [CVF10, RCLWW10]. Parkinson [WLL17]. Parsimonious [COM22, Ros09]. part [WCL23]. Partial [SEX21, ARK^{+}18, BL11, GB16, JD18, RKM^{+}23, WDSJ23, YL13, dCdCAGM16]. Partial-mastery [SEX21]. Partially [SW17, AMR16, BMT13, CR13, DS14, XKG^{+}19]. Particle [Lyo08, DS14, FZZW17, HFH10, KP15, KS17, LHF^{+}20, LLS^{+}22, WMA^{+}14]. particle-based [LHF^{+}20]. particles [CHH^{+}14, GL18]. particulate [DBF^{+}16, PYP^{+}09, PD20, XBS23]. partition [CAL^{+}23, MZI18, PQR21, Zan15]. partitioned [SBSH18]. Partitioning [GPZ^{+}22, ST11, JGF08, Sad14, SHH22]. partitions [CLZ09]. partnership [AH16, KHLB22]. passive [vdHWC^{+}12]. past [BLTV14, LR20b, MDP21]. pathology [LWLX19]. paths [ZW15]. pathway [FMA18, LZ11]. pathways [CC19, LWSP17, SCTV11]. Pathwise [FHHT07]. Patient [DTL^{+}23, GG19, QDN^{+}21, SMC^{+}20]. Patient-centered [DTL^{+}23]. patient-specific [SMC^{+}20]. patients [CFH^{+}14, LTL19, LHH19, YLH17]. pattern [AV15, AT15, SRH16, YD23, YLG15]. patterns [AM07, DMGJ20, FGS^{+}10, GHO23, HJ18, HST19, LCG09, MC17, MZA22, SFPS^{+}21, SPH17, WYL^{+}22]. Paving [JSR16]. payments [WD10]. PCA [CMR18]. PCR [SKZ14]. peak [GWZ19, HCW11, KOJ^{+}14]. peaks [DGM^{+}08, KN20, SJGM13, SD10]. peaks-over-thresholds [KN20]. pediatric [Bro09, Hat14, Hav14, NSS14a, NSS14b, Pad14, Wal14, ZS18]. pedigree [HPB23, ZW12]. Penalized [LYY^{+}23, SM13, BK20, BH11, FSM^{+}19, HWF15, KSW^{+}21, MBYWX19, MK21, MDGM21, RZC^{+}18, SK22, WLG17, WL08b, ZLOS10]. penalties [FFW09]. penalty [LRZ08]. People [vdHWC^{+}12, LBK^{+}23, LNR^{+}22]. peptide [LES12]. percentages [KSP16]. percentiles [WTCW10]. performance [BLM09, CAS20, GTW13, HSD^{+}22, KZ16, LPT^{+}11, PS12, Sco09, YH20]. performer [HSFP11]. period [BBG^{+}12, HLY^{+}21, JMJ^{+}21, LCB16, WMKG19]. period-luminosity [HLY^{+}21]. periodic [LRHF12, LCB16]. periodically [FPL10]. periodontal [HS14, Ros16]. periods [HLY^{+}21]. permanent [MB22]. Permutation [CLTZ22, GBST19, SW10]. permutational [MAB^{+}14]. permutational-splitting [MAB^{+}14]. persistence [Mas22]. Persistent [BMM^{+}16, KM17, MLX23]. person [YLH07]. person-to-person [YLH07]. personal [Bic10, ZSFS22]. personalized [AZM11, YGAT20]. perspective [Lie13]. Perspectives [Dav17, HK17a, Kra17, NZ17a, NZ17b, Sch17, Zho17a]. perturbation [DB22]. Perturbed [RLHD21, MLCW13]. PET [LS18, OMM^{+}14]. phalanxes [TWZ15]. pharmaceutical [GM15]. pharmacogenetic [TLH14]. pharmacokinetics [DPHL10]. phase [AMGG13, CK14, FFM^{+}21, KOB^{+}20, LSY^{+}22, MSJ14, YY11, ZBLC17, URZF21]. Phenomenological [JGC^{+}18]. phenotype [LBL20]. phenotypes [ABNG14, NKAY10, SS15a]. Phenotypic [RSH12, CSB^{+}15, KHDV20, KHDV22]. Philadelphia [BJ19, HGSJ23]. photo [PGL^{+}19]. photo-switching [PGL^{+}19]. Photodegradation [DHM^{+}17]. phylogenetic [BHB^{+}21, CSB^{+}15, HHN^{+}20, KRC23, LL09, MMY^{+}16, Pur11, TMN18, ZNB^{+}21]. Phylogenetically [CGN22]. phylogenetics [FIM^{+}21]. Phylogeny [ZWZ19, ZSMJ19]. Phylogeny-based [ZWZ19]. physical [BCA18, BE23, CLZ16, LYH^{+}16, MAE^{+}08, WSK^{+}21, YSR22]. Physics [Lyo08]. piecewise [LvdVvWvdW13]. pigeonhole [Owe07]. pipeline [WSK^{+}21]. PIRLS [GPRR16]. pitcher [SC14]. place [Zan15]. place-names [Zan15]. placebo [JPTO17]. placement [AC12, SIS^{+}20]. Plackett [CTM14]. plaid [CMR15]. plan [CLTZ22]. planar [CKM21]. Planet [Gne12]. planning [YHX13]. plant [FSPWWE18, KDH^{+}19, PPB11, VCC22, WFC^{+}22]. plants [LC10]. player [HGM15, SFDMM22, TVJM13]. playing [BLM^{+}23]. plots [CZM10]. Point [LCRM21, TLF^{+}19, XS11, YBL^{+}17, AK12, BWBS14, CLZ16, CT18, CDM18, CSS11, CDB11, DAAM22, EKW20, ERM15, FJK10, FSG16, GDJR20, HWK21, HKT12, ISR12, JWH22, KNWJ14, KKL11, KPDO23, LCG09, LW17, MSSS^{+}10, SGC23, SYZ15, SZ12, SJHJD20, TMY17, TCW21, TEF22, UH20, WS10a, WS10b, WJT^{+}21, XKS15, ZGM23, ZX22]. point-source [MSSS^{+}10]. points [GBC21, KÓ14a, LLR09, MBGDS11, Mas22, SMZ16]. Poisson [CMR18, CBvdHvdH08, ELD09, PTH22, SRC15, SH08, SZ12, WS10a, WS10b, Wit11, XCCL20, ZK10, ZYXS16, ZRCC21]. Poisson-lognormal [SRC15]. Poisson-type [ZRCC21]. policies [KGGQ15]. policy [GC08, Gel13, GRS16, WME17]. political [FHI18, LBA11, ZPBW^{+}18]. politicians [NMW^{+}21]. poll [JLS23]. pollen [HISV15]. polling [GQ10]. polls [DPT22]. pollutant [MBL^{+}17]. pollutants [WCW^{+}22]. pollution [BC23, BJS^{+}22, KDL^{+}17, Lee18, OSL^{+}14, SLZS08, WMT^{+}21, WBK^{+}19, XQ23, OSL^{+}14]. Pólya [GH12, HGRS17]. polymer [SSH^{+}11]. polymer-ZnO [SSH^{+}11]. polymeric [DHM^{+}17]. polynomial [PL08]. pooled [WS10c]. popularity [JLLK20, ZFB14]. popularity-based [JLLK20]. Population [RBB11, SRC15, BBG^{+}12, CT07, DPR11, DPHL10, Far22, HHHV17, HEHM23, LSS^{+}12, MRMB15, MAM17, NMD19, NCHJ13, NECS17, NS20, PNB22, Sad18, SX09, SPF20, TL11, ZBC16, vdHWC^{+}12]. population-specific [CT07]. populations [BvdH09, CR13, DPR^{+}20, FFM^{+}21, GGFG^{+}18, HSSF21, HCRB23, Men18, PMQW14, RHR12, Sha16, WBA^{+}14]. portal [DH10]. portfolio [LXC11]. ports [DH10]. positive [TON20, WLML23]. positive-unlabelled [WLML23]. possible [STJ^{+}07]. post [LQNM19, Sta08a, Sta08b, VFH16]. post-election [Sta08a, Sta08b]. post-selection [LQNM19]. post-stratification [VFH16]. postal [MGRG^{+}23]. postdisaster [ACS^{+}23]. Postelection [JLS23]. posterior [CCD22]. postural [CD12]. potential [KSD11, NZRC13, SPsLC16, SF11, ZNSL14, ZB11]. potentials [FPLM18, Fin13]. poverty [IHJ16, MNR14, TB22]. Power [MJ16, MH14, VC14, ERM15, Gil17, HMM09, JND12, JSF^{+}22, KDH^{+}19, LM10a, LM10b, LSS^{+}20, MBR09, PLM^{+}16, SWLS14, YSR22, ZCGC21, ZLJW23]. Power-law [MH14, VC14]. Power-weighted [MJ16]. Powerful [LX18, ZTLW20]. Practical [PYP^{+}09]. pre [HT08]. pre-validation [HT08]. Precinct [GRS16]. precipitation [BRG08, MSSS^{+}10, RS12, RTB22, SHSZ19, SYZ15, SKS12, SS15b, WFHZ23]. precise [WR12]. precision [Loh09, SHAB22]. predict [QGM^{+}14]. predictability [HGSJ23]. Predicting [FW21, THN^{+}19, ZM16, ABNG14, CFW17, EFDS20, ERM15, FBM09, LZK^{+}15, LZTB16, MRM12, ZFB14]. Prediction [BL19, FAL^{+}10, GPBT22, HMM09, LLZ19, MDGM21, AS10a, BK21, Bro08, Chi12, FMB^{+}12, FLP^{+}15, HLK18, HCP^{+}17, JL19, JWZBC19, JPK21, KHDV20, KHDV22, KB23, LDV^{+}10, LRMM15, LWFW16, LPKP22, LCMJ11, LHH19, MHG18, MRSA19, MH19, PTH22, PSL^{+}16, RDL23, RCBB19, SGCT17, SHM20, SKS12, SCW^{+}23, THN^{+}19, TWHP15, WSM^{+}16, WLL17, YLH17, ZWS08, ZCG^{+}09, ZSG11, ZSG14, ZZXL23]. Prediction-based [FAL^{+}10]. predictions [DHM^{+}17, Tin11, YHX13, YTHY18]. Predictive [FP08, KLH^{+}16, ZCP14, BCA18, BPSC14, CSGD16, RHZ^{+}15, RGSB^{+}18, ZNSL14]. predictor [GL08]. predictors [BMT13, GSC^{+}20, HSVF09, PZB^{+}10, Sch15, ST11, SXC^{+}20, WS10c]. Preelectoral [DPT22]. preferences [CTM14, GY23, LRM17, WME17]. preferential [CdVM^{+}22, DSB19, WZS19, ZST14]. pregnancy [FBW^{+}17, MGTZ21]. prejudice [GRS16]. premature [YLS14]. preoperative [SKBL23]. preprocessing [RHHH13]. presence [DLM14, ENH^{+}18, Fin13, FH13, Gho10, GMB15, LK13, MG22, NMD19, PZSW23, SKAL19, WS10a, WS10b, WCD23, WZLP20, ZZ18]. presence-only [FH13, MG22, WS10a, WS10b]. Preserving [IHJ16, HSFP11]. presidential [JLS23, Men18]. pressure [LT12, LCSZ15, SFGLR15]. preterm [SXC^{+}20, ZSG^{+}13]. Prevalence [VFH16, ARC07, JMJ^{+}21, JD18, KM17, ZS18, vdBN09]. prevention [LBK^{+}23]. preventive [ZCD^{+}20]. previously [Sta23]. price [CMZ19, CHJCK18, Dup17, MB22, NBZ11, RFB17, Yua09, dCP10]. prices [CGM17, Lie13, Lie19, PPM14, VIF13]. pricing [GH12, RB10a, RB11]. primary [DGCT10, Hat14, Hav14, NSS14a, NSS14b, Pad14, Wal14]. Principal [DKLL19, JFM11, LAS16, SP19, ZD13, AXEC18, APW^{+}09, DCCP09, EHM18, GMM08, HZG22b, JEK^{+}22, KDH^{+}19, LHH10, LMM15, LBL20, MLM13, PZ19, RC23b, WT08, YLH17, YFM19, ZW15, ZB11, ZGS^{+}14]. printing [SDHZ14]. prior [DMGJ20, GJPS08]. prioritization [CFRW19]. Prioritizing [KSW^{+}21]. priors [AMB^{+}20, LdGK^{+}17, MHC15, NWJ20, PHWM11, PS12, URZF21, VH14, WLM^{+}21, WMA^{+}14]. prison [ENH^{+}18]. private [PL11, SBSH18, dCP10]. private-values [dCP10]. Probabilistic [ARC07, BRG08, DRB21, FMB^{+}12, SM15, SLBL23, YWB^{+}23, AY12, CMR18, DB22, KLCM20, LG20, LR20b, RAB20, SHSZ19, XLS^{+}19]. probabilities [CW13a, CW13b, CRZ13, FBW^{+}17, Gil13, GNCS22, Moh13a, RP13, RBB11, SB20, Sch13, TSG17, Whi13]. Probability [SJM^{+}14, MKN22, TACH21]. probation [KB23]. probe [WI07]. probe-level [WI07]. probes [DSB19]. probing [HS14]. Probit [MBDL14, MMGC22, ZNB^{+}21]. problem [BK20, BdHZ08, BFF^{+}09, LPT^{+}11, LFMM23, LLL10, MM08, SCDG17, WS10a, WS10b, YE14, ZCS13, ZKS15]. problems [AS10a, Efr08, GS11, TL11, WOH23, YSR22]. procedural [SPH17]. procedure [BG09, MAB^{+}14, STG21, WTCW10, ZTLW20, ZWW13]. procedures [CSZK14, ZZD11]. process [BJ12, BPSC14, BWBS14, CT18, CMJZ22, CDM18, DBF^{+}16, DMGJ20, FHSJ14, FSG16, GGMG23, GB16, HHHV17, ISR12, JGVM18, JWH22, JSF^{+}22, KNWJ14, KKL11, Lee18, LDV^{+}10, LW17, LCRM21, MGMB19, OIHH09, PKGG23, SX09, SJHJD20, TAC^{+}16, TVJM13, THN^{+}19, URZF21, WS10a, WS10b, WJT^{+}21, WKR21, XQ23, XS11, YBL^{+}17, ZK10]. process-convolution [Lee18]. processes [BMAF^{+}23, BSNP16, CKHP15, CGW^{+}10, CW10, ÇL12, CSS11, CWS15, DAAM22, DSH^{+}13, ERM15, FD20, FLP23, HKT12, JGC^{+}18, JLGJL12, KPDO23, LYH^{+}16, LL09, MRW09, MKS^{+}14, MAL^{+}22, PM08, PPM14, RAY14, SH18, SB20, SYZ15, SH08, SZ12, SG17, SCA13, UH20, WWCZ22, XKS15, XKG^{+}19, ZX22]. procurement [GDJR20]. product [CAL^{+}23, HMT12, JFM11]. production [SCL^{+}13, SJHJD20]. products [MB22]. professional [FMBG15]. professor [SMZ21]. Profile [WZL12, MZ12, RGT13]. profiles [LKTJ^{+}15, RLHD21, RSD22, Wei07]. Profiling [CCJ^{+}09, MIP22, PTGN12, SZ12, SPS20]. profit [ZPR14]. program [BMAF^{+}23, IR13, MLM13, YKHS21]. programmes [CTM14]. programming [ZSG^{+}13]. programs [GMM08, ME18]. progression [AXEC18, LGL^{+}18, YLH17]. project [BL19]. projections [ARC07, CGCA21, LR20b, PK11, Ros16]. promote [ME18]. promotion [MKM23, PA23]. prone [GMB15, SR09a, SW17, WBK^{+}19]. propagation [Sad18, SCL^{+}13]. Propensity [LL19, BLM^{+}23, CK14, FHI18]. Properties [BFM12, LKTJ^{+}15]. proportional [FSM^{+}19, PZSW23, SW17, WLG17, WK10, ZHJZ15]. proportions [BBL22, FZCV22, HGB21]. prosecutors [BSLL10]. prospective [RG23]. prostate [LCMJ11]. prosthesis [SIS^{+}20]. protection [SS10b]. protective [CBvdHvdH08]. protects [PLM^{+}16]. protein [BTJ^{+}14, BFF^{+}09, DK12, GPR^{+}22, HTM^{+}13, JJRZ21, LDV^{+}10, LMS10, MPT12, PK11, RS14, WLK18]. proteins [HBW17, Kou08]. proteomic [BBM20, FKSBS19]. proteomics [CWWW17, CLGK22, KPA^{+}10, LLM20, MBH^{+}11, OGP^{+}18, WKG^{+}15]. protocols [LGL^{+}18]. Prototype [BT11, RLSF12]. provided [SS10b]. Providing [BBB^{+}18, SBSH18]. proxies [BLTV14, Ber11, CR11, DL11a, HU11, Hol11, HK23b, Kap11, MM11, MW11b, NL11, Rou11, SMR11, Sme11, Tin11, WA11]. proximal [CLK^{+}12]. proximity [JHMC16]. Proxy [AT15]. Pseudo [AL16, TFB14, WS10a, WS10b]. pseudo-absence [WS10a, WS10b]. Pseudo-value [AL16]. PSF [BHP10]. psychiatric [AD22, RHZ^{+}15]. PTEM [JLLK20]. PTSD [ACS^{+}23]. public [PL11, WR12, WME17]. pulmonary [CHS^{+}16, FS14, SRZ^{+}15]. purchase [BLM^{+}23]. purchases [LMMS21]. purposive [CDF^{+}18].

Qini [BMGN21]. Qini-based [BMGN21]. qRT [SKZ14]. qRT-PCR [SKZ14]. QTL [Wen16]. quality [BGC20, DLM14, FND09, KBMF^{+}23, LYH^{+}16]. Quantification [CDP^{+}17, DGM^{+}08, FSG16, KP15, KS17, LZZL18, LBBM21, WZ16]. quantify [MAB^{+}14, WFHZ23]. Quantifying [FWK^{+}13, LL16, RAKS14, YFHE20, SDH18, RAKS15]. Quantile [SFGLR15, WILW22, AL16, BL19, BGC20, CGI08, CMAC^{+}23, HCYH20, LZ21, LZCW21, LLM20]. quantiles [BPS22, BL19, GGPM16]. Quantitative [ZJLC08, BRG08, CWWW17, MAZM13, MBH^{+}11, WGL^{+}18b]. Quantum [Wan11]. question [JLLK20]. questionnaire [MM08]. queueing [SJ11]. queues [RCLWW10].

R [GQ10]. race [JL19, KPC^{+}19]. race-independent [JL19]. racial [GRS16, GQ10, KM17]. radar [FRL08, LGK18]. radial [HCKFZ21, XDM15]. radiation [CM09, DH10]. radiative [GBH^{+}15]. radio [CVF10, ZMA^{+}19]. Radio-iBAG [ZMA^{+}19]. radioactive [DH10]. Radiogenomic [MBK^{+}21]. RADIOHEAD [MBK^{+}21]. Radiomics [ZMA^{+}19]. Radiomics-based [ZMA^{+}19]. rainfall [BCR^{+}19, BdHZ08, DT23, FRL08, HGRS17]. Random [IKBL08, NQdB^{+}07, URZF21, VFMD17, WNRZ11, Zan15, BLM^{+}23, BZC^{+}19, BFF^{+}09, CFW17, CDB11, DPHL10, FPLM18, FCC15, FTE^{+}21, GRL^{+}13, Goe11, GPZ^{+}22, GREG15, JLL^{+}19, JLB^{+}14, JWL^{+}19, JYB16, LA22, LSL^{+}15, LSY^{+}22, MN15, PQR21, PK11, PC20, Tin11, WTCW10, WL08a, ZPGO21, ZNSL14, ZYXS16]. random-coefficient [JWL^{+}19]. Random-effects [URZF21, CFW17, LSY^{+}22]. Random-set [NQdB^{+}07]. Randomization [JFRS17, WDSJ23, ZHYS23, ZZ08]. randomized [CA22, CCS18, CBvdHvdH08, DLKM20, DK18, HZF22, IR13, JFRS17, STD13, WLM^{+}21, WDSJ23, YY11, ZZTL22, ZHFN23]. randomness [LMB18]. RAnk [WILW22, ABNG14, DK18, FH09, GZB^{+}11, GM08, HSFP11, HH21, Hof07, LZCW21, OSL^{+}14, Tal15, WME17, WG23, WCL23]. rank-based [WCL23]. rank-dependent [Tal15]. rank-score [LZCW21]. ranked [FDH10]. Ranking [SHGA10, CD20, DTL^{+}23, GY23, MV12, NZ12, SMZ16]. rape [YSL08]. Rapid [YJD21, MN15]. rare [FCGA^{+}13, WSH^{+}14]. rarity [Wei07]. Rasch [BMH16]. rate [AS10a, BYZ18, BG09, FYB^{+}15, FSG16, GPR^{+}22, JCJ20, KXC09, KH13, LR20b, MKS^{+}14, Mur10, PHCM^{+}10, RHC23, Sch08, SDT08, TT09, TWH13, WFH^{+}22, Wen16, ZLS^{+}17, ZYFF19]. rates [CdVM^{+}22, GBC21, HE15, LL16, Mar08, MMWH11, TP11, YSL08, vdKvEW17]. ratings [LBD18b, TVJM13]. ratio [CGCA21, PSL^{+}16, ZLJW23]. ratios [BWS19, RBB11, Sch23, SC16]. ray [DGM^{+}08]. rays [FDKP13, SCL^{+}13]. RCRnorm [JWL^{+}19]. RCT [NS20, NECS17]. RCT-to-target-population [NS20, NECS17]. reach [Far22, MZ12]. reaction [DK12, YHE20]. Reactive [ERM15]. read [WZ16]. reading [PPLK18, WRSS15]. Real [FCPL19, GRS23, PL08, BWT^{+}20, DB15, MM15]. Real-Time [FCPL19, GRS23, BWT^{+}20, DB15]. realistic [HGG13]. reassessment [LY16, LYY13]. recapture [ARK^{+}18, CA18, DPR^{+}20, FS13a, KMKB16, LK13, MC17, MBDL14, MLKQ22, RBB11, WMKG19]. received [ZB11]. recency [SLBL23]. receptor [QDN^{+}21]. recidivism [JL19]. recollections [Sto10]. recombination [FYB^{+}15]. recommendations [AZM11, YKHS21]. reconcile [EOB21]. Reconstructing [BLTV14, WYH^{+}14, WOK^{+}16, ZLOS10, ZW08, HISV15]. Reconstruction [ZSMJ19, BB11, LdGK^{+}17, SX09, SHR^{+}22, THSL12, ZSMJ20]. reconstructions [Ber11, CR11, DL11a, EHKW12, GREG15, HU11, Hol11, Kap11, KÓ14a, MM11, MW11b, NL11, Rou11, SMR11, Sme11, Tin11, WA11]. record [Sad18, STG21, TL11, XLS^{+}19]. records [GSC^{+}20, HISV15, HCP^{+}17, JG23, RFWE22, SPS20, TACH21, WSM^{+}16, WLML23, ZCD^{+}20]. records-based [TACH21]. recovery [LK13]. rectum [LCG09]. recurrence [HAFFH21, LCMJ11, WZLP20]. recurrent [BC23, ÇL12, LTL19, SHM20]. recycle [TSS10]. Red [HH21]. redesign [vdBR10]. reduce [NZRC13]. Reduced [OSL^{+}14, CHOK14]. Reduced-rank [OSL^{+}14]. Reducing [JCCG18, ZST14]. reduction [AN14, Fuk19, TJW10, YKHS21]. reductions [MB22]. Reef [GKP^{+}16]. Reexamining [Lin13]. reference [Tal13]. reference-invariant [Tal13]. referenced [LSS^{+}20]. Refining [CLK^{+}13, FMA18]. reflectances [WFC^{+}22]. refreshment [SRH16]. refugees [Far22]. regime [EJD19, FLS16]. regime-switching [EJD19]. regimens [SML^{+}21]. regimes [JLS^{+}17, TWA18, ZZ18]. Region [LSS^{+}20, RHC23, JDP^{+}13]. region-level [RHC23]. Region-referenced [LSS^{+}20]. regional [GEC13, KKLS15, KKLS16, PDM19, SFC11]. regionally [QBC13]. regions [FS13b, HK17b, ZMC^{+}21]. registration [SMW^{+}22, WDL22]. registry [RHR12, Sad14, WNX^{+}23]. Regression [LZ13, SZL16, SZ20, ZLR19, ZW15, ACG13, AS23, APC23, AMB^{+}20, BMGN21, BL19, BH11, CLK^{+}12, CL13, CHS^{+}16, CMJZ22, CGM10, DFN08, DVF13, DT23, FW21, Fre08, GM16, GJPS08, GKM23, GKS17, GS11, GV14, GL08, HZL^{+}15, Hof15, HHH10b, HCYH20, HPF13, JB21, JWL^{+}19, JG23, KGGQ15, KX12, KKLS15, KKLS16, LvdVvWvdW13, LZK^{+}15, LTL19, LL10, LMM15, LZW^{+}15, LMMS21, LZ21, Lin13, LT12, LBL20, LLM20, LKTJ^{+}15, LRS12, LRS15, MBYWX19, MWW20, MBGDS11, MK21, Mey08, MMGC22, MTZZ21, MZI18, MHK22, MDGM21, NDRF17, PAS23, PZB^{+}10, PRRW11, PWP^{+}21, PT12, PL08, QTL^{+}22, QGM^{+}14, QDN^{+}21, RZC^{+}18, RHZ^{+}15, RDH^{+}20, RLH^{+}13, SPI^{+}23, SHSZ19, SS10a, SLBL23, SFGLR15, ST11, SXC^{+}20, SKBL23, SRL10, Tad15, TLH14, VKG12, WD10, WNZK14, WZ17, WEWX21, WFH^{+}22, WZL12, WLA^{+}21, WL08b, YLL12, ZMC^{+}21, ZS17]. regressions [HWF15, HCRB23, ZCP14]. regularization [LBD^{+}18a, NV18, THSL12, VVSK18]. Regularized [LCYZ23, PZB^{+}10, WNZK14, WRSS15, AT10, LCMJ11, LRS12, LRS15, WG23, ZSH13]. regulation [BKGJ14, Dav17, HK17a, Kra17, MLCW13, NZ17a, NZ17b, Sch17, Zho17a]. regulatory [JSZZ10, JCS07, SCV^{+}10, WL22, ZW07]. rehospitalization [YKHS21]. reinforcement [SWH22, TWA18, SC14]. Rejoinder [CW13b, Feu08a, JJ16b, LNW08a, Mac20b, MW11a, NSS14a, NZ17b, SR09c, WLPP16]. relapse [YLG15]. related [FWGS11, KMMS13, OKGM14, RHR12, REFT18, SWLS14, SPsLC16, WRNR14]. relatedness [OSB15]. Relating [KDH^{+}19]. relational [CB10, Hof15, WH11]. relations [HLY^{+}21, SHGA10]. relationship [Sto10, ZHYS23]. relationships [CLK^{+}13, EFDS20, NMW^{+}21]. relative [FYB^{+}15, HMP22]. relaxing [HZG22b]. relevant [GM16]. reliability [LKB21, MR15, Sin09, ZCGC21]. reliable [Ber11, CR11, DL11a, HU11, Hol11, Kap11, MM11, MW11b, NL11, Rou11, SMR11, Sme11, Tin11, WA11]. rely [NZRC13]. remaining [HMM09]. Remembering [Cut10, Fri10, Ols10, Yu10]. Remembrance [Büh10]. remittances [TB22]. remodeling [BBM20]. remote [LRHF12]. remotely [EKW20]. removal [MMM^{+}16, vdBN09]. removing [JL19, YRY17]. remuneration [SMZ21]. renewal [MKS^{+}14]. Rényi [Tal13, Tal15]. repeated [BSLL10, BDC^{+}11, CGM17, WLL17, ZMLS22]. repeated-measures [BDC^{+}11]. replacement [YSR21]. Replicability [HY14, HF20]. replicated [AMR16, FRBT13]. replicates [CSC^{+}12]. replication [HMP22, SFPS^{+}21]. reported [CWS15, GMB15, WSGH12]. reporting [HCD^{+}21, VFMD17]. reports [CGCN22, Kan20]. representations [ZPGO21]. Representatives [YR21]. reproducibility [LBHB11]. reproducible [BC09]. reproduction [vdBN09]. reproductive [RD14]. rerandomization [BDR16]. resampling [LPKP22, YLH07]. resampling-based [YLH07]. research [BC09, BBB^{+}18, SRZ^{+}15]. resident [ZST16]. residing [vdHWC^{+}12]. Residual [CSS11, DGM^{+}08, AL16, BWBS14, RDL23, ZSG11, ZSG14]. Residual-based [DGM^{+}08]. resistance [HWHWA11, PRRW11]. Resnick [TAC^{+}16]. resolution [FPLM18, FCPL19, GMNP^{+}21, GSD^{+}18, HH21, OW11, OBHL22, SFDMM22]. resolutions [KS19, WBKJ22]. resolved [SR23]. resonance [BM11, HBW17, WBKJ22, ZJLC08]. resource [CCS18]. respondent [BGM17, GGFG^{+}18]. respondent-driven [BGM17, GGFG^{+}18]. responders [JPTO17]. Response [BBDP11, SST^{+}08, CCS18, CBvdHvdH08, DHG19, FFJJ14, FS13a, FLHA15, HK23a, HWHWA11, KH23, KX12, KPC^{+}19, MDWH21, PD20, PTGN12, RDH^{+}20, SLZS08, TTB22, WD10, WBB13, WZF^{+}13, YSL08, ZHYS23]. Response-adaptive [BBDP11]. responses [AWL13, CBvdHvdH08, DLS^{+}17, FDH10, JL09, VRN^{+}11, XBS23]. restricted [Mey08, PG13, QDN^{+}21]. restrictions [CCJ^{+}09, LSM15, PSD13]. results [HEHM23]. retention [DZ23]. retrieval [GGG^{+}12]. retrospective [WSGH12]. retrovirus [HBP17]. return [CA22]. returns [FPC20b, FH19, JSX16]. Reuse [TSS10]. reveal [LLR15]. revealed [IWG13, LRM17]. reveals [CLZ16]. review [BW18, Cox07, ZSP19, dGC07]. reviews [MHG18]. Revisiting [DJ11]. reweigh [TSS10]. reweighting [NZRC13]. rhesus [MMGC22]. rhythm [GGG^{+}12]. rhythms [HKP^{+}19]. richness [CDF^{+}18]. Ricker [HHHV17]. Ricker-type [HHHV17]. Riemannian [SKKS14]. right [HMM09, TMY17]. rights [Sad18]. rill [BK21]. Risk [LCMJ11, ASX13, BD22, Bro09, CFLP15, DHG19, FK22, FCGA^{+}13, Goe11, HE14, HZG22b, HCP^{+}17, KH23, KBMF^{+}23, Mar08, PG13, QGM^{+}14, RLH^{+}15, RWK17, RSD22, SMC^{+}20, SCW^{+}23, SKBL23, WSM^{+}16, WACY20, WCH^{+}23, YGAT20, YKLK23, YTHY18, ZLDR17]. risk-based [RLH^{+}15]. risk-prediction [WSM^{+}16]. risks [AL16, CSL^{+}08, LGL^{+}18, MGTZ21]. risky [JAZ15]. river [ADE15, JSR16]. rivers [BPS22]. Rizzo [New09]. RNA [ABNG14, CGFT15, EHM18, LZZL18, Lia19, LRI21, LWZ19, MKKN21, NvdBCR23, RHL^{+}22, RAKS14, RAKS15, SSD15, SKZ14, WZ16, XZX18, ZRCC21, ZSMJ20, ZLDR18]. RNA-Seq [WZ16, XZX18, RHL^{+}22, SKZ14, CGFT15, LZZL18, LRI21, MKKN21, RAKS15, SSD15, ZRCC21]. RNA-sequencing [LWZ19, RAKS14]. RNAi [CSC^{+}12]. RNDClone [ZSMJ20]. robotic [SIS^{+}20]. robots [AFS07]. Robust [CA22, DVF13, FD11, FKSBS19, GMMW17, KW23, PLM^{+}16, RMP17, RFWE22, TACH21, YL13, YWQG23, ZSH13, ZW18, Dup17, JLS^{+}17, MSG^{+}20, QM23, ST14, SRL10, ZMA^{+}20]. ROC [dCdCAGM16]. rockslide [ZBT^{+}20]. role [BLM^{+}23, HE14, HGB21, KBG21, SP19, SFDMM22, DPT22, vdHWC^{+}12]. role-playing [BLM^{+}23]. roll [GCL^{+}15]. rolling [PHWM11]. root [Joh09]. rooted [ZBC16]. rotating [CLTZ22]. rotation [COC23, HRP10]. rotation-based [COC23]. routine [DW21]. routing [Sin11, TF11]. rover [BML^{+}20]. roving [LCZ^{+}17]. rubric [LBD18b]. Rugby [GNCS22]. rule [FP08, MRM12, NV18]. rules [LRMM15]. run [GEF22]. running [LMMS21]. runs [AM07, HS10]. rural [KN13]. Ryan [CW13b, CRZ13, Gil13, Moh13a, RP13, Sch13, Whi13].

S&P500 [RWK17]. SAFE [SIS^{+}20]. safety [HCD^{+}21, KHLB22, KSD11]. Sample [DLKM20, SSD15, CWE18, CDF^{+}18, FH13, KZ16, MAB^{+}14, MNB^{+}12, SMW^{+}22, SS20, SF11, ZMLS22]. sample-based [CDF^{+}18]. sampled [Ger09, HG10, KM17, LCB16, MLKQ22, SM10, WOK^{+}16]. samples [CK14, FWGS11, GSC^{+}20, LZZL18, SDP22, SRH16, SMZ16, WLA^{+}21, ZW08]. sampling [BFM12, BGM17, BK21, CP20, CK14, CLTZ22, CA18, DSB19, FFM^{+}21, FTE^{+}21, GGFG^{+}18, HHK^{+}16, JL10, LKB21, LYBA22, MGRG^{+}23, MP11, PHJ22, QGFL08, RTB^{+}21, RG23, SRZ^{+}15, WZS19, WMKG19, YBL^{+}17, ZKS15]. sandwich [SRL10]. satellite [AYJ^{+}09, GGPM16, MVW^{+}23, ZZD22]. Saudi [LG20]. Saxon [Zan15]. SCAD [FFW09]. Scalable [FCPL19, MK21, TEF22, RMP17, WMGB23]. Scalar [MHK22, LZW^{+}15, LLM20]. scalar-on-image [LZW^{+}15]. scale [BR08, CGW^{+}10, DTL^{+}23, GG19, GS11, HGB21, JL09, JJRZ21, JG23, LNW08a, LNW08b, MRMB15, MB08, MKM23, Mur08, PYP^{+}09, PDS13, QTL^{+}22, Qiu08, SPsLC16, Ste07, Tib08, TvdL08, WTJ10, YBL^{+}17, ZNB^{+}21, ZS17]. scaled [EFDS20]. scales [BMH16]. scaling [DGH08, LCYZ23, LSZL22, YH20]. SCALPEL [PSW18]. Scan [ZYXS16, LZ07, TMN18]. scenario [CGCA21]. scenario-based [CGCA21]. scenarios [BCA18]. Scheffé [Ros16]. schemes [FLP^{+}15]. schizophrenia [CSGD16, ZLDR17]. School [BDR16, DZ23, PHLH12]. schools [PL11, ZPR14, YFM19]. science [BC23, Fie07, Kad18, BL07a, BL07b]. sciences [FGS08, JB21]. scientific [FDH10, TCZ16]. sclerosis [GMLB^{+}14, PSL^{+}16, ZGS^{+}14]. score [BMH16, CK14, FHI18, HZF22, LL19, LZCW21, SCA13, TFB14]. scores [AS10a, NSMM23]. scoring [HS10, HYS19, LES12, YWL^{+}12]. screening [BDC^{+}11, BR16, CLR16, DAL^{+}23, DB22, FLRZ23, KDS20, KAvdW^{+}23, KHZK23, LX18, NZ12, SML^{+}11, SMZ16, WZF18, ZW15, ZCD^{+}20, PS12]. scRNA [LZCW21]. scRNA-seq [LZCW21]. Sea [HH21, CKHP15, DCHP21, DRB21, TETJ17]. sea-level [CKHP15]. search [GL18, LRS12, LRS15, MBR09, TJW10, WIC^{+}10]. season [Bro08, CMAC^{+}23]. Seasonal [SHM20, OHC^{+}17, XKS15]. seasonality [AABC^{+}19, CMZ19]. Seber [WHAW21]. second [GTZ^{+}21, STJ^{+}07]. Section [Gne12, Ano18, FGS08, Kaf11a, Kaf11b, Kaf12]. sectional [MGTZ21]. sections [CGM17]. Secure [SBSH18]. security [CBvdHvdH08, Moh13b]. SEER [HZG^{+}22a, WNX^{+}23]. SEER-Medicare [HZG^{+}22a]. segmental [YH13]. segmentation [CCH19, FHSJ14, LLL10]. seismic [CDM18, ZCS13]. seizure [QW08, WCH^{+}23]. seizures [WJL16].

SELC [MRW09]. Selected [Sto10, CK14]. selecting [ZKY14]. selection [AS10a, AS23, BYZ18, BTA20, BG09, BT11, BKGJ14, BE23, BvdBS^{+}15, BH11, BJ09, CC19, CL13, CLLR20, CWH20, DM18, DVF13, FLS16, FSM17, FND09, FLP23, FKSBS19, GGG^{+}12, GPRZ17, GTW13, GS11, GL08, Hun12, JCS07, KS19, KSW^{+}21, LL10, LZW^{+}15, LRI21, LWFW16, LQNM19, MAM17, MDR10, NWJ20, PWP^{+}21, RJ11, RLSF12, SzCT10, SBS14, SCTV11, WZ17, WCH^{+}23, WCW^{+}22, WLA^{+}21, YJZ09, ZZ18, ZZ22, ZJBS21, ZHM^{+}19, ZST14]. selective [AH16, BW18]. SELEX [ABB^{+}12]. Self [PW12, CK14, CT18, CD18, CWS15, CBvdHvdH08, GMB15, SGL^{+}08, WSGH12]. Self-exciting [PW12, CT18, CD18]. self-organizing [SGL^{+}08]. self-protective [CBvdHvdH08]. self-reported [CWS15, GMB15, WSGH12]. self-selected [CK14]. Semi [CLGK22, RS09, HS14, JPS21, KMKB16, PG13, WFS19, XDO10]. semi-competing [PG13]. semi-complete [KMKB16]. semi-Markov [JPS21]. semi-nonnegative [WFS19]. Semi-parametric [RS09, HS14]. Semi-supervised [CLGK22, XDO10]. semicompeting [MGTZ21]. semicontinuous [CKK^{+}22]. Semiparametric [GMB15, JWH22, LMKC12, MRSA19, MIP22, PPB11, RK22, VKG12, YGAT20, ZLS^{+}17, AMGG13, AMB^{+}20, CTB17, COD22, GEC13, GQ11, HHLC16, HMT12, HH21, Hof07, JLDQ10, JLA16, JCJ20, LTL19, LT12, ML14, MTZZ21, PA23, RF07, WHAW21, ZLR19]. semireference [TTH21]. semireference-based [TTH21]. Senate [LRM17]. sensible [KW23, WYT^{+}20]. sensing [LRHF12, LPKP22]. sensitive [CCS18, JL19, KB10]. Sensitivity [GM09, HZG22b, NPM12, NECS17, Ros18, FH19, GTW13, HBHM13, HHH10b, HZG^{+}22a, PHM^{+}23, Ros12, SVYP11, NS20]. sensor [LCZ^{+}17, RFWE22]. sensor-based [RFWE22]. sentenced [ENH^{+}18]. Separable [FH14]. separation [WG23]. Seq [SJGM13, SKZ14, CGFT15, LZZL18, LRI21, LZCW21, MKKN21, RAKS15, SSD15, ZRCC21, WZ16, XZX18, RHL^{+}22]. sequence [CCH19, MVW^{+}23, Thi11, ZW12, ZSMJ20]. sequence-based [ZW12]. sequences [CLFC23, FM17, FJK10, SYZ11, ZASM12]. sequencing [CJM^{+}17, CDP^{+}17, HSSF21, JTLE22, Kaf12, Lia19, LWZ19, MM08, MNB^{+}12, NvdBCR23, RSI16, RAKS14, SZ12, SS15a, Wit11, ZWW13, ZSMJ19, ZLDR18]. Sequential [CP20, FLS16, IGA22, JGF08, JWZBC19, RB10a, RB11, RGSB^{+}18, RG23, ABB^{+}12, CB22, MRW09, PACB21, SIS^{+}20, TSS10, WLK18, ZGJ^{+}22]. Sequentially [AHZ23]. serial [CMZ19, DT19]. series [AMR18, BvdB22, BHIK09, BGK^{+}15, CGCA21, DGL13, EKO22, ESO19, FZZW17, FZCV22, FHSJ14, GV14, Hun12, KH13, LLKP18, MJ16, MB22, QW08, RS09, SJM^{+}14, SM20a, STA18, SH08, SBS14, TSY22, Tin11]. service [MMWH11, Ser11]. services [SJ11]. set [COC23, DB15, Efr09, HE14, KSW^{+}21, LCB16, MN15, NQdB^{+}07, WHLN15, ZLJW23]. sets [ACG13, BZN18, CDB11, DVF13, ET07, FTE^{+}21, KB23, LSAR12, PDS13, RC23b, SJH11, SWM13]. setting [ST14]. several [EHKW12, Fre08, Thi11]. severity [PPB^{+}14, SXZ23]. Sex [CWS15, BSLL10, CGCA21]. sex-offenders [BSLL10]. sexually [AH16]. Shafer [ELD09]. Shape [KS17, CCJ^{+}09, DKLL19, DKS18, GBNS22, LA22, Mey08, QDN^{+}21, SSD^{+}19, VDP08, ZS09]. Shape-constrained [KS17]. shape-restricted [Mey08, QDN^{+}21]. shapes [CKM21, MFB^{+}13, WSK^{+}21]. shared [ENF14, LNC^{+}19, WWL22]. sharing [BCJ15, JSR16, WR12]. sharper [Sta08b]. sheet [BK21, CHAP16, DAL^{+}23, LHF^{+}20]. shielding [CLEB14]. shock [GBH^{+}15]. shocks [HKP23]. shooter [STJ^{+}07]. shooting [SB20]. SHOPPER [RAB20]. short [BLTV14, DB15, EJD19, RDL23, SKS12, WYKH07, ZBG14]. short- [BLTV14]. short-term [DB15, EJD19, RDL23, ZBG14]. shortfall [HHH10a]. Should [GC08, Gel13, Efr08, FRL18]. Shrinkage [RJP16, BMT13, FO11, LdGK^{+}17, RJ11, WRSS15]. shuffle [BY13]. sibling [FH19]. side [RC23a]. sigmoidoscopy [KDS20]. sign [CGC12]. sign-coherent [CGC12]. Signal [ZLJW23, Big13, NQdB^{+}07, WWL22]. Signal-noise [ZLJW23]. signaling [CC19, LZ11]. signalling [WYH^{+}14]. signals [FO11, FRBT13, HST19, JPTO17, LSS^{+}20, MHB^{+}09, WBC15, WOC18]. significance [ET07, Hua18, WT08]. significant [RAY14, WWM^{+}14]. silo [LFMM23]. similarities [AZM11]. similarity [MFB^{+}13, TLH14]. simple [AICV11, BG09, FRL18, SSZT19]. simplified [DKS18]. simulate [CTB17]. simulated [NCHJ13]. Simulating [RC23b, GBH^{+}15]. Simulation [HS09, BCA18, GGFG^{+}18, KKR13, LYBA22, Wan11, WYT^{+}20]. simulation-based [GGFG^{+}18]. simulations [FC20, PMMS16]. Simultaneous [CA18, Efr08, HLY^{+}21, LGL^{+}12, RG21, Thi11, AFS07, FM17, KHDV20, KHDV22, LCMJ11, SYZ11, SML^{+}21]. SING [RG21]. Single [MHG18, BBM20, CDF^{+}20, DK12, FYB^{+}15, FWK^{+}13, GMNP^{+}21, LWZ19, LZ11, MKKN21, MGSD19, NvdBCR23, WOC18, WLM^{+}21, WL22, YHE20, ZLDR18]. single-arm [WLM^{+}21]. single-cell [BBM20, CDF^{+}20, LWZ19, MKKN21, NvdBCR23, WL22]. single-locus [FYB^{+}15]. single-molecule [GMNP^{+}21]. single-trial [WOC18]. singular [MHB^{+}09, ZSH13]. SIR [OHC^{+}17]. site [YFM19, ZWS08, ZST14]. sites [HZL^{+}15, HBP17, MZA22]. Size [DPR^{+}20, BvdH09, DLKM20, Far22, HMP22, LFMM23, MRMB15, MKN22, Mur10, PHCM^{+}10, PNB22, RBB11, Sad18, SSD15, TL11, WMA^{+}14, vdHWC^{+}12]. sizes [FCC15]. skeptical [Bir08]. skewed [AT15, Goe11]. skill [FMBG15]. Skip [MM08]. skipping [RJP16]. slab [MRG21]. sleep [HAFFH21, JEAS09]. SLOPE [BvdBS^{+}15]. SLOPE-Adaptive [BvdBS^{+}15]. Small [KB10, MWP^{+}15, MNR14, ALC09, BL19, BK21, BvdB22, IHJ16, MRSA19, SF11]. small-area [ALC09]. smelt [ZGJ^{+}22]. smoking [LL16, LR20a, LR21, Ros16, Ros18]. smoking-attributable [LR20a]. Smooth [LAS16, SML^{+}21]. Smoothed [ZHB09]. Smoothing [CLK^{+}12, DPR11, CFH^{+}14, CLK^{+}13, DW21, MWP^{+}15, SMC^{+}20, WZF^{+}13, WFH^{+}22, WRSS15, WLP^{+}16, WLPP16, ZHB09, ZDL10]. SNIP [HPB23]. snow [BD11, WKR21]. SNP [DLL^{+}18, HVL14, KSW^{+}21, LGL^{+}12, LBL20, SPPR08, SSZT19, WIC^{+}10, ZLJW23]. SNP-set [KSW^{+}21, ZLJW23]. Social [GY23, ZHYS23, AS17, CBvdHvdH08, HG10, HWPH10, KM17, KBG21, NMW^{+}21, NS17, OV17, SHF^{+}16, SKS10, WWCZ22, ZKS15, vdKvEW17]. socioeconomic [Hua18, Tal15]. soil [CN07, MHH17, RCBB19]. Solar [HLK18, SSH^{+}11]. solution [FCPL19]. solutions [LPT^{+}11]. solve [WS10a, WS10b]. somatic [DTZP13, FGS^{+}10, OSB15, TP11]. Some [FRL18]. song [DMA19]. Sorafenib [SC16]. sorted [HPB23]. sound [DLS^{+}17, SJHJD20]. sounds [CAV^{+}19]. source [MSSS^{+}10, Moh13b, THSL12, WG23, WOH23, WBA^{+}14]. sources [BPSC14, CGCN22, EKW20, FMA18, HS22, HLK18, MHC15, PNB22, WESVS23, WKLvD16, YFHE20, YOZC23]. South [BJS^{+}22, MGM^{+}14]. Southeast [LGK18]. Southern [WTB16, KN13, UH20]. Space [DW21, MWP^{+}15, SG17, ASX13, BGH10, BCA18, CS13, CTB17, CSS11, DMA19, DKLL19, FZSI^{+}08, HHA15, HGS23, HS09, LPT^{+}11, MMGR21, NMW^{+}21, OW11, OHC^{+}17, STM17, SGL^{+}08, Ser11, SS15b, WCW15, WLK18, WYKH07, XFS10, ZBG14, ZCH^{+}16, vdBR10]. Space-time [DW21, MWP^{+}15, BCA18, CTB17, SS15b]. spaced [LCSZ15]. spaces [CD20, YR21]. spare [LNW08a]. Sparse [ACG13, GT10, HHLC16, JSZZ10, LHH10, LRZ08, LS22, ML13, PK11, PZ19, SGCW07, SWM13, ST11, WWL22, BTJ^{+}14, BR08, BC23, CLK^{+}12, CL13, FCPL19, GPR^{+}22, Hua19, HCRB23, JEK^{+}22, KG11, LNW08b, LCSZ15, MB08, Mur08, PRRW11, PDM19, QTL^{+}22, Qiu08, RHL^{+}22, RFB17, SJM^{+}14, Tib08, TvdL08, VRN^{+}11, WJF^{+}15, WFH^{+}22, YL11, ZRA^{+}21, ZW15, ZSCL23, ZZD22]. sparsely [PM08]. Sparsity [CGC12, AMB^{+}20, BKS21, KX12, LBD^{+}18a]. Spatial [ARK^{+}18, BJ19, BD11, BL11, CMAC^{+}23, FCGA^{+}13, FRL08, Hat14, Hav14, JLGJL12, LZW^{+}15, NSS14a, NSS14b, Pad14, Ste07, Ste09, SSH^{+}11, Wal14, WMT^{+}21, WFC^{+}22, YR21, BFM12, BZC^{+}19, BRG08, BWBS14, BdHZ08, CGW^{+}10, CKK^{+}22, CZM10, CLGK22, DSCS19, DJ11, FFM^{+}21, Fin13, FBM09, FMBG15, GGMG23, GS13, HHK^{+}16, HH21, HJS22, HBW17, HRFS19, ISR12, JRHM22, JYB16, KNWJ14, LYRR22, LL16, LCG09, LSL^{+}15, LBD18b, LZTB16, LZ07, MHH17, PBSVS23, RMP17, RF07, RB10b, RS12, RTB22, RCBB19, SFC11, SJH11, SRC15, SP20, SSD^{+}19, VIF13, WL08a, WKR21, WBKJ22, WTB16, ZHB09, ZSLH23, ZDL10]. spatial-temporal [WL08a]. Spatially [CR13, FSG16, GMLB^{+}14, GKZS12, LSS^{+}12, MHH17, RHHH18, SR23, ZBT^{+}20]. spatially-aided [ZBT^{+}20]. Spatio [ESF14, EJD19, FS13b, GKP^{+}16, BPSC14, BHW15, CD18, DAAM22, DBF^{+}16, LGK18, LCZ^{+}17, MLP^{+}19, OSL^{+}14, PYP^{+}09, SKS12, YBL^{+}17]. Spatio-temporal [ESF14, EJD19, FS13b, GKP^{+}16, BPSC14, BHW15, CD18, DAAM22, DBF^{+}16, LGK18, LCZ^{+}17, MLP^{+}19, OSL^{+}14, PYP^{+}09, SKS12, YBL^{+}17]. Spatiotemporal [KPDO23, LG20, PKGG23, ZZD22, ZX22, FCPL19, GDG^{+}16, HWK21, HISV15, JSH^{+}22, MK21, QWC17, REG^{+}11, WJT^{+}21, ZGV^{+}16, BC23, FC20, RK22]. Spatiotemporal-textual [ZX22]. SPDE [FC20]. speaker [FSJW11]. Special [Ano18, FGS08, Kaf11a, Kaf11b, Kaf12, CLEB14, JRHM22, Gne12]. speciated [SCRS^{+}20]. species [AMR16, BZC^{+}19, CFW17, CGW^{+}10, CDF^{+}18, FBM09, HWF15, JDP^{+}13, ZW07]. specific [CJM^{+}17, CT07, GL08, KAGK^{+}23, LN12, MAM17, PAS23, QWC17, RSI16, RSD22, SMW^{+}22, SM13, SMC^{+}20, YKHS21, vdKvEW17]. specificity [dCdCAGM16]. specificity-ROC [dCdCAGM16]. spectra [KP15, KS17, WKLvD16]. spectral [EKO22, GGMG23, HKP^{+}19, KH13, LLR10, LSS^{+}20, QW08, SBS14]. spectrometric [PHT15]. spectrometry [KPA^{+}10, KOJ^{+}14, LMS10, LES12, LLM20, MAZM13, OGP^{+}18, Ryu22, WKG^{+}15]. spectrometry-based [KPA^{+}10]. spectroscopy [COM22, HCW11]. speech [SM20a]. speed [BCA18, HMT12, SM15, ZBG14]. spend [CA22]. spending [ML14]. sphere [FPLM18]. spherical [FDKP13]. Spike [MRG21, KKL11, LL11, MP11, VVSK18, WCW15]. Spike-and-slab [MRG21]. spiking [GTZ^{+}21]. spills [AMR16]. splice [GZB^{+}11]. splice-junction [GZB^{+}11]. splicing [RAKS14, RAKS15]. spline [AN14, LBND13]. splines [LvdVvWvdW13, Mey08, SK22]. Split [SHW18]. Split-door [SHW18]. splitting [GBMRR20, MAB^{+}14]. sporadic [WWCZ22]. sport [LMB18, MV12]. sports [SBD23]. spot [Lie13]. spread [CD17, MH14]. Spurious [Tin11]. sputum [SRA^{+}15]. square [SPF20]. squares [ACG13, DMVT23]. stability [AS23, GGQY07]. Stabilizing [PWP^{+}21]. stable [RS12, SYZ15]. stacking [CPG^{+}21]. stage [BRG08, JB21, MHG18, SIL^{+}11]. stagnation [HCYH20]. standard [WZ18]. Standardization [LS18, KBMF^{+}23, Mar08]. standardized [HPF13]. star [LCB16]. stars [HLY^{+}21]. State [FZSI^{+}08, LPT^{+}11, Yua09, ASX13, GFW^{+}09, HS09, JAZ15, KAvdW^{+}23, MMGR21, MLCW13, MR15, NDRF17, OHC^{+}17, SM13, WMKG19, WYKH07, XFS10, ZCH^{+}16, vdBR10, Dup17]. State-space [FZSI^{+}08, LPT^{+}11, OHC^{+}17, XFS10, ZCH^{+}16, vdBR10]. state-specific [SM13]. States [RY11, HE15, RHC23, ZS18]. Static [LCZ^{+}17, KH13, SJA^{+}13]. station [NMD19]. stationarity [AK12]. stationary [XDO10]. statistic [LT11, LZ07, ST14]. Statistical [Ben08, Bir08, CM09, CW20, CN07, DKS18, Feu08a, Feu08b, Feu13, FGS^{+}10, Fuc08, GREG15, GBNS22, HW08, HF20, Ing08, Kaf12, KP15, Men18, MV08, NPM12, PK18, QYP09, Sti08, SKKS14, SKZ14, TETJ17, XKG^{+}19, YWL^{+}12, ZMB23, ZCG^{+}09, vDDS^{+}09, AFS07, BNW08, Ber11, BFF^{+}09, CLW20, CAS20, CR11, DL11a, ESO19, FMB^{+}12, FH13, HU11, Hol11, Kap11, KY07, LC10, LYH^{+}16, Lyo08, MAE^{+}08, MM11, MW11b, NL11, OMM^{+}14, RSI16, RGF^{+}20, Rou11, SFPS^{+}21, SMR11, SGCW07, Sme11, Sta08a, Tin11, WA11, WSK^{+}21, WYKH07, WI07, YE14, ZOZ17, ZMA^{+}20, ZLJW23, ZLDR18, PK19]. statistical-physical [MAE^{+}08]. statistically [DSC^{+}23]. statisticians [Cra16, JJ16a, JJ16b, KP16, KT16, RRS16, Sil16, WR16]. Statistics [Fie07, Fie08, Bic10, Cox07, DKZ09, EOB21, FGS08, Goe14, GY23, HTP14, Kaf11b, RLH^{+}13, SM15, WCW15, ZYXS16, ZZ08, Zho17b, ZS17]. status [HCYH20, MGM^{+}14, SBD23]. steady [MLCW13]. steady-state [MLCW13]. steel [MSJ14]. steeply [KS17]. stellar [JSF^{+}22, vDDS^{+}09]. stem [FGA09, PBS^{+}23]. step [LM10a, LM10b, SYZ15]. Stephen [Ano18, Rub18]. stepped [JFRS17, WDSJ23]. stepped-wedge [JFRS17, WDSJ23]. steps [RAY14]. stereological [MSJ14]. stereotypes [DFGY23]. sticky [FSJW11]. stillbirth [SMC^{+}20, WFH^{+}22, WCW^{+}22]. stimuli [GTZ^{+}21]. stimuli-bundle [GTZ^{+}21]. stimulus [KPC^{+}19]. Stochastic [BCA18, CD17, Kou08, SCA13, SAV^{+}14, AY12, BL11, CGM17, CW10, CGCN22, DS14, FGA09, HCS18, JCCG18, JLL09, KB10, LBA11, LRDD22, LKB21, LYBA22, NMW^{+}21, PC20, PM08, RB10a, RB11, RSH12, RHHH18, SS15b, TMvD^{+}17, TDBM23, TJW10, URZF21, XZC17, ZJLC08, ZK10, KN17]. stochasticity [GTZ^{+}21]. stock [CCdCW18, FFJJ14, JSX16, RCBB19]. stop [GRS16]. stop-and-frisk [GRS16]. stopover [WMKG19, WHAW21]. stopping [GEF22]. stops [DAAM22]. storage [JCCG18]. storms [ESF14]. Strategies [ZPMA10, CH14, DMVT23, FFM^{+}21, LN12, RCF^{+}13, SB20, ZZTL22, ZLR20]. strategy [HGM15]. stratification [CGI08, GMM08, KDH^{+}19, LMM15, MLM13, SP19, Sco09, VFH16]. stratified [BFM12, DLKM20, LYBA22, Ros18]. Streaming [SBW^{+}09, HAL21, IGA22]. streams [RFB17]. street [FS13a]. strength [RHR12]. Strengthening [KM16]. strike [ZTH19]. stroke [LRMM15, QWC17]. strong [KM16]. Stronger [ZSG^{+}13]. structural [BGK^{+}15, CB22, HSFP11, LBD^{+}18a, MB22, RGPC19, STA18, SzCT10, VIF13]. Structure [IHJ16, NvdBCR23, CHS^{+}16, EHM18, FMBG15, LDV^{+}10, MBYWX19, MRV10, MDWH21, NCHJ13, OW11, PK11, RS14, SRC15, SHM15, WZD19, WG23, ZHM^{+}19]. Structured [BML^{+}20, DB22, PRRW11, WZ17, YJZ09, CLK^{+}12, Fuk19, HSSF21, JND12, KX12, KKLS15, KKLS16, LL10, LS22, MM15, NV18, ZYC^{+}17, ZHYS23]. structures [COM22, GCC^{+}11, KY07, MTZZ21]. student [GPRR16, LMM15]. students [MIP22]. studied [RSH12]. studies [AMR18, AMGG13, AD22, BHC^{+}20, BAH22, BDL^{+}16, CW20, DHG19, FH19, GBST19, GM08, GMM08, GS11, HY14, HVL14, JCJ20, JL11, JHMC16, KH23, KSD11, LBK^{+}23, LT11, LGL^{+}12, LZLW14, LWLW15, LBL20, ML23, MCCW09, MSH21, MLX23, OMM^{+}14, PL11, PDS13, RN14, RD14, RG23, SRA^{+}15, SSZT19, SRH16, SL20, SL19, ST11, SCK19, SH11, SW10, TMY17, TTB22, TP11, URZF21, WDSJ23, WIC^{+}10, WHC^{+}22, WHNW15, XCS11, YSR21, YWQG23, ZLS^{+}17, ZYC^{+}17, ZCRC18, ZLZB18, ZMA^{+}20, ZRCC21, ZZTL22, ZSS23, ZLD12, Zho17b, ZS17]. Study [BDR16, OSL^{+}14, AXEC18, ANFM09, AS23, BHC^{+}20, BDC^{+}11, CFRW19, DK18, FWGS11, FW21, FLHA15, HT08, HL08, HEHM23, Hun12, JEAS09, JLS^{+}17, JD18, JGC^{+}18, KDS20, KZ16, LHPW13, LMKC12, LRM17, PZB^{+}10, RRSM18, Ros12, Ros16, Ros18, SC16, TACH21, TWHP15, WACY20, YSR22, YR21, YKLK23, ZY12, ZASM12, ZSLH23, ZSCL23, ZMB23, ZSG^{+}13, ZPR14]. studying [HCD^{+}21, RGPC19]. stunting [KKMS16]. style [CD12, RB10a, RB11]. styles [YH20]. Subbotin [CA23]. subclassification [ZMB23]. subclonal [HSSF21]. subclone [ZWZ19, ZSMJ19, ZSMJ20]. subcomposition [WZ17]. Subdiffusion [Kou08]. subgraph [JLB^{+}14]. Subgroup [STMC17, ZZ22, ZSFS22, SHH22]. Subgroup-effects [ZSFS22]. subgroups [SF11, WS14]. Subject [PAS23, MBDL14, MAM17, SVYP11, SCDD18, ZGV^{+}16]. Subject-specific [PAS23, MAM17]. subjective [FW21, YMP11]. submatrices [SWPN09]. subnational [BNMG23, DW21]. subnetworks [ZKY14]. subpopulations [SKAL19]. Subsampling [BBB^{+}10, WEWX21]. substance [GMM08]. substitutes [RAB20]. substitution [LL09]. subtype [LRI21]. success [HMP22]. suicide [WACY20]. sum [CQ09, JL11]. summaries [JMY^{+}14]. summary [CD20, WS10c, WCW15, YOZC23, Zho17b, ZS17]. summer [PPLK18]. SUP [CHH^{+}14]. super [GMNP^{+}21]. super-resolution [GMNP^{+}21]. superhighways [BMLG21]. superior [PS12, Ros12]. supermarket [SP20]. supernova [GFW^{+}09]. supervised [CLGK22, WBC15, XDO10]. supplemental [CT07]. supply [Gho10]. Support [Deb09, HS13, KZS23, MVV13, PA23]. supported [KBH^{+}11]. Supreme [RY11]. Surface [GDJR20, Ber11, CR11, DCHP21, DL11a, HU11, HH21, Hol11, Kap11, MM11, MW11b, NL11, PTGN12, RF07, Rou11, SMR11, Sme11, Tin11, WA11]. surge [PACB21]. surgery [TACH21]. surgical [LHH19]. Surrogate [HYL23, HZG22b, QGFL08, ZGJ^{+}22]. surveillance [LGL^{+}18, LQNM19, SKKS14, XQ23]. Survey [McE09, AT15, ACS^{+}23, BBL22, CBvdHvdH08, DW21, DH18, Gau11, HTP14, JGF08, MV14, ML14, MWP^{+}15, RFWE22, SGC23, SS10b, SS20, YSL08, YBL^{+}17, vdBR10, BBL22, JRHM22, Sav16]. survey-based [BBL22]. surveys [CA18, EOB21, FFM^{+}21, FT18, PKP16, TSG17, WLA^{+}21, ZTCS20]. Survival [FSM^{+}19, JL11, AL16, HSFP11, IKBL08, JLL^{+}19, JLS^{+}17, JLRK23, KXC09, LTL19, MHC15, NWJ20, QDN^{+}21, Sin09, SCW^{+}23, TLH14, WACY20, YWQG23, ZHJZ15, ZZXL23]. survivor [HSFP11]. susceptibility [WHC^{+}22, ZMC^{+}21, ZCG^{+}09]. sustained [PS12]. SVD [OP09]. switching [CFMR18, EJD19, MMGR21, PGL^{+}19, SRCK16]. symmetric [FLP23]. symmetry [BHP10]. synchrony [KKL11, OIHH09]. syndrome [dCdCAGM16]. Synthesis [YMP11, BBB^{+}18, CPvV^{+}11, FK22]. Synthesising [PPB^{+}14]. Syrian [CSS18]. system [HCD^{+}21, JWL^{+}19, MLCW13, MVV13, SP19, WD10]. systematic [RWK17]. systems [BCJ15, CCS18, DT23, ERM15, NKAY10, PPB11]. Székely [New09].

T [WYKH07, ZS09]. T-wave [ZS09]. Table [Ano16d, Ano16e, Ano16f, Ano23b]. tables [JGF08, Kip22, OMM^{+}14, Thi11]. tabulations [JRHM22]. tail [GH22, SCDD18]. tailed [PS12, VDP08].

Tajima [CP20]. taking [DGCT10]. tantrum [QYP09]. tapering [CCD22]. target [FZSI^{+}08, NECS17, NS20, WP12]. Targeted [SML^{+}21, SvdLMP14, SMC^{+}20]. Targeting [RLH^{+}13]. tax [KGGQ15]. teacher [DK18, ML11]. teaching [LSM15]. team [LMB18, PLCX23]. technique [CCS18]. techniques [DPR11]. technological [KPA^{+}10]. teenage [JAZ15]. Teesta [JSR16]. Telecom [ZST16]. temperature [Ber11, BB11, CMAC^{+}23, CR11, DCHP21, DL11a, GDJR20, GS13, HU11, HH21, Hol11, Kap11, KÓ14a, KKR13, MM11, MW11b, NL11, Rou11, SMR11, Sme11, Tin11, WA11, WRNR14, WTB16, ZHO22]. Temperatures [PMMS16, BLTV14, Ber11, CR11, DL11a, DSB19, HU11, Hol11, Kap11, MM11, MW11b, NL11, Rou11, SMR11, Sme11, TAC^{+}16, Tin11, WA11]. temporal [BPSC14, BHW15, CD18, DAAM22, DBF^{+}16, ESF14, EJD19, FS13b, FRL08, GKP^{+}16, GS13, HHHV17, HCD^{+}21, DFGY23, KÓ14a, LL16, LSL^{+}15, LGK18, LCZ^{+}17, MLP^{+}19, OSL^{+}14, PYP^{+}09, PMMS16, QBC13, SKS12, WFH^{+}22, WL08a, YBL^{+}17]. Tensor [LZ21, MZI18, SMZ21, DKZ09, Hof15, ML23, WFS19, ZGS^{+}14]. Tensor-variate [SMZ21]. tensors [YZS^{+}13]. tenures [PG13]. term [DB15, EJD19, HL08, RDL23, SKS12, ZBG14]. terminal [WZLP20]. terminology [SPH17]. terrain [KKR13, SHSZ19]. terrorism [SM20a]. terrorist [CD18, CW13a, CW13b, CRZ13, Gil13, Moh13a, PW12, RGT13, RVW20, RP13, Sch13, Whi13]. tertiary [ST11]. Test [WILW22, BBM20, BMH16, Bro08, CWE18, DHM^{+}17, GBST19, LZCW21, LX18, MKS^{+}14, OSB15, Ros12, SSD^{+}19, SPF20, SHH22, TMN18, WDSJ23, WGL^{+}18b, YLH07, YHX13]. test-negative [WDSJ23]. testicular [VKG12]. Testing [BAH22, BFF^{+}09, CLEB14, FDKP13, KBB^{+}11, PG14, WT08, ZLDR17, dCP10, AXEC18, BW18, BNW08, BST15, BZN18, CdVM^{+}22, CSZK14, DLKM20, ET07, Efr08, FFR^{+}08, FS14, GBST19, GLB^{+}17, GGQY07, GKZS12, HKP^{+}19, JEAS09, Lie19, LZP16, PLCX23, SJGM13, Sco09, SPsLC16, SDH18, SZO12, WBB13, WWM^{+}14, ZSS23, ZWW13, ZCG^{+}09]. tests [AABC^{+}19, AHZ23, BYZ18, CLM22, CLTZ22, DK18, Hua18, JND12, KM16, KHZK23, NPM12, OW11, SCDG17, STD13, WCL23, YZAD13, ZMLS22, ZLJW23]. text [JMY^{+}14]. texts [GGG^{+}12]. textual [ZX22]. TFisher [ZTLW20]. their [CGT^{+}14, DCHP21, DL11b, FDH10, KY07, LLKP18, MGM^{+}14, YOZC23]. theme [YD23]. theoretic [YH13]. Theory [CWE18, KN20, KHZK23, WZS19]. therapeutic [SHAB22]. therapy [PHWM11, YSR21]. threads [ZPBW^{+}18]. Three [WFS19, HBW17, KAvdW^{+}23]. three-cube [HBW17]. three-state [KAvdW^{+}23]. Three-way [WFS19]. threshold [BYZ18, SD10]. thresholds [DHG19, KN20]. throughput [BC09, DBTP21, LBHB11, LS22, SGLB10, SPPR08, SS15a]. tiling [JLL09]. Time [BHIK09, CCdCW18, FCPL19, HGS23, PPM14, AMR18, AS10b, BJ19, BLM^{+}23, BPS22, BKS21, BGH10, BCA18, BWT^{+}20, BvdB22, BGK^{+}15, CS13, CTB17, CW10, CGCA21, CCJ^{+}09, CLR16, CSS11, DB15, DLKM20, DGL13, DW21, EKO22, ENF14, ENH^{+}18, ESO19, FS13a, FZZW17, FZCV22, FHSJ14, FSPWWE18, FRBT13, GMMW17, GRS23, GSC^{+}20, GMB15, GV14, HSFP11, HHA15, HS09, Hun12, JLL09, JPS21, KAvdW^{+}23, KSAX10, KHBV20, KH13, LZK^{+}15, LMW10, LLKP18, MJ16, MB22, MWP^{+}15, PA23, PDM19, PL08, QW08, QYP09, RAY14, RHR12, RS09, SGL^{+}08, SJM^{+}14, SM20a, SRZ^{+}15, STA18, SMW^{+}22, Ser11, SH08, SPsLC16, SG17, SW17, SHM15, SBS14, SH11, SS15b, TMvD^{+}17, TSY22, TCW21, Tin11, TFB14, TB22, WLL17, WLG17, WL08a, WZLP20, WWMH13, WK10, WMA^{+}14, WYKH07, XZC17, YFHE20, YLG15, YLC^{+}17, ZD13, ZW19]. time [ZBG14]. time-course [FRBT13, SHM15, TCW21, ZD13]. time-dependent [SW17]. Time-discretization [HGS23]. time-series [BvdB22, BGK^{+}15, SJM^{+}14, STA18]. time-to-event [AS10b, GSC^{+}20, TFB14]. Time-varying [CCdCW18, BPS22, BKS21, CLR16, DLKM20, ENF14, FSPWWE18, KSAX10, KHBV20, LMW10, PDM19, SW17, TB22, WLG17, WK10, YFHE20]. Time-warped [PPM14]. timeliness [Kan20]. timeout [GEF22]. times [CMPR22, KPC^{+}19, TFB14, WACY20]. TIMSS [GPRR16]. tissue [OMM^{+}14, PHT15, WFS19, YWL^{+}12]. TOF [HCW11]. tomb [Feu13, Fie08]. tomography [DLZL16, Haz15, SSH^{+}11, XFS10, ZCS13]. tool [MIP22]. toolbox [ISR12]. top [HLK18, RVW20]. top-down [HLK18]. Topic [TCZ16, BL07a, BL07b, MHG18, MMBL20, RHHH13]. Topic-adjusted [TCZ16]. topical [JLLK20]. topics [ZPBW^{+}18]. Topological [KF10, SC23, WOC18, MLX23, WSU^{+}19]. topology [DLZL16]. Torus [EHM18, KOB^{+}20, JPK21]. total [Far22, LR20b, RRSM18, Ste07, SHR^{+}22]. tournaments [MV12]. toxicants [BZS19]. Toxicity [PTGN12, LYY13, LKTJ^{+}15]. TPRM [MZI18]. trace [PK18, PK19]. tracer [OMM^{+}14]. traces [SAV^{+}14]. Tracking [DH18, MN15, PBS^{+}23, DB15, LLS^{+}22, MGMB19, QHPD19, SFDMM22, SKKS14, WLP^{+}16, WLPP16, XKG^{+}19]. tracks [LZTB16]. tracts [YZS^{+}13]. trade [FK22, WH11]. trade-off [FK22]. traditional [RFWE22]. traffic [BMLG21, Chi12, CCH19, Haz15, IGA22, LLR09, MGRG^{+}23, PS15, ZGLH13]. train [WCW15]. training [DK18, MLM13, PHT15, SSD15, SLBL23, YWL^{+}12]. trains [KKL11, LL11, MP11]. trait [TLH14, WWL22]. traits [LC10, WGL^{+}18b, ZLZB18, ZNB^{+}21]. trajectories [ACS^{+}23, Ger09, SKKS14, SXC^{+}20]. trajectory [MV14, MGMB19, WNX^{+}23]. transcript [WZ16]. Transcription [ZWS08, FWK^{+}13, LW18, WP12]. transcriptome [LSL^{+}15, MSH21, TCW21]. transcriptome-wide [MSH21]. transcriptomic [HST19, LT11, SR23, TTH21]. transect [JL10]. transfer [HBW17, JGVM18, KN17, PSL^{+}16, SH18]. transformers [HMM09]. transgenerational [HZL^{+}15]. transient [ÇL12, PMMS16]. transition [HCYH20, SB20]. transitional [YKHS21]. translate [MM15]. transmission [CD17, FGS^{+}10, PHLH12, RTB^{+}21, WOK^{+}16, YLH07]. transmitted [AH16]. transplant [AL16, ZLR20]. transport [FBH23, PKGG23]. transportation [KSD11, LSAR12]. Transposable [AT10]. traps [ARK^{+}18]. Travel [WWMH13]. treatment [CAS20, DLL^{+}18, DMVT23, DTL^{+}23, FLS16, FHI18, GMM08, HBHM13, HHH10a, HZF22, IR13, JLS^{+}17, JCK22, LN12, MSG^{+}20, NECS17, NS20, STMC17, SSL^{+}10, SGNM22, SML^{+}21, STD13, SHH22, TWA18, VGH14, YWB^{+}23, YLG15, ZLR19, ZZ18, ZZ22, ZHFN23, ZSFS22, ZB11]. treatments [DLKM20, Fre08, HZG^{+}22a, LL19, Ros12, STMC17, TB22]. treaty [JSR16]. Tree [KX12, NV18, TWA18, WZHC12, BZC^{+}19, FBM09, GGG^{+}12, GH12, HR22, KHZK23, LNC^{+}19, LKTJ^{+}15, MM22, PT12, PLCX23, Pur11, SCW^{+}23, TMN18]. Tree-based [TWA18, KHZK23]. Tree-guided [KX12]. TreeClone [ZSMJ19]. treelet [CLK^{+}13]. Treelets [BR08, LNW08a, LNW08b, MB08, Mur08, Qiu08, Tib08, TvdL08]. Trees [ZLR19, APW^{+}09, BHB^{+}21, BMM^{+}16, BFF^{+}09, CGM10, DKS18, FIM^{+}21, GTW13, HTM^{+}13, Loh09, LZ13, RLH^{+}13, RHHH13, SPI^{+}23, WOK^{+}16, YWB^{+}23, ZBC16]. trend [BGC20, DB15, GKS17, JMJ^{+}21, PL08, VFH16, ZD13, ZGM23]. trend-cycle [PL08]. Trends [GV14, BJ19, Bro09, GKZS12, IWG13, LL16, McE09, RHR12, SFGLR15]. Trial [SW17, DLKM20, EKO22, PBS^{+}23, QM23, STMC17, SSL^{+}10, SHC12, STD13, WOC18, YFM19]. trials [DHL18, DK18, HZF22, JFRS17, LY16, LSY^{+}22, NZRC13, QGFL08, WLM^{+}21, XLDO13, YY11, ZZTL22, ZHFN23, ZZ08]. tribal [CGM17]. trigger [OBHL22, SM20a]. triggered [SCW^{+}23]. trimmed [ACG13]. tropical [LGK18, MVW^{+}23, WFHZ23]. trumps [Rub08]. Truncated [WCL23, BvdH09, BvdH19, CGN22, CVF10, HMM09, TDS^{+}14]. truncation [ZTLW20]. Trust [MSS09]. Truth [WSGH12]. trying [Far22]. tsunami [GSD^{+}18]. tube [GCC^{+}11]. tube-fitting [GCC^{+}11]. tuberculosis [SRA^{+}15]. Tumor [ZSMJ20, CPP^{+}14, CDP^{+}17, LMGJ15, LWLX19, MBK^{+}21, XZX18, ZWZ19, ZSMJ19]. tumors [OSB15]. tuning [GTW13, KSH^{+}13, RMP17, SWHO11]. turbine [LKB21, MR15, SDH18]. turbines [LBND13]. Tuscany [AM16]. tutoring [SP19]. tweets [ZFB14]. Twitter [CT18, MM15, MMBL20]. Two [BKS21, CK14, FFM^{+}21, JB21, LBD^{+}18a, SIL^{+}11, ZMLS22, AMGG13, BRG08, CSC^{+}12, CWE18, CLM22, FGA09, GM15, JAZ15, KOJ^{+}14, LAS16, SYZ15, SCDD18, THSL12, WCL23, WBKJ22, ZBLC17]. two-channel [CSC^{+}12]. two-compartment [FGA09]. two-dimensional [KOJ^{+}14, LAS16]. Two-level [LBD^{+}18a]. two-part [WCL23]. Two-phase [CK14, FFM^{+}21, AMGG13, ZBLC17]. Two-sample [ZMLS22, CWE18]. Two-stage [JB21, SIL^{+}11, BRG08]. two-state [JAZ15]. two-step [SYZ15]. Two-way [BKS21, CLM22, THSL12]. type [CFH^{+}14, FGMP16, HHHV17, HGB21, KNWJ14, LFMM23, NPM12, YLS14, ZZL11, ZRCC21, ZNB^{+}21, GFW^{+}09]. types [HGM15, JD18, LHMN13, WJT^{+}21].

U.S [LRM17]. U.S. [BBB^{+}18, YR21]. UK [DMGJ20, QTL^{+}22]. ultra [SCL^{+}13]. ultra-high [SCL^{+}13]. ultrafine [FZZW17]. ultrahigh [CLR16]. ultrahigh-dimensional [CLR16]. ultralong [CLFC23]. Unbiased [WDSJ23]. uncertain [WACY20]. uncertainties [DCHP21, YOZC23, YTHY18]. Uncertainty [CFMR18, LBBM21, AS23, BM22, BBDP11, DGCT10, FSG16, KP15, KS17, LR20b, MSS09, Sad18, SHC12, WFHZ23]. Uncovering [MRV10, MMGR21]. underground [ERM15]. underlying [LC10]. understand [PHLH12]. Understanding [ZST16, KM17, LMB18, Sun22, GRS16]. unemployment [MRSA19]. unfolding [KP15, KS17]. unified [LMB18, NKAY10, Zho17b, ZLDR18]. uninsurance [RHC23]. Unique [CSS18]. unit [PHJ22]. unit-level [PHJ22]. United [CHJCK18, RY11, ZS18]. univariate [BD22]. university [LMM15, MIP22, SMZ21]. unknown [LXC11, MBGDS11, WZ18, ZGM23]. unlabeled [CDB11, MPT12, TON20]. unlabelled [WLML23]. unmarked [CR13]. unmatched [DK18]. unmeasured [FH19, HBHM13, HZG^{+}22a, Mar08]. Unmixing [BMH16]. unobserved [NECS17, NS20, OBHL22, SP20]. unordered [BR08, LNW08a, LNW08b, MB08, Mur08, Qiu08, Tib08, TvdL08]. unstable [BJ09]. Unsupervised [FFR^{+}08, HAL21, LRI21]. update [Feu13]. updating [IHJ16, MDR10]. upgrades [SDH18]. uplift [BMGN21]. upon [RBB11]. urban [BC23, GF19, HGSJ23]. Uruguayan [LSS^{+}12]. Usage [Goe14, ZST16]. Use [MHB^{+}09, CWE18, DMVT23, FS13a, MK21, MMGC22, SKBL23, WR12, ZNSL14]. used [Sch15]. uses [WHLN15]. Using [HK23b, HGSJ23, JD18, MLX23, OSB15, Ros16, WS10c, WMA^{+}14, vdBN09, AS10a, ARC07, AN14, AMB^{+}20, BLM^{+}23, BLTV14, BL19, BRG08, BCA18, BK20, BGC20, BGK^{+}15, CKHP15, CJMF18, CHAP16, CKK^{+}22, CGCA21, CQ09, CDP^{+}17, CFH^{+}14, CCD22, CZM10, CB22, CSL^{+}08, CH14, CAL^{+}23, CLK^{+}13, CBvdHvdH08, CDB11, DBG21, DPR^{+}20, DZ23, DH18, EKW20, ELD09, EFDS20, FDKP13, FYB^{+}15, FD20, FD11, FWK^{+}13, FDR16, FND09, FMA18, FSPWWE18, FRL08, GGPM16, GWZ19, GGMG23, GFW^{+}09, GH12, GSC^{+}20, Gil17, GTW13, GKS17, GPBT22, GM15, HAFFH21, HH21, HS13, HCYH20, HCD^{+}21, HYL23, HWF15, HBP17, HCRB23, ISR12, JPTO17, JYB16, JGC^{+}18, JLGJL12, JLRK23, KKL11, KBH^{+}11, KMKB16, KN20, LdGK^{+}17, LRMM15, LMS10, LES12, LMM15, LSL^{+}15, LWZ19, LPT^{+}11, MAE^{+}08, MRMB15, MRW09, MC17, MGM^{+}14]. using [MB22, MCCW09, Mey08, MAB^{+}14, MKN22, MBH^{+}11, MKM23, MNB^{+}12, MHC15, NCHJ13, NMW^{+}21, NWJ20, PHWM11, PQR21, PKGG23, PTH22, PG14, PT12, PPLK18, PACB21, PS12, QW08, RKM^{+}23, RSI16, RTB^{+}21, RDL23, RTB22, RC23b, RY11, RODC19, Sad14, SPPR08, SGCT17, SRC15, SLZS08, SG16, SFB16, SHC12, SC14, SHGA10, SS20, SJHJD20, SCW^{+}23, Tal15, TDBM23, THN^{+}19, TLH14, UH20, URZF21, WJF^{+}15, WSM^{+}16, WSU^{+}19, WFS19, WLM^{+}21, WFH^{+}22, Wen16, WWMH13, Wit11, WOK^{+}16, XZX18, XDM15, YL13, YLS14, YLLS21, YLG15, YLL12, ZLR19, ZWZ19, ZBC16, ZSP19, ZS18, ZSG11, ZHJZ15, ZM16, ZSMJ19, ZZD22, dCP10]. uterine [FCC15]. utero [BZS19]. utility [FK22, KY07, Sin09]. utility-risk [FK22]. utilization [SPH17, WJT^{+}21].

V1 [VRN^{+}11]. vaccination [DHL18]. vaccine [BHC^{+}20, DHL18, DHG19, HCD^{+}21, KH23, QGFL08, YGLH08, YWQG23]. Vaccines [OV17]. valid [AHZ23]. validation [GSC^{+}20, HT08, NCHJ13, OP09, RGSB^{+}18, SWHO11, TT09, TWHP15]. value [AL16, BYZ18, CCdCW18, CA23, DT23, MSSS^{+}10, MHB^{+}09, ML11, RCF^{+}13, REFT18, SW10, WTJ10, WD10, ZSH13, ZBLC17]. value-added [ML11]. valued [AMR18, FTE^{+}21, Haz15, LLKP18, MRV10, SDP22]. values [CDN12, LR20b, RAY14, SPS20, TDS^{+}14, ZTLW20, dCP10]. variability [KH13, RD14, DPT22]. Variable [BKGJ14, CL13, CFRW19, FSM17, GPRZ17, GTW13, LL10, LWFW16, MDR10, ZZ18, BZN18, BvdBS^{+}15, DBG21, DVF13, ENH^{+}18, FND09, FKSBS19, GEC13, GS11, HLY^{+}21, HYL23, JCS07, JTLE22, KS19, KZS23, LZW^{+}15, LCB16, MGM^{+}14, NWJ20, PWP^{+}21, RJ11, SP19, SGNM22, VFH16, WYT^{+}20, WCH^{+}23, WCW^{+}22, YJZ09, ZZ22, ZHM^{+}19]. variables [CKHP15, CLZ09, CGC12, GT10, Goe11, JCK22, KGGQ15, KKMS16, KBG21, LMMS21, SH18, Sch15, SZO12, TWZ15, YLS14, YKLK23]. Variance [KFB11, MAZM13, AXEC18, BY13, GDJR20, KSW^{+}21, LXC11, MKS^{+}14, RS10, SGC23, VGH14, Zho17b]. variant [SYZ11]. variants [CGT^{+}14, FYB^{+}15, ZWW13]. variate [SMZ21]. variation [HCD^{+}21, LHMN13, Ste07, ZJLC08, FGMP16]. Variational [CLFC23, CMR18, CWE18, FCGA^{+}13, MRV10, WME17, WYW^{+}23, ZYFF19]. variations [NZ12]. Varying [YZS^{+}13, BPS22, BKS21, CCdCW18, CLR16, CLLR20, DLM14, DLKM20, ENF14, FSPWWE18, GPRZ17, GMLB^{+}14, KSAX10, KHBV20, LGL^{+}18, LWLW15, LMW10, MKS^{+}14, MHH17, PDM19, RHHH18, Ser11, SW17, TB22, WLG17, WK10, YFHE20]. varying-coefficient [GPRZ17, LWLW15]. VCSEL [KSW^{+}21]. Vector [Deb09, FTE^{+}21, HS22, HR22, HKP23, LG20, LSS^{+}12, PA23, ZPGO21]. vector-based [ZPGO21]. vector-borne [LSS^{+}12]. vector-valued [FTE^{+}21]. vegetation [CSZK14]. velocity [HCKFZ21]. verbal [KLCM20]. verification [BBB^{+}18]. Vertex [FLP^{+}15, WL10]. via [ASX13, AK12, BYZ18, BPS22, BvdBS^{+}15, BDR16, BHIK09, CFLP15, CHS^{+}16, CGC12, CDN12, CWS15, DTL^{+}23, FFW09, FHSJ14, FP08, GDJR20, GMLB^{+}14, GREG15, HZL^{+}15, HGRS17, Hua18, HIH^{+}21, JND12, KBMF^{+}23, LRZ08, LZP16, LSS^{+}12, MAM17, MG22, MHK22, PTGN12, SH18, SPH17, SHH22, WNZK14, WYT^{+}20, WMKG19, YJD21, Yua09, ZJLC08, ZK10, ZYFF19, ZBT^{+}20, ZSG^{+}13]. victims [PG13]. video [SKKS14, SHR^{+}22]. view [CMJ09]. viewers [YSG16]. views [Kad08]. VIF [DVF13]. violations [Sad18]. violence [FLHA15, YSL08]. viral [BBM20, HJS22]. virological [CGCN22]. virtual [FFJJ14, SML^{+}11]. Virus [XDM15, CCJ^{+}09]. visibility [TCZ16]. visitor [DAAM22]. visual [SF11]. visualization [ESO19]. Visualizing [GHK^{+}13]. vivo [XKG^{+}19]. volatility [CGM17, RB10a, RB11]. volumes [GEC13]. voluntary [VFMD17]. Voronoi [BWBS14, PG14]. voters [RLH^{+}13]. votes [DMA19]. voting [GQ10, GCL^{+}15, YR21]. Voxel [OMM^{+}14]. Voxel-level [OMM^{+}14]. vs [SML^{+}21]. vulnerability [LSS^{+}12].

wage [HCYH20]. wake [LYBA22]. war [RHHH13, Gil17]. warped [PPM14]. warping [CMZ19, ZM16]. warranty [SHM20]. water [JSR16, MHH17, REFT18]. water-related [REFT18]. wave [JLGJL12, RAY14, TETJ17, ZS09]. Wavelet [HKP^{+}19, RHZ^{+}15, SS15a]. Wavelet-based [SS15a]. Wavelet-domain [RHZ^{+}15]. wavelets [WNZK14]. waves [SRCK16]. way [AWL13, BKS21, CLM22, JGF08, JSR16, THSL12, WFS19]. weak [ENH^{+}18, FRBT13, KM16, WSH^{+}14]. Weakly [WBC15, GJPS08]. wealth [Gau11]. weapon [CLEB14]. wearable [ZMLS22]. weather [CLW20, LGK18, Sun22]. weathering [DHM^{+}17]. web [IGA22]. wedge [JFRS17, WDSJ23]. weight [LCYZ23, WGL^{+}18b]. weighted [DMVT23, LT11, MJ16, XDM15]. weighting [LL19, TACH21, ZTLW20]. weights [GM16, KBMF^{+}23, MKN22, RHC23, VFH16]. Weinberg [ZSS23]. West [WKR21]. Western [MBD11]. whale [CWH20, DLS^{+}17]. whales [YBL^{+}17]. which [NZRC13]. white [YZS^{+}13]. who [LBK^{+}23]. whole [CJM^{+}17, WYW^{+}23]. whole-brain [WYW^{+}23]. wide [BDL^{+}16, GS11, HY14, Hua19, JHMC16, LZLW14, LWLW15, LX18, MSH21, SSZT19, ZLS^{+}17, Zho17b, ZS17]. WikiLeaks [RHHH13]. wildfire [HJS22, KPDO23, UH20, XS11]. wildland [LYRR22]. win [GNCS22, LMB18]. wind [BCA18, EJD19, JCCG18, LM10a, LM10b, LBND13, LG20, LKB21, MR15, RF07, SM15, SDH18, ZCGC21, ZBG14]. window [WCW^{+}22]. windows [WHC^{+}22, ZMC^{+}21]. winner [XCS11]. winning [FCPL19, HSFP11]. Wiskott [PBS^{+}23]. within [CMAC^{+}23, EKO22, Kou08, PHLH11, PHLH12, SCK19]. within-household [PHLH11, SCK19]. within-school [PHLH12]. without [DHL18, LSM15, Mac20a, Mac20b, SLBL23]. women [vdKvEW17, GMB15]. Woodard [CW13b, CRZ13, Gil13, Moh13a, RP13, Sch13, Whi13]. word [Kip22]. word-frequency [Kip22]. words [RY11]. workforce [SH08]. Workload [ANFM09]. world [FT18, MM15]. wrapped [JLGJL12]. written [GGG^{+}12]. WTCCC [ZZL11].

X [DGM^{+}08]. X-ray [DGM^{+}08].

Yahoo [AZM11]. Year [Gne12, CMAC^{+}23, McE09]. years [Ber11, CR11, DL11a, HU11, Hol11, Kap11, MM11, MW11b, NL11, Rou11, SMR11, Sme11, Tin11, WA11]. yeast [FPL10]. yield [HSH12]. York [BDR16, Dup17, GRS16]. youth [KHZK23].

Zero [CVF10, LZCW21, BvdH09, BvdH19, CBvdHvdH08, FLHA15, JB21, KZS23, LMKC12, MGSD19, NvdBCR23, RBF^{+}20, WL22, XCCL20, ZCM^{+}11]. Zero-inflated [CVF10, LZCW21, CBvdHvdH08, LMKC12, MGSD19, NvdBCR23, RBF^{+}20, WL22, XCCL20, ZCM^{+}11]. zero-one-truncated [BvdH19]. zerotruncated [BvdH09, BvdH19]. zeros [WCL23]. ZIQRank [LZCW21]. ZnO [SSH^{+}11]. zone [ZTH19]. zones [Dup17].

References

Ameijeiras-Alonso:2019:FSI

[AABC^{+}19] Jose Ameijeiras-Alonso, Akli Benali, Rosa M. Crujeiras, Alberto Rodríguez-Casal, and José M. C. Pereira. Fire seasonality identification with multimodality tests. Annals of Applied Statistics, 13(4):2120–2139, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1574910038.

Atherton:2012:MSE

[ABB^{+}12] Juli Atherton, Nathan Boley, Ben Brown, Nobuo Ogawa, Stuart M. Davidson, Michael B. Eisen, Mark D. Biggin, and Peter Bickel. A model for sequential evolution of ligands by exponential enrichment (SELEX) data. Annals of Applied Statistics, 6(3):928–949, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1346418568.

Afsari:2014:RDP

[ABNG14] Bahman Afsari, Ulisses M. Braga-Neto, and Donald Geman. Rank discriminants for predicting phenotypes from RNA expression. Annals of Applied Statistics, 8(3):1469–1491, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091221.

Aksakalli:2012:OOP

[AC12] Vural Aksakalli and Elvan Ceyhan. Optimal obstacle placement with disambiguations. Annals of Applied Statistics, 6(4):1730–1774, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629058.

Alfons:2013:SLT

[ACG13] Andreas Alfons, Christophe Croux, and Sarah Gelper. Sparse least trimmed squares regression for analyzing highdimensional large data sets. Annals of Applied Statistics, 7(1):226–248, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1365527197.

Anthopolos:2023:BGM

[ACS^{+}23] Rebecca Anthopolos, Qixuan Chen, Joseph Sedransk, Mary Thompson, Gang Meng, and Sandro Galea. A Bayesian growth mixture model for complex survey data: Clustering postdisaster PTSD trajectories. Annals of Applied Statistics, 17(3):2494–2514, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-3/A-
Bayesian-growth-mixture-model-for-complex-survey-
data/10.1214/23-AOAS1729.full.

Aliverti:2022:CML

[AD22] Emanuele Aliverti and David B. Dunson. Composite mixture of log–linear models with application to psychiatric studies. Annals of Applied Statistics, 16(2):765– 790, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Composite-mixture-of-log-linear-models-
with-application-to-psychiatric/10.1214/21-AOAS1515.
full.

Asadi:2015:ERN

[ADE15] Peiman Asadi, Anthony C. Davison, and Sebastian Engelke. Extremes on river networks. Annals of Applied Statistics, 9(4):2023–2050, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453994189.

Araneda:2007:SAS

[AFS07] Anita Araneda, Stephen E. Fienberg, and Alvaro Soto. A statistical approach to simultaneous mapping and localization for mobile robots. Annals of Applied Statistics, 1(1): 66–84, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143729.

Admiraal:2016:MCS

[AH16] Ryan Admiraal and Mark S. Handcock. Modeling concurrency and selective mixing in heterosexual partnership networks with applications to sexually transmitted diseases. Annals of Applied Statistics, 10(4):2021–2046, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606850.

Arnold:2023:SVT

[AHZ23] Sebastian Arnold, Alexander Henzi, and Johanna F. Ziegel. Sequentially valid tests for forecast calibration. Annals of Applied Statistics, 17(3):1909–1935, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Sequentially-valid-tests-for-forecast-
calibration/10.1214/22-AOAS1697.full.

Anaya-Izquierdo:2011:OSC

[AICV11] Karim Anaya-Izquierdo, Frank Critchley, and Karen Vines. Orthogonal simple component analysis: a new, exploratory approach. Annals of Applied Statistics, 5(1):486–522, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1300715200.

Aston:2012:ESC

[AK12] John A. D. Aston and Claudia Kirch. Evaluating stationarity via change-point alternatives with applications to fMRI data. Annals of Applied Statistics, 6(4):1906–1948, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1356629065.

Ahn:2016:PVA

[AL16] Kwang Woo Ahn and Brent R. Logan. Pseudo-value approach for conditional quantile residual lifetime analysis for clustered survival and competing risks data with applications to bone marrow transplant data. Annals of Applied Statistics, 10(2):618–637, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199887.

Aitkin:2009:BMC

[ALC09] Murray Aitkin, Charles C. Liu, and Tom Chadwick. Bayesian model comparison and model averaging for small-area estimation. Annals of Applied Statistics, 3(1):199–221, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888368.

Aston:2007:DAG

[AM07] John A. D. Aston and Donald E. K. Martin. Distributions associated with general runs and patterns in hidden Markov models. Annals of Applied Statistics, 1(2):585–611, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1196438032.

Arpino:2016:ACE

[AM16] Bruno Arpino and Alessandra Mattei. Assessing the causal effects of financial aids to firms in Tuscany allowing for interference. Annals of Applied Statistics, 10(3):1170–1194, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069604.

Antonelli:2020:EHE

[AMB^{+}20] Joseph Antonelli, Maitreyi Mazumdar, David Bellinger, David Christiani, Robert Wright, and Brent Coull. Estimating the health effects of environmental mixtures using Bayesian semiparametric regression and sparsity inducing priors. Annals of Applied Statistics, 14(1):257–275, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002674.

Ahn:2013:BSA

[AMGG13] Jaeil Ahn, Bhramar Mukherjee, Stephen B. Gruber, and Malay Ghosh. Bayesian semiparametric analysis for twophase studies of gene-environment interaction. Annals of Applied Statistics, 7(1):543–569, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527210.

Arbel:2016:BND

[AMR16] Julyan Arbel, Kerrie Mengersen, and Judith Rousseau. Bayesian nonparametric dependent model for partially replicated data: The influence of fuel spills on species diversity. Annals of Applied Statistics, 10(3):1496–1516, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1475069616.

Adrian:2018:CVT

[AMR18] Daniel W. Adrian, Ranjan Maitra, and Daniel B. Rowe. Complex-valued time series modeling for improved activation detection in fMRI studies. Annals of Applied Statistics, 12(3):1451–1478, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652961.

Alkema:2014:GEC

[AN14] Leontine Alkema and Jin Rou New. Global estimation of child mortality using a Bayesian B-spline bias-reduction model. Annals of Applied Statistics, 8(4):2122–2149, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001737.

Ane:2008:ACD

[Ané08] Cécile Ané. Analysis of comparative data with hierarchical autocorrelation. Annals of Applied Statistics, 2 (3):1078–1102, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1223908053.

Aldor-Noiman:2009:WFC

[ANFM09] Sivan Aldor-Noiman, Paul D. Feigin, and Avishai Mandelbaum. Workload forecasting for a call center: Methodology and a case study. Annals of Applied Statistics, 3(4):1403– 1447, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453946.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial board. Annals of Applied Statistics, 10(2):??, June 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1469199883.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial board. Annals of Applied Statistics, 10(3):??, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1475069598.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial board. Annals of Applied Statistics, 10(4):??, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606834.

Anonymous:2016:TCa

[Ano16d] Anonymous. Table of contents. Annals of Applied Statistics, 10(2):??, June 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1469199882.

Anonymous:2016:TCb

[Ano16e] Anonymous. Table of contents. Annals of Applied Statistics, 10(3):??, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1475069597.

Anonymous:2016:TCc

[Ano16f] Anonymous. Table of contents. Annals of Applied Statistics, 10(4):??, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606833.

Anonymous:2018:SSM

[Ano18] Anonymous. Special section in memory of Stephen E. Fienberg (1942–2016) AOAS Editor-in-Chief 2013–2015. Annals of Applied Statistics, 12(2):iii–x, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1532743469.

Anonymous:2023:EBb

[Ano23a] Anonymous. Editorial board. Annals of Applied Statistics, 17(3):??, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-17/issue-3/Editorial-Board/aoas173p1.full.

Anonymous:2023:TCa

[Ano23b] Anonymous. Table of contents. Annals of Applied Statistics, 17(3):??, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-17/issue-3/Table-of-Contents/aoas173c1.full.

Alonso-Pena:2023:AAE

[APC23] María Alonso-Pena and Rosa M. Crujeiras. Analyzing animal escape data with circular nonparametric multimodal regression. Annals of Applied Statistics, 17(1):130–152, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Analyzing-animal-escape-data-with-circular-
nonparametric-multimodal-regression/10.1214/22-AOAS1619.
full.

Aydin:2009:PCA

[APW^{+}09] Burcu Aydın, Gábor Pataki, Haonan Wang, Elizabeth Bullitt, and J. S. Marron. A principal component analysis for trees. Annals of Applied Statistics, 3(4):1597–1615, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453955.

Azose:2018:ELC

[AR18] Jonathan J. Azose and Adrian E. Raftery. Estimating large correlation matrices for international migration. Annals of Applied Statistics, 12(2):940–970, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1532743482.

Alkema:2007:PPH

[ARC07] Leontine Alkema, Adrian E. Raftery, and Samuel J. Clark. Probabilistic projections of HIV prevalence using Bayesian melding. Annals of Applied Statistics, 1(1):229–248, June 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1183143737.

Augustine:2018:SCR

[ARK^{+}18] Ben C. Augustine, J. Andrew Royle, Marcella J. Kelly, Christopher B. Satter, Robert S. Alonso, Erin E. Boydston, and Kevin R. Crooks. Spatial capture–recapture with partial identity: An application to camera traps. Annals of Applied Statistics, 12(1):67–95, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564465.

Ahdesmaki:2010:FSO

[AS10a] Miika Ahdesmäki and Korbinian Strimmer. Feature selection in omics prediction problems using cat scores and false nondiscovery rate control. Annals of Applied Statistics, 4(1): 503–519, March 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1273584465.

Albert:2010:AJM

[AS10b] Paul S. Albert and Joanna H. Shih. An approach for jointly modeling multivariate longitudinal measurements and discrete time-to-event data. Annals of Applied Statistics, 4 (3):1517–1532, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409384.

Aronow:2017:EAC

[AS17] Peter M. Aronow and Cyrus Samii. Estimating average causal effects under general interference, with application to a social network experiment. Annals of Applied Statistics, 11(4):1912–1947, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1514430272.

Allenbrand:2023:MSU

[AS23] Corban Allenbrand and Ben Sherwood. Model selection uncertainty and stability in beta regression models: a study of bootstrap-based model averaging with an empirical application to clickstream data. Annals of Applied Statistics, 17(1): 680–710, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Model-selection-uncertainty-and-stability-
in-beta-regression-models/10.1214/22-AOAS1647.full.

Aktekin:2013:AMD

[ASX13] Tevfik Aktekin, Refik Soyer, and Feng Xu. Assessment of mortgage default risk via Bayesian state space models. Annals of Applied Statistics, 7(3):1450–1473, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1380804802.

Ait-Sahalia:2009:HFM

[ASY09] Yacine Aït-Sahalia and Jialin Yu. High frequency market microstructure noise estimates and liquidity measures. Annals of Applied Statistics, 3(1):422–457, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1239888377.

Allen:2010:TRC

[AT10] Genevera I. Allen and Robert Tibshirani. Transposable regularized covariance models with an application to missing data imputation. Annals of Applied Statistics, 4(2):764–790, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1280842139.

Andridge:2015:ANB

[AT15] Rebecca Andridge and Katherine Jenny Thompson. Assessing nonresponse bias in a business survey: Proxy pattern– mixture analysis for skewed data. Annals of Applied Statistics, 9(4):2237–2265, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453994199.

Anderlucci:2015:CPM

[AV15] Laura Anderlucci and Cinzia Viroli. Covariance pattern mixture models for the analysis of multivariate heterogeneous longitudinal data. Annals of Applied Statistics, 9(2):777– 800, June 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1437397111.

Airoldi:2013:MWB

[AWL13] Edoardo M. Airoldi, Xiaopei Wang, and Xiaodong Lin. Multi-way blockmodels for analyzing coordinated highdimensional responses. Annals of Applied Statistics, 7 (4):2431–2457, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1387823326.

Agniel:2018:FPV

[AXEC18] Denis Agniel, Wen Xie, Myron Essex, and Tianxi Cai. Functional principal variance component testing for a genetic association study of HIV progression. Annals of Applied Statistics, 12(3):1871–1893, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652978.

Allassonniere:2012:SAP

[AY12] Stéphanie Allassonnière and Laurent Younes. A stochastic algorithm for probabilistic independent component analysis. Annals of Applied Statistics, 6(1):125–160, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1331043391.

Anderes:2009:MLE

[AYJ^{+}09] E. Anderes, B. Yu, V. Jovanovic, C. Moroney, M. Garay, A. Braverman, and E. Clothiaux. Maximum likelihood estimation of cloud height from multi-angle satellite imagery. Annals of Applied Statistics, 3(3):902–921, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1254773271.

Alkema:2017:BAG

[AZC^{+}17] Leontine Alkema, Sanqian Zhang, Doris Chou, Alison Gemmill, Ann-Beth Moller, Doris Ma Fat, Lale Say, Colin Mathers, and Daniel Hogan. A Bayesian approach to the global estimation of maternal mortality. Annals of Applied Statistics, 11(3):1245–1274, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168829.

Agarwal:2011:MII

[AZM11] Deepak Agarwal, Liang Zhang, and Rahul Mazumder. Modeling item–item similarities for personalized recommendations on Yahoo! front page. Annals of Applied Statistics, 5(3):1839–1875, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514287.

Brill:2022:TDA

[BAH22] Barak Brill, Amnon Amir, and Ruth Heller. Testing for differential abundance in compositional counts data, with application to microbiome studies. Annals of Applied Statistics, 16(4):2648–2671, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Testing-for-
differential-abundance-in-compositional-counts-data-
with-application/10.1214/22-AOAS1607.full.

Brynjarsdottir:2011:BHM

[BB11] Jenný Brynjarsdóttir and L. Mark Berliner. Bayesian hierarchical modeling for temperature reconstruction from geothermal data. Annals of Applied Statistics, 5(2B):1328– 1359, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562723.

Bickel:2010:SMG

[BBB^{+}10] Peter J. Bickel, Nathan Boley, James B. Brown, Haiyan Huang, and Nancy R. Zhang. Subsampling methods for genomic inference. Annals of Applied Statistics, 4(4):1660– 1697, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167794.

Barrientos:2018:PAC

[BBB^{+}18] Andrés F. Barrientos, Alexander Bolton, Tom Balmat, Jerome P. Reiter, John M. de Figueiredo, Ashwin Machanavajjhala, Yan Chen, Charley Kneifel, and Mark DeLong. Providing access to confidential research data through synthesis and verification: an application to data on employees of the U.S. federal government. Annals of Applied Statistics, 12(2): 1124–1156, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743488.

Bornkamp:2011:RAD

[BBDP11] Björn Bornkamp, Frank Bretz, Holger Dette, and José Pinheiro. Response-adaptive dose-finding under model uncertainty. Annals of Applied Statistics, 5(2B):1611–1631, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562735.

Bilgrau:2018:ECC

[BBE^{+}18] Anders Ellern Bilgrau, Rasmus Froberg Brøndum, Poul Svante Eriksen, Karen Dybkær, and Martin Bøgsted. Estimating a common covariance matrix for network meta-analysis of gene expression datasets in diffuse large B-cell lymphoma. Annals of Applied Statistics, 12(3):1894–1913, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1536652979.

Baggaley:2012:IPD

[BBG^{+}12] Andrew W. Baggaley, Richard J. Boys, Andrew Golightly, Graeme R. Sarson, and Anvar Shukurov. Inference for population dynamics in the Neolithic period. Annals of Applied Statistics, 6(4):1352–1376, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1356629043.

Benedetti:2022:ASD

[BBL22] Marco H. Benedetti, Veronica J. Berrocal, and Roderick J. Little. Accounting for survey design in Bayesian disaggregation of survey-based areal estimates of proportions: An application to the American Community Survey. Annals of Applied Statistics, 16(4):2201–2230, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Accounting-for-survey-design-in-Bayesian-
disaggregation-of-survey-based/10.1214/21-AOAS1585.
full.

Banerjee:2020:NNB

[BBM20] Trambak Banerjee, Bhaswar B. Bhattacharya, and Gourab Mukherjee. A nearest-neighbor based nonparametric test for viral remodeling in heterogeneous single-cell proteomic data. Annals of Applied Statistics, 14(4):1777–1805, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

14/issue-4/A-nearest-neighbor-based-nonparametric-
test-for-viral-remodeling-in/10.1214/20-AOAS1362.
full.

Baggerly:2009:DCC

[BC09] Keith A. Baggerly and Kevin R. Coombes. Deriving chemosensitivity from cell lines: Forensic bioinformatics and reproducible research in high-throughput biology. Annals of Applied Statistics, 3(4):1309–1334, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453942.

Bonas:2023:CSF

[BC23] Matthew Bonas and Stefano Castruccio. Calibration of SpatioTemporal forecasts from citizen science urban air pollution data with sparse recurrent neural networks. Annals of Applied Statistics, 17(3):1820–1840, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Calibration-
of-SpatioTemporal-forecasts-from-citizen-science-
urban-air-pollution/10.1214/22-AOAS1683.full.

Bessac:2018:SSP

[BCA18] Julie Bessac, Emil Constantinescu, and Mihai Anitescu. Stochastic simulation of predictive space-time scenarios of wind speed using observations and physical model outputs. Annals of Applied Statistics, 12(1):432–458, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564479.

Bouveyron:2015:DFM

[BCJ15] Charles Bouveyron, Etienne Côme, and Julien Jacques. The discriminative functional mixture model for a comparative analysis of bike sharing systems. Annals of Applied Statistics, 9(4):1726–1760, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453993092.

Bertolacci:2019:CID

[BCR^{+}19] Michael Bertolacci, Edward Cripps, Ori Rosen, John W. Lau, and Sally Cripps. Climate inference on daily rainfall across the Australian continent, 1876-2015. Annals of Applied Statistics, 13(2):683–712, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1560758424.

Blanchet:2011:SME

[BD11] Juliette Blanchet and Anthony C. Davison. Spatial modeling of extreme snow depth. Annals of Applied Statistics, 5(3):1699–1725, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514282.

Belzile:2022:IIR

[BD22] Léo R. Belzile and Anthony C. Davison. Improved inference on risk measures for univariate extremes. Annals of Applied Statistics, 16(3):1524–1549, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Improved-inference-
on-risk-measures-for-univariate-extremes/10.1214/
21-AOAS1555.full.

Brentnall:2011:MER

[BDC^{+}11] Adam R. Brentnall, Stephen W. Duffy, Martin J. Crowder, Maureen G. C. Gillan, Susan M. Astley, Matthew G. Wallis, Jonathan James, Caroline R. M. Boggis, and Fiona J. Gilbert. A method for exploratory repeated-measures analysis applied to a breast-cancer screening study. Annals of Applied Statistics, 5(4):2448–2469, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1324399602.

Buishand:2008:SEA

[BdHZ08] T. A. Buishand, L. de Haan, and C. Zhou. On spatial extremes: With application to a rainfall problem. Annals of Applied Statistics, 2(2):624–642, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1215118531.

Briollais:2016:BGM

[BDL^{+}16] Laurent Briollais, Adrian Dobra, Jinnan Liu, Matt Friedlander, Hilmi Ozcelik, and Hélène Massam. A Bayesian graphical model for genome-wide association studies (GWAS). Annals of Applied Statistics, 10(2):786–811, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1469199893.

Branson:2016:ICB

[BDR16] Zach Branson, Tirthankar Dasgupta, and Donald B. Rubin. Improving covariate balance in 2^{K} factorial designs via rerandomization with an application to a New York City Department of Education high school study. Annals of Applied Statistics, 10(4):1958–1976, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1483606847.

Bodnar:2023:BMS

[BE23] Olha Bodnar and Viktor Eriksson. Bayesian model selection: Application to the adjustment of fundamental physical constants. Annals of Applied Statistics, 17(3):2118–2138, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Bayesian-model-selection--Application-
to-the-adjustment-of-fundamental/10.1214/22-AOAS1710.
full.

Bentley:2008:DSA

[Ben08] Donald L. Bentley. Discussion of: Statistical analysis of an archaeological find. Annals of Applied Statistics, 2(1):66– 73, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367808. See [Feu08b].

Berliner:2011:DSA

[Ber11] L. Mark Berliner. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):45–46, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715171. See [MW11b].

Busch:2009:TSH

[BFF^{+}09] Jorge R. Busch, Pablo A. Ferrari, Ana Georgina Flesia, Ricardo Fraiman, Sebastian P. Grynberg, and Florencia Leonardi. Testing statistical hypothesis on random trees and applications to the protein classification problem. Annals of Applied Statistics, 3(2):542–563, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676185.

Barabesi:2012:PDB

[BFM12] Lucio Barabesi, Sara Franceschi, and Marzia Marcheselli. Properties of design-based estimation under stratified spatial sampling with application to canopy coverage estimation. Annals of Applied Statistics, 6(1):210–228, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1331043394.

Benjamini:2009:SFS

[BG09] Yoav Benjamini and Yulia Gavrilov. A simple forward selection procedure based on false discovery rate control. Annals of Applied Statistics, 3(1):179–198, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1239888367.

Brantley:2020:BDE

[BGC20] Halley L. Brantley, Joseph Guinness, and Eric C. Chi. Baseline drift estimation for air quality data using quantile trend filtering. Annals of Applied Statistics, 14(2):585–604, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1593449317.

Bridle:2009:HGC

[BGH^{+}09] Sarah Bridle, Mandeep Gill, Alan Heavens, Catherine Heymans, F. William High, Henk Hoekstra, Mike Jarvis, Donnacha Kirk, Thomas Kitching, Jean-Paul Kneib, Konrad Kuijken, John Shawe-Taylor, David Lagatutta, Rachel Mandelbaum, Richard Massey, Yannick Mellier, Baback Moghaddam, Yassir Moudden, Reiko Nakajima, Stephane Paulin-Henriksson, Sandrine Pires, Anais Rassat, Adam Amara, Alexandre Refregier, Jason Rhodes, Tim Schrabback, Elisabetta Semboloni, Marina Shmakova, Ludovic van Waerbeke, Dugan Witherick, Lisa Voigt, David Wittman, Douglas Applegate, Sreekumar T. Balan, Joel Berge, Gary Bernstein, Hakon Dahle, and Thomas Erben. Handbook for the GREAT08 challenge: an image analysis competition for cosmological lensing. Annals of Applied Statistics, 3(1):6– 37, March 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1239888361.

Berrocal:2010:BST

[BGH10] Veronica J. Berrocal, Alan E. Gelfand, and David M. Holland. A bivariate space–time downscaler under space and time misalignment. Annals of Applied Statistics, 4(4):1942– 1975, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167805.

Brodersen:2015:ICI

[BGK^{+}15] Kay H. Brodersen, Fabian Gallusser, Jim Koehler, Nicolas Remy, and Steven L. Scott. Inferring causal impact using Bayesian structural time-series models. Annals of Applied Statistics, 9(1):247–274, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226092.

Beaudry:2017:IRD

[BGM17] Isabelle S. Beaudry, Krista J. Gile, and Shruti H. Mehta. Inference for respondent-driven sampling with misclassification. Annals of Applied Statistics, 11(4):2111–2141, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1514430279.

Breheny:2011:CDA

[BH11] Patrick Breheny and Jian Huang. Coordinate descent algorithms for nonconvex penalized regression, with applications to biological feature selection. Annals of Applied Statistics, 5(1):232–253, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1300715189.

Bastide:2021:EBI

[BHB^{+}21] Paul Bastide, Lam Si Tung Ho, Guy Baele, Philippe Lemey, and Marc A. Suchard. Efficient Bayesian inference of general Gaussian models on large phylogenetic trees. Annals of Applied Statistics, 15(2):971–997, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Efficient-Bayesian-inference-of-general-
Gaussian-models-on-large-phylogenetic/10.1214/20-
AOAS1419.full.

Barkley:2020:CIO

[BHC^{+}20] Brian G. Barkley, Michael G. Hudgens, John D. Clemens, Mohammad Ali, and Michael E. Emch. Causal inference from observational studies with clustered interference, with application to a cholera vaccine study. Annals of Applied Statistics, 14(3):1432–1448, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Causal-inference-from-observational-studies-
with-clustered-interference-with-application/10.1214/
19-AOAS1314.full.

Breto:2009:TSA

[BHIK09] Carles Bretó, Daihai He, Edward L. Ionides, and Aaron A. King. Time series analysis via mechanistic models. Annals of Applied Statistics, 3(1):319–348, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1239888373.

Bissantz:2010:IPC

[BHP10] Nicolai Bissantz, Hajo Holzmann, and Mirosław Pawlak. Improving PSF calibration in confocal microscopic imaging — estimating and exploiting bilateral symmetry. Annals of Applied Statistics, 4(4):1871–1891, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167802.

Bovy:2011:EDI

[BHR11] Jo Bovy, David W. Hogg, and Sam T. Roweis. Extreme deconvolution: Inferring complete distribution functions from noisy, heterogeneous and incomplete observations. Annals of Applied Statistics, 5(2B):1657–1677, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1310562737.

Bradley:2015:MST

[BHW15] Jonathan R. Bradley, Scott H. Holan, and Christopher K. Wikle. Multivariate spatio-temporal models for highdimensional areal data with application to Longitudinal Employer–Household Dynamics. Annals of Applied Statistics, 9(4):1761–1791, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453993093.

Bickel:2010:LBI

[Bic10] Peter J. Bickel. Leo Breiman: an important intellectual and personal force in statistics, my life and that of many others. Annals of Applied Statistics, 4(4):1634–1637, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167786.

Bigot:2013:FMC

[Big13] Jérémie Bigot. Fréchet means of curves for signal averaging and application to ECG data analysis. Annals of Applied Statistics, 7(4):2384–2401, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1387823324.

Bird:2008:DSA

[Bir08] Sheila M. Bird. Discussion of: “Statistical analysis of an archeological find” — skeptical counting challenges to an archaeological find. Annals of Applied Statistics, 2(1):74– 76, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367809. See [Feu08b].

Buzbas:2009:MLE

[BJ09] Erkan Ozge Buzbas and Paul Joyce. Maximum likelihood estimates under k-allele models with selection can be numerically unstable. Annals of Applied Statistics, 3(3):1147–1162, September 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1254773282.

Ba:2012:CGP

[BJ12] Shan Ba and V. Roshan Joseph. Composite Gaussian process models for emulating expensive functions. Annals of Applied Statistics, 6(4):1838–1860, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629062.

Balocchi:2019:SMT

[BJ19] Cecilia Balocchi and Shane T. Jensen. Spatial modeling of trends in crime over time in Philadelphia. Annals of Applied Statistics, 13(4):2235–2259, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910043.

Bouveyron:2022:CCM

[BJS^{+}22] Charles Bouveyron, Julien Jacques, Amandine Schmutz, Fanny Simões, and Silvia Bottini. Co-clustering of multivariate functional data for the analysis of air pollution in the South of France. Annals of Applied Statistics, 16(3):1400–1422, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Co-clustering-
of-multivariate-functional-data-for-the-analysis-
of/10.1214/21-AOAS1547.full.

Boonstra:2020:ICP

[BK20] Philip S. Boonstra and John C. Krauss. Inferring a consensus problem list using penalized multistage models for ordered data. Annals of Applied Statistics, 14(3):1557–1580, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Inferring-a-consensus-problem-list-using-
penalized-multistage-models-for/10.1214/20-AOAS1361.
full.

Berg:2021:ABP

[BK21] Emily Berg and Jae-Kwang Kim. An approximate best prediction approach to small area estimation for sheet and rill erosion under informative sampling. Annals of Applied Statistics, 15(1):102–125, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-1/An-
approximate-best-prediction-approach-to-small-area-
estimation-for/10.1214/20-AOAS1388.full.

Bleich:2014:VSB

[BKGJ14] Justin Bleich, Adam Kapelner, Edward I. George, and Shane T. Jensen. Variable selection for BART: an application to gene regulation. Annals of Applied Statistics, 8(3):1750–1781, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1414091233.

Bartlett:2021:TWS

[BKS21] Thomas E. Bartlett, Ioannis Kosmidis, and Ricardo Silva. Two-way sparsity for time-varying networks with applications in genomics. Annals of Applied Statistics, 15(2):856– 879, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Two-way-sparsity-for-time-varying-networks-
with-applications-in/10.1214/20-AOAS1416.full.

Bonvini:2022:CIE

[BKVW22] Matteo Bonvini, Edward H. Kennedy, Valerie Ventura, and Larry Wasserman. Causal inference for the effect of mobility on COVID-19 deaths. Annals of Applied Statistics, 16(4):2458–2480, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Causal-inference-
for-the-effect-of-mobility-on-COVID-19/10.1214/22-
AOAS1599.full.

Blei:2007:CTMa

[BL07a] David M. Blei and John D. Lafferty. A correlated topic model of Science. Annals of Applied Statistics, 1(1):17–35, June 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1183143727.

Blei:2007:CTMb

[BL07b] David M. Blei and John D. Lafferty. A correlated topic model of Science. Annals of Applied Statistics, 1(2):634, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1196438034.

Bolin:2011:SMG

[BL11] David Bolin and Finn Lindgren. Spatial models generated by nested stochastic partial differential equations, with an application to global ozone mapping. Annals of Applied Statistics, 5(1):523–550, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1300715201.

Berg:2019:PSA

[BL19] Emily Berg and Danhyang Lee. Prediction of small area quantiles for the conservation effects assessment project using a mixed effects quantile regression model. Annals of Applied Statistics, 13(4):2158–2188, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910040.

Bartolucci:2009:LMM

[BLM09] Francesco Bartolucci, Monia Lupparelli, and Giorgio E. Montanari. Latent Markov model for longitudinal binary data: An application to the performance evaluation of nursing homes. Annals of Applied Statistics, 3(2):611–636, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1245676188.

Banerjee:2023:JMP

[BLM^{+}23] Trambak Banerjee, Peng Liu, Gourab Mukherjee, Shantanu Dutta, and Hai Che. Joint modeling of playing time and purchase propensity in massively multiplayer online role-playing games using crossed random effects. Annals of Applied Statistics, 17(3):2533–2554, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Joint-modeling-of-playing-time-and-purchase-
propensity-in-massively/10.1214/23-AOAS1731.full.

Bachl:2015:BME

[BLTG15] Fabian E. Bachl, Alex Lenkoski, Thordis L. Thorarinsdottir, and Christoph S. Garbe. Bayesian motion estimation for dust aerosols. Annals of Applied Statistics, 9(3):1298–1327, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488740.

Barboza:2014:RPT

[BLTV14] Luis Barboza, Bo Li, Martin P. Tingley, and Frederi G. Viens. Reconstructing past temperatures from natural proxies and estimated climate forcings using short- and longmemory models. Annals of Applied Statistics, 8(4):1966– 2001, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001732.

Banerjee:2008:OFD

[BM08] Tathagata Banerjee and Rahul Mukerjee. Optimal factorial designs for cDNA microarray experiments. Annals of Applied Statistics, 2(1):366–385, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1206367825.

Bhattacharya:2011:NNB

[BM11] Sourabh Bhattacharya and Ranjan Maitra. A nonstationary nonparametric Bayesian approach to dynamically modeling effective connectivity in functional magnetic resonance imaging experiments. Annals of Applied Statistics, 5(2B):1183– 1206, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562718.

Baugh:2022:HBM

[BM22] Samuel Baugh and Karen McKinnon. Hierarchical Bayesian modeling of ocean heat content and its uncertainty. Annals of Applied Statistics, 16(4):2603–2625, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Hierarchical-
Bayesian-modeling-of-ocean-heat-content-and-its-uncertainty/
10.1214/22-AOAS1605.full.

Ben-Michael:2023:EEC

[BMAF^{+}23] Eli Ben-Michael, David Arbour, Avi Feller, Alexander Franks, and Steven Raphael. Estimating the effects of a California gun control program with multitask Gaussian processes. Annals of Applied Statistics, 17(2):985– 1016, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Estimating-the-effects-of-a-California-
gun-control-program-with/10.1214/22-AOAS1654.full.

Belbahri:2021:QBU

[BMGN21] Mouloud Belbahri, Alejandro Murua, Olivier Gandouet, and Vahid Partovi Nia. Qini-based uplift regression. Annals of Applied Statistics, 15(3):1247–1272, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Qini-based-uplift-regression/10.1214/21-
AOAS1465.full.

Bolsinova:2016:URS

[BMH16] Maria Bolsinova, Gunter Maris, and Herbert Hoijtink. Unmixing Rasch scales: How to score an educational test. Annals of Applied Statistics, 10(2):925–945, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1469199899.

Bhat:2020:SDB

[BML^{+}20] K. Sham Bhat, Kary Myers, Earl Lawrence, James Colgan, and Elizabeth Judge. Structured discrepancy in Bayesian model calibration for ChemCam on the Mars Curiosity rover. Annals of Applied Statistics, 14(4):2020–2036, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Structured-discrepancy-in-Bayesian-model-
calibration-for-ChemCam-on-the/10.1214/20-AOAS1373.
full.

Bhuyan:2021:ACE

[BMLG21] Prajamitra Bhuyan, Emma J. McCoy, Haojie Li, and Daniel J. Graham. Analysing the causal effect of London cycle superhighways on traffic congestion. Annals of Applied Statistics, 15(4):1999–2022, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Analysing-the-
causal-effect-of-London-cycle-superhighways-on-traffic/
10.1214/21-AOAS1450.full.

Bendich:2016:PHA

[BMM^{+}16] Paul Bendich, J. S. Marron, Ezra Miller, Alex Pieloch, and Sean Skwerer. Persistent homology analysis of brain artery trees. Annals of Applied Statistics, 10(1):198–218, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909913.

Boonstra:2013:BSM

[BMT13] Philip S. Boonstra, Bhramar Mukherjee, and Jeremy M. G. Taylor. Bayesian shrinkage methods for partially observed data with many predictors. Annals of Applied Statistics, 7(4):2272–2292, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1387823319.

Bao:2023:EHE

[BNMG23] Le Bao, Xiaoyue Niu, Mary Mahy, and Peter D. Ghys. Estimating HIV epidemics for subnational areas. Annals of Applied Statistics, 17(3):2515–2532, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Estimating-HIV-
epidemics-for-subnational-areas/10.1214/23-AOAS1730.
full.

Barry:2008:SFT

[BNW08] William T. Barry, Andrew B. Nobel, and Fred A. Wright. A statistical framework for testing functional categories in microarray data. Annals of Applied Statistics, 2(1):286– 315, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367822.

Barata:2022:FIT

[BPS22] Raquel Barata, Raquel Prado, and Bruno Sansó. Fast inference for time-varying quantiles via flexible dynamic models with application to the characterization of atmospheric rivers. Annals of Applied Statistics, 16(1):247– 271, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

16/issue-1/Fast-inference-for-time-varying-quantiles-
via-flexible-dynamic-models/10.1214/21-AOAS1497.full.

Bliznyuk:2014:NPL

[BPSC14] Nikolay Bliznyuk, Christopher J. Paciorek, Joel Schwartz, and Brent Coull. Nonlinear predictive latent process models for integrating spatio-temporal exposure data from multiple sources. Annals of Applied Statistics, 8(3):1538–1560, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091224.

Bickel:2008:DTA

[BR08] Peter J. Bickel and Ya’acov Ritov. Discussion of: “Treelets — An adaptive multi-scale basis for sparse unordered data”. Annals of Applied Statistics, 2(2):474–477, June 2008. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1215118520. See [LNW08b].

Brown:2016:LSD

[BR16] Philip J. Brown and Martin S. Ridout. Level-screening designs for factors with many levels. Annals of Applied Statistics, 10(2):864–883, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199896.

Berrocal:2008:PQP

[BRG08] Veronica J. Berrocal, Adrian E. Raftery, and Tilmann Gneiting. Probabilistic quantitative precipitation field forecasting using a two-stage spatial model. Annals of Applied Statistics, 2(4):1170–1193, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1231424205.

Brown:2008:SPB

[Bro08] Lawrence D. Brown. In-season prediction of batting averages: a field test of empirical Bayes and Bayes methodologies. Annals of Applied Statistics, 2(1):113–152, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1206367815.

Brown:2009:AAB

[Bro09] Elizabeth R. Brown. Assessing the association between trends in a biomarker and risk of event with an application in pediatric HIV/AIDS. Annals of Applied Statistics, 3(3):1163–1182, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1254773283.

Bargagli-Stoffi:2022:HCE

[BSDG22] Falco J. Bargagli-Stoffi, Kristof De Witte, and Giorgio Gnecco. Heterogeneous causal effects with imperfect compliance: a Bayesian machine learning approach. Annals of Applied Statistics, 16(3):1986–2009, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Heterogeneous-causal-effects-with-imperfect-
compliance--A-Bayesian-machine/10.1214/21-AOAS1579.
full.

Bandyopadhyay:2010:CAP

[BSLL10] Dipankar Bandyopadhyay, Debajyoti Sinha, Stuart Lipsitz, and Elizabeth Letourneau. Changing approaches of prosecutors towards juvenile repeated sex-offenders: a Bayesian evaluation. Annals of Applied Statistics, 4(2):805–829, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1280842141.

Bonner:2016:ELM

[BSNP16] Simon J. Bonner, Matthew R. Schofield, Patrik Noren, and Steven J. Price. Extending the latent multinomial model with complex error processes and dynamic Markov bases. Annals of Applied Statistics, 10(1):246–263, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1458909915.

Bien:2015:CHT

[BST15] Jacob Bien, Noah Simon, and Robert Tibshirani. Convex hierarchical testing of interactions. Annals of Applied Statistics, 9(1):27–42, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1430226083.

Bien:2011:PSI

[BT11] Jacob Bien and Robert Tibshirani. Prototype selection for interpretable classification. Annals of Applied Statistics, 5(4):2403–2424, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1324399600.

Baker:2020:FSD

[BTA20] Yulia Baker, Tiffany M. Tang, and Genevera I. Allen. Feature selection for data integration with mixed multiview data. Annals of Applied Statistics, 14(4):1676–1698, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Feature-selection-for-data-integration-
with-mixed-multiview-data/10.1214/20-AOAS1389.full.

Baladandayuthapani:2014:BSG

[BTJ^{+}14] Veerabhadran Baladandayuthapani, Rajesh Talluri, Yuan Ji, Kevin R. Coombes, Yiling Lu, Bryan T. Hennessy, Michael A. Davies, and Bani K. Mallick. Bayesian sparse graphical models for classification with application to protein expression data. Annals of Applied Statistics, 8 (3):1443–1468, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1414091220.

Buhlmann:2010:RLB

[Büh10] Peter Bühlmann. Remembrance of Leo Breiman. Annals of Applied Statistics, 4(4):1638–1641, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167787.

Boonstra:2022:MTS

[BvdB22] Harm Jan Boonstra and Jan van den Brakel. Multilevel time-series models for small area estimation at different frequencies and domain levels. Annals of Applied Statistics, 16(4):2314–2338, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Multilevel-time-
series-models-for-small-area-estimation-at-different/
10.1214/21-AOAS1592.full.

Bogdan:2015:SAV

[BvdBS^{+}15] Małgorzata Bogdan, Ewout van den Berg, Chiara Sabatti, Weijie Su, and Emmanuel J. Candès. SLOPE-adaptive variable selection via convex optimization. Annals of Applied Statistics, 9(3):1103–1140, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1446488733.

Bohning:2009:CAZ

[BvdH09] Dankmar Böhning and Peter G. M. van der Heijden. A covariate adjustment for zero-truncated approaches to estimating the size of hidden and elusive populations. Annals of Applied Statistics, 3(2):595–610, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676187.

Bohning:2019:IZT

[BvdH19] Dankmar Böhning and Peter G. M. van der Heijden. The identity of the zero-truncated, one-inflated likelihood and the zero-one-truncated likelihood for general count densities with an application to drink-driving in Britain. Annals of Applied Statistics, 13(2):1198–1211, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758443.

Balakrishnan:2018:HTH

[BW18] Sivaraman Balakrishnan and Larry Wasserman. Hypothesis testing for high-dimensional multinomials: A selective review. Annals of Applied Statistics, 12(2):727–749, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743474.

Bray:2014:VRA

[BWBS14] Andrew Bray, Ka Wong, Christopher D. Barr, and Frederic Paik Schoenberg. Voronoi residual analysis of spatial point process models with applications to California earthquake forecasts. Annals of Applied Statistics, 8(4):2247–2267, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001742.

Berger:2019:MBR

[BWS19] Moritz Berger, Michael Wagner, and Matthias Schmid. Modeling biomarker ratios with gamma distributed components. Annals of Applied Statistics, 13(1):548–572, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1554861660. See corrigendum [Sch23].

Birrell:2020:ERT

[BWT^{+}20] Paul J. Birrell, Lorenz Wernisch, Brian D. M. Tom, Leonhard Held, Gareth O. Roberts, Richard G. Pebody, and Daniela De Angelis. Efficient real-time monitoring of an emerging influenza pandemic: How feasible? Annals of Applied Statistics, 14(1):74–93, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1587002665.

Bickel:2009:DBD

[BX09] Peter J. Bickel and Ying Xu. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1266– 1269, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453934. See [SR09b].

Benjamini:2013:SEE

[BY13] Yuval Benjamini and Bin Yu. The shuffle estimator for explainable variance in fMRI experiments. Annals of Applied Statistics, 7(4):2007–2033, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1387823308.

Bader:2018:ATS

[BYZ18] Brian Bader, Jun Yan, and Xuebin Zhang. Automated threshold selection for extreme value analysis via ordered goodness-of-fit tests with adjustment for false discovery rate. Annals of Applied Statistics, 12(1):310–329, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564474.

Bartlett:2016:DEN

[BZ16] Thomas E. Bartlett and Alexey Zaikin. Detection of epigenomic network community oncomarkers. Annals of Applied Statistics, 10(3):1373–1396, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1475069611.

Berg:2019:LDM

[BZC^{+}19] Stephen Berg, Jun Zhu, Murray K. Clayton, Monika E. Shea, and David J. Mladenoff. A latent discrete Markov random field approach to identifying and classifying historical forest communities based on spatial multivariate tree species counts. Annals of Applied Statistics, 13(4):2312–2340, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910046.

Bodwin:2018:TBA

[BZN18] Kelly Bodwin, Kai Zhang, and Andrew Nobel. A testing based approach to the discovery of differentially correlated variable sets. Annals of Applied Statistics, 12(2):1180–1203, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1532743490.

Baek:2019:BAI

[BZS19] Jonggyu Baek, Bin Zhu, and Peter X. K. Song. Bayesian analysis of infant’s growth dynamics with in utero exposure to environmental toxicants. Annals of Applied Statistics, 13(1):297–320, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861650.

Conn:2018:SMM

[CA18] Paul B. Conn and Ray T. Alisauskas. Simultaneous modelling of movement, measurement error, and observer dependence in mark-recapture distance sampling: an application to Arctic bird surveys. Annals of Applied Statistics, 12(1): 96–122, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564466.

Chen:2022:RCI

[CA22] Aiyou Chen and Timothy C. Au. Robust causal inference for incremental return on ad spend with randomized paired geo experiments. Annals of Applied Statistics, 16(1):

1–20, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Robust-causal-inference-for-incremental-
return-on-ad-spend-with/10.1214/21-AOAS1493.full.

Chang:2023:SGM

[CA23] Andersen Chang and Genevera I. Allen. Subbotin graphical models for extreme value dependencies with applications to functional neuronal connectivity. Annals of Applied Statistics, 17(3):2364–2386, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Subbotin-graphical-
models-for-extreme-value-dependencies-with-applications-
to/10.1214/22-AOAS1723.full.

Criscuolo:2023:HCF

[CAL^{+}23] Tulio L. Criscuolo, Renato M. Assunção, Rosangela H. Loschi, Wagner Meira, Jr., and Danna Cruz-Reyes. Handling categorical features with many levels using a product partition model. Annals of Applied Statistics, 17(1):786– 814, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Handling-categorical-features-with-many-
levels-using-a-product-partition/10.1214/22-AOAS1651.
full.

Cheng:2020:SPM

[CAS20] David Cheng, Rajeev Ayyagari, and James Signorovitch. The statistical performance of matching-adjusted indirect comparisons: Estimating treatment effects with aggregate external control data. Annals of Applied Statistics, 14(4):1806– 1833, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/The-statistical-performance-of-matching-
adjusted-indirect-comparisons--Estimating/10.1214/
20-AOAS1359.full.

Crispino:2019:BMA

[CAV^{+}19] Marta Crispino, Elja Arjas, Valeria Vitelli, Natasha Barrett, and Arnoldo Frigessi. A Bayesian Mallows approach to nontransitive pair comparison data: How human are sounds? Annals of Applied Statistics, 13(1):492–519, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1554861658.

Chang:2010:HRM

[CB10] Jonathan Chang and David M. Blei. Hierarchical relational models for document networks. Annals of Applied Statistics, 4(1):124–150, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1273584450.

Clarke:2022:EME

[CB22] Paul S. Clarke and Yanchun Bao. Estimating mode effects from a sequential mixed-mode experiment using structural moment models. Annals of Applied Statistics, 16 (3):1563–1585, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-16/issue-3/Estimating-mode-effects-from-a-
sequential-mixed-mode-experiment-using/10.1214/21-
AOAS1557.full.

Cruyff:2008:ASP

[CBvdHvdH08] Maarten J. L. F. Cruyff, Ulf Böckenholt, Ardo van den Hout, and Peter G. M. van der Heijden. Accounting for selfprotective responses in randomized response data from a social security survey using the zero-inflated Poisson model. Annals of Applied Statistics, 2(1):316–331, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1206367823.

Chang:2017:MLF

[CBZG17] Lo-Bin Chang, Eran Borenstein, Wei Zhang, and Stuart Geman. Maximum likelihood features for generative image models. Annals of Applied Statistics, 11(3):1275–1308, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1507168830.

Castelletti:2019:OBM

[CC19] Federico Castelletti and Guido Consonni. Objective Bayes model selection of Gaussian interventional essential graphs for the identification of signaling pathways. Annals of Applied Statistics, 13(4):2289–2311, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910045.

Cheung:2022:NFE

[CCD22] Ying Kuen Cheung, Thevaa Chandereng, and Keith M. Diaz. A novel framework to estimate multidimensional minimum effective doses using asymmetric posterior gain and ϵ-tapering. Annals of Applied Statistics, 16(3):1445–1458, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/A-novel-framework-to-estimate-multidimensional-
minimum-effective-doses-using/10.1214/21-AOAS1549.
full.

Castro-Camilo:2018:TVE

[CCdCW18] Daniela Castro-Camilo, Miguel de Carvalho, and Jennifer Wadsworth. Time-varying extreme value dependence with application to leading European stock markets. Annals of Applied Statistics, 12(1):283–309, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1520564473.

Chiou:2019:IMC

[CCH19] Jeng-Min Chiou, Yu-Ting Chen, and Tailen Hsing. Identifying multiple changes for a functional data sequence with application to freeway traffic segmentation. Annals of Applied Statistics, 13(3):1430–1463, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1571277759.

Carmichael:2021:JIA

[CCH^{+}21] Iain Carmichael, Benjamin C. Calhoun, Katherine A. Hoadley, Melissa A. Troester, Joseph Geradts, Heather D. Couture, Linnea Olsson, Charles M. Perou, Marc Niethammer, Jan Hannig, and J. S. Marron. Joint and individual analysis of breast cancer histologic images and genomic covariates. Annals of Applied Statistics, 15(4):1697–1722, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Joint-and-individual-analysis-of-breast-
cancer-histologic-images-and/10.1214/20-AOAS1433.
full.

Chien:2009:PTC

[CCJ^{+}09] Li-Chu Chien, I-Shou Chang, Shih Sheng Jiang, Pramod K. Gupta, Chi-Chung Wen, Yuh-Jenn Wu, and Chao A. Hsiung. Profiling time course expression of virus genes — an illustration of Bayesian inference under shape restrictions. Annals of Applied Statistics, 3(4):1542–1565, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453952.

Chung:2018:BRR

[CCS18] Ray S. W. Chung, Amanda M. Y. Chu, and Mike K. P. So. Bayesian randomized response technique with multiple sensitive attributes: The case of information systems resource misuse. Annals of Applied Statistics, 12(3):1969–1992, September 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1536652982.

Chambaz:2012:CPS

[CD12] Antoine Chambaz and Christophe Denis. Classification in postural style. Annals of Applied Statistics, 6(3):977–993, September 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1346418570.

Cule:2017:SMI

[CD17] Madeleine Cule and Peter Donnelly. Stochastic modelling and inference in electronic hospital databases for the spread of infections: Clostridium difficile transmission in Oxfordshire hospitals 2007–2010. Annals of Applied Statistics, 11(2): 655–679, June 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1500537718.

Clark:2018:MES

[CD18] Nicholas J. Clark and Philip M. Dixon. Modeling and estimation for self-exciting spatio-temporal models of terrorist activity. Annals of Applied Statistics, 12(1):633–653, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564487.

Chen:2020:MHA

[CD20] Yen-Chi Chen and Adrian Dobra. Measuring human activity spaces from GPS data with density ranking and summary curves. Annals of Applied Statistics, 14(1):409–432, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002680.

Czogiel:2011:BMU

[CDB11] Irina Czogiel, Ian L. Dryden, and Christopher J. Brignell. Bayesian matching of unlabeled marked point sets using random fields, with an application to molecular alignment. Annals of Applied Statistics, 5(4):2603–2629, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399608.

Chiarucci:2018:JIE

[CDF^{+}18] Alessandro Chiarucci, Rosa Maria Di Biase, Lorenzo Fattorini, Marzia Marcheselli, and Caterina Pisani. Joining the incompatible: Exploiting purposive lists for the sample-based estimation of species richness. Annals of Applied Statistics, 12(3):1679–1699, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652970.

Camerlenghi:2020:NBM

[CDF^{+}20] Federico Camerlenghi, Bianca Dumitrascu, Federico Ferrari, Barbara E. Engelhardt, and Stefano Favaro. Nonparametric Bayesian multiarmed bandits for single-cell experiment design. Annals of Applied Statistics, 14(4):2003–2019, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Nonparametric-Bayesian-multiarmed-bandits-

for-single-cell-experiment-design/10.1214/20-AOAS1370.
full.

Cheng:2018:CEG

[CDM18] Yicheng Cheng, Murat Dundar, and George Mohler. A coupled ETAS-I^{2} GMM point process with applications to seismic fault detection. Annals of Applied Statistics, 12 (3):1853–1870, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652977.

Cooley:2012:ACD

[CDN12] Daniel Cooley, Richard A. Davis, and Philippe Naveau. Approximating the conditional density given large observed values via a multivariate extremes framework, with application to environmental data. Annals of Applied Statistics, 6(4):1406–1429, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629045.

Cheng:2017:QMT

[CDP^{+}17] Yichen Cheng, James Y. Dai, Thomas G. Paulson, Xiaoyu Wang, Xiaohong Li, Brian J. Reid, and Charles Kooperberg. Quantification of multiple tumor clones using gene array and sequencing data. Annals of Applied Statistics, 11(2):967– 991, June 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1500537731.

Campbell:2022:BAP

[CdVM^{+}22] Harlan Campbell, Perry de Valpine, Lauren Maxwell, Valentijn M. T. de Jong, Thomas P. A. Debray, Thomas Jaenisch, and Paul Gustafson. Bayesian adjustment for preferential testing in estimating infection fatality rates, as motivated by the COVID-19 pandemic. Annals of Applied Statistics, 16(1): 436–459, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Bayesian-adjustment-for-preferential-testing-
in-estimating-infection-fatality-rates/10.1214/21-
AOAS1499.full.

Chervoneva:2014:END

[CFH^{+}14] Inna Chervoneva, Boris Freydin, Brian Hipszer, Tatiyana V. Apanasovich, and Jeffrey I. Joseph. Estimation of nonlinear differential equation model for glucose-insulin dynamics in type I diabetic patients using generalized smoothing. Annals of Applied Statistics, 8(2):886–904, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229518.

Carota:2015:BND

[CFLP15] Cinzia Carota, Maurizio Filippone, Roberto Leombruni, and Silvia Polettini. Bayesian nonparametric disclosure risk estimation via mixed effects log-linear models. Annals of Applied Statistics, 9(1):525–546, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226103.

Casarin:2018:UTL

[CFMR18] Roberto Casarin, Claudia Foroni, Massimiliano Marcellino, and Francesco Ravazzolo. Uncertainty through the lenses of a mixed-frequency Bayesian panel Markov-switching model. Annals of Applied Statistics, 12(4):2559–2586, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078056.

Crawford:2019:VPN

[CFRW19] Lorin Crawford, Seth R. Flaxman, Daniel E. Runcie, and Mike West. Variable prioritization in nonlinear black box methods: a genetic association case study. Annals of Applied Statistics, 13(2):958–989, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758434.

Cantoni:2017:REH

[CFW17] E. Cantoni, J. Mills Flemming, and A. H. Welsh. A randomeffects hurdle model for predicting bycatch of endangered marine species. Annals of Applied Statistics, 11(4):2178–2199, December 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1514430282.

Chiquet:2012:SSC

[CGC12] Julien Chiquet, Yves Grandvalet, and Camille Charbonnier. Sparsity with sign-coherent groups of variables via the cooperative-Lasso. Annals of Applied Statistics, 6(2):795– 830, June 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1339419617.

Chao:2021:GES

[CGCA21] Fengqing Chao, Patrick Gerland, Alex R. Cook, and Leontine Alkema. Global estimation and scenario-based projections of sex ratio at birth and missing female births using a Bayesian hierarchical time series mixture model. Annals of Applied Statistics, 15(3):1499–1528, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Global-estimation-and-scenario-based-projections-
of-sex-ratio-at/10.1214/20-AOAS1436.full.

Chkrebtii:2022:ISK

[CGCN22] Oksana A. Chkrebtii, Yury E. García, Marcos A. Capistrán, and Daniel E. Noyola. Inference for stochastic kinetic models from multiple data sources for joint estimation of infection dynamics from aggregate reports and virological data. Annals of Applied Statistics, 16(2):959–981, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Inference-for-stochastic-kinetic-models-
from-multiple-data-sources-for/10.1214/21-AOAS1527.
full.

Cui:2015:HHM

[CGFT15] Shiqi Cui, Subharup Guha, Marco A. R. Ferreira, and Allison N. Tegge. hmmSeq: a hidden Markov model for detecting differentially expressed genes from RNA-seq data. Annals of Applied Statistics, 9(2):901–925, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1437397117.

Cannamela:2008:CSQ

[CGI08] Claire Cannamela, Josselin Garnier, and Bertrand Iooss. Controlled stratification for quantile estimation. Annals of Applied Statistics, 2(4):1554–1580, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1231424222.

Chipman:2010:BBA

[CGM10] Hugh A. Chipman, Edward I. George, and Robert E. McCulloch. BART: Bayesian additive regression trees. Annals of Applied Statistics, 4(1):266–298, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584455.

Cagnone:2017:MMS

[CGM17] Silvia Cagnone, Simone Giannerini, and Lucia Modugno. Multilevel models with stochastic volatility for repeated cross-sections: an application to tribal art prices. Annals of Applied Statistics, 11(2):1040–1062, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1500537734.

Chung:2022:PIB

[CGN22] Hee Cheol Chung, Irina Gaynanova, and Yang Ni. Phylogenetically informed Bayesian truncated copula graphical models for microbial association networks. Annals of Applied Statistics, 16(4):2437–2457, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Phylogenetically-
informed-Bayesian-truncated-copula-graphical-models-
for-microbial-association/10.1214/21-AOAS1598.full.

Cassese:2014:HBM

[CGT^{+}14] Alberto Cassese, Michele Guindani, Mahlet G. Tadesse, Francesco Falciani, and Marina Vannucci. A hierarchical Bayesian model for inference of copy number variants and their association to gene expression. Annals of Applied Statistics, 8(1):148–175, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966282.

Chakraborty:2010:MLS

[CGW^{+}10] Avishek Chakraborty, Alan E. Gelfand, Adam M. Wilson, Andrew M. Latimer, and John A. Silander, Jr. Modeling large scale species abundance with latent spatial processes. Annals of Applied Statistics, 4(3):1403–1429, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1287409379.

Cotton:2014:EED

[CH14] Cecilia A. Cotton and Patrick J. Heagerty. Evaluating epoetin dosing strategies using observational longitudinal data. Annals of Applied Statistics, 8(4):2356–2377, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1419001747.

Chakrabarti:2017:MNI

[Cha17] Deepayan Chakrabarti. Modeling node incentives in directed networks. Annals of Applied Statistics, 11(4):2298–2331, December 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1514430287.

Chang:2016:IIS

[CHAP16] Won Chang, Murali Haran, Patrick Applegate, and David Pollard. Improving ice sheet model calibration using paleoclimate and modern data. Annals of Applied Statistics, 10(4):2274–2302, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606860.

Chen:2014:SCA

[CHH^{+}14] Ting-Li Chen, Dai-Ni Hsieh, Hung Hung, I-Ping Tu, Pei-Shien Wu, Yi-Ming Wu, Wei-Hau Chang, and Su-Yun Huang. γ-SUP: a clustering algorithm for cryo-electron microscopy images of asymmetric particles. Annals of Applied Statistics, 8(1):259–285, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966286.

Chiou:2012:DFP

[Chi12] Jeng-Min Chiou. Dynamical functional prediction and classification, with application to traffic flow prediction. Annals of Applied Statistics, 6(4):1588–1614, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629052.

Chu:2018:EUK

[CHJCK18] Ba Chu, Kim Huynh, David Jacho-Chávez, and Oleksiy Kryvtsov. On the evolution of the United Kingdom price distributions. Annals of Applied Statistics, 12(4):2618–2646, December 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1542078058.

Chang:2014:FDR

[CHOK14] Won Chang, Murali Haran, Roman Olson, and Klaus Keller. Fast dimension-reduced climate model calibration and the effect of data aggregation. Annals of Applied Statistics, 8(2): 649–673, June 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1404229509.

Chen:2016:LLA

[CHS^{+}16] Kun Chen, Eric A. Hoffman, Indu Seetharaman, Feiran Jiao, Ching-Long Lin, and Kung-Sik Chan. Linking lung airway structure to pulmonary function via composite bridge regression. Annals of Applied Statistics, 10(4):1880–1906, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606844.

Chen:2017:ASC

[CJM^{+}17] Hao Chen, Yuchao Jiang, Kara N. Maxwell, Katherine L. Nathanson, and Nancy Zhang. Allele-specific copy number estimation by whole exome sequencing. Annals of Applied Statistics, 11(2):1169–1192, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537739.

Caye:2018:FII

[CJMF18] Kevin Caye, Flora Jay, Olivier Michel, and Olivier François. Fast inference of individual admixture coefficients using geographic data. Annals of Applied Statistics, 12(1):586– 608, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564485.

Chen:2014:TPS

[CK14] Sixia Chen and Jae-Kwang Kim. Two-phase sampling experiment for propensity score estimation in self-selected samples. Annals of Applied Statistics, 8(3):1492–1515, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1414091222.

Cahill:2015:MSL

[CKHP15] Niamh Cahill, Andrew C. Kemp, Benjamin P. Horton, and Andrew C. Parnell. Modeling sea-level change using errorsin-variables integrated Gaussian processes. Annals of Applied Statistics, 9(2):547–571, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1437397101.

Chang:2022:IMC

[CKK^{+}22] Won Chang, Bledar A. Konomi, Georgios Karagiannis, Yawen Guan, and Murali Haran. Ice model calibration using semicontinuous spatial data. Annals of Applied Statistics, 16(3):1937–1961, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Ice-model-calibration-
using-semicontinuous-spatial-data/10.1214/21-AOAS1577.
full.

Cho:2021:APC

[CKM21] Min Ho Cho, Sebastian Kurtek, and Steven N. MacEachern. Aggregated pairwise classification of elastic planar shapes. Annals of Applied Statistics, 15(2):619–637, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Aggregated-pairwise-classification-of-
elastic-planar-shapes/10.1214/21-AOAS1452.full.

Cigsar:2012:ATC

[ÇL12] Candemir Çiğşar and Jerald F. Lawless. Assessing transient carryover effects in recurrent event processes, with application to chronic health conditions. Annals of Applied Statistics, 6(4):1641–1663, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1356629054.

Chen:2013:VSS

[CL13] Jun Chen and Hongzhe Li. Variable selection for sparse Dirichlet-multinomial regression with an application to microbiome data analysis. Annals of Applied Statistics, 7(1): 418–442, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527205.

Chan:2014:TSS

[CLEB14] Kung-Sik Chan, Jinzheng Li, William Eichinger, and Erwei Bai. Testing for shielding of special nuclear weapon materials. Annals of Applied Statistics, 8(1):553–576, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1396966298.

Chen:2023:VBA

[CLFC23] Xinyuan Chen, Yiwei Li, Xiangnan Feng, and Joseph T. Chang. Variational Bayesian analysis of nonhomogeneous hidden Markov models with long and ultralong sequences. Annals of Applied Statistics, 17(2):1615–1640, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Variational-Bayesian-analysis-of-nonhomogeneous-
hidden-Markov-models-with-long/10.1214/22-AOAS1685.
full.

Crook:2022:SSN

[CLGK22] Oliver M. Crook, Kathryn S. Lilley, Laurent Gatto, and Paul D. W. Kirk. Semi-supervised nonparametric Bayesian modelling of spatial proteomics. Annals of Applied Statistics, 16(4):2554–2576, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Semi-supervised-
nonparametric-Bayesian-modelling-of-spatial-proteomics/
10.1214/22-AOAS1603.full.

Chen:2012:SPG

[CLK^{+}12] Xi Chen, Qihang Lin, Seyoung Kim, Jaime G. Carbonell, and Eric P. Xing. Smoothing proximal gradient method for general structured sparse regression. Annals of Applied Statistics, 6(2):719–752, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419614.

Crossett:2013:RGI

[CLK^{+}13] Andrew Crossett, Ann B. Lee, Lambertus Klei, Bernie Devlin, and Kathryn Roeder. Refining genetically inferred relationships using treelet covariance smoothing. Annals of Applied Statistics, 7(2):669–690, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338463.

Chu:2020:FSG

[CLLR20] Wanghuan Chu, Runze Li, Jingyuan Liu, and Matthew Reimherr. Feature selection for generalized varying coefficient mixed-effect models with application to obesity GWAS. Annals of Applied Statistics, 14(1):276–298, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002675.

Chen:2022:DTW

[CLM22] Yunxiao Chen, Yan Lu, and Irini Moustaki. Detection of two-way outliers in multivariate data and application to cheating detection in educational tests. Annals of Applied Statistics, 16(3):1718–1746, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Detection-of-
two-way-outliers-in-multivariate-data-and-application/
10.1214/21-AOAS1564.full.

Chu:2016:FST

[CLR16] Wanghuan Chu, Runze Li, and Matthew Reimherr. Feature screening for time-varying coefficient models with ultrahighdimensional longitudinal data. Annals of Applied Statistics, 10(2):596–617, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199886.

Chen:2022:PTU

[CLTZ22] Jiahua Chen, Yukun Liu, Carilyn G. Taylor, and James V. Zidek. Permutation tests under a rotating sampling plan with clustered data. Annals of Applied Statistics, 16(2):936– 958, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Permutation-tests-under-a-rotating-sampling-
plan-with-clustered-data/10.1214/21-AOAS1526.full.

Chakraborty:2020:SAN

[CLW20] Arnab Chakraborty, Soumendra Nath Lahiri, and Alyson Wilson. A statistical analysis of noisy crowdsourced weather data. Annals of Applied Statistics, 14(1):116–142, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002667.

Chernoff:2009:DIV

[CLZ09] Herman Chernoff, Shaw-Hwa Lo, and Tian Zheng. Discovering influential variables: a method of partitions. Annals of Applied Statistics, 3(4):1335–1369, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453943.

Chen:2016:CPA

[CLZ16] Mengjie Chen, Haifan Lin, and Hongyu Zhao. Change point analysis of histone modifications reveals epigenetic blocks linking to physical domains. Annals of Applied Statistics, 10(1):506–526, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1458909925.

Cabella:2009:SCA

[CM09] Paolo Cabella and Domenico Marinucci. Statistical challenges in the analysis of cosmic microwave background radiation. Annals of Applied Statistics, 3(1):61–95, March 2009.

CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888363.

Castillo-Mateo:2023:SQA

[CMAC^{+}23] Jorge Castillo-Mateo, Jesús Asín, Ana C. Cebrián, Alan E. Gelfand, and Jesús Abaurrea. Spatial quantile autoregression for season within year daily maximum temperature data. Annals of Applied Statistics, 17(3):2305–2325, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Spatial-quantile-autoregression-for-season-
within-year-daily-maximum-temperature/10.1214/22-
AOAS1719.full.

Culp:2009:MVL

[CMJ09] Mark Culp, George Michailidis, and Kjell Johnson. On multi-view learning with additive models. Annals of Applied Statistics, 3(1):292–318, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1239888372.

Chen:2022:ADG

[CMJZ22] Jialei Chen, Simon Mak, V. Roshan Joseph, and Chuck Zhang. Adaptive design for Gaussian process regression under censoring. Annals of Applied Statistics, 16(2):744– 764, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Adaptive-design-for-Gaussian-process-regression-
under-censoring/10.1214/21-AOAS1512.full.

Coulombe:2022:EME

[CMPR22] Janie Coulombe, Erica E. M. Moodie, Robert W. Platt, and Christel Renoux. Estimation of the marginal effect of antidepressants on body mass index under confounding and endogenous covariate-driven monitoring times. Annals of Applied Statistics, 16(3):1868–1890, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Estimation-of-the-marginal-effect-of-antidepressants-
on-body-mass/10.1214/21-AOAS1570.full.

Chekouo:2015:GPB

[CMR15] Thierry Chekouo, Alejandro Murua, and Wolfgang Raffelsberger. The Gibbs-plaid biclustering model. Annals of Applied Statistics, 9(3):1643–1670, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1446488755.

Chiquet:2018:VIP

[CMR18] Julien Chiquet, Mahendra Mariadassou, and Stéphane Robin. Variational inference for probabilistic Poisson PCA. Annals of Applied Statistics, 12(4):2674–2698, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078060.

Chen:2019:MSS

[CMZ19] Ying Chen, J. S. Marron, and Jiejie Zhang. Modeling seasonality and serial dependence of electricity price curves with warping functional autoregressive dynamics. Annals of Applied Statistics, 13(3):1590–1616, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1571277765.

Cook:2007:ESL

[CN07] R. Dennis Cook and Liqiang Ni. Elevated soil lead: Statistical modeling and apportionment of contributions from lead-based paint and leaded gasoline. Annals of Applied Statistics, 1(1):130–151, June 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1183143732.

Crawford:2023:RBF

[COC23] Amy M. Crawford, Danica M. Ommen, and Alicia L. Carriquiry. A rotation-based feature and Bayesian hierarchical model for the forensic evaluation of handwriting evidence in a closed set. Annals of Applied Statistics, 17(2):1127–1151, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-2/A-
rotation-based-feature-and-Bayesian-hierarchical-
model-for-the/10.1214/22-AOAS1662.full.

Chakraborty:2022:BSL

[COD22] Antik Chakraborty, Otso Ovaskainen, and David B. Dunson. Bayesian semiparametric long memory models for discretized event data. Annals of Applied Statistics, 16(3):1380–1399, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Bayesian-semiparametric-long-memory-models-
for-discretized-event-data/10.1214/21-AOAS1546.full.

Casa:2022:PBF

[COM22] Alessandro Casa, Tom F. O’Callaghan, and Thomas Brendan Murphy. Parsimonious Bayesian factor analysis for modelling latent structures in spectroscopy data. Annals of Applied Statistics, 16(4):2417–2436, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Parsimonious-Bayesian-factor-analysis-
for-modelling-latent-structures-in-spectroscopy/10.
1214/21-AOAS1597.full.

Cope:2009:DBD

[Cop09] Leslie Cope. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1279–1281, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1267453936. See [SR09b].

Cox:2007:ASR

[Cox07] D. R. Cox. Applied statistics: a review. Annals of Applied Statistics, 1(1):1–16, June 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1183143726.

Cappello:2020:SIS

[CP20] Lorenzo Cappello and Julia A. Palacios. Sequential importance sampling for multiresolution Kingman–Tajima coalescent counting. Annals of Applied Statistics, 14(2):727–751, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1593449323.

Capezza:2021:ASD

[CPG^{+}21] Christian Capezza, Biagio Palumbo, Yannig Goude, Simon N. Wood, and Matteo Fasiolo. Additive stacking for disaggregate electricity demand forecasting. Annals of Applied Statistics, 15(2):727–746, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Additive-stacking-
for-disaggregate-electricity-demand-forecasting/10.
1214/20-AOAS1417.full.

Cervone:2014:LMA

[CPP^{+}14] Daniel Cervone, Natesh S. Pillai, Debdeep Pati, Ross Berbeco, and John Henry Lewis. A location-mixture autoregressive model for online forecasting of lung tumor motion. Annals of Applied Statistics, 8(3):1341–1371, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1414091216.

Conti:2011:MHI

[CPvV^{+}11] Stefano Conti, Anne M. Presanis, Maaike G. van Veen, Maria Xiridou, Martin C. Donoghoe, Annemarie Rinder Stengaard, and Daniela De Angelis. Modeling of the HIV infection epidemic in the Netherlands: a multi-parameter evidence synthesis approach. Annals of Applied Statistics, 5(4):2359– 2384, December 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1324399598.

Chatterjee:2009:DFC

[CQ09] Snigdhansu Chatterjee and Peihua Qiu. Distribution-free cumulative sum control charts using bootstrap-based control limits. Annals of Applied Statistics, 3(1):349–369, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888374.

Craigmile:2011:DSA

[CR11] Peter Craigmile and Bala Rajaratnam. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):88–90, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715181. See [MW11b].

Chandler:2013:SEM

[CR13] Richard B. Chandler and J. Andrew Royle. Spatially explicit models for inference about density in unmarked or partially marked populations. Annals of Applied Statistics, 7(2):936– 954, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338474.

Crawford:2016:DCC

[Cra16] Forrest W. Crawford. Discussion of “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1842–1845, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606841. See [JJ16a,

JJ16b].

Cui:2013:DEH

[CRZ13] Qiurong Cui, Karl Rohe, and Zhengjun Zhang. Discussion of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1891–1894, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823301. See [SR09b].

Castruccio:2013:GST

[CS13] Stefano Castruccio and Michael L. Stein. Global space– time models for climate ensembles. Annals of Applied Statistics, 7(3):1593–1611, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804808.

Cybis:2015:APC

[CSB^{+}15] Gabriela B. Cybis, Janet S. Sinsheimer, Trevor Bedford, Alison E. Mather, Philippe Lemey, and Marc A. Suchard. Assessing phenotypic correlation through the multivariate phylogenetic latent liability model. Annals of Applied Statistics, 9(2):969–991, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1437397120.

Chen:2012:BME

[CSC^{+}12] Chung-Hsing Chen, Wen-Chi Su, Chih-Yu Chen, Jing-Ying Huang, Fang-Yu Tsai, Wen-Chang Wang, Chao A. Hsiung, King-Song Jeng, and I-Shou Chang. A Bayesian measurement error model for two-channel cell-based RNAi data with replicates. Annals of Applied Statistics, 6(1):356–382, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1331043400.

Chekouo:2016:BPM

[CSGD16] Thierry Chekouo, Francesco C. Stingo, Michele Guindani, and Kim-Anh Do. A Bayesian predictive model for imaging genetics with application to schizophrenia. Annals of Applied Statistics, 10(3):1547–1571, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1475069618.

Commenges:2008:EDK

[CSL^{+}08] D. Commenges, A. Sayyareh, L. Letenneur, J. Guedj, and A. Bar-Hen. Estimating a difference of Kullback–Leibler risks using a normalized difference of AIC. Annals of Applied Statistics, 2(3):1123–1142, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1223908055.

Clements:2011:RAM

[CSS11] Robert Alan Clements, Frederic Paik Schoenberg, and Danijel Schorlemmer. Residual analysis methods for space–time point processes with applications to earthquake forecast models in California. Annals of Applied Statistics, 5(4):2549– 2571, December 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1324399606.

Chen:2018:UEE

[CSS18] Beidi Chen, Anshumali Shrivastava, and Rebecca C. Steorts. Unique entity estimation with application to the Syrian conflict. Annals of Applied Statistics, 12(2):1039–1067, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743485.

Clements:2014:AMT

[CSZK14] Nicolle Clements, Sanat K. Sarkar, Zhigen Zhao, and Dong-Yun Kim. Applying multiple testing procedures to detect change in East African vegetation. Annals of Applied Statistics, 8(1):286–308, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966287.

Coram:2007:IPS

[CT07] Marc Coram and Hua Tang. Improving population-specific allele frequency estimates by adapting supplemental data: an empirical Bayes approach. Annals of Applied Statistics, 1(2):459–479, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1196438027.

Chen:2018:MSE

[CT18] Feng Chen and Wai Hong Tan. Marked self-exciting point process modelling of information diffusion on Twitter. Annals of Applied Statistics, 12(4):2175–2196, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078041.

Chailan:2017:SMS

[CTB17] Romain Chailan, Gwladys Toulemonde, and Jean-Noel Bacro. A semiparametric method to simulate bivariate spacetime extremes. Annals of Applied Statistics, 11(3):1403–1428, September 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1507168834.

Caron:2014:BNP

[CTM14] François Caron, Yee Whye Teh, and Thomas Brendan Murphy. Bayesian nonparametric Plackett–Luce models for the analysis of preferences for college degree programmes. Annals of Applied Statistics, 8(2):1145–1181, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1404229529.

Cutler:2010:RLB

[Cut10] Adele Cutler. Remembering Leo Breiman. Annals of Applied Statistics, 4(4):1621–1633, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167785.

Couturier:2010:ZIT

[CVF10] Dominique-Laurent Couturier and Maria-Pia Victoria-Feser. Zero-inflated truncated generalized Pareto distribution for the analysis of radio audience data. Annals of Applied Statistics, 4(4):1824–1846, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167800.

Chan:2010:BES

[CW10] Kwun Chuen Gary Chan and Mei-Cheng Wang. Backward estimation of stochastic processes with failure events as time origins. Annals of Applied Statistics, 4(3):1602–1620, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409388.

Clauset:2013:EHF

[CW13a] Aaron Clauset and Ryan Woodard. Estimating the historical and future probabilities of large terrorist events. Annals of Applied Statistics, 7(4):1838–1865, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823295. See discussion [Moh13a, RP13, Whi13, Gil13, Sch13, CRZ13] and rejoinder [CW13b].

Clauset:2013:REH

[CW13b] Aaron Clauset and Ryan Woodard. Rejoinder of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1895–1897, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823302. See [SR09b].

Cheng:2020:SMA

[CW20] Chao Cheng and Molin Wang. Statistical methods for analysis of combined categorical biomarker data from multiple studies. Annals of Applied Statistics, 14(3):1146–1163, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Statistical-methods-for-analysis-of-combined-
categorical-biomarker-data-from/10.1214/20-AOAS1337.
full.

Chen:2018:UBV

[CWE18] Yen-Chi Chen, Y. Samuel Wang, and Elena A. Erosheva. On the use of bootstrap with variational inference: Theory, interpretation, and a two-sample test example. Annals of Applied Statistics, 12(2):846–876, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1532743479.

Cunen:2020:FMS

[CWH20] Céline Cunen, Lars Walløe, and Nils Lid Hjort. Focused model selection for linear mixed models with an application to whale ecology. Annals of Applied Statistics, 14(2):872– 904, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449330.

Crawford:2015:SLS

[CWS15] Forrest W. Crawford, Robert E. Weiss, and Marc A. Suchard. Sex, lies and self-reported counts: Bayesian mixture models for heaping in longitudinal count data via birth-death processes. Annals of Applied Statistics, 9(2):572–596, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1437397102.

Chen:2017:MEM

[CWWW17] Lin S. Chen, Jiebiao Wang, Xianlong Wang, and Pei Wang. A mixed-effects model for incomplete data from labelingbased quantitative proteomics experiments. Annals of Applied Statistics, 11(1):114–138, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1491616874.

Choudhury:2010:ASD

[CZM10] Kingshuk Roy Choudhury, Limian Zheng, and John J. Mackrill. Analysis of spatial distribution of marker expression in cells using boundary distance plots. Annals of Applied Statistics, 4(3):1365–1382, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1287409377.

DAngelo:2022:IST

[DAAM22] Nicoletta D’Angelo, Giada Adelfio, Antonino Abbruzzo, and Jorge Mateu. Inhomogeneous spatio-temporal point processes on linear networks for visitors’ stops data. Annals of Applied Statistics, 16(2):791–815, June 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-2/Inhomogeneous-
spatio-temporal-point-processes-on-linear-networks-
for-visitors/10.1214/21-AOAS1519.full.

Denti:2023:HMM

[DAL^{+}23] Francesco Denti, Ricardo Azevedo, Chelsie Lo, Damian G. Wheeler, Sunil P. Gandhi, Michele Guindani, and Babak Shahbaba. A horseshoe mixture model for Bayesian screening with an application to light sheet fluorescence microscopy in brain imaging. Annals of Applied Statistics, 17(3):2639–2658, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/A-horseshoe-mixture-model-for-Bayesian-
screening-with-an-application/10.1214/23-AOAS1736.
full.

Davis:2017:DEB

[Dav17] Mark H. A. Davis. Discussion of “Elicitability and backtesting: Perspectives for banking regulation”. Annals of Applied Statistics, 11(4):1886–1887, December 2017. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1514430268. See [NZ17a].

Dagum:2015:NSA

[DB15] Estela Bee Dagum and Silvia Bianconcini. A new set of asymmetric filters for tracking the short-term trend in real-time. Annals of Applied Statistics, 9(3):1433–1458, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488746.

Dirmeier:2022:SHM

[DB22] Simon Dirmeier and Niko Beerenwinkel. Structured hierarchical models for probabilistic inference from perturbation screening data. Annals of Applied Statistics, 16 (3):2010–2029, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-16/issue-3/Structured-hierarchical-models-
for-probabilistic-inference-from-perturbation-screening-
data/10.1214/21-AOAS1580.full.

Datta:2016:NDN

[DBF^{+}16] Abhirup Datta, Sudipto Banerjee, Andrew O. Finley, Nicholas A. S. Hamm, and Martijn Schaap. Nonseparable dynamic nearest neighbor Gaussian process models for large spatio-temporal data with an application to particulate matter analysis. Annals of Applied Statistics, 10(3):1286–1316, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069608.

DAngelo:2021:IFI

[DBG21] Silvia D’Angelo, Lorraine Brennan, and Isobel Claire Gormley. Inferring food intake from multiple biomarkers using a latent variable model. Annals of Applied Statistics, 15(4):2043– 2060, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Inferring-food-intake-from-multiple-biomarkers-
using-a-latent-variable/10.1214/21-AOAS1478.full.

DeVito:2021:BMF

[DBTP21] Roberta De Vito, Ruggero Bellio, Lorenzo Trippa, and Giovanni Parmigiani. Bayesian multistudy factor analysis for high-throughput biological data. Annals of Applied Statistics, 15(4):1723–1741, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Bayesian-multistudy-
factor-analysis-for-high-throughput-biological-data/
10.1214/21-AOAS1456.full.

Di:2009:MFP

[DCCP09] Chong-Zhi Di, Ciprian M. Crainiceanu, Brian S. Caffo, and Naresh M. Punjabi. Multilevel functional principal component analysis. Annals of Applied Statistics, 3(1):458– 488, March 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1239888378.

deCarvalho:2016:FCA

[dCdCAGM16] Vanda Inácio de Carvalho, Miguel de Carvalho, Todd A. Alonzo, and Wenceslao González-Manteiga. Functional covariate-adjusted partial area under the specificity-ROC curve with an application to metabolic syndrome diagnosis. Annals of Applied Statistics, 10(3):1472–1495, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1475069615.

Dai:2021:LCN

[DCHP21] Chenguang Dai, Duo Chan, Peter Huybers, and Natesh Pillai. Late 19th century navigational uncertainties and their influence on sea surface temperature estimates. Annals of Applied Statistics, 15(1):22–40, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Late-19th-century-
navigational-uncertainties-and-their-influence-on-
sea/10.1214/20-AOAS1367.full.

deCastro:2010:TAP

[dCP10] Luciano I. de Castro and Harry J. Paarsch. Testing affiliation in private-values models of first-price auctions using grid distributions. Annals of Applied Statistics, 4(4):2073– 2098, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167810.

Durante:2016:LAD

[DD16] Daniele Durante and David B. Dunson. Locally adaptive dynamic networks. Annals of Applied Statistics, 10(4):2203– 2232, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606857.

Debruyne:2009:OMS

[Deb09] Michiel Debruyne. An outlier map for support vector machine classification. Annals of Applied Statistics, 3(4):1566– 1580, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453953.

Dupuis:2023:MPE

[DET23] Debbie J. Dupuis, Sebastian Engelke, and Luca Trapin. Modeling panels of extremes. Annals of Applied Statistics, 17(1): 498–517, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Modeling-panels-of-extremes/10.1214/22-
AOAS1639.full.

Iorio:2023:BNM

[DFGY23] Maria De Iorio, Stefano Favaro, Alessandra Guglielmi, and Lifeng Ye. Bayesian nonparametric mixture modeling for temporal dynamics of gender stereotypes. Annals of Applied Statistics, 17(3):2256–2278, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Bayesian-nonparametric-
mixture-modeling-for-temporal-dynamics-of-gender-
stereotypes/10.1214/22-AOAS1717.full.

Drignei:2008:PEC

[DFN08] Dorin Drignei, Chris E. Forest, and Doug Nychka. Parameter estimation for computationally intensive nonlinear regression with an application to climate modeling. Annals of Applied Statistics, 2(4):1217–1230, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1231424207.

deGonzalez:2007:IIR

[dGC07] Amy Berrington de González and D. R. Cox. Interpretation of interaction: a review. Annals of Applied Statistics, 1(2):371–385, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1196438023.

Drylewicz:2010:MDB

[DGCT10] J. Drylewicz, J. Guedj, D. Commenges, and R. Thiébaut. Modeling the dynamics of biomarkers during primary HIV infection taking into account the uncertainty of infection date. Annals of Applied Statistics, 4(4):1847–1870, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167801.

Diaconis:2008:HMS

[DGH08] Persi Diaconis, Sharad Goel, and Susan Holmes. Horseshoes in multidimensional scaling and local kernel methods. Annals of Applied Statistics, 2(3):777–807, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1223908041.

Deng:2013:LLD

[DGL13] Wanlu Deng, Zhi Geng, and Hongzhe Li. Learning local directed acyclic graphs based on multivariate time series data. Annals of Applied Statistics, 7(3):1663–1683, September 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1380804811.

Davies:2008:RBL

[DGM^{+}08] P. L. Davies, U. Gather, M. Meise, D. Mergel, and T. Mildenberger. Residual-based localization and quantification of peaks in X-ray diffractograms. Annals of Applied Statistics, 2(3):861–886, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1223908044.

Dalal:2010:DRM

[DH10] Siddhartha R. Dalal and Bing Han. Detection of radioactive material entering national ports: a Bayesian approach to radiation portal data. Annals of Applied Statistics, 4 (3):1256–1271, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409372.

DelaCruz:2011:DDD

[DH11] Omar De la Cruz and Susan Holmes. The duality diagram in data analysis: Examples of modern applications. Annals of Applied Statistics, 5(4):2266–2277, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1324399594.

Donnat:2018:TND

[DH18] Claire Donnat and Susan Holmes. Tracking network dynamics: a survey using graph distances. Annals of Applied Statistics, 12(2):971–1012, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1532743483.

Donovan:2019:NII

[DHG19] Kevin M. Donovan, Michael G. Hudgens, and Peter B. Gilbert. Nonparametric inference for immune response thresholds of risk in vaccine studies. Annals of Applied Statistics, 13(2):1147–1165, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758441.

Dean:2018:DVT

[DHL18] Natalie E. Dean, M. Elizabeth Halloran, and Ira M. Longini. Design of vaccine trials during outbreaks with and without a delayed vaccination comparator. Annals of Applied Statistics, 12(1):330–347, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564475.

Duan:2017:PMB

[DHM^{+}17] Yuanyuan Duan, Yili Hong, William Q. Meeker, Deborah L. Stanley, and Xiaohong Gu. Photodegradation modeling based on laboratory accelerated test data and predictions under outdoor weathering for polymeric materials. Annals of Applied Statistics, 11(4):2052–2079, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1514430277.

Dray:2011:RGD

[DJ11] Stéphane Dray and Thibaut Jombart. Revisiting Guerry’s data: Introducing spatial constraints in multivariate analysis. Annals of Applied Statistics, 5(4):2278–2299, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399595.

Du:2012:CAE

[DK12] Chao Du and S. C. Kou. Correlation analysis of enzymatic reaction of a single protein molecule. Annals of Applied Statistics, 6(3):950–976, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1346418569.

Ding:2018:RTU

[DK18] Peng Ding and Luke Keele. Rank tests in unmatched clustered randomized trials applied to a study of teacher training. Annals of Applied Statistics, 12(4):2151–2174, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078040.

Dryden:2019:PNS

[DKLL19] Ian L. Dryden, Kwang-Rae Kim, Charles A. Laughton, and Huiling Le. Principal nested shape space analysis of molecular dynamics data. Annals of Applied Statistics, 13(4):2213– 2234, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910042.

Duncan:2018:SSA

[DKS18] Adam Duncan, Eric Klassen, and Anuj Srivastava. Statistical shape analysis of simplified neuronal trees. Annals of Applied Statistics, 12(3):1385–1421, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1536652959.

Dryden:2009:NES

[DKZ09] Ian L. Dryden, Alexey Koloydenko, and Diwei Zhou. Non-Euclidean statistics for covariance matrices, with applications to diffusion tensor imaging. Annals of Applied Statistics, 3(3):1102–1123, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1254773280.

Dass:2009:HMM

[DL09] Sarat C. Dass and Mingfei Li. Hierarchical mixture models for assessing fingerprint individuality. Annals of Applied Statistics, 3(4):1448–1466, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1267453947.

Davis:2011:DSA

[DL11a] Richard A. Davis and Jingchen Liu. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):52–55, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715173. See [MW11b].

Dobra:2011:CGG

[DL11b] Adrian Dobra and Alex Lenkoski. Copula Gaussian graphical models and their application to modeling functional disability data. Annals of Applied Statistics, 5(2A):969–993, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562213.

Dempsey:2020:SMR

[DLKM20] Walter Dempsey, Peng Liao, Santosh Kumar, and Susan A. Murphy. The stratified micro-randomized trial design: Sample size considerations for testing nested causal effects of time-varying treatments. Annals of Applied Statistics, 14(2): 661–684, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449320.

Ding:2018:CIS

[DLL^{+}18] Ying Ding, Ying Grace Li, Yushi Liu, Stephen J. Ruberg, and Jason C. Hsu. Confident inference for SNP effects on treatment efficacy. Annals of Applied Statistics, 12(3):1727–1748, September 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1536652972.

Dass:2014:GMM

[DLM14] Sarat C. Dass, Chae Young Lim, and Tapabrata Maiti. A generalized mixed model framework for assessing fingerprint individuality in presence of varying image quality. Annals of Applied Statistics, 8(3):1314–1340, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1414091215.

DeRuiter:2017:MMH

[DLS^{+}17] Stacy L. DeRuiter, Roland Langrock, Tomas Skirbutas, Jeremy A. Goldbogen, John Calambokidis, Ari S. Friedlaender, and Brandon L. Southall. A multivariate mixed hidden Markov model for blue whale behaviour and responses to sound exposure. Annals of Applied Statistics, 11(1):362– 392, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616885.

Dou:2010:MHO

[DLZ10] Yiping Dou, Nhu D. Le, and James V. Zidek. Modeling hourly ozone concentration fields. Annals of Applied Statistics, 4(3):1183–1213, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409369.

Deng:2016:FPE

[DLZL16] Ke Deng, Yang Li, Weiping Zhu, and Jun S. Liu. Fast parameter estimation in loss tomography for networks of general topology. Annals of Applied Statistics, 10(1):144–164, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909911.

Dobra:2018:LMS

[DM18] Adrian Dobra and Reza Mohammadi. Loglinear model selection and human mobility. Annals of Applied Statistics, 12(2): 815–845, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743478.

DAngelo:2019:LSM

[DMA19] Silvia D’Angelo, Thomas Brendan Murphy, and Marco Alfò. Latent space modelling of multidimensional networks with application to the exchange of votes in Eurovision song contest. Annals of Applied Statistics, 13(2):900–930, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758432.

Diana:2020:HDD

[DMGJ20] Alex Diana, Eleni Matechou, Jim Griffin, and Alison Johnston. A hierarchical dependent Dirichlet process prior for modelling bird migration patterns in the UK. Annals of Applied Statistics, 14(1):473–493, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1587002683.

deChaumaray:2020:MHM

[DMN20] Marie Du Roy de Chaumaray, Matthieu Marbac, and Fabien Navarro. Mixture of hidden Markov models for accelerometer data. Annals of Applied Statistics, 14(4):1834–1855, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Mixture-of-hidden-Markov-models-for-accelerometer-
data/10.1214/20-AOAS1375.full.

Dong:2023:EUG

[DMVT23] Larry Dong, Erica E. M. Moodie, Laura Villain, and Rodolphe Thiébaut. Evaluating the use of generalized dynamic weighted ordinary least squares for individualized HIV treatment strategies. Annals of Applied Statistics, 17(3):2432–2451, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Evaluating-the-
use-of-generalized-dynamic-weighted-ordinary-least-
squares/10.1214/22-AOAS1726.full.

Dette:2010:ODR

[DPHL10] Holger Dette, Andrey Pepelyshev, and Tim Holland-Letz. Optimal designs for random effect models with correlated errors with applications in population pharmacokinetics. Annals of Applied Statistics, 4(3):1430–1450, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1287409380.

DAmato:2011:MIP

[DPR11] Valeria D’Amato, Gabriella Piscopo, and Maria Russolillo. The mortality of the Italian population: Smoothing techniques on the Lee–Carter model. Annals of Applied Statistics, 5(2A):705–724, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1310562202.

Datta:2020:SEK

[DPR^{+}20] Abhirup Datta, Andrew Pita, Amrita Rao, Bhekie Sithole, Zandile Mnisi, and Stefan Baral. Size estimation of key populations in the HIV epidemic in eSwatini using incomplete and misaligned capture-recapture data. Annals of Applied Statistics, 14(3):1207–1241, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Size-estimation-of-key-populations-in-
the-HIV-epidemic-in/10.1214/20-AOAS1327.full.

Stefano:2022:PPV

[DPT22] Domenico De Stefano, Francesco Pauli, and Nicola Torelli. Preelectoral polls variability: a hierarchical Bayesian model to assess the role of house effects with application to Italian elections. Annals of Applied Statistics, 16(1):460–476, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Preelectoral-polls-variability--A-hierarchical-
Bayesian-model-to-assess/10.1214/21-AOAS1507.full.

Director:2021:PFA

[DRB21] Hannah M. Director, Adrian E. Raftery, and Cecilia M. Bitz. Probabilistic forecasting of the Arctic sea ice edge with contour modeling. Annals of Applied Statistics, 15(2):711– 726, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Probabilistic-forecasting-of-the-Arctic-
sea-ice-edge-with-contour/10.1214/20-AOAS1405.full.

Ditlevsen:2014:EPO

[DS14] Susanne Ditlevsen and Adeline Samson. Estimation in the partially observed stochastic Morris–Lecar neuronal model with particle filter and stochastic approximation methods. Annals of Applied Statistics, 8(2):674–702, June 2014. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229510.

Dinsdale:2019:MOT

[DSB19] Daniel Dinsdale and Matias Salibian-Barrera. Modelling ocean temperatures from bio-probes under preferential sampling. Annals of Applied Statistics, 13(2):713–745, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758425.

Dai:2023:DAD

[DSC^{+}23] Ben Dai, Xiaotong Shen, Lin Yee Chen, Chunlin Li, and Wei Pan. Data-adaptive discriminative feature localization with statistically guaranteed interpretation. Annals of Applied Statistics, 17(3):2019–2038, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Data-adaptive-discriminative-feature-localization-
with-statistically-guaranteed-interpretation/10.1214/
22-AOAS1705.full.

Davies:2019:MMS

[DSCS19] Tilman M. Davies, Matthew R. Schofield, Jon Cornwall, and Philip W. Sheard. Modelling multilevel spatial behaviour in binary-mark muscle fibre configurations. Annals of Applied Statistics, 13(3):1329–1347, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1571277755.

Doss:2013:FBD

[DSH^{+}13] Charles R. Doss, Marc A. Suchard, Ian Holmes, Midori Kato-Maeda, and Vladimir N. Minin. Fitting birth–death processes to panel data with applications to bacterial DNA fingerprinting. Annals of Applied Statistics, 7(4):2315–2335, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823321.

Dupuis:2019:GLO

[DT19] Debbie J. Dupuis and Luca Trapin. Ground-level ozone: Evidence of increasing serial dependence in the extremes. Annals of Applied Statistics, 13(1):34–59, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1554861640.

Dupuis:2023:MFE

[DT23] Debbie J. Dupuis and Luca Trapin. Mixed-frequency extreme value regression: Estimating the effect of mesoscale convective systems on extreme rainfall intensity. Annals of Applied Statistics, 17(2):1398–1418, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Mixed-frequency-
extreme-value-regression--Estimating-the-effect-of/
10.1214/22-AOAS1675.full.

Duan:2023:PPC

[DTL^{+}23] Rui Duan, Jiayi Tong, Lifeng Lin, Lisa Levine, Mary Sammel, Joel Stoddard, Tianjing Li, Christopher H Schmid, Haitao Chu, and Yong Chen. PALM: Patient-centered treatment ranking via large-scale multivariate network metaanalysis. Annals of Applied Statistics, 17(1):815–837, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/PALM--Patient-centered-treatment-ranking-
via-large-scale-multivariate/10.1214/22-AOAS1652.
full.

Ding:2013:HBA

[DTZP13] Jie Ding, Lorenzo Trippa, Xiaogang Zhong, and Giovanni Parmigiani. Hierarchical Bayesian analysis of somatic mutation data in cancer. Annals of Applied Statistics, 7(2): 883–903, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338472.

Dupuis:2017:EPD

[Dup17] Debbie J. Dupuis. Electricity price dependence in New York State zones: a robust detrended correlation approach. Annals of Applied Statistics, 11(1):248–273, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1491616880.

Dobra:2019:MAM

[DVA^{+}19] Adrian Dobra, Camilo Valdes, Dragana Ajdic, Bertrand Clarke, and Jennifer Clarke. Modeling association in microbial communities with clique loglinear models. Annals of Applied Statistics, 13(2):931–957, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1560758433.

Dupuis:2013:RVR

[DVF13] Debbie J. Dupuis and Maria-Pia Victoria-Feser. Robust VIF regression with application to variable selection in large data sets. Annals of Applied Statistics, 7(1):319–341, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1365527201.

Dong:2021:STS

[DW21] Tracy Qi Dong and Jon Wakefield. Space-time smoothing models for subnational measles routine immunization coverage estimation with complex survey data. Annals of Applied Statistics, 15(4):1959–1979, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Space-time-smoothing-
models-for-subnational-measles-routine-immunization-
coverage/10.1214/21-AOAS1474.full.

Diaz:2023:CDD

[DZ23] Juan D. Díaz and José R. Zubizarreta. Complex discontinuity designs using covariates: Impact of school grade retention on later life outcomes in Chile. Annals of Applied Statistics, 17(1):67–88, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Complex-discontinuity-
designs-using-covariates--Impact-of-school-grade/
10.1214/22-AOAS1616.full.

Elmasri:2020:HBM

[EFDS20] Mohamad Elmasri, Maxwell J. Farrell, T. Jonathan Davies, and David A. Stephens. A hierarchical Bayesian model for predicting ecological interactions using scaled evolutionary relationships. Annals of Applied Statistics, 14(1):221– 240, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002672.

Erosheva:2007:DDT

[EFJ07] Elena A. Erosheva, Stephen E. Fienberg, and Cyrille Joutard. Describing disability through individual-level mixture models for multivariate binary data. Annals of Applied Statistics, 1(2):502–537, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1196438029.

Efron:2008:SIW

[Efr08] Bradley Efron. Simultaneous inference: When should hypothesis testing problems be combined? Annals of Applied Statistics, 2(1):197–223, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1206367818.

Efron:2009:SMI

[Efr09] Bradley Efron. Are a set of microarrays independent of each other? Annals of Applied Statistics, 3(3):922–942, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1254773272.

Efron:2012:BIP

[Efr12] Bradley Efron. Bayesian inference and the parametric bootstrap. Annals of Applied Statistics, 6(4):1971–1997, December 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1356629067.

Erasto:2012:FCC

[EHKW12] Panu Erästö, Lasse Holmström, Atte Korhola, and Jan Weckström. Finding a consensus on credible features among several paleoclimate reconstructions. Annals of Applied Statistics, 6(4):1377–1405, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629044.

Eltzner:2018:TPC

[EHM18] Benjamin Eltzner, Stephan Huckemann, and Kanti V. Mardia. Torus principal component analysis with applications to RNA structure. Annals of Applied Statistics, 12(2):1332– 1359, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743497.

Ezzat:2019:STS

[EJD19] Ahmed Aziz Ezzat, Mikyoung Jun, and Yu Ding. Spatiotemporal short-term wind forecast: a calibrated regimeswitching method. Annals of Applied Statistics, 13(3): 1484–1510, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277761.

Embleton:2022:MSM

[EKO22] Jonathan Embleton, Marina I. Knight, and Hernando Ombao. Multiscale spectral modelling for nonstationary time series within an ordered multiple-trial experiment. Annals of Applied Statistics, 16(4):2774–2803, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Multiscale-spectral-modelling-for-nonstationary-
time-series-within-an-ordered/10.1214/22-AOAS1614.
full.

Eckley:2020:NCP

[EKW20] Idris Eckley, Claudia Kirch, and Silke Weber. A novel change-point approach for the detection of gas emission sources using remotely contained concentration data. Annals of Applied Statistics, 14(3):1258–1284, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-14/issue-3/A-
novel-change-point-approach-for-the-detection-of-
gas/10.1214/20-AOAS1345.full.

Edlefsen:2009:ELP

[ELD09] Paul T. Edlefsen, Chuanhai Liu, and Arthur P. Dempster. Estimating limits from Poisson counting data using Dempster–Shafer analysis. Annals of Applied Statistics, 3(2): 764–790, June 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1245676194.

Enki:2014:TVS

[ENF14] Doyo G. Enki, Angela Noufaily, and C. Paddy Farrington. A time-varying shared frailty model with application to infectious diseases. Annals of Applied Statistics, 8(1):430– 447, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966293.

Ertefaie:2018:IVA

[ENH^{+}18] Ashkan Ertefaie, Anh Nguyen, David J. Harding, Jeffrey D. Morenoff, and Wei Yang. Instrumental variable analysis with censored data in the presence of many weak instruments: Application to the effect of being sentenced to prison on time to employment. Annals of Applied Statistics, 12(4):2647– 2673, December 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1542078059.

Erciulescu:2021:BMR

[EOB21] Andreea L. Erciulescu, Jean D. Opsomer, and F. Jay Breidt. A bridging model to reconcile statistics based on data from multiple surveys. Annals of Applied Statistics, 15(2):1068– 1079, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/A-bridging-model-to-reconcile-statistics-
based-on-data-from/10.1214/20-AOAS1437.full.

Ertekin:2015:RPP

[ERM15]  Şeyda Ertekin, Cynthia Rudin, and Tyler H. McCormick. Re-

active point processes: a new approach to predicting power failures in underground electrical systems. Annals of Applied Statistics, 9(1):122–144, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226087.

Economou:2014:STM

[ESF14] Theodoros Economou, David B. Stephenson, and Christopher A. T. Ferro. Spatio-temporal modelling of extreme storms. Annals of Applied Statistics, 8(4):2223–2246, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001741.

Euan:2019:CBT

[ESO19] Carolina Euán, Ying Sun, and Hernando Ombao. Coherencebased time series clustering for statistical inference and visualization of brain connectivity. Annals of Applied Statistics, 13(2):990–1015, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758435.

Efron:2007:TSS

[ET07] Bradley Efron and Robert Tibshirani. On testing the significance of sets of genes. Annals of Applied Statistics, 1(1): 107–129, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143731.

Foulkes:2010:PBC

[FAL^{+}10] Andrea S. Foulkes, Livio Azzoni, Xiaohong Li, Margaret A. Johnson, Colette Smith, Karam Mounzer, and Luis J. Montaner. Prediction-based classification for longitudinal biomarkers. Annals of Applied Statistics, 4(3):1476–1497, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409382.

Farcomeni:2022:HMR

[Far22] Alessio Farcomeni. How many refugees and migrants died trying to reach Europe? Joint population size and total estimation. Annals of Applied Statistics, 16(4):2339–2351, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/How-many-refugees-and-migrants-died-trying-
to-reach-Europe/10.1214/21-AOAS1593.full.

Freulon:2023:COT

[FBH23] Paul Freulon, Jérémie Bigot, and Boris P. Hejblum. CytOpT: Optimal transport with domain adaptation for interpreting flow cytometry data. Annals of Applied Statistics, 17(2): 1086–1104, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/CytOpT--Optimal-transport-with-domain-

adaptation-for-interpreting-flow/10.1214/22-AOAS1660.
full.

Finley:2009:HSM

[FBM09] Andrew O. Finley, Sudipto Banerjee, and Ronald E. McRoberts. Hierarchical spatial models for predicting tree species assemblages across large domains. Annals of Applied Statistics, 3(3):1052–1079, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1254773278.

Friedrich:2017:NEP

[FBW^{+}17] Sarah Friedrich, Jan Beyersmann, Ursula Winterfeld, Martin Schumacher, and Arthur Allignol. Nonparametric estimation of pregnancy outcome probabilities. Annals of Applied Statistics, 11(2):840–867, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537726.

Fuglstad:2020:CCS

[FC20] Geir-Arne Fuglstad and Stefano Castruccio. Compression of climate simulations with a nonstationary global SpatioTemporal SPDE model. Annals of Applied Statistics, 14(2):542– 559, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449315.

Fanshawe:2015:LCU

[FCC15] T. R. Fanshawe, C. M. Chapman, and T. Crick. Lymphangiogenesis and carcinoma in the uterine cervix: Joint and hierarchical models for random cluster sizes and continuous outcomes. Annals of Applied Statistics, 9(4):1932–1949, December 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1453993099.

Forbes:2013:SRM

[FCGA^{+}13] Florence Forbes, Myriam Charras-Garrido, Lamiae Azizi, Senan Doyle, and David Abrial. Spatial risk mapping for rare disease with hidden Markov fields and variational EM. Annals of Applied Statistics, 7(2):1192–1216, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1372338484.

Flaxman:2019:SHR

[FCPL19] Seth Flaxman, Michael Chirico, Pau Pereira, and Charles Loeffler. Scalable high-resolution forecasting of sparse spatiotemporal events with kernel methods: a winning solution to the NIJ “Real-Time Crime Forecasting Challenge”. Annals of Applied Statistics, 13(4):2564–2585, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1574910055.

Finegold:2011:RGM

[FD11] Michael Finegold and Mathias Drton. Robust graphical modeling of gene networks using classical and alternative t-distributions. Annals of Applied Statistics, 5(2A):1057– 1080, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562216.

Ferrari:2020:IME

[FD20] Federico Ferrari and David B. Dunson. Identifying main effects and interactions among exposures using Gaussian processes. Annals of Applied Statistics, 14(4):1743–1758, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Identifying-main-effects-and-interactions-
among-exposures-using-Gaussian-processes/10.1214/
20-AOAS1363.full.

Francis:2010:MHR

[FDH10] Brian Francis, Regina Dittrich, and Reinhold Hatzinger. Modeling heterogeneity in ranked responses by nonparametric maximum likelihood: How do Europeans get their scientific knowledge? Annals of Applied Statistics, 4(4):2181– 2202, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167815.

Fay:2013:TIH

[FDKP13] Gilles Faÿ, Jacques Delabrouille, Gérard Kerkyacharian, and Dominique Picard. Testing the isotropy of high energy cosmic rays using spherical needlets. Annals of Applied Statistics, 7(2):1040–1073, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1372338478.

Fosdick:2016:CDF

[FDR16] Bailey K. Fosdick, Maria DeYoreo, and Jerome P. Reiter. Categorical data fusion using auxiliary information. Annals of Applied Statistics, 10(4):1907–1929, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606845.

Feldman:2010:LM

[Fel10] Jacob Feldman. Leo and me. Annals of Applied Statistics, 4(4):1656, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167792.

Feuerverger:2008:RSA

[Feu08a] Andrey Feuerverger. Rejoinder of: “Statistical analysis of an archeological find”. Annals of Applied Statistics, 2(1):99– 112, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367814. See [Feu08b].

Feuerverger:2008:SAA

[Feu08b] Andrey Feuerverger. Statistical analysis of an archeological find. Annals of Applied Statistics, 2(1):3–54, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1206367805. See discussion [Sti08, Fuc08, Ben08, Bir08, HW08, Ing08, MV08, Kad08] and rejoinder [Feu08a].

Feuerverger:2009:DBD

[Feu09] Andrey Feuerverger. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1282–1284, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453937. See [SR09b].

Feuerverger:2013:TND

[Feu13] Andrey Feuerverger. The tomb next door: an update to “statistical analysis of an archeological find”. Annals of Applied Statistics, 7(4):2081–2105, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1387823311.

Fan:2014:FRA

[FFJJ14] Yingying Fan, Natasha Foutz, Gareth M. James, and Wolfgang Jank. Functional response additive model estimation with online virtual stock markets. Annals of Applied Statistics, 8(4):2435–2460, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1419001750.

Fattorini:2021:TPS

[FFM^{+}21] Lorenzo Fattorini, Sara Franceschi, Marzia Marcheselli, Caterina Pisani, and Luca Pratelli. Two-phase sampling strategies for design-based mapping of continuous spatial populations in environmental surveys. Annals of Applied Statistics, 15(1):287–303, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Two-phase-sampling-
strategies-for-design-based-mapping-of-continuous/
10.1214/20-AOAS1392.full.

Ferkingstad:2008:UEB

[FFR^{+}08] Egil Ferkingstad, Arnoldo Frigessi, Håvard Rue, Gudmar Thorleifsson, and Augustine Kong. Unsupervised empirical Bayesian multiple testing with external covariates. Annals of Applied Statistics, 2(2):714–735, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1215118535.

Fan:2009:NEA

[FFW09] Jianqing Fan, Yang Feng, and Yichao Wu. Network exploration via the adaptive LASSO and SCAD penalties. Annals of Applied Statistics, 3(2):521–541, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676184.

Fong:2009:BIM

[FGA09] Youyi Fong, Peter Guttorp, and Janis Abkowitz. Bayesian inference and model choice in a hidden stochastic twocompartment model of hematopoietic stem cell fate decisions. Annals of Applied Statistics, 3(4):1695–1709, December 2009.

CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1267453960.

Feller:2016:CWV

[FGMP16] Avi Feller, Todd Grindal, Luke Miratrix, and Lindsay C. Page. Compared to what? Variation in the impacts of early childhood education by alternative care type. Annals of Applied Statistics, 10(3):1245–1285, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1475069607.

Fuentes:2008:SSS

[FGS08] Montserrat Fuentes, Peter Guttorp, and Michael L. Stein. Special section on statistics in the atmospheric sciences. Annals of Applied Statistics, 2(4):1143–1147, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1231424203.

Fu:2010:SIT

[FGS^{+}10] Audrey Qiuyan Fu, Diane P. Genereux, Reinhard Stöger, Charles D. Laird, and Matthew Stephens. Statistical inference of transmission fidelity of DNA methylation patterns over somatic cell divisions in mammals. Annals of Applied Statistics, 4(2):871–892, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1280842144.

Feng:2009:ILR

[FH09] Xingdong Feng and Xuming He. Inference on low-rank data matrices with applications to microarray data. Annals of Applied Statistics, 3(4):1634–1654, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453957.

Fithian:2013:FSE

[FH13] William Fithian and Trevor Hastie. Finite-sample equivalence in statistical models for presence-only data. Annals of Applied Statistics, 7(4):1917–1939, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1387823304.

Fosdick:2014:SFA

[FH14] Bailey K. Fosdick and Peter D. Hoff. Separable factor analysis with applications to mortality data. Annals of Applied Statistics, 8(1):120–147, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966281.

Fogarty:2019:ESA

[FH19] Colin B. Fogarty and Raiden B. Hasegawa. Extended sensitivity analysis for heterogeneous unmeasured confounding with an application to sibling studies of returns to education. Annals of Applied Statistics, 13(2):767–796, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758427.

Friedman:2007:PCO

[FHHT07] Jerome Friedman, Trevor Hastie, Holger Höfling, and Robert Tibshirani. Pathwise coordinate optimization. Annals of Applied Statistics, 1(2):302–332, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1196438020.

Fong:2018:CBP

[FHI18] Christian Fong, Chad Hazlett, and Kosuke Imai. Covariate balancing propensity score for a continuous treatment: Application to the efficacy of political advertisements. Annals of Applied Statistics, 12(1):156–177, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1520564468.

Fox:2014:JMM

[FHSJ14] Emily B. Fox, Michael C. Hughes, Erik B. Sudderth, and Michael I. Jordan. Joint modeling of multiple time series via the beta process with application to motion capture segmentation. Annals of Applied Statistics, 8(3):1281–1313, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091214.

Fienberg:2007:ESF

[Fie07] Stephen E. Fienberg. Editorial: Statistics and forensic science. Annals of Applied Statistics, 1(2):285–286, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1196438018.

Fienberg:2008:ESL

[Fie08] Stephen E. Fienberg. Editorial: Statistics and “The lost tomb of Jesus”. Annals of Applied Statistics, 2(1):1–2, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1206367804.

Fienberg:2010:IPMa

[Fie10a] Stephen E. Fienberg. Introduction to papers on the modeling and analysis of network data. Annals of Applied Statistics, 4 (1):1–4, March 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1273584444.

Fienberg:2010:IPMb

[Fie10b] Stephen E. Fienberg. Introduction to papers on the modeling and analysis of network data — II. Annals of Applied Statistics, 4(2):533–534, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1280842129.

Feng:2021:ECL

[FIM^{+}21] Jean Feng, William S. DeWitt III, Aaron McKenna, Noah Simon, Amy D. Willis, and Frederick A. Matsen IV. Estimation of cell lineage trees by maximum-likelihood phylogenetics. Annals of Applied Statistics, 15(1):343–362, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Estimation-of-cell-lineage-trees-by-maximum-
likelihood-phylogenetics/10.1214/20-AOAS1400.full.

Finazzi:2013:GMP

[Fin13] Francesco Finazzi. Geostatistical modeling in the presence of interaction between the measuring instruments, with an application to the estimation of spatial market potentials. Annals of Applied Statistics, 7(1):81–101, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1365527191.

Fiorentino:2008:LE

[Fio08] John E. Fiorentino. Letter to the Editor. Annals of Applied Statistics, 2(1):430–431, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1206367828.

Fotopoulos:2010:EAD

[FJK10] Stergios B. Fotopoulos, Venkata K. Jandhyala, and Elena Khapalova. Exact asymptotic distribution of change-point mle for change in the mean of Gaussian sequences. Annals of Applied Statistics, 4(2):1081–1104, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1280842153.

Feldman:2022:BDS

[FK22] Joseph Feldman and Daniel R. Kowal. Bayesian data synthesis and the utility-risk trade-off for mixed epidemiological data. Annals of Applied Statistics, 16(4):2577–2602, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Bayesian-data-synthesis-and-the-utility-
risk-trade-off-for/10.1214/22-AOAS1604.full.

Freue:2019:REN

[FKSBS19] Gabriela V. Cohen Freue, David Kepplinger, Matías Salibián-Barrera, and Ezequiel Smucler. Robust elastic net estimators for variable selection and identification of proteomic biomarkers. Annals of Applied Statistics, 13(4):2065–2090, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1574910036.

Fulton:2015:MME

[FLHA15] Kara A. Fulton, Danping Liu, Denise L. Haynie, and Paul S. Albert. Mixed model and estimating equation approaches for zero inflation in clustered binary response data with application to a dating violence study. Annals of Applied Statistics, 9(1):275–299, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1430226093.

Fishkind:2015:VNS

[FLP^{+}15] D. E. Fishkind, V. Lyzinski, H. Pao, L. Chen, and C. E. Priebe. Vertex nomination schemes for membership prediction. Annals of Applied Statistics, 9(3):1510–1532, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1446488749.

Franzolini:2023:MSM

[FLP23] Beatrice Franzolini, Antonio Lijoi, and Igor Prünster. Model selection for maternal hypertensive disorders with symmetric hierarchical Dirichlet processes. Annals of Applied Statistics, 17(1):313–332, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Model-selection-
for-maternal-hypertensive-disorders-with-symmetric-
hierarchical-Dirichlet/10.1214/22-AOAS1628.full.

Fan:2023:DSA

[FLRZ23] Yimei Fan, Yuan Liao, Ilya O. Ryzhov, and Kunpeng Zhang. A dynamic screening algorithm for hierarchical binary marketing data. Annals of Applied Statistics, 17(3):2326–2344, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/A-dynamic-screening-algorithm-for-hierarchical-
binary-marketing-data/10.1214/22-AOAS1720.full.

Fan:2016:SAS

[FLS16] Ailin Fan, Wenbin Lu, and Rui Song. Sequential advantage selection for optimal treatment regime. Annals of Applied Statistics, 10(1):32–53, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1458909906.

Fan:2017:EBA

[FM17] Zhou Fan and Lester Mackey. Empirical Bayesian analysis of simultaneous changepoints in multiple data sequences. Annals of Applied Statistics, 11(4):2200–2221, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1514430283.

Franks:2018:RCP

[FMA18] Alexander M. Franks, Florian Markowetz, and Edoardo M. Airoldi. Refining cellular pathway models using an ensemble of heterogeneous data sources. Annals of Applied Statistics, 12(3):1361–1384, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652958.

Filippone:2012:PPN

[FMB^{+}12] M. Filippone, A. F. Marquand, C. R. V. Blain, S. C. R. Williams, J. Mourão-Miranda, and M. Girolami. Probabilistic prediction of neurological disorders with a statistical assessment of neuroimaging data modalities. Annals of Applied Statistics, 6(4):1883–1905, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629064.

Franks:2015:CSS

[FMBG15] Alexander Franks, Andrew Miller, Luke Bornn, and Kirk Goldsberry. Characterizing the spatial structure of defensive skill in professional basketball. Annals of Applied Statistics, 9(1):94–121, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1430226086.

Fouskakis:2009:BVS

[FND09] D. Fouskakis, I. Ntzoufras, and D. Draper. Bayesian variable selection using cost-adjusted BIC, with application to costeffective measurement of quality of health care. Annals of Applied Statistics, 3(2):663–690, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676190.

Fiecas:2011:GSE

[FO11] Mark Fiecas and Hernando Ombao. The generalized shrinkage estimator for the analysis of functional connectivity of brain signals. Annals of Applied Statistics, 5(2A):1102– 1125, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562218.

Friedman:2008:PLR

[FP08] Jerome H. Friedman and Bogdan E. Popescu. Predictive learning via rule ensembles. Annals of Applied Statistics, 2(3):916–954, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1223908046.

Fisher:2020:OAA

[FPC20a] Jared D. Fisher, Davide Pettenuzzo, and Carlos M. Carvalho. Optimal asset allocation with multivariate Bayesian dynamic linear models. Annals of Applied Statistics, 14(1): 299–338, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002676.

Fisher:2020:MEC

[FPC20b] Jared D. Fisher, David W. Puelz, and Carlos M. Carvalho. Monotonic effects of characteristics on returns. Annals of Applied Statistics, 14(4):1622–1650, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Monotonic-effects-of-characteristics-on-
returns/10.1214/20-AOAS1351.full.

Fan:2010:BMA

[FPL10] Xiaodan Fan, Saumyadipta Pyne, and Jun S. Liu. Bayesian meta-analysis for identifying periodically expressed genes in fission yeast cell cycle. Annals of Applied Statistics, 4(2): 988–1013, June 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1280842149.

Fan:2018:MRM

[FPLM18] Minjie Fan, Debashis Paul, Thomas C. M. Lee, and Tomoko Matsuo. A multi-resolution model for non-Gaussian random fields on a sphere with application to ionospheric electrostatic potentials. Annals of Applied Statistics, 12(1):459– 489, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564480.

Fu:2013:BCR

[FRBT13] Audrey Qiuyan Fu, Steven Russell, Sarah J. Bray, and Simon Tavaré. Bayesian clustering of replicated time-course gene expression data with weak signals. Annals of Applied Statistics, 7(3):1334–1361, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804798.

Freedman:2008:RAE

[Fre08] David A. Freedman. On regression adjustments in experiments with several treatments. Annals of Applied Statistics, 2(1):176–196, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1206367817. See com-

ments [Lin13].

Friedman:2010:RL

[Fri10] Jerome H. Friedman. Remembering Leo. Annals of Applied Statistics, 4(4):1649–1651, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167790.

Fuentes:2008:STM

[FRL08] Montserrat Fuentes, Brian Reich, and Gyuwon Lee. Spatial– temporal mesoscale modeling of rainfall intensity using gage and radar data. Annals of Applied Statistics, 2(4):1148–1169, December 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1231424204.

Fraser:2018:WSM

[FRL18] D. A. S. Fraser, N. Reid, and Wei Lin. When should modes of inference disagree? some simple but challenging examples. Annals of Applied Statistics, 12(2):750–770, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743475.

Farcomeni:2013:HBR

[FS13a] Alessio Farcomeni and Daria Scacciatelli. Heterogeneity and behavioral response in continuous time capture–recapture, with application to street cannabis use in Italy. Annals of Applied Statistics, 7(4):2293–2314, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1387823320.

French:2013:STE

[FS13b] Joshua P. French and Stephan R. Sain. Spatio-temporal exceedance locations and confidence regions. Annals of Applied Statistics, 7(3):1421–1449, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804801.

Fogarty:2014:ETF

[FS14] Colin B. Fogarty and Dylan S. Small. Equivalence testing for functional data with an application to comparing pulmonary function devices. Annals of Applied Statistics, 8(4):2002– 2026, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001733.

Fox:2016:SIB

[FSG16] Eric Warren Fox, Frederic Paik Schoenberg, and Joshua Seth Gordon. Spatially inhomogeneous background rate estimators and uncertainty quantification for nonparametric Hawkes point process models of earthquake occurrences. Annals of Applied Statistics, 10(3):1725–1756, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1475069625.

Fox:2011:SHH

[FSJW11] Emily B. Fox, Erik B. Sudderth, Michael I. Jordan, and Alan S. Willsky. A sticky HDP-HMM with application to speaker diarization. Annals of Applied Statistics, 5(2A): 1020–1056, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562215.

Fop:2017:VSL

[FSM17] Michael Fop, Keith M. Smart, and Thomas Brendan Murphy. Variable selection for latent class analysis with application to low back pain diagnosis. Annals of Applied Statistics, 11(4):2080–2110, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1514430278.

Feng:2019:SAD

[FSM^{+}19] Jean Feng, David A. Shaw, Vladimir N. Minin, Noah Simon, and Frederick A. Matsen IV. Survival analysis of DNA mutation motifs with penalized proportional hazards. Annals of Applied Statistics, 13(2):1268–1294, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1560758446.

Fruhwirth-Schnatter:2018:APC

[FSPWWE18] Sylvia Frühwirth-Schnatter, Stefan Pittner, Andrea Weber, and Rudolf Winter-Ebmer. Analysing plant closure effects using time-varying mixture-of-experts Markov chain clustering. Annals of Applied Statistics, 12(3):1796–1830, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1536652975.

Fienberg:2018:IWS

[FT18] Stephen E. Fienberg and Judith M. Tanur. The interlocking world of surveys and experiments. Annals of Applied Statistics, 12(2):1157–1179, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1532743489.

Fossum:2021:LES

[FTE^{+}21] Trygve Olav Fossum, Cédric Travelletti, Jo Eidsvik, David Ginsbourger, and Kanna Rajan. Learning excursion sets of vector-valued Gaussian random fields for autonomous ocean sampling. Annals of Applied Statistics, 15(2):597– 618, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Learning-excursion-sets-of-vector-valued-
Gaussian-random-fields-for/10.1214/21-AOAS1451.full.

Fuchs:2008:DSA

[Fuc08] Camil Fuchs. Discussion of: “Statistical analysis of an archeological find”. Annals of Applied Statistics, 2(1):57– 65, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367807. See [Feu08b].

Fukuyama:2019:AGM

[Fuk19] Julia Fukuyama. Adaptive gPCA: a method for structured dimensionality reduction with applications to microbiome data. Annals of Applied Statistics, 13(2):1043–1067, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758437.

Franck:2021:PCC

[FW21] Christopher T. Franck and Christopher E. Wilson. Predicting competitions by combining conditional logistic regression and subjective Bayes: an Academy Awards case study. Annals of Applied Statistics, 15(4):2083–2100, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Predicting-competitions-by-combining-conditional-
logistic-regression-and-subjective-Bayes/10.1214/
21-AOAS1464.full.

Feng:2011:GGA

[FWGS11] Zeny Feng, William W. L. Wong, Xin Gao, and Flavio Schenkel. Generalized genetic association study with samples of related individuals. Annals of Applied Statistics, 5(3):2109–2130, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514297.

Finkenstadt:2013:QIE

[FWK^{+}13] Bärbel Finkenstädt, Dan J. Woodcock, Michal Komorowski, Claire V. Harper, Julian R. E. Davis, Mike R. H. White, and David A. Rand. Quantifying intrinsic and extrinsic noise in gene transcription using the linear noise approximation: An application to single cell data. Annals of Applied Statistics, 7(4):1960–1982, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1387823306.

Friedman:2010:LB

[FY10] Jerry Friedman and Bin Yu. Leo Breiman (1929–2005). Annals of Applied Statistics, 4(4):??, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167784.

Fearnhead:2015:ERR

[FYB^{+}15] Paul Fearnhead, Shoukai Yu, Patrick Biggs, Barbara Holland, and Nigel French. Estimating the relative rate of recombination to mutation in bacteria from single-locus variants using composite likelihood methods. Annals of Applied Statistics, 9(1):200–224, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226090.

Fisher:2022:DMC

[FZCV22] Thomas J. Fisher, Jing Zhang, Stephen P. Colegate, and Michael J. Vanni. Detecting and modeling changes in a time series of proportions. Annals of Applied Statistics, 16(1):477– 494, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Detecting-and-modeling-changes-in-a-time-
series-of-proportions/10.1214/21-AOAS1509.full.

Fushing:2008:SSB

[FZSI^{+}08] Hsieh Fushing, Li Zhu, David I. Shapiro-Ilan, James F. Campbell, and Edwin E. Lewis. State-space based mass event-history model I: Many decision-making agents with one target. Annals of Applied Statistics, 2(4):1503–1522, December 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1231424220.

Fischer:2017:FTS

[FZZW17] Heidi J. Fischer, Qunfang Zhang, Yifang Zhu, and Robert E. Weiss. Functional time series models for ultrafine particle distributions. Annals of Applied Statistics, 11(1):297– 319, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616882.

Gautier:2011:HBE

[Gau11] Eric Gautier. Hierarchical Bayesian estimation of inequality measures with nonrectangular censored survey data with an application to wealth distribution of French households. Annals of Applied Statistics, 5(2B):1632–1656, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562736.

Gu:2016:PPG

[GB16] Mengyang Gu and James O. Berger. Parallel partial Gaussian process emulation for computer models with massive output. Annals of Applied Statistics, 10(3):1317–1347, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1475069609.

Glynn:2021:IPC

[GBC21] Chris Glynn, Thomas H. Byrne, and Dennis P. Culhane. Inflection points in community-level homeless rates. Annals of Applied Statistics, 15(2):1037–1053, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Inflection-points-in-community-level-homeless-
rates/10.1214/20-AOAS1414.full.

Gramacy:2015:CLC

[GBH^{+}15] Robert B. Gramacy, Derek Bingham, James Paul Holloway, Michael J. Grosskopf, Carolyn C. Kuranz, Erica Rutter, Matt Trantham, and R. Paul Drake. Calibrating a large computer experiment simulating radiative shock hydrodynamics. Annals of Applied Statistics, 9(3):1141–1168, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1446488734.

Godichon-Baggioni:2020:MCA

[GBMRR20] Antoine Godichon-Baggioni, Cathy Maugis-Rabusseau, and Andrea Rau. Multiview cluster aggregation and splitting, with an application to multiomic breast cancer data. Annals of Applied Statistics, 14(2):752–767, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1593449324.

Guo:2022:SSA

[GBNS22] Xiaoyang Guo, Aditi Basu Bal, Tom Needham, and Anuj Srivastava. Statistical shape analysis of brain arterial networks (BAN). Annals of Applied Statistics, 16(2):1130– 1150, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Statistical-shape-analysis-of-brain-arterial-
networks-BAN/10.1214/21-AOAS1536.full.

Gagnon-Bartsch:2019:CPT

[GBST19] Johann Gagnon-Bartsch and Yotam Shem-Tov. The classification permutation test: a flexible approach to testing for covariate imbalance in observational studies. Annals of Applied Statistics, 13(3):1464–1483, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1571277760.

Gelman:2008:SDM

[GC08] Andrew Gelman and Cexun Jeffrey Cai. Should the Democrats move to the left on economic policy? Annals of Applied Statistics, 2(2):536–549, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1215118527. See cor-

rection [Gel13].

Goldsmith:2011:NTF

[GCC^{+}11] Jeff Goldsmith, Brian Caffo, Ciprian Crainiceanu, Daniel Reich, Yong Du, and Craig Hendrix. Nonlinear tube-fitting for the analysis of anatomical and functional structures. Annals of Applied Statistics, 5(1):337–363, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1300715193.

Guo:2015:EHG

[GCL^{+}15] Jian Guo, Jie Cheng, Elizaveta Levina, George Michailidis, and Ji Zhu. Estimating heterogeneous graphical models for discrete data with an application to roll call voting. Annals of Applied Statistics, 9(2):821–848, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1437397113.

George:2016:ASM

[GDG^{+}16] Brandon George, Jr., Thomas Denney, Himanshu Gupta, Louis Dell’Italia, and Inmaculada Aban. Applying a spatiotemporal model for longitudinal cardiac imaging data. Annals of Applied Statistics, 10(1):527–548, March 2016.

CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909926.

Gao:2020:STM

[GDJR20] Zhenguo Gao, Pang Du, Ran Jin, and John L. Robertson. Surface temperature monitoring in liver procurement via functional variance change-point analysis. Annals of Applied Statistics, 14(1):143–159, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1587002668.

Grabski:2023:BCM

[GDTP23] Isabella N. Grabski, Roberta De Vito, Lorenzo Trippa, and Giovanni Parmigiani. Bayesian combinatorial MultiStudy factor analysis. Annals of Applied Statistics, 17 (3):2212–2235, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-17/issue-3/Bayesian-combinatorial-MultiStudy-
factor-analysis/10.1214/22-AOAS1715.full.

Gruhl:2013:SAM

[GEC13] Jonathan Gruhl, Elena A. Erosheva, and Paul K. Crane. A semiparametric approach to mixed outcome latent variable models: Estimating the association between cognition and regional brain volumes. Annals of Applied Statistics, 7(4):2361– 2383, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823323.

Gibbs:2022:CET

[GEF22] Connor P. Gibbs, Ryan Elmore, and Bailey K. Fosdick. The causal effect of a timeout at stopping an opposing run in the NBA. Annals of Applied Statistics, 16(3):1359–1379, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/The-causal-effect-of-a-timeout-at-stopping-
an-opposing/10.1214/21-AOAS1545.full.

Gelman:2013:CSD

[Gel13] Andrew Gelman. Correction: Should the Democrats move to the left on economic policy? Annals of Applied Statistics, 7 (2):1248, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338488. See [GC08].

Genovese:2009:DBD

[Gen09] Christopher R. Genovese. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1299–1302, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453940. See [SR09b].

Gervini:2009:DHO

[Ger09] Daniel Gervini. Detecting and handling outlying trajectories in irregularly sampled functional datasets. Annals of Applied Statistics, 3(4):1758–1775, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453963.

Glynn:2019:DHU

[GF19] Chris Glynn and Emily B. Fox. Dynamics of homelessness in urban America. Annals of Applied Statistics, 13(1):573– 605, March 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1554861661.

Gretton:2009:DBD

[GFS09] Arthur Gretton, Kenji Fukumizu, and Bharath K. Sriperumbudur. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1285–1294, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1267453938. See [SR09b].

Genovese:2009:IDE

[GFW^{+}09] Christopher Genovese, Peter Freeman, Larry Wasserman, Robert Nichol, and Christopher Miller. Inference for the dark energy equation of state using Type IA supernova data. Annals of Applied Statistics, 3(1):144–178, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888366.

Gu:2019:DCP

[GG19] Chenyang Gu and Roee Gutman. Development of a common patient assessment scale across the continuum of care: a nested multiple imputation approach. Annals of Applied Statistics, 13(1):466–491, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861657.

Gao:2020:MED

[GGCM20] Yuanjun Gao, Jack Goetz, Matthew Connelly, and Rahul Mazumder. Mining events with declassified diplomatic documents. Annals of Applied Statistics, 14(4):1699–1723, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Mining-events-with-declassified-diplomatic-
documents/10.1214/20-AOAS1344.full.

Griffin:2018:SBF

[GGFG^{+}18] Maryclare Griffin, Krista J. Gile, Karen I. Fredricksen-Goldsen, Mark S. Handcock, and Elena A. Erosheva. A simulation-based framework for assessing the feasibility of respondent-driven sampling for estimating characteristics in populations of lesbian, gay and bisexual older adults. Annals of Applied Statistics, 12(4):2252–2278, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078044.

Galves:2012:CTS

[GGG^{+}12] Antonio Galves, Charlotte Galves, Jesús E. García, Nancy L. Garcia, and Florencia Leonardi. Context tree selection and linguistic rhythm retrieval from written texts. Annals of Applied Statistics, 6(1):186–209, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1331043393.

Gelsinger:2023:LGC

[GGMG23] Megan L. Gelsinger, Maryclare Griffin, David Matteson, and Joseph Guinness. Log-Gaussian Cox process modeling of large spatial lightning data using spectral and Laplace approximations. Annals of Applied Statistics, 17(3):2078–2094, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Log-Gaussian-
Cox-process-modeling-of-large-spatial-lightning-data/
10.1214/22-AOAS1708.full.

Gaetan:2016:CCS

[GGPM16] Carlo Gaetan, Paolo Girardi, Roberto Pastres, and Antoine Mangin. Clustering chlorophyll-a satellite data using quantiles. Annals of Applied Statistics, 10(2):964–988, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1469199901.

Gordon:2007:CMN

[GGQY07] Alexander Gordon, Galina Glazko, Xing Qiu, and Andrei Yakovlev. Control of the mean number of false discoveries, Bonferroni and stability of multiple testing. Annals of Applied Statistics, 1(1):179–190, June 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1183143734.

George:2012:OPU

[GH12] Edward I. George and Sam K. Hui. Optimal pricing using online auction experiments: A Pólya tree approach. Annals of Applied Statistics, 6(1):55–82, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1331043388.

Gong:2022:ATD

[GH22] Yan Gong and Raphaël Huser. Asymmetric tail dependence modeling, with application to cryptocurrency market data. Annals of Applied Statistics, 16(3):1822–1847, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Asymmetric-tail-dependence-modeling-with-
application-to-cryptocurrency-market-data/10.1214/
21-AOAS1568.full.

Gaydos:2013:VGC

[GHK^{+}13] Travis L. Gaydos, Nancy E. Heckman, Mark Kirkpatrick, J. R. Stinchcombe, Johanna Schmitt, Joel Kingsolver, and J. S. Marron. Visualizing genetic constraints. Annals of Applied Statistics, 7(2):860–882, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338471.

Ghosh:2010:IBA

[Gho10] Samiran Ghosh. An imputation-based approach for parameter estimation in the presence of ambiguous censoring with application in industrial supply chain. Annals of Applied Statistics, 4(4):1976–1999, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1294167806.

Guerrero:2023:CCL

[GHO23] Matheus B. Guerrero, Raphaël Huser, and Hernando Ombao. Conex-Connect: Learning patterns in extremal brain connectivity from MultiChannel EEG data. Annals of Applied Statistics, 17(1):178–198, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/ConexConnect-
-Learning-patterns-in-extremal-brain-connectivity-
from-MultiChannel/10.1214/22-AOAS1621.full.

Gill:2013:DEH

[Gil13] Jeff Gill. Discussion of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1881– 1887, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823299. See [SR09b].

Gillespie:2017:ENC

[Gil17] Colin S. Gillespie. Estimating the number of casualties in the American Indian War: a Bayesian analysis using the power law distribution. Annals of Applied Statistics, 11(4):2357– 2374, December 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1514430289.

Gelman:2008:WID

[GJPS08] Andrew Gelman, Aleks Jakulin, Maria Grazia Pittau, and Yu-Sung Su. A weakly informative default prior distribution for logistic and other regression models. Annals of Applied Statistics, 2(4):1360–1383, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1231424214.

Ghosh:2023:BNL

[GKM23] Satyajit Ghosh, Kshitij Khare, and George Michailidis. The Bayesian nested lasso for mixed frequency regression models. Annals of Applied Statistics, 17(3):2279–2304, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/The-Bayesian-nested-lasso-for-mixed-frequency-
regression-models/10.1214/22-AOAS1718.full.

Gladish:2016:STA

[GKP^{+}16] Daniel W. Gladish, Petra M. Kuhnert, Daniel E. Pagendam, Christopher K. Wikle, Rebecca Bartley, Ross D. Searle, Robin J. Ellis, Cameron Dougall, Ryan D. R. Turner, Stephen E. Lewis, Zoë T. Bainbridge, and Jon E. Brodie. Spatio-temporal assimilation of modelled catchment loads with monitoring data in the Great Barrier Reef. Annals of Applied Statistics, 10(3):1590–1618, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1475069620.

Gromenko:2017:ECT

[GKS17] Oleksandr Gromenko, Piotr Kokoszka, and Jan Sojka. Evaluation of the cooling trend in the ionosphere using functional regression with incomplete curves. Annals of Applied Statistics, 11(2):898–918, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537728.

Gromenko:2012:ETS

[GKZS12] Oleksandr Gromenko, Piotr Kokoszka, Lie Zhu, and Jan Sojka. Estimation and testing for spatially indexed curves with application to ionospheric and magnetic field trends. Annals of Applied Statistics, 6(2):669–696, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1339419612.

Gustafson:2008:BMR

[GL08] Paul Gustafson and Geneviève Lefebvre. Bayesian multinomial regression with class-specific predictor selection. Annals of Applied Statistics, 2(4):1478–1502, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1231424219.

Golchi:2018:FCB

[GL18] Shirin Golchi and Richard Lockhart. A frequency-calibrated Bayesian search for new particles. Annals of Applied Statistics, 12(3):1939–1968, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652981.

Ginestet:2017:HTN

[GLB^{+}17] Cedric E. Ginestet, Jun Li, Prakash Balachandran, Steven Rosenberg, and Eric D. Kolaczyk. Hypothesis testing for network data in functional neuroimaging. Annals of Applied Statistics, 11(2):725–750, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537721.

Gormley:2008:MEM

[GM08] Isobel Claire Gormley and Thomas Brendan Murphy. A mixture of experts model for rank data with applications in election studies. Annals of Applied Statistics, 2(4):1452–1477, December 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1231424218.

Green:2009:SIF

[GM09] Peter J. Green and Julia Mortera. Sensitivity of inferences in forensic genetics to assumptions about founding genes. Annals of Applied Statistics, 3(2):731–763, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1245676193.

Guseo:2015:MCB

[GM15] Renato Guseo and Cinzia Mortarino. Modeling competition between two pharmaceutical drugs using innovation diffusion models. Annals of Applied Statistics, 9(4):2073–2089, December 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1453994192.

Garcia:2016:CRE

[GM16] Tanya P. Garcia and Samuel Müller. Cox regression with exclusion frequency-based weights to identify neuroimaging markers relevant to Huntington’s disease onset. Annals of Applied Statistics, 10(4):2130–2156, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606854.

Gu:2015:STE

[GMB15] Xiangdong Gu, Yunsheng Ma, and Raji Balasubramanian. Semiparametric time to event models in the presence of errorprone, self-reported outcomes — with application to the Women’s Health Initiative. Annals of Applied Statistics, 9(2): 714–730, June 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1437397108.

Ge:2014:AMS

[GMLB^{+}14] Tian Ge, Nicole Müller-Lenke, Kerstin Bendfeldt, Thomas E. Nichols, and Timothy D. Johnson. Analysis of multiple sclerosis lesions via spatially varying coefficients. Annals of Applied Statistics, 8(2):1095–1118, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229527.

Griffin:2008:APS

[GMM08] Beth Ann Griffin, Daniel F. McCaffrey, and Andrew R. Morral. An application of principal stratification to control for institutionalization at follow-up in studies of substance abuse treatment programs. Annals of Applied Statistics, 2(3):1034–1055, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1223908051.

Garcia:2017:RME

[GMMW17] Tanya P. Garcia, Yanyuan Ma, Karen Marder, and Yuanjia Wang. Robust mixed effects model for clustered failure time data: Application to Huntington’s disease event measures. Annals of Applied Statistics, 11(2):1085–1116, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1500537736.

Gabitto:2021:BNA

[GMNP^{+}21] Mariano I. Gabitto, Herve Marie-Nelly, Ari Pakman, Andras Pataki, Xavier Darzacq, and Michael I. Jordan. A Bayesian nonparametric approach to superresolution single-molecule localization. Annals of Applied Statistics, 15(4):1742–1766, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-4/A-
Bayesian-nonparametric-approach-to-super-resolution-
single-molecule-localization/10.1214/21-AOAS1441.
full.

Guan:2022:GWP

[GNCS22] Tianyu Guan, Robert Nguyen, Jiguo Cao, and Tim Swartz. In-game win probabilities for the National Rugby League. Annals of Applied Statistics, 16(1):349–367, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/In-game-win-
probabilities-for-the-National-Rugby-League/10.1214/
21-AOAS1514.full.

Gneiting:2012:SSY

[Gne12] Tilmann Gneiting. Section on the Special Year for Mathematics of Planet Earth (MPE 2013). Annals of Applied Statistics, 6(4):1349–1351, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1356629042.

Goerg:2011:LRV

[Goe11] Georg M. Goerg. Lambert W random variables — a new family of generalized skewed distributions with applications to risk estimation. Annals of Applied Statistics, 5(3):2197–2230, September 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1318514301.

Goerg:2014:ULW

[Goe14] Georg M. Goerg. Usage of the Lambert W function in statistics. Annals of Applied Statistics, 8(4):2567, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1419001755.

Guan:2022:PHC

[GPBT22] Zoe Guan, Giovanni Parmigiani, Danielle Braun, and Lorenzo Trippa. Prediction of hereditary cancers using neural networks. Annals of Applied Statistics, 16(1):495– 520, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Prediction-of-hereditary-cancers-using-
neural-networks/10.1214/21-AOAS1510.full.

Gauran:2022:BLF

[GPR^{+}22] Iris Ivy M. Gauran, Junyong Park, Ilia Rattsev, Thomas A. Peterson, Maricel G. Kann, and DoHwan Park. Bayesian local false discovery rate for sparse count data with application to the discovery of hotspots in protein domains. Annals of Applied Statistics, 16(3):1459–1475, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Bayesian-local-false-discovery-rate-for-
sparse-count-data-with/10.1214/21-AOAS1551.full.

Grilli:2016:ETP

[GPRR16] Leonardo Grilli, Fulvia Pennoni, Carla Rampichini, and Isabella Romeo. Exploiting TIMSS and PIRLS combined data: Multivariate multilevel modelling of student achievement. Annals of Applied Statistics, 10(4):2405–2426, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606865.

Gao:2017:VSC

[GPRZ17] Jiti Gao, Bin Peng, Zhao Ren, and Xiaohui Zhang. Variable selection for a categorical varying-coefficient model with identifications for determinants of body mass index. Annals of Applied Statistics, 11(2):1117–1145, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1500537737.

Gu:2022:PAM

[GPZ^{+}22] Yu Gu, John S. Preisser, Donglin Zeng, Poojan Shrestha, Molina Shah, Miguel A. Simancas-Pallares, Jeannie Ginnis, and Kimon Divaris. Partitioning around medoids clustering and random forest classification for GIS-informed imputation of fluoride concentration data. Annals of Applied Statistics, 16(1):551–572, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Partitioning-
around-medoids-clustering-and-random-forest-classification-
for-GIS/10.1214/21-AOAS1516.full.

Greiner:2010:EPR

[GQ10] D. James Greiner and Kevin M. Quinn. Exit polling and racial bloc voting: Combining individual-level and r × c ecological data. Annals of Applied Statistics, 4(4):1774–1796, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167798.

Gutierrez:2011:MBS

[GQ11] Luis Gutiérrez and Fernando A. Quintana. Multivariate Bayesian semiparametric models for authentication of food and beverages. Annals of Applied Statistics, 5(4):2385–2402, December 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1324399599.

Guillot:2015:SPR

[GREG15] Dominique Guillot, Bala Rajaratnam, and Julien Emile-Geay. Statistical paleoclimate reconstructions via Markov random fields. Annals of Applied Statistics, 9(1):324–352, March 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1430226095.

Geyer:2013:LAG

[GRL^{+}13] Charles J. Geyer, Caroline E. Ridley, Robert G. Latta, Julie R. Etterson, and Ruth G. Shaw. Local adaptation and genetic effects on fitness: Calculations for exponential family models with random effects. Annals of Applied Statistics, 7(3):1778–1795, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804816.

Goel:2016:PPU

[GRS16] Sharad Goel, Justin M. Rao, and Ravi Shroff. Precinct or prejudice? Understanding racial disparities in New York City’s stop-and-frisk policy. Annals of Applied Statistics, 10(1):365–394, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1458909920.

Gibson:2023:RTM

[GRS23] Graham C. Gibson, Nicholas G. Reich, and Daniel Sheldon. Real-time mechanistic Bayesian forecasts of COVID-19 mortality. Annals of Applied Statistics, 17(3):1801–1819, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Real-time-mechanistic-Bayesian-forecasts-
of-COVID-19-mortality/10.1214/22-AOAS1671.full.

Guan:2011:BVS

[GS11] Yongtao Guan and Matthew Stephens. Bayesian variable selection regression for genome-wide association studies and other large-scale problems. Annals of Applied Statistics, 5(3):1780–1815, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514285.

Guinness:2013:INH

[GS13] Joseph Guinness and Michael L. Stein. Interpolation of nonstationary high frequency spatial–temporal temperature data. Annals of Applied Statistics, 7(3):1684–1708, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL http://projecteuclid.org/euclid.
aoas/1380804812.

Giganti:2020:ADE

[GSC^{+}20] Mark J. Giganti, Pamela A. Shaw, Guanhua Chen, Sally S. Bebawy, Megan M. Turner, Timothy R. Sterling, and Bryan E. Shepherd. Accounting for dependent errors in predictors and time-to-event outcomes using electronic health records, validation samples and multiple imputation. Annals of Applied Statistics, 14(2):1045–1061, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1593449337.

Guillas:2018:FEH

[GSD^{+}18] Serge Guillas, Andria Sarri, Simon J. Day, Xiaoyu Liu, and Frederic Dias. Functional emulation of high resolution tsunami modelling over Cascadia. Annals of Applied Statistics, 12(4):2023–2053, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078035.

Gertheiss:2010:SMC

[GT10] Jan Gertheiss and Gerhard Tutz. Sparse modeling of categorial explanatory variables. Annals of Applied Statistics, 4(4):2150–2180, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167814.

Gramacy:2013:VSS

[GTW13] Robert B. Gramacy, Matt Taddy, and Stefan M. Wild. Variable selection and sensitivity analysis using dynamic trees, with an application to computer code performance tuning. Annals of Applied Statistics, 7(1):51–80, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1365527190.

Glynn:2021:ASO

[GTZ^{+}21] Chris Glynn, Surya T. Tokdar, Azeem Zaman, Valeria C. Caruso, Jeff T. Mohl, Shawn M. Willett, and Jennifer M. Groh. Analyzing second order stochasticity of neural spiking under stimuli-bundle exposure. Annals of Applied Statistics, 15(1):41–63, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Analyzing-second-
order-stochasticity-of-neural-spiking-under-stimuli-
bundle/10.1214/20-AOAS1383.full.

Guolo:2014:BRT

[GV14] Annamaria Guolo and Cristiano Varin. Beta regression for time series analysis of bounded data, with application to Canada Google^{(}R) flu trends. Annals of Applied Statistics, 8(1):74–88, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966279.

Gao:2019:EDN

[GWZ19] Fei Gao, Yuanjia Wang, and Donglin Zeng. Early diagnosis of neurological disease using peak degeneration ages of multiple biomarkers. Annals of Applied Statistics, 13(2):1295– 1318, June 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1560758447.

Gu:2023:SOS

[GY23] Jiaqi Gu and Philip L. H. Yu. Social order statistics models for ranking data with analysis of preferences in social networks. Annals of Applied Statistics, 17(1):89–107, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Social-order-statistics-models-for-ranking-
data-with-analysis-of/10.1214/22-AOAS1617.full.

Gelfond:2011:LRC

[GZB^{+}11] Jonathan Gelfond, Lee Ann Zarzabal, Tarea Burton, Suzanne Burns, Mari Sogayar, and Luiz O. F. Penalva. Latent rank change detection for analysis of splice-junction microarrays with nonlinear effects. Annals of Applied Statistics, 5(1):364– 380, March 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1300715194.

Hadj-Amar:2021:IRS

[HAFFH21] Beniamino Hadj-Amar, Bärbel Finkenstädt, Mark Fiecas, and Robert Huckstepp. Identifying the recurrence of sleep apnea using a harmonic hidden Markov model. Annals of Applied Statistics, 15(3):1171–1193, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-

applied-statistics/volume-15/issue-3/Identifying-
the-recurrence-of-sleep-apnea-using-a-harmonic-hidden/
10.1214/21-AOAS1455.full.

Hoeltgebaum:2021:USA

[HAL21] Henrique Hoeltgebaum, Niall Adams, and F. Din-Houn Lau. Unsupervised streaming anomaly detection for instrumented infrastructure. Annals of Applied Statistics, 15(3):1101–1125, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Unsupervised-streaming-anomaly-detection-
for-instrumented-infrastructure/10.1214/20-AOAS1424.
full.

Hatfield:2014:DSA

[Hat14] Laura A. Hatfield. Discussion of “Spatial accessibility of pediatric primary healthcare: Measurement and inference”. Annals of Applied Statistics, 8(4):1947–1951, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1419001728. See [NSS14b].

Haviland:2014:DSA

[Hav14] Amelia M. Haviland. Discussion of “Spatial accessibility of pediatric primary healthcare: Measurement and inference”. Annals of Applied Statistics, 8(4):1952–1955, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1419001729. See [NSS14b].

Hazelton:2015:NTI

[Haz15] Martin L. Hazelton. Network tomography for integer-valued traffic. Annals of Applied Statistics, 9(1):474–506, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226101.

Handorf:2013:ECU

[HBHM13] Elizabeth A. Handorf, Justin E. Bekelman, Daniel F. Heitjan, and Nandita Mitra. Evaluating costs with unmeasured confounding: A sensitivity analysis for the treatment effect. Annals of Applied Statistics, 7(4):2062–2080, December 2013.

CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1387823310.

Hunter:2017:AER

[HBP17] David R. Hunter, Le Bao, and Mary Poss. Assignment of endogenous retrovirus integration sites using a mixture model. Annals of Applied Statistics, 11(2):751–770, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1500537722.

Hooghoudt:2017:TBI

[HBW17] Jan-Otto Hooghoudt, Margarida Barroso, and Rasmus Waagepetersen. Toward Bayesian inference of the spatial distribution of proteins from three-cube Förster resonance energy transfer data. Annals of Applied Statistics, 11(3):1711–1737, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168845.

Huang:2021:MVS

[HCD^{+}21] Jing Huang, Yi Cai, Jingcheng Du, Ruosha Li, Susan S. Ellenberg, Sean Hennessy, Cui Tao, and Yong Chen. Monitoring vaccine safety by studying temporal variation of adverse events using vaccine adverse event reporting system. Annals of Applied Statistics, 15(1):252–269, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Monitoring-vaccine-safety-by-studying-
temporal-variation-of-adverse-events/10.1214/20-AOAS1393.
full.

Holzer:2021:HGB

[HCKFZ21] Parker H. Holzer, Jessi Cisewski-Kehe, Debra Fischer, and Lily Zhao. A Hermite–Gaussian based exoplanet radial velocity estimation method. Annals of Applied Statistics, 15(2): 527–555, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/A-HermiteGaussian-based-exoplanet-radial-
velocity-estimation-method/10.1214/20-AOAS1406.full.

Hyun:2017:FRP

[HCP^{+}17] Noorie Hyun, Li C. Cheung, Qing Pan, Mark Schiffman, and Hormuzd A. Katki. Flexible risk prediction models for left or interval-censored data from electronic health records. Annals of Applied Statistics, 11(2):1063–1084, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1500537735.

Hyun:2023:MCP

[HCRB23] Sangwon Hyun, Mattias Rolf Cape, Francois Ribalet, and Jacob Bien. Modeling cell populations measured by flow cytometry with covariates using sparse mixture of regressions. Annals of Applied Statistics, 17(1):357–377, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Modeling-cell-populations-measured-by-
flow-cytometry-with-covariates-using/10.1214/22-AOAS1631.
full.

Ho:2018:DLB

[HCS18] Lam Si Tung Ho, Forrest W. Crawford, and Marc A. Suchard. Direct likelihood-based inference for discretely observed stochastic compartmental models of infectious disease. Annals of Applied Statistics, 12(3):1993–2021, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1536652983.

House:2011:BNM

[HCW11] Leanna L. House, Merlise A. Clyde, and Robert L. Wolpert. Bayesian nonparametric models for peak identification in MALDI-TOF mass spectroscopy. Annals of Applied Statistics, 5(2B):1488–1511, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1310562730.

Hsu:2020:AWS

[HCYH20] Chih-Yuan Hsu, Yi-Hau Chen, Ruoh-Rong Yu, and Tsung-Wei Hung. Assessing wage status transition and stagnation using quantile transition regression. Annals of Applied Statistics, 14(1):160–177, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1587002669.

Holzmann:2014:RIS

[HE14] Hajo Holzmann and Matthias Eulert. The role of the information set for forecasting — with applications to risk management. Annals of Applied Statistics, 8(1):595–621, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1396966300.

Hooker:2015:GFN

[HE15] Giles Hooker and Stephen P. Ellner. Goodness of fit in nonlinear dynamics: Misspecified rates or misspecified states? Annals of Applied Statistics, 9(2):754–776, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1437397110.

Huang:2023:LPO

[HEHM23] Melody Huang, Naoki Egami, Erin Hartman, and Luke Miratrix. Leveraging population outcomes to improve the generalization of experimental results: Application to the JTPA study. Annals of Applied Statistics, 17(3):2139–2164, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Leveraging-population-outcomes-to-improve-
the-generalization-of-experimental-results/10.1214/
22-AOAS1712.full.

Hung:2020:SMR

[HF20] Kenneth Hung and William Fithian. Statistical methods for replicability assessment. Annals of Applied Statistics, 14(3):1063–1087, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Statistical-
methods-for-replicability-assessment/10.1214/20-AOAS1336.
full.

Hughes:2010:LIP

[HFH10] John Hughes, John Fricks, and William Hancock. Likelihood inference for particle location in fluorescence microscopy. Annals of Applied Statistics, 4(2):830–848, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1280842142.

Handcock:2010:MSN

[HG10] Mark S. Handcock and Krista J. Gile. Modeling social networks from sampled data. Annals of Applied Statistics, 4(1): 5–25, March 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1273584445.

Hunt:2021:RSE

[HGB21] Gregory J. Hunt and Johann A. Gagnon-Bartsch. The role of scale in the estimation of cell-type proportions. Annals of Applied Statistics, 15(1):270–286, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/The-role-of-
scale-in-the-estimation-of-cell-type/10.1214/20-AOAS1395.
full.

Hardin:2013:MGR

[HGG13] Johanna Hardin, Stephan Ramon Garcia, and David Golan. A method for generating realistic correlation matrices. Annals of Applied Statistics, 7(3):1733–1762, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1380804814.

Hahn:2015:BHM

[HGM15] P. Richard Hahn, Indranil Goswami, and Carl F. Mela. A Bayesian hierarchical model for inferring player strategy types in a number guessing game. Annals of Applied Statistics, 9(3):1459–1483, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1446488747.

Holsclaw:2017:BNM

[HGRS17] Tracy Holsclaw, Arthur M. Greene, Andrew W. Robertson, and Padhraic Smyth. Bayesian nonhomogeneous Markov models via Pólya–Gamma data augmentation with applications to rainfall modeling. Annals of Applied Statistics, 11(1): 393–426, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616886.

Hewitt:2023:TDA

[HGS23] Joshua Hewitt, Alan E. Gelfand, and Robert S. Schick. Time-discretization approximation enriches continuous-time discrete-space models for animal movement. Annals of Applied Statistics, 17(1):740–760, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Time-discretization-
approximation-enriches-continuous-time-discrete-space-
models-for/10.1214/22-AOAS1649.full.

Humphrey:2023:UPI

[HGSJ23] Colman Humphrey, Ryan Gross, Dylan S. Small, and Shane T. Jensen. Using predictability to improve matching of urban locations in Philadelphia. Annals of Applied Statistics, 17(3):2659–2679, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Using-predictability-
to-improve-matching-of-urban-locations-in-Philadelphia/
10.1214/23-AOAS1739.full.

Hazra:2021:EHR

[HH21] Arnab Hazra and Raphaël Huser. Estimating high-resolution red sea surface temperature hotspots, using a low-rank semiparametric spatial model. Annals of Applied Statistics, 15(2): 572–596, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Estimating-high-resolution-Red-Sea-surface-
temperature-hotspots-using-a/10.1214/20-AOAS1418.
full.

Hanks:2015:CTD

[HHA15] Ephraim M. Hanks, Mevin B. Hooten, and Mat W. Alldredge. Continuous-time discrete-space models for animal movement. Annals of Applied Statistics, 9(1):145–165, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226088.

Hare:2017:AMB

[HHC17] Eric Hare, Heike Hofmann, and Alicia Carriquiry. Automatic matching of bullet land impressions. Annals of Applied Statistics, 11(4):2332–2356, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1514430288.

He:2010:DTE

[HHH10a] Xuming He, Ya-Hui Hsu, and Mingxiu Hu. Detection of treatment effects by covariate-adjusted expected shortfall. Annals of Applied Statistics, 4(4):2114–2125, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167812.

Hosman:2010:SLR

[HHH10b] Carrie A. Hosman, Ben B. Hansen, and Paul W. Holland. The sensitivity of linear regression coefficients’ confidence limits to the omission of a confounder. Annals of Applied Statistics, 4(2):849–870, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1280842143.

Hartmann:2017:GPF

[HHHV17] Marcelo Hartmann, Geoffrey R. Hosack, Richard M. Hillary, and Jarno Vanhatalo. Gaussian process framework for temporal dependence and discrepancy functions in Ricker-type population growth models. Annals of Applied Statistics, 11(3):1375–1402, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168833.

Hanks:2016:LSM

[HHK^{+}16] Ephraim M. Hanks, Mevin B. Hooten, Steven T. Knick, Sara J. Oyler-McCance, Jennifer A. Fike, Todd B. Cross, and Michael K. Schwartz. Latent spatial models and sampling design for landscape genetics. Annals of Applied Statistics, 10(2):1041–1062, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199904.

Han:2016:SMG

[HHLC16] Fang Han, Xiaoyan Han, Han Liu, and Brian Caffo. Sparse median graphs estimation in a high-dimensional semiparametric model. Annals of Applied Statistics, 10(3):1397–1426, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069612.

Hannaford:2020:ICH

[HHN^{+}20] Naomi E. Hannaford, Sarah E. Heaps, Tom M. W. Nye, Tom A. Williams, and T. Martin Embley. Incorporating compositional heterogeneity into Lie Markov models for phylogenetic inference. Annals of Applied Statistics, 14(4):1964– 1983, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Incorporating-compositional-heterogeneity-
into-Lie-Markov-models-for-phylogenetic-inference/
10.1214/20-AOAS1369.full.

Huang:2021:EMG

[HIH^{+}21] Theodore Huang, Gregory Idos, Christine Hong, Stephen B. Gruber, Giovanni Parmigiani, and Danielle Braun. Extending models via gradient boosting: an application to Mendelian models. Annals of Applied Statistics, 15 (3):1126–1146, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-15/issue-3/Extending-models-via-gradient-boosting-
-An-application-to-Mendelian/10.1214/21-AOAS1482.
full.

Holmstrom:2015:BSM

[HISV15] Lasse Holmström, Liisa Ilvonen, Heikki Seppä, and Siim Veski. A Bayesian spatiotemporal model for reconstructing climate from multiple pollen records. Annals of Applied Statistics, 9(3):1194–1225, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1446488736.

Heggeseth:2018:HGM

[HJ18] Brianna C. Heggeseth and Nicholas P. Jewell. How Gaussian mixture models might miss detecting factors that impact growth patterns. Annals of Applied Statistics, 12(1): 222–245, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564471.

Holbrook:2022:BMS

[HJS22] Andrew J. Holbrook, Xiang Ji, and Marc A. Suchard. Bayesian mitigation of spatial coarsening for a Hawkes model applied to gunfire, wildfire and viral contagion. Annals of Applied Statistics, 16(1):573–595, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Bayesian-mitigation-
of-spatial-coarsening-for-a-Hawkes-model-applied/
10.1214/21-AOAS1517.full.

Holzmann:2017:DEB

[HK17a] Hajo Holzmann and Bernhard Klar. Discussion of “Elicitability and backtesting: Perspectives for banking regulation”. Annals of Applied Statistics, 11(4):1875–1882, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1514430266. See [NZ17a].

Holzmann:2017:FRI

[HK17b] Hajo Holzmann and Bernhard Klar. Focusing on regions of interest in forecast evaluation. Annals of Applied Statistics, 11(4):2404–2431, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1514430291.

Hoffmann:2023:MCR

[HK23a] Clara Hoffmann and Nadja Klein. Marginally calibrated response distributions for end-to-end learning in autonomous driving. Annals of Applied Statistics, 17(2):1740–1763, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Marginally-calibrated-response-distributions-
for-end-to-end-learning-in/10.1214/22-AOAS1693.full.

Holzmann:2023:UPI

[HK23b] Hajo Holzmann and Bernhard Klar. Using proxies to improve forecast evaluation. Annals of Applied Statistics, 17(3):2236–2255, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-17/issue-3/Using-proxies-to-improve-forecast-
evaluation/10.1214/22-AOAS1716.full.

Hargreaves:2019:WST

[HKP^{+}19] Jessica K. Hargreaves, Marina I. Knight, Jon W. Pitchford, Rachael J. Oakenfull, Sangeeta Chawla, Jack Munns, and Seth J. Davis. Wavelet spectral testing: Application to nonstationary circadian rhythms. Annals of Applied Statistics, 13(3):1817–1846, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277775.

Huber:2023:BPV

[HKP23] Florian Huber, Tamás Krisztin, and Michael Pfarrhofer. A Bayesian panel vector autoregression to analyze the impact of climate shocks on high-income economies. Annals of Applied Statistics, 17(2):1543–1573, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-2/A-
Bayesian-panel-vector-autoregression-to-analyze-the-
impact-of/10.1214/22-AOAS1681.full.

Hill:2012:FGP

[HKT12] Bryony J. Hill, Wilfrid S. Kendall, and Elke Thönnes. Fibregenerated point processes and fields of orientations. Annals of Applied Statistics, 6(3):994–1020, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418571.

Huang:2008:MLT

[HL08] Yangxin Huang and Tao Lu. Modeling long-term longitudinal HIV dynamics with application to an AIDS clinical study. Annals of Applied Statistics, 2(4):1384–1408, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1231424215.

Hwang:2018:BET

[HLK18] Youngdeok Hwang, Siyuan Lu, and Jae-Kwang Kim. Bottom-up estimation and top-down prediction: Solar energy prediction combining information from multiple sources. Annals of Applied Statistics, 12(4):2096–2120, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078038.

He:2021:SIP

[HLY^{+}21] Shiyuan He, Zhenfeng Lin, Wenlong Yuan, Lucas M. Macri, and Jianhua Z. Huang. Simultaneous inference of periods and period-luminosity relations for Mira variable stars. Annals of Applied Statistics, 15(2):662–687, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Simultaneous-inference-of-periods-and-
period-luminosity-relations-for-Mira/10.1214/21-AOAS1440.
full.

Hong:2009:PRL

[HMM09] Yili Hong, William Q. Meeker, and James D. McCalley. Prediction of remaining life of power transformers based on left truncated and right censored lifetime data. Annals of Applied Statistics, 3(2):857–879, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676198.

Held:2022:ARS

[HMP22] Leonhard Held, Charlotte Micheloud, and Samuel Pawel. The assessment of replication success based on relative effect size. Annals of Applied Statistics, 16(2):706–720, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/The-assessment-of-replication-success-
based-on-relative-effect-size/10.1214/21-AOAS1502.
full.

Hartman:2012:IIN

[HMT12] Brian M. Hartman, Bani K. Mallick, and Debabrata Talukdar. Investigating international new product diffusion speed: a semiparametric approach. Annals of Applied Statistics, 6(2):625–651, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419610.

Hoff:2007:ERL

[Hof07] Peter D. Hoff. Extending the rank likelihood for semiparametric copula estimation. Annals of Applied Statistics, 1(1): 265–283, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143739.

Hoff:2015:MTR

[Hof15] Peter D. Hoff. Multilinear tensor regression for longitudinal relational data. Annals of Applied Statistics, 9(3):1169–1193, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488735.

Holmstrom:2011:DSA

[Hol11] Lasse Holmström. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):71–75, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715177. See [MW11b].

Huang:2023:SAS

[HPB23] Theodore Huang, Matthew Ploenzke, and Danielle Braun. SNIP: An adaptation of sorted neighborhood methods for deduplicating pedigree data. Annals of Applied Statistics, 17(3):2619–2638, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/SNIP--An-adaptation-
of-sorted-neighborhood-methods-for-deduplicating/
10.1214/23-AOAS1735.full.

Huang:2013:LRA

[HPF13] Ying Huang, Margaret S. Pepe, and Ziding Feng. Logistic regression analysis with standardized markers. Annals of Applied Statistics, 7(3):1640–1662, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1380804810.

Huber:2022:IBA

[HR22] Florian Huber and Luca Rossini. Inference in Bayesian additive vector autoregressive tree models. Annals of Applied Statistics, 16(1):104–123, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Inference-in-
Bayesian-additive-vector-autoregressive-tree-models/
10.1214/21-AOAS1488.full.

Huang:2019:CSM

[HRFS19] Yen-Ning Huang, Brian J. Reich, Montserrat Fuentes, and A. Sankarasubramanian. Complete spatial model calibration. Annals of Applied Statistics, 13(2):746–766, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758426.

Haddou:2010:NME

[HRP10] Mohammed Haddou, Louis-Paul Rivest, and Michael Pierrynowski. A nonlinear mixed effects directional model for the estimation of the rotation axes of the human ankle. Annals of Applied Statistics, 4(4):1892–1912, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167803.

Hobolth:2009:SEC

[HS09] Asger Hobolth and Eric A. Stone. Simulation from endpointconditioned, continuous-time Markov chains on a finite state space, with applications to molecular evolution. Annals of Applied Statistics, 3(3):1204–1231, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1254773285.

Huber:2010:BMS

[HS10] Michael R. Huber and Rodney X. Sturdivant. Building a model for scoring 20 or more runs in a baseball game. Annals of Applied Statistics, 4(2):791–804, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1280842140.

Hill:2013:ALC

[HS13] Jennifer Hill and Yu-Sung Su. Assessing lack of common support in causal inference using Bayesian nonparametrics: Implications for evaluating the effect of breastfeeding on children’s cognitive outcomes. Annals of Applied Statistics, 7(3):1386–1420, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804800.

Hill:2014:SPB

[HS14] E. G. Hill and E. H. Slate. A semi-parametric Bayesian model of inter- and intra-examiner agreement for periodontal probing depth. Annals of Applied Statistics, 8(1):331– 351, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966289.

Hector:2022:JIA

[HS22] Emily C. Hector and Peter X.-K. Song. Joint integrative analysis of multiple data sources with correlated vector outcomes. Annals of Applied Statistics, 16(3):1700–1717, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Joint-integrative-analysis-of-multiple-
data-sources-with-correlated-vector/10.1214/21-AOAS1563.
full.

Haneuse:2022:MPE

[HSD^{+}22] Sebastien Haneuse, Deborah Schrag, Francesca Dominici, Sharon-Lise Normand, and Kyu Ha Lee. Measuring performance for end-of-life care. Annals of Applied Statistics, 16(3):1586–1607, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Measuring-performance-
for-end-of-life-care/10.1214/21-AOAS1558.full.

Han:2011:EWO

[HSFP11] Xu Han, Dylan S. Small, Dean P. Foster, and Vishal Patel. The effect of winning an Oscar award on survival: Correcting for healthy performer survivor bias with a rank preserving structural accelerated failure time model. Annals of Applied Statistics, 5(2A):746–772, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562204.

Hays:2012:FDF

[HSH12] Spencer Hays, Haipeng Shen, and Jianhua Z. Huang. Functional dynamic factor models with application to yield curve forecasting. Annals of Applied Statistics, 6(3):870–894, September 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1346418566.

He:2021:BNM

[HSSF21] Shai He, Aaron Schein, Vishal Sarsani, and Patrick Flaherty. A Bayesian nonparametric model for inferring subclonal populations from structured DNA sequencing data. Annals of Applied Statistics, 15(2):925–951, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-2/A-
Bayesian-nonparametric-model-for-inferring-subclonal-
populations-from-structured/10.1214/20-AOAS1434.full.

Huo:2019:BLH

[HST19] Zhiguang Huo, Chi Song, and George Tseng. Bayesian latent hierarchical model for transcriptomic meta-analysis to detect biomarkers with clustered meta-patterns of differential expression signals. Annals of Applied Statistics, 13(1): 340–366, March 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1554861652.

Huber:2009:AMP

[HSVF09] Philippe Huber, Olivier Scaillet, and Maria-Pia Victoria-Feser. Assessing multivariate predictors of financial market movements: a latent factor framework for ordinal data. Annals of Applied Statistics, 3(1):249–271, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888370.

Hofling:2008:SPV

[HT08] Holger Höfling and Robert Tibshirani. A study of prevalidation. Annals of Applied Statistics, 2(2):643–664, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1215118532.

Hu:2013:LE

[HTL13] Yuao Hu, Ye Tian, and Heng Lian. Letter to the Editor. Annals of Applied Statistics, 7(2):1244–1246, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1372338486.

Henao:2013:LPT

[HTM^{+}13] Ricardo Henao, J. Will Thompson, M. Arthur Moseley, Geoffrey S. Ginsburg, Lawrence Carin, and Joseph E. Lucas. Latent protein trees. Annals of Applied Statistics, 7(2):691– 713, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338464.

Horton:2014:AMO

[HTP14] Nicholas J. Horton, Daniell Toth, and Polly Phipps. Adjusting models of ordered multinomial outcomes for nonignorable nonresponse in the occupational employment statistics survey. Annals of Applied Statistics, 8(2):956–973, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229521.

Haran:2011:DSA

[HU11] Murali Haran and Nathan M. Urban. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):61–64, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715175. See [MW11b].

Huang:2018:JST

[Hua18] Yen-Tsung Huang. Joint significance tests for mediation effects of socioeconomic adversity on adiposity via epigenetics. Annals of Applied Statistics, 12(3):1535–1557, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1536652964.

Huang:2019:GWA

[Hua19] Yen-Tsung Huang. Genome-wide analyses of sparse mediation effects under composite null hypotheses. Annals of Applied Statistics, 13(1):60–84, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1554861641.

Hung:2012:OSN

[Hun12] Ying Hung. Order selection in nonlinear time series models with application to the study of cell memory. Annals of Applied Statistics, 6(3):1256–1279, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418582.

Huang:2014:JAS

[HVL14] Yen-Tsung Huang, Tyler J. VanderWeele, and Xihong Lin. Joint analysis of SNP and gene expression data in genetic association studies of complex diseases. Annals of Applied Statistics, 8(1):352–376, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966290.

Hofling:2008:DSA

[HW08] Holger Höfling and Larry Wasserman. Discussion of: “Statistical analysis of an archeological find”. Annals of Applied Statistics, 2(1):77–83, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1206367810.  See

[Feu08b].

Hui:2015:MSD

[HWF15] Francis K. C. Hui, David I. Warton, and Scott D. Foster. Multi-species distribution modeling using penalized mixture of regressions. Annals of Applied Statistics, 9(2):866– 882, June 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1437397115.

Huang:2011:DBN

[HWHWA11] Yangxin Huang, Hulin Wu, Jeanne Holden-Wiltse, and Edward P. Acosta. A dynamic Bayesian nonlinear mixedeffects model of HIV response incorporating medication adherence, drug resistance and covariates. Annals of Applied Statistics, 5(1):551–577, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1300715202.

Hepler:2021:MSC

[HWK21] Staci A. Hepler, Lance A. Waller, and David M. Kline. A multivariate spatiotemporal change-point model of opioid overdose deaths in Ohio. Annals of Applied Statistics, 15(3):1329–1342, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-3/A-multivariate-
spatiotemporal-change-point-model-of-opioid-overdose-
deaths/10.1214/20-AOAS1415.full.

Heard:2010:BAD

[HWPH10] Nicholas A. Heard, David J. Weston, Kiriaki Platanioti, and David J. Hand. Bayesian anomaly detection methods for social networks. Annals of Applied Statistics, 4(2):645– 662, June 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1280842134.

Heller:2014:RAG

[HY14] Ruth Heller and Daniel Yekutieli. Replicability analysis for genome-wide association studies. Annals of Applied Statistics, 8(1):481–498, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966295.

Huang:2023:SMA

[HYL23] Yen-Tsung Huang, Jih-Chang Yu, and Jui-Hsiang Lin. Surrogate marker assessment using mediation and instrumental variable analyses in a case-cohort design. Annals of Applied Statistics, 17(2):1239–1259, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Surrogate-marker-
assessment-using-mediation-and-instrumental-variable-
analyses-in/10.1214/22-AOAS1667.full.

Huling:2019:FCI

[HYS19] Jared D. Huling, Menggang Yu, and Maureen Smith. Fused comparative intervention scoring for heterogeneity of longitudinal intervention effects. Annals of Applied Statistics, 13(2): 824–847, June 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1560758429.

Hu:2022:ATE

[HZF22] Zonghui Hu, Zhiwei Zhang, and Dean Follmann. Assessing treatment effect through compliance score in randomized trials with noncompliance. Annals of Applied Statistics, 16(4):2279–2290, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Assessing-treatment-
effect-through-compliance-score-in-randomized-trials-
with/10.1214/21-AOAS1590.full.

Hu:2022:FSA

[HZG^{+}22a] Liangyuan Hu, Jungang Zou, Chenyang Gu, Jiayi Ji, Michael Lopez, and Minal Kale. A flexible sensitivity analysis approach for unmeasured confounding with multiple treatments and a binary outcome with application to SEER-Medicare lung cancer data. Annals of Applied Statistics, 16(2):1014–1037, June 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-16/issue-2/A-
flexible-sensitivity-analysis-approach-for-unmeasured-
confounding-with-multiple/10.1214/21-AOAS1530.full.

Huang:2022:SAE

[HZG22b] Ying Huang, Yingying Zhuang, and Peter Gilbert. Sensitivity analysis for evaluating principal surrogate endpoints relaxing the equal early clinical risk assumption. Annals of Applied Statistics, 16(3):1774–1794, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Sensitivity-
analysis-for-evaluating-principal-surrogate-endpoints-
relaxing-the-equal/10.1214/21-AOAS1566.full.

Han:2015:IHT

[HZL^{+}15] Shengtong Han, Hongmei Zhang, Gabrielle A. Lockett, Nandini Mukherjee, John W. Holloway, and Wilfried Karmaus.

Identifying heterogeneous transgenerational DNA methylation sites via clustering in beta regression. Annals of Applied Statistics, 9(4):2052–2072, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453994191.

Hu:2015:BDE

[HZY^{+}15] Jie Hu, Zhongying Zhao, Hari Krishna Yalamanchili, Junwen Wang, Kenny Ye, and Xiaodan Fan. Bayesian detection of embryonic gene expression onset in C. elegans. Annals of Applied Statistics, 9(2):950–968, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1437397119.

Irie:2022:SMM

[IGA22] Kaoru Irie, Chris Glynn, and Tevfik Aktekin. Sequential modeling, monitoring, and forecasting of streaming web traffic data. Annals of Applied Statistics, 16(1):300– 325, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Sequential-modeling-monitoring-and-forecasting-
of-streaming-web-traffic-data/10.1214/21-AOAS1505.
full.

Isidro:2016:ESP

[IHJ16] Marissa Isidro, Stephen Haslett, and Geoff Jones. Extended Structure Preserving Estimation (ESPREE) for updating small area estimates of poverty. Annals of Applied Statistics, 10(1):451–476, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1458909923.

Ishwaran:2008:RSF

[IKBL08] Hemant Ishwaran, Udaya B. Kogalur, Eugene H. Blackstone, and Michael S. Lauer. Random survival forests. Annals of Applied Statistics, 2(3):841–860, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1223908043.

Ingermanson:2008:DSA

[Ing08] Randall Ingermanson. Discussion of: “Statistical analysis of an archaeological find”. Annals of Applied Statistics, 2(1):

84–90, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367811. See [Feu08b].

Imai:2013:ETE

[IR13] Kosuke Imai and Marc Ratkovic. Estimating treatment effect heterogeneity in randomized program evaluation. Annals of Applied Statistics, 7(1):443–470, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527206.

Illian:2012:TFC

[ISR12] Janine B. Illian, Sigrunn H. Sørbye, and Håvard Rue. A toolbox for fitting complex spatial point process models using integrated nested Laplace approximation (INLA). Annals of Applied Statistics, 6(4):1499–1530, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629049.

Ionides:2013:MEM

[IWG13] Edward L. Ionides, Zhen Wang, and José A. Tapia Granados. Macroeconomic effects on mortality revealed by panel analysis with nonlinear trends. Annals of Applied Statistics, 7(3):1362–1385, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804799.

James:2007:CAM

[Jam07] Gareth M. James. Curve alignment by moments. Annals of Applied Statistics, 1(2):480–501, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1196438028.

Jackson:2015:TSM

[JAZ15] John C. Jackson, Paul S. Albert, and Zhiwei Zhang. A twostate mixed hidden Markov model for risky teenage driving behavior. Annals of Applied Statistics, 9(2):849–865, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1437397114.

Jha:2021:TSC

[JB21] Jayant Jha and Prajamitra Bhuyan. Two-stage circularcircular regression with zero inflation: Application to medical sciences. Annals of Applied Statistics, 15(3):1343–1365, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Two-stage-circular-circular-regression-
with-zero-inflation--Application/10.1214/20-AOAS1429.
full.

Jeong:2018:RSG

[JCCG18] Jaehong Jeong, Stefano Castruccio, Paola Crippa, and Marc G. Genton. Reducing storage of global wind ensembles with stochastic generators. Annals of Applied Statistics, 12(1):490–509, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564481.

Jeong:2020:SMM

[JCJ20] Seok-Oh Jeong, Dongseok Choi, and Woncheol Jang. A semiparametric mixture method for local false discovery rate estimation from multiple studies. Annals of Applied Statistics, 14(3):1242–1257, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-14/issue-3/A-
semiparametric-mixture-method-for-local-false-discovery-
rate-estimation/10.1214/20-AOAS1341.full.

Johnson:2022:DHT

[JCK22] Michael Johnson, Jiongyi Cao, and Hyunseung Kang. Detecting heterogeneous treatment effects with instrumental variables and application to the Oregon health insurance experiment. Annals of Applied Statistics, 16(2):1111– 1129, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Detecting-heterogeneous-treatment-effects-
with-instrumental-variables-and-application-to/10.
1214/21-AOAS1535.full.

Jensen:2007:BVS

[JCS07] Shane T. Jensen, Guang Chen, and Christian J. Stoeckert, Jr. Bayesian variable selection and data integration for biological regulatory networks. Annals of Applied Statistics, 1(2):612– 633, December 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1196438033.

Jiang:2018:UMT

[JD18] Zhichao Jiang and Peng Ding. Using missing types to improve partial identification with application to a study of HIV prevalence in Malawi. Annals of Applied Statistics, 12(3):1831–1852, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652976.

Jiang:2013:NCF

[JDP^{+}13] Xun Jiang, Dipak K. Dey, Rachel Prunier, Adam M. Wilson, and Kent E. Holsinger. A new class of flexible link functions with application to species co-occurrence in Cape Floristic Region. Annals of Applied Statistics, 7(4):2180– 2204, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823315.

Jensen:2009:BTM

[JEAS09] Shane T. Jensen, Ibrahim Erkan, Erna S. Arnardottir, and Dylan S. Small. Bayesian testing of many hypotheses × many genes: a study of sleep apnea. Annals of Applied Statistics, 3(3):1080–1101, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1254773279.

Jiang:2022:BMS

[JEK^{+}22] Lingjing Jiang, Chris Elrod, Jane J. Kim, Austin D. Swafford, Rob Knight, and Wesley K. Thompson. Bayesian multivariate sparse functional principal components analysis with application to longitudinal microbiome multiomics data. Annals of Applied Statistics, 16(4):2231–2249, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

16/issue-4/Bayesian-multivariate-sparse-functional-
principal-components-analysis-with-application-to/
10.1214/21-AOAS1587.full.

Jung:2011:PAA

[JFM11] Sungkyu Jung, Mark Foskey, and J. S. Marron. Principal arc analysis on direct product manifolds. Annals of Applied Statistics, 5(1):578–603, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1300715203.

Ji:2017:RIS

[JFRS17] Xinyao Ji, Gunther Fink, Paul Jacob Robyn, and Dylan S. Small. Randomization inference for stepped-wedge clusterrandomized trials: an application to community-based health insurance. Annals of Applied Statistics, 11(1):1–20, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1491616869.

Jung:2023:BCR

[JG23] Alexander Wolfgang Jung and Moritz Gerstung. Bayesian Cox regression for large-scale inference with applications to electronic health records. Annals of Applied Statistics, 17(2): 1064–1085, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Bayesian-Cox-regression-for-large-scale-
inference-with-applications-to/10.1214/22-AOAS1658.
full.

Johnson:2018:PFD

[JGC^{+}18] Leah R. Johnson, Robert B. Gramacy, Jeremy Cohen, Erin Mordecai, Courtney Murdock, Jason Rohr, Sadie J. Ryan, Anna M. Stewart-Ibarra, and Daniel Weikel. Phenomenological forecasting of disease incidence using heteroskedastic Gaussian processes: a dengue case study. Annals of Applied Statistics, 12(1):27–66, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564464.

Jackson:2008:SCA

[JGF08] L. Fraser Jackson, Alistair G. Gray, and Stephen E. Fienberg. Sequential category aggregation and partitioning approaches for multi-way contingency tables based on survey and census data. Annals of Applied Statistics, 2(3):955–981, September 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1223908047.

Jarvenpaa:2018:GPM

[JGVM18] Marko Järvenpää, Michael U. Gutmann, Aki Vehtari, and Pekka Marttinen. Gaussian process modelling in approximate Bayesian computation to estimate horizontal gene transfer in bacteria. Annals of Applied Statistics, 12(4):2228–2251, December 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1542078043.

Johnston:2016:GPM

[JHMC16] Ian Johnston, Timothy Hancock, Hiroshi Mamitsuka, and Luis Carvalho. Gene-proximity models for genome-wide association studies. Annals of Applied Statistics, 10(3): 1217–1244, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1475069606.

Ji:2016:CCN

[JJ16a] Pengsheng Ji and Jiashun Jin. Coauthorship and citation networks for statisticians. Annals of Applied Statistics, 10(4):1779–1812, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606836. See discus-

sion [RRS16, WR16, KP16, KT16, Cra16] and rejoinder [JJ16b].

Ji:2016:RCC

[JJ16b] Pengsheng Ji and Jiashun Jin. Rejoinder: “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1846–1853, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1483606842.

Jernigan:2021:LSM

[JJRZ21] Robert Jernigan, Kejue Jia, Zhao Ren, and Wen Zhou. Large-scale multiple inference of collective dependence with applications to protein function. Annals of Applied Statistics, 15(2):902–924, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Large-scale-
multiple-inference-of-collective-dependence-with-
applications-to/10.1214/20-AOAS1431.full.

Jenney:2009:EDM

[JL09] Brenda Jenney and Sharon Lohr. Experimental designs for multiple-level responses, with application to a large-scale educational intervention. Annals of Applied Statistics, 3(2): 691–709, June 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1245676191.

Jang:2010:DEG

[JL10] Woncheol Jang and Ji Meng Loh. Density estimation for grouped data with application to line transect sampling. Annals of Applied Statistics, 4(2):893–915, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1280842145.

Johnson:2011:SES

[JL11] Brent A. Johnson and Qi Long. Survival ensembles by the sum of pairwise differences with application to lung cancer microarray studies. Annals of Applied Statistics, 5(2A):1081– 1101, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562217.

Johndrow:2019:ARS

[JL19] James E. Johndrow and Kristian Lum. An algorithm for removing sensitive information: Application to raceindependent recidivism prediction. Annals of Applied Statistics, 13(1):189–220, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861646.

Jaspers:2016:BAS

[JLA16] Stijn Jaspers, Philippe Lambert, and Marc Aerts. A Bayesian approach to the semiparametric estimation of a minimum inhibitory concentration distribution. Annals of Applied Statistics, 10(2):906–924, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199898.

Jernite:2014:RSM

[JLB^{+}14] Yacine Jernite, Pierre Latouche, Charles Bouveyron, Patrick Rivera, Laurent Jegou, and Stéphane Lamassé. The random subgraph model for the analysis of an ecclesiastical network in Merovingian Gaul. Annals of Applied Statistics, 8(1):377– 405, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966291.

Jara:2010:BSI

[JLDQ10] Alejandro Jara, Emmanuel Lesaffre, Maria De Iorio, and Fernando Quintana. Bayesian semiparametric inference for multivariate doubly-interval-censored data. Annals of Applied Statistics, 4(4):2126–2149, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1294167813.

Jona-Lasinio:2012:SAW

[JLGJL12] Giovanna Jona-Lasinio, Alan Gelfand, and Mattia Jona-Lasinio. Spatial analysis of wave direction data using wrapped Gaussian processes. Annals of Applied Statistics, 6(4):1478–1498, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629048.

Johnson:2009:DSC

[JLL09] W. Evan Johnson, X. Shirley Liu, and Jun S. Liu. Doubly stochastic continuous-time hidden Markov approach for analyzing genome tiling arrays. Annals of Applied Statistics, 3(3):1183–1203, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1254773284.

Jaeger:2019:ORS

[JLL^{+}19] Byron C. Jaeger, D. Leann Long, Dustin M. Long, Mario Sims, Jeff M. Szychowski, Yuan-I Min, Leslie A. Mcclure, George Howard, and Noah Simon. Oblique random survival forests. Annals of Applied Statistics, 13(3):1847–1883, September 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1571277776.

Jung:2020:PPB

[JLLK20] Hohyun Jung, Jae-Gil Lee, Namgil Lee, and Sung-Ho Kim. PTEM: a popularity-based topical expertise model for community question answering. Annals of Applied Statistics, 14(3):1304–1325, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/PTEM--A-popularity-
based-topical-expertise-model-for-community/10.1214/
20-AOAS1346.full.

Josephs:2023:BCA

[JLRK23] Nathaniel Josephs, Lizhen Lin, Steven Rosenberg, and Eric D. Kolaczyk. Bayesian classification, anomaly detection, and survival analysis using network inputs with application to the microbiome. Annals of Applied Statistics, 17(1):199– 224, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Bayesian-classification-anomaly-detection-
and-survival-analysis-using-network-inputs/10.1214/
22-AOAS1623.full.

Jiang:2017:DRE

[JLS^{+}17] Runchao Jiang, Wenbin Lu, Rui Song, Michael G. Hudgens, and Sonia Naprvavnik. Doubly robust estimation of optimal treatment regimes for survival data-with application to an HIV/AIDS study. Annals of Applied Statistics, 11(3):1763–1786, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168847.

Jiang:2023:PAP

[JLS23] Jiming Jiang, Yuanyuan Li, and Peter X. K. Song. Postelection analysis of presidential election/poll data. Annals of Applied Statistics, 17(3):2059–2077, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Postelection-analysis-of-presidential-
electionpoll-data/10.1214/22-AOAS1707.full.

Jadhav:2021:PDC

[JMJ^{+}21] Sneha Jadhav, Chenjin Ma, Yefei Jiang, Ben-Chang Shia, and Shuangge Ma. Pan-disease clustering analysis of the trend of period prevalence. Annals of Applied Statistics, 15(4):1945–1958, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Pan-disease-
clustering-analysis-of-the-trend-of-period-prevalence/
10.1214/21-AOAS1470.full.

Jia:2014:CCS

[JMY^{+}14] Jinzhu Jia, Luke Miratrix, Bin Yu, Brian Gawalt, Laurent El Ghaoui, Luke Barnesmoore, and Sophie Clavier. Concise comparative summaries (CCS) of large text corpora with a human experiment. Annals of Applied Statistics, 8(1):499– 529, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966296.

Jacob:2012:MPG

[JND12] Laurent Jacob, Pierre Neuvial, and Sandrine Dudoit. More power via graph-structured tests for differential expression of gene networks. Annals of Applied Statistics, 6(2):561– 600, June 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1339419608.

Johnstone:2009:AND

[Joh09] Iain M. Johnstone. Approximate null distribution of the largest root in multivariate analysis. Annals of Applied Statistics, 3(4):1616–1633, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1267453956.

Jordan:2010:LB

[Jor10] Michael I. Jordan. Leo Breiman. Annals of Applied Statistics, 4(4):1642–1643, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167788.

Jung:2021:CTC

[JPK21] Sungkyu Jung, Kiho Park, and Byungwon Kim. Clustering on the torus by conformal prediction. Annals of Applied Statistics, 15(4):1583–1603, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Clustering-on-
the-torus-by-conformal-prediction/10.1214/21-AOAS1459.
full.

Johnson:2021:CTS

[JPS21] Devin Johnson, Noel Pelland, and Jeremy Sterling. A continuous-time semi-Markov model for animal movement in a dynamic environment. Annals of Applied Statistics, 15(2): 797–812, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/A-continuous-time-semi-Markov-model-for-
animal-movement-in/10.1214/20-AOAS1408.full.

Jiang:2017:LCM

[JPTO17] Bei Jiang, Eva Petkova, Thaddeus Tarpey, and R. Todd Ogden. Latent class modeling using matrix covariates with application to identifying early placebo responders based on EEG signals. Annals of Applied Statistics, 11(3):1513–1536, September 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1507168838.

Janicki:2022:BNM

[JRHM22] Ryan Janicki, Andrew M. Raim, Scott H. Holan, and Jerry J. Maples. Bayesian nonparametric multivariate spatial mixture mixed effects models with application to American Community Survey special tabulations. Annals of Applied Statistics, 16(1):144–168, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Bayesian-nonparametric-multivariate-spatial-
mixture-mixed-effects-models-with-application/10.
1214/21-AOAS1494.full.

Jun:2008:NCM

[JS08] Mikyoung Jun and Michael L. Stein. Nonstationary covariance models for global data. Annals of Applied Statistics, 2(4):1271–1289, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1231424210.

Jones:2022:IED

[JSF^{+}22] David E. Jones, David C. Stenning, Eric B. Ford, Robert L. Wolpert, Thomas J. Loredo, Christian Gilbertson, and Xavier Dumusque. Improving exoplanet detection power: Multivariate Gaussian process models for stellar activity. Annals of Applied Statistics, 16(2):652– 679, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Improving-exoplanet-detection-power--Multivariate-
Gaussian-process-models-for/10.1214/21-AOAS1471.full.

Johannesson:2022:ABI

[JSH^{+}22] Árni V. Jóhannesson, Stefan Siegert, Raphaël Huser, Haakon Bakka, and Birgir Hrafnkelsson. Approximate Bayesian inference for analysis of spatiotemporal flood frequency data. Annals of Applied Statistics, 16(2):905–935, June 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-2/Approximate-
Bayesian-inference-for-analysis-of-spatiotemporal-
flood-frequency-data/10.1214/21-AOAS1525.full.

Jana:2016:CBR

[JSR16] Kaushik Jana, Debasis Sengupta, and Kalyan Rudra. Correction of bifurcated river flow measurements from historical data: Paving the way for the Teesta water sharing treaty. Annals of Applied Statistics, 10(3):1757–1775, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1475069626.

Jensen:2009:BBH

[JSW09] Shane T. Jensen, Kenneth E. Shirley, and Abraham J. Wyner. Bayesball: a Bayesian hierarchical model for evaluating fielding in major league baseball. Annals of Applied Statistics, 3(2):491–520, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1245676183.

Jiang:2016:ACC

[JSX16] Hui Jiang, Patrick W. Saart, and Yingcun Xia. Asymmetric conditional correlations in stock returns. Annals of Applied Statistics, 10(2):989–1018, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1469199902.

James:2010:SRN

[JSZZ10] Gareth M. James, Chiara Sabatti, Nengfeng Zhou, and Ji Zhu. Sparse regulatory networks. Annals of Applied Statistics, 4(2):663–686, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1280842135.

Jones:2022:CLV

[JTLE22] Andrew Jones, F. William Townes, Didong Li, and Barbara E. Engelhardt. Contrastive latent variable modeling with application to case-control sequencing experiments. Annals of Applied Statistics, 16(3):1268–1291, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Contrastive-latent-variable-modeling-with-
application-to-case-control-sequencing/10.1214/21-
AOAS1534.full.

Jensen:2022:SPP

[JWH22] Louis G. Jensen, David J. Williamson, and Ute Hahn. Semiparametric point process modeling of blinking artifacts in PALM. Annals of Applied Statistics, 16(3):1500–1523, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Semiparametric-point-process-modeling-
of-blinking-artifacts-in-PALM/10.1214/21-AOAS1553.
full.

Jia:2019:RIS

[JWL^{+}19] Gaoxiang Jia, Xinlei Wang, Qiwei Li, Wei Lu, Ximing Tang, Ignacio Wistuba, and Yang Xie. RCRnorm: an integrated system of random-coefficient hierarchical regression models for normalizing NanoString nCounter data. Annals of Applied Statistics, 13(3):1617–1647, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1571277766.

Jun:2019:SDM

[JWZBC19] Seong-Hwan Jun, Samuel W. K. Wong, James V. Zidek, and Alexandre Bouchard-Côté. Sequential decision model for inference and prediction on nonuniform hypergraphs with application to knot matching from computational forestry. Annals of Applied Statistics, 13(3):1678–1707, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1571277768.

Jank:2010:LNO

[JY10] Wolfgang Jank and Inbal Yahav. E-loyalty networks in online auctions. Annals of Applied Statistics, 4(1):151–178, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1273584451.

Jin:2016:BHS

[JYB16] Ick Hoon Jin, Ying Yuan, and Dipankar Bandyopadhyay. A Bayesian hierarchical spatial model for dental caries assessment using non-Gaussian Markov random fields. Annals of Applied Statistics, 10(2):884–905, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1469199897.

Kadane:2008:DBV

[Kad08] Joseph B. Kadane. Discussion of: Bayesian views of an archaeological find. Annals of Applied Statistics, 2(1):97– 98, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367813. See [Feu08b].

Kadane:2018:FS

[Kad18] Joseph B. Kadane. Fingerprint science. Annals of Applied Statistics, 12(2):771–787, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1532743476.

Kafadar:2011:SSM

[Kaf11a] Karen Kafadar. Special section on modern multivariate analysis. Annals of Applied Statistics, 5(4):2265, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399593.

Kafadar:2011:SSS

[Kaf11b] Karen Kafadar. Special section on statistics in neuroscience. Annals of Applied Statistics, 5(2B):1127–1131, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562715.

Kafadar:2012:SSS

[Kaf12] Karen Kafadar. Special section: Statistical methods for next-generation gene sequencing data. Annals of Applied Statistics, 6(2):429–431, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419601.

Knutson:2023:EGC

[KAGK^{+}23] Victoria Knutson, Serge Aleshin-Guendel, Ariel Karlinsky, William Msemburi, and Jon Wakefield. Estimating global and country-specific excess mortality during the Covid-19 pandemic. Annals of Applied Statistics, 17(2):1353– 1374, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Estimating-global-and-country-specific-
excess-mortality-during-the-Covid/10.1214/22-AOAS1673.
full.

Kang:2020:MTA

[Kan20] Yicheng Kang. Measuring timeliness of annual reports filing by jump additive models. Annals of Applied Statistics, 14(4):1604–1621, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-4/Measuring-timeliness-
of-annual-reports-filing-by-jump-additive-models/
10.1214/20-AOAS1365.full.

Kaplan:2011:DSA

[Kap11] Alexey Kaplan. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):47–51, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715172. See [MW11b].

Klausch:2023:BAF

[KAvdW^{+}23] Thomas Klausch, Eddymurphy U. Akwiwu, Mark A. van de Wiel, Veerle M. H. Coupé, and Johannes Berkhof. A Bayesian accelerated failure time model for interval censored three-state screening outcomes. Annals of Applied Statistics, 17(2):1285–1306, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-2/A-
Bayesian-accelerated-failure-time-model-for-interval-
censored-three/10.1214/22-AOAS1669.full.

Kriegler:2010:SAE

[KB10] Brian Kriegler and Richard Berk. Small area estimation of the homeless in Los Angeles: an application of costsensitive stochastic gradient boosting. Annals of Applied Statistics, 4(3):1234–1255, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1287409371.

Kuchibhotla:2023:NCP

[KB23] Arun K. Kuchibhotla and Richard A. Berk. Nested conformal prediction sets for classification with applications to probation data. Annals of Applied Statistics, 17(1):761– 785, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Nested-conformal-prediction-sets-for-classification-
with-applications-to-probation/10.1214/22-AOAS1650.
full.

Kitching:2011:GLA

[KBB^{+}11] Thomas Kitching, Sreekumar Balan, Gary Bernstein, Matthias Bethge, Sarah Bridle, Frederic Courbin, Marc Gentile, Alan Heavens, Michael Hirsch, Reshad Hosseini, Alina Kiessling, Adam Amara, Donnacha Kirk, Konrad Kuijken, Rachel Mandelbaum, Baback Moghaddam, Guldariya Nurbaeva, Stephane Paulin-Henriksson, Anais Rassat, Jason Rhodes, Bernhard Schölkopf, John Shawe-Taylor, Mandeep Gill, Marina Shmakova, Andy Taylor, Malin Velander, Ludovic van Waerbeke, Dugan Witherick, David Wittman, Stefan Harmeling, Catherine Heymans, Richard Massey, Barnaby Rowe, Tim Schrabback, and Lisa Voigt. Gravitational lensing accuracy testing 2010 (GREAT10) challenge handbook. Annals of Applied Statistics, 5(3):2231–2263, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL http://projecteuclid.org/euclid.
aoas/1318514302.

Kormaksson:2012:IMB

[KBFM12] Matthias Kormaksson, James G. Booth, Maria E. Figueroa, and Ari Melnick. Integrative model-based clustering of microarray methylation and expression data. Annals of Applied Statistics, 6(3):1327–1347, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418585.

Kuha:2021:MAA

[KBG21] Jouni Kuha, Erzsébet Bukodi, and John H. Goldthorpe. Mediation analysis for associations of categorical variables: The role of education in social class mobility in Britain. Annals of Applied Statistics, 15(4):2061–2082, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Mediation-analysis-for-associations-of-
categorical-variables--The-role/10.1214/21-AOAS1467.
full.

Kaufman:2011:EEC

[KBH^{+}11] Cari G. Kaufman, Derek Bingham, Salman Habib, Katrin Heitmann, and Joshua A. Frieman. Efficient emulators of computer experiments using compactly supported correlation functions, with an application to cosmology. Annals of Applied Statistics, 5(4):2470–2492, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1324399603.

Keele:2023:HQR

[KBMF^{+}23] Luke J. Keele, Eli Ben-Michael, Avi Feller, Rachel Kelz, and Luke Miratrix. Hospital quality risk standardization via approximate balancing weights. Annals of Applied Statistics, 17(2):901–928, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Hospital-quality-
risk-standardization-via-approximate-balancing-weights/
10.1214/22-AOAS1629.full.

Konomi:2013:BOC

[KDH^{+}13] Bledar A. Konomi, Soma S. Dhavala, Jianhua Z. Huang, Subrata Kundu, David Huitink, Hong Liang, Yu Ding, and Bani K. Mallick. Bayesian object classification of gold nanoparticles. Annals of Applied Statistics, 7(2):640– 668, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338462.

Kim:2019:BMM

[KDH^{+}19] Chanmin Kim, Michael J. Daniels, Joseph W. Hogan, Christine Choirat, and Corwin M. Zigler. Bayesian methods for multiple mediators: Relating principal stratification and causal mediation in the analysis of power plant emission controls. Annals of Applied Statistics, 13(3):1927–1956, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1571277778.

Keller:2017:CAC

[KDL^{+}17] Joshua P. Keller, Mathias Drton, Timothy Larson, Joel D. Kaufman, Dale P. Sandler, and Adam A. Szpiro. Covariateadaptive clustering of exposures for air pollution epidemiology cohorts. Annals of Applied Statistics, 11(1):93–113, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616873.

Karmakar:2020:EFC

[KDS20] Bikram Karmakar, Chyke A. Doubeni, and Dylan S. Small. Evidence factors in a case-control study with application to the effect of flexible sigmoidoscopy screening on colorectal cancer. Annals of Applied Statistics, 14(2):829–849, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1593449328.

Kilner:2010:TIE

[KF10] James M. Kilner and Karl J. Friston. Topological inference for EEG and MEG. Annals of Applied Statistics, 4(3):1272–1290, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409373.

Kim:2011:VEN

[KFB11] Jae Kwang Kim, Wayne A. Fuller, and William R. Bell. Variance estimation for nearest neighbor imputation for US Census long form data. Annals of Applied Statistics, 5(2A):824– 842, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562207.

Knowles:2011:NBS

[KG11] David Knowles and Zoubin Ghahramani. Nonparametric Bayesian sparse factor models with application to gene expression modeling. Annals of Applied Statistics, 5(2B):1534– 1552, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562732.

Kang:2015:JDG

[KGGQ15] Yicheng Kang, Xiaodong Gong, Jiti Gao, and Peihua Qiu. Jump detection in generalized error-in-variables regression with an application to Australian health tax policies. Annals of Applied Statistics, 9(2):883–900, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1437397116.

Krafty:2013:CCA

[KH13] Robert T. Krafty and Martica Hall. Canonical correlation analysis between time series and static outcomes, with application to the spectral analysis of heart rate variability. Annals of Applied Statistics, 7(1):570–587, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1365527211.

Kang:2023:IIR

[KH23] Chaeryon Kang and Ying Huang. Identification of immune response combinations associated with heterogeneous infection risk in the immune correlates analysis of HIV vaccine studies. Annals of Applied Statistics, 17(2):1199– 1219, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Identification-of-immune-response-combinations-
associated-with-heterogeneous-infection-risk/10.1214/
22-AOAS1665.full.

Koslovsky:2020:BTV

[KHBV20] Matthew D. Koslovsky, Emily T. Hébert, Michael S. Businelle, and Marina Vannucci. A Bayesian time-varying effect model for behavioral mHealth data. Annals of Applied Statistics, 14(4):1878–1902, December 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-14/issue-4/A-
Bayesian-time-varying-effect-model-for-behavioral-
mHealth-data/10.1214/20-AOAS1402.full.

Koslovsky:2020:BMM

[KHDV20] Matthew D. Koslovsky, Kristi L. Hoffman, Carrie R. Daniel, and Marina Vannucci. A Bayesian model of microbiome data for simultaneous identification of covariate associations and prediction of phenotypic outcomes. Annals of Applied Statistics, 14(3):1471–1492, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-14/issue-3/A-
Bayesian-model-of-microbiome-data-for-simultaneous-
identification-of/10.1214/20-AOAS1354.full. See cor-

rection [KHDV22].

Koslovsky:2022:CBM

[KHDV22] Matthew D. Koslovsky, Kristi L. Hoffman, Carrie R. Daniel, and Marina Vannucci. Correction to: A Bayesian model of microbiome data for simultaneous identification of covariate associations and prediction of phenotypic outcomes. Annals of Applied Statistics, 16(1):625, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Correction-to--A-Bayesian-model-of-microbiome-
data-for/10.1214/21-AOAS1573.full. See [KHDV20].

Kahmann:2022:IEL

[KHLB22] Sydney Kahmann, Erin Hartman, Jorja Leap, and P. Jeffrey Brantingham. Impact evaluation of the LAPD community safety partnership. Annals of Applied Statistics, 16(2):1215– 1235, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Impact-evaluation-of-the-LAPD-community-
safety-partnership/10.1214/21-AOAS1543.full.

Krantsevich:2023:BDT

[KHZK23] Chelsea Krantsevich, P. Richard Hahn, Yi Zheng, and Charles Katz. Bayesian decision theory for tree-based adaptive screening tests with an application to youth delinquency. Annals of Applied Statistics, 17(2):1038–1063, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Bayesian-decision-theory-for-tree-based-
adaptive-screening-tests-with/10.1214/22-AOAS1657.
full.

Kipnis:2022:HCD

[Kip22] Alon Kipnis. Higher criticism for discriminating wordfrequency tables and authorship attribution. Annals of Applied Statistics, 16(2):1236–1252, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-2/Higher-criticism-
for-discriminating-word-frequency-tables-and-authorship-
attribution/10.1214/21-AOAS1544.full.

Katenka:2012:ICM

[KK12] Natallia Katenka and Eric D. Kolaczyk. Inference and characterization of multi-attribute networks with application to computational biology. Annals of Applied Statistics, 6(3):1068–1094, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1346418574.

Komarek:2013:CMC

[KK13] Arnošt Komárek and Lenka Komárková. Clustering for multivariate continuous and discrete longitudinal data. Annals of Applied Statistics, 7(1):177–200, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1365527195.

Kass:2011:ASM

[KKL11] Robert E. Kass, Ryan C. Kelly, and Wei-Liem Loh. Assessment of synchrony in multiple neural spike trains using loglinear point process models. Annals of Applied Statistics, 5(2B): 1262–1292, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562721.

Kim:2023:GAM

[KKL23] Yura Kim, Daniel Kessler, and Elizaveta Levina. Graphaware modeling of brain connectivity networks. Annals of Applied Statistics, 17(3):2095–2117, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Graph-aware-
modeling-of-brain-connectivity-networks/10.1214/22-
AOAS1709.full.

Klein:2015:BSA

[KKLS15] Nadja Klein, Thomas Kneib, Stefan Lang, and Alexander Sohn. Bayesian structured additive distributional regression with an application to regional income inequality in Germany. Annals of Applied Statistics, 9(2):1024–1052, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1437397122. See correction [KKLS16].

Klein:2016:CBS

[KKLS16] Nadja Klein, Thomas Kneib, Stefan Lang, and Alexander Sohn. Correction: Bayesian structured additive distributional regression with an application to regional income inequality in Germany. Annals of Applied Statistics, 10(2): 1135–1136, June 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1469199908. See [KKLS15].

Kang:2016:FMA

[KKMS16] Hyunseung Kang, Benno Kreuels, Jürgen May, and Dylan S. Small. Full matching approach to instrumental variables estimation with application to the effect of malaria on stunting. Annals of Applied Statistics, 10(1):335–364, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909919.

Kleiber:2013:DMM

[KKR13] William Kleiber, Richard W. Katz, and Balaji Rajagopalan. Daily minimum and maximum temperature simulation over complex terrain. Annals of Applied Statistics, 7(1):588– 612, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527212.

Kang:2016:DFD

[KL16] Jian Kang and Lexin Li. Discussion of “Fiber direction estimation in diffusion MRI”. Annals of Applied Statistics, 10(3):1162–1165, September 2016. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1475069602. See [WLP^{+}16].

Kunihama:2020:BFM

[KLCM20] Tsuyoshi Kunihama, Zehang Richard Li, Samuel J. Clark, and Tyler H. McCormick. Bayesian factor models for probabilistic cause of death assessment with verbal autopsies. Annals of Applied Statistics, 14(1):241–256, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002673.

Koepke:2016:PMC

[KLH^{+}16] Amanda A. Koepke, Ira M. Longini, Jr., M. Elizabeth Halloran, Jon Wakefield, and Vladimir N. Minin. Predictive modeling of cholera outbreaks in Bangladesh. Annals of Applied Statistics, 10(2):575–595, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1469199885.

Keele:2016:HSS

[KM16] Luke Keele and Jason W. Morgan. How strong is strong enough? Strengthening instruments through matching and weak instrument tests. Annals of Applied Statistics, 10(2): 1086–1106, June 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1469199906.

Krivitsky:2017:ISN

[KM17] Pavel N. Krivitsky and Martina Morris. Inference for social network models from egocentrically sampled data, with application to understanding persistent racial disparities in HIV prevalence in the US. Annals of Applied Statistics, 11(1): 427–455, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616887.

King:2016:CRA

[KMKB16] Ruth King, Brett T. McClintock, Darren Kidney, and David Borchers. Capture-recapture abundance estimation using a semi-complete data likelihood approach. Annals of Applied Statistics, 10(1):264–285, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1458909916.

Kao:2013:MME

[KMMS13] Ming-Hung Kao, Dibyen Majumdar, Abhyuday Mandal, and John Stufken. Maximin and maximin-efficient event-related fMRI designs under a nonlinear model. Annals of Applied Statistics, 7(4):1940–1959, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1387823305.

Kadane:2013:NKS

[KN13] Joseph B. Kadane and Ferdinand L. Næshagen. The number of killings in southern rural Norway, 1300–1569. Annals of Applied Statistics, 7(2):846–859, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338470.

Kelly:2017:LTS

[KN17] Luke J. Kelly and Geoff K. Nicholls. Lateral transfer in Stochastic Dollo models. Annals of Applied Statistics, 11(2):

1146–1168, June 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1500537738.

Kiriliouk:2020:CEE

[KN20] Anna Kiriliouk and Philippe Naveau. Climate extreme event attribution using multivariate peaks-over-thresholds modeling and counterfactual theory. Annals of Applied Statistics, 14(3):1342–1358, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Climate-extreme-
event-attribution-using-multivariate-peaks-over-thresholds-
modeling/10.1214/20-AOAS1355.full.

Kang:2014:BHS

[KNWJ14] Jian Kang, Thomas E. Nichols, Tor D. Wager, and Timothy D. Johnson. A Bayesian hierarchical spatial point process model for multi-type neuroimaging meta-analysis. Annals of Applied Statistics, 8(3):1800–1824, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1414091235.

Kelly:2014:CPT

[KÓ14a] Morgan Kelly and Cormac Ó Gráda. Change points and temporal dependence in reconstructions of annual temperature: Did Europe experience a Little Ice Age? Annals of Applied Statistics, 8(3):1372–1394, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1414091217.

Kumazawa:2014:NEM

[KO14b] Takao Kumazawa and Yosihiko Ogata. Nonstationary ETAS models for nonstandard earthquakes. Annals of Applied Statistics, 8(3):1825–1852, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1414091236.

Klein:2020:TGM

[KOB^{+}20] Natalie Klein, Josue Orellana, Scott L. Brincat, Earl K. Miller, and Robert E. Kass. Torus graphs for multivariate phase coupling analysis. Annals of Applied Statistics, 14(2): 635–660, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449319.

Kim:2014:NMP

[KOJ^{+}14] Seongho Kim, Ming Ouyang, Jaesik Jeong, Changyu Shen, and Xiang Zhang. A new method of peak detection for analysis of comprehensive two-dimensional gas chromatography mass spectrometry data. Annals of Applied Statistics, 8(2): 1209–1231, June 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1404229531.

Kosorok:2009:DBD

[Kos09] Michael R. Kosorok. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1270–1278, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1267453935. See [SR09b].

Kosorok:2013:CDB

[Kos13] Michael R. Kosorok. Correction: Discussion of Brownian distance covariance. Annals of Applied Statistics, 7(2): 1247, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338487. See [SR09b].

Kou:2008:SMN

[Kou08] S. C. Kou. Stochastic modeling in nanoscale biophysics: Subdiffusion within proteins. Annals of Applied Statistics, 2(2): 501–535, June 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1215118526.

Kuusela:2015:SUE

[KP15] Mikael Kuusela and Victor M. Panaretos. Statistical unfolding of elementary particle spectra: Empirical Bayes estimation and bias-corrected uncertainty quantification. Annals of Applied Statistics, 9(3):1671–1705, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1446488756.

Karwa:2016:DCC

[KP16] Vishesh Karwa and Sonja Petrović. Discussion of “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1827–1834, December 2016. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1483606839. See [JJ16a, JJ16b].

Karpievitch:2010:LCM

[KPA^{+}10] Yuliya V. Karpievitch, Ashoka D. Polpitiya, Gordon A. Anderson, Richard D. Smith, and Alan R. Dabney. Liquid chromatography mass spectrometry-based proteomics: Biological and technological aspects. Annals of Applied Statistics, 4(4):1797–1823, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167799.

Kunkel:2019:BRM

[KPC^{+}19] Deborah Kunkel, Kevin Potter, Peter F. Craigmile, Mario Peruggia, and Trisha Van Zandt. A Bayesian race model for response times under cyclic stimulus discriminability. Annals of Applied Statistics, 13(1):271–296, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1554861649.

Koh:2023:SWM

[KPDO23] Jonathan Koh, François Pimont, Jean-Luc Dupuy, and Thomas Opitz. Spatiotemporal wildfire modeling through point processes with moderate and extreme marks. Annals of Applied Statistics, 17(1):560–582, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Spatiotemporal-
wildfire-modeling-through-point-processes-with-moderate-
and-extreme/10.1214/22-AOAS1642.full.

Kratz:2017:DEB

[Kra17] Marie Kratz. Discussion of “Elicitability and backtesting: Perspectives for banking regulation”. Annals of Applied Statistics, 11(4):1894–1900, December 2017. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1514430270. See [NZ17a].

Kelly:2023:LCD

[KRC23] Luke J. Kelly, Robin J. Ryder, and Grégoire Clarté. Lagged couplings diagnose Markov chain Monte Carlo phylogenetic inference. Annals of Applied Statistics, 17(2):1419– 1443, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Lagged-couplings-diagnose-Markov-chain-
Monte-Carlo-phylogenetic-inference/10.1214/22-AOAS1676.
full.

Kuusela:2017:SCU

[KS17] Mikael Kuusela and Philip B. Stark. Shape-constrained uncertainty quantification in unfolding steeply falling elementary particle spectra. Annals of Applied Statistics, 11 (3):1671–1710, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168844.

Katsevich:2019:MKF

[KS19] Eugene Katsevich and Chiara Sabatti. Multilayer knockoff filter: Controlled variable selection at multiple resolutions. Annals of Applied Statistics, 13(1):1–33, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1554861639.

Kolar:2010:ETV

[KSAX10] Mladen Kolar, Le Song, Amr Ahmed, and Eric P. Xing. Estimating time-varying networks. Annals of Applied Statistics, 4(1):94–123, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1273584449.

Karwa:2011:CIT

[KSD11] Vishesh Karwa, Aleksandra B. Slavković, and Eric T. Donnell. Causal inference in transportation safety studies: Comparison of potential outcomes and causal diagrams. Annals of Applied Statistics, 5(2B):1428–1455, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1310562728.

Kleiber:2013:PTM

[KSH^{+}13] William Kleiber, Stephan R. Sain, Matthew J. Heaton, Michael Wiltberger, C. Shane Reese, and Derek Bingham. Parameter tuning for a multi-fidelity dynamical model of the magnetosphere. Annals of Applied Statistics, 7(3):1286–1310, September 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1380804796.

Kobak:2016:IPE

[KSP16] Dmitry Kobak, Sergey Shpilkin, and Maxim S. Pshenichnikov. Integer percentages as electoral falsification fingerprints. Annals of Applied Statistics, 10(1):54–73, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909907.

Kim:2021:VPS

[KSW^{+}21] Juhyun Kim, Judong Shen, Anran Wang, Devan V. Mehrotra, Seyoon Ko, Jin J. Zhou, and Hua Zhou. VCSEL: Prioritizing SNP-set by penalized variance component selection. Annals of Applied Statistics, 15(4):1652–1672, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/VCSEL-Prioritizing-SNP-set-by-penalized-
variance-component-selection/10.1214/21-AOAS1491.
full.

Kolar:2016:DCC

[KT16] Mladen Kolar and Matt Taddy. Discussion of “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1835–1841, December 2016. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1483606840. See [JJ16a, JJ16b].

Kim:2023:RSA

[KW23] Jungeum Kim and Xiao Wang. Robust sensible adversarial learning of deep neural networks for image classification. Annals of Applied Statistics, 17(2):961–984, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Robust-sensible-adversarial-learning-of-
deep-neural-networks-for-image/10.1214/22-AOAS1637.
full.

Kim:2012:TGG

[KX12] Seyoung Kim and Eric P. Xing. Tree-guided group lasso for multi-response regression with structured sparsity, with an application to eQTL mapping. Annals of Applied Statistics, 6(3):1095–1117, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1346418575.

Kim:2009:NLC

[KXC09] Sungduk Kim, Yingmei Xi, and Ming-Hui Chen. A new latent cure rate marker model for survival data. Annals of Applied Statistics, 3(3):1124–1146, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1254773281.

Klebanov:2007:DCS

[KY07] Lev Klebanov and Andrei Yakovlev. Diverse correlation structures in gene expression data and their utility in improving statistical inference. Annals of Applied Statistics, 1(2):538–559, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1196438030.

Kilcioglu:2016:MIC

[KZ16] Cinar Kilcioglu and José R. Zubizarreta. Maximizing the information content of a balanced matched sample in a study of the economic performance of green buildings. Annals of Applied Statistics, 10(4):1997–2020, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1483606849.

Kuha:2023:LVM

[KZS23] Jouni Kuha, Siliang Zhang, and Fiona Steele. Latent variable models for multivariate dyadic data with zero inflation: Analysis of intergenerational exchanges of family support. Annals of Applied Statistics, 17(2):1521– 1542, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Latent-variable-models-for-multivariate-
dyadic-data-with-zero-inflation/10.1214/22-AOAS1680.
full.

Lila:2022:FRE

[LA22] Eardi Lila and John A. D. Aston. Functional random effects modeling of brain shape and connectivity. Annals of Applied Statistics, 16(4):2122–2144, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Functional-random-effects-modeling-of-
brain-shape-and-connectivity/10.1214/21-AOAS1572.
full.

Lila:2016:SPC

[LAS16] Eardi Lila, John A. D. Aston, and Laura M. Sangalli. Smooth principal component analysis over two-dimensional manifolds with an application to neuroimaging. Annals of Applied Statistics, 10(4):1854–1879, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1483606843.

Lazar:2016:DFD

[Laz16] Nicole A. Lazar. Discussion of “Fiber direction estimation in diffusion MRI”. Annals of Applied Statistics, 10 (3):1160–1161, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1475069601.  See

[WLP^{+}16].

Latouche:2011:OSB

[LBA11] Pierre Latouche, Etienne Birmelé, and Christophe Ambroise. Overlapping stochastic block models with application to the French political blogosphere. Annals of Applied Statistics, 5(1):309–336, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1300715192.

Lin:2021:UQC

[LBBM21] Luyao Lin, Derek Bingham, Floor Broekgaarden, and Ilya Mandel. Uncertainty quantification of a computer model for binary black hole formation. Annals of Applied Statistics, 15(4):1604–1627, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Uncertainty-
quantification-of-a-computer-model-for-binary-black-
hole/10.1214/21-AOAS1484.full.

Li:2018:TLS

[LBD^{+}18a] Xin Li, Alex Belianinov, Ondrej Dyck, Stephen Jesse, and Chiwoo Park. Two-level structural sparsity regularization for identifying lattices and defects in noisy images. Annals of Applied Statistics, 12(1):348–377, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1520564476.

Linero:2018:MRM

[LBD18b] Antonio R. Linero, Jonathan R. Bradley, and Apurva Desai. Multi-rubric models for ordinal spatial data with application to online ratings data. Annals of Applied Statistics, 12(4):2054–2074, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078036.

Li:2011:MRH

[LBHB11] Qunhua Li, James B. Brown, Haiyan Huang, and Peter J. Bickel. Measuring reproducibility of high-throughput experiments. Annals of Applied Statistics, 5(3):1752–1779, September 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1318514284.

Lee:2023:ECE

[LBK^{+}23] TingFang Lee, Ashley L. Buchanan, Natallia V. Katenka, Laura Forastiere, M. Elizabeth Halloran, Samuel R. Friedman, and Georgios Nikolopoulos. Estimating causal effects of HIV prevention interventions with interference in network-based studies among people who inject drugs. Annals of Applied Statistics, 17(3):2165–2191, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Estimating-causal-effects-of-HIV-prevention-

interventions-with-interference-in/10.1214/22-AOAS1713.
full.

Liu:2020:CPC

[LBL20] Zhonghua Liu, Ian Barnett, and Xihong Lin. A comparison of principal component methods between multiple phenotype regression and multiple SNP regression in genetic association studies. Annals of Applied Statistics, 14(1):433–451, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002681.

Lee:2013:BSM

[LBND13] Giwhyun Lee, Eunshin Byon, Lewis Ntaimo, and Yu Ding. Bayesian spline method for assessing extreme loads on wind turbines. Annals of Applied Statistics, 7(4):2034–2061, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823309.

Li:2010:GSF

[LC10] Gengxin Li and Yuehua Cui. A general statistical framework for dissecting parent-of-origin effects underlying endosperm traits in flowering plants. Annals of Applied Statistics, 4(3):1214–1233, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409370.

Li:2012:GCG

[LC12] Shaoyu Li and Yuehua Cui. Gene-centric gene–gene interaction: a model-based kernel machine method. Annals of Applied Statistics, 6(3):1134–1161, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418577.

Long:2016:ECP

[LCB16] James P. Long, Eric C. Chi, and Richard G. Baraniuk. Estimating a common period for a set of irregularly sampled functions with applications to periodic variable star data. Annals of Applied Statistics, 10(1):165–197, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1458909912.

Liang:2009:AMC

[LCG09] Shengde Liang, Bradley P. Carlin, and Alan E. Gelfand. Analysis of Minnesota colon and rectum cancer point patterns with spatial and nonspatial covariate information. Annals of Applied Statistics, 3(3):943–962, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1254773273.

Long:2011:RPP

[LCMJ11] Qi Long, Matthias Chung, Carlos S. Moreno, and Brent A. Johnson. Risk prediction for prostate cancer recurrence through regularized estimation with simultaneous adjustment for nonlinear clinical effects. Annals of Applied Statistics, 5(3):2003–2023, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514293.

Liu:2021:PPM

[LCRM21] Xueying Liu, Jeremy Carter, Brad Ray, and George Mohler. Point process modeling of drug overdoses with heterogeneous and missing data. Annals of Applied Statistics, 15(1):88– 101, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Point-process-modeling-of-drug-overdoses-
with-heterogeneous-and-missing/10.1214/20-AOAS1384.
full.

Lu:2015:BAA

[LCSZ15] Zhao-Hua Lu, Sy-Miin Chow, Andrew Sherwood, and Hongtu Zhu. Bayesian analysis of ambulatory blood pressure dynamics with application to irregularly spaced sparse data. Annals of Applied Statistics, 9(3):1601–1620, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1446488753.

Li:2023:RFD

[LCYZ23] Yan Li, Kun Chen, Jun Yan, and Xuebin Zhang. Regularized fingerprinting in detection and attribution of climate change with weight matrix optimizing the efficiency in scaling factor estimation. Annals of Applied Statistics, 17(1):225–239, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Regularized-
fingerprinting-in-detection-and-attribution-of-climate-
change-with/10.1214/22-AOAS1624.full.

Ludwig:2017:SRS

[LCZ^{+}17] Guilherme Ludwig, Tingjin Chu, Jun Zhu, Haonan Wang, and Kirsten Koehler. Static and roving sensor data fusion for spatio-temporal hazard mapping with application to occupational exposure assessment. Annals of Applied Statistics, 11(1):139–160, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616875.

Leday:2017:GNR

[LdGK^{+}17] Gwenaël G. R. Leday, Mathisca C. M. de Gunst, Gino B. Kpogbezan, Aad W. van der Vaart, Wessel N. van Wieringen, and Mark A. van de Wiel. Gene network reconstruction using global-local shrinkage priors. Annals of Applied Statistics, 11(1):41–68, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616871.

Lennox:2010:DPM

[LDV^{+}10] Kristin P. Lennox, David B. Dahl, Marina Vannucci, Ryan Day, and Jerry W. Tsai. A Dirichlet process mixture of hidden Markov models for protein structure prediction. Annals of Applied Statistics, 4(2):916–942, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1280842146.

Lee:2018:LAP

[Lee18] Duncan Lee. A locally adaptive process-convolution model for estimating the health impact of air pollution. Annals of Applied Statistics, 12(4):2540–2558, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078055.

Li:2012:LBS

[LES12] Qunhua Li, Jimmy K. Eng, and Matthew Stephens. A likelihood-based scoring method for peptide identification using mass spectrometry. Annals of Applied Statistics, 6 (4):1775–1794, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629059.

Lopez-Fidalgo:2023:DEE

[LFMM23] Jesus Lopez-Fidalgo, Caterina May, and Jose Antonio Moler. Designing experiments for estimating an appropriate outlet size for a silo type problem. Annals of Applied Statistics, 17(1):606–620, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Designing-experiments-
for-estimating-an-appropriate-outlet-size-for-a/10.
1214/22-AOAS1644.full.

Li:2018:GFA

[LG18] Gen Li and Irina Gaynanova. A general framework for association analysis of heterogeneous data. Annals of Applied Statistics, 12(3):1700–1726, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1536652971.

Lenzi:2020:SPW

[LG20] Amanda Lenzi and Marc G. Genton. Spatiotemporal probabilistic wind vector forecasting over Saudi Arabia. Annals of Applied Statistics, 14(3):1359–1378, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Spatiotemporal-probabilistic-wind-vector-
forecasting-over-Saudi-Arabia/10.1214/20-AOAS1347.
full.

Liu:2018:STM

[LGK18] Xiao Liu, Vikneswaran Gopal, and Jayant Kalagnanam. A spatio-temporal modeling framework for weather radar image data in tropical Southeast Asia. Annals of Applied Statistics, 12(1):378–407, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564477.

Li:2012:SSI

[LGL^{+}12] Zhen Li, Vikneswaran Gopal, Xiaobo Li, John M. Davis, and George Casella. Simultaneous SNP identification in association studies with missing data. Annals of Applied Statistics, 6(2):432–456, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419602.

Lange:2018:ECC

[LGL^{+}18] Jane M. Lange, Roman Gulati, Amy S. Leonardson, Daniel W. Lin, Lisa F. Newcomb, Bruce J. Trock, H. Ballentine Carter, Peter R. Carroll, Matthew R. Cooperberg, Janet E. Cowan, Lawrence H. Klotz, and Ruth Etzioni. Estimating and comparing cancer progression risks under varying surveillance protocols. Annals of Applied Statistics, 12(3):1773–1795, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652974.

Lee:2020:FPB

[LHF^{+}20] Ben Seiyon Lee, Murali Haran, Robert W. Fuller, David Pollard, and Klaus Keller. A fast particle-based approach for calibrating a 3-D model of the Antarctic ice sheet. Annals of Applied Statistics, 14(2):605–634, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1593449318.

Lee:2010:SLP

[LHH10] Seokho Lee, Jianhua Z. Huang, and Jianhua Hu. Sparse logistic principal components analysis for binary data. Annals of Applied Statistics, 4(3):1579–1601, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1287409387.

Lorenzi:2019:HIF

[LHH19] Elizabeth Lorenzi, Ricardo Henao, and Katherine Heller. Hierarchical infinite factor models for improving the prediction of surgical complications for geriatric patients. Annals of Applied Statistics, 13(4):2637–2661, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1574910058.

Lock:2013:JIV

[LHMN13] Eric F. Lock, Katherine A. Hoadley, J. S. Marron, and Andrew B. Nobel. Joint and individual variation explained (JIVE) for integrated analysis of multiple data types. Annals of Applied Statistics, 7(1):523–542, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1365527209.

Li:2013:BIN

[LHPW13] Shuang Li, Li Hsu, Jie Peng, and Pei Wang. Bootstrap inference for network construction with an application to a breast cancer microarray study. Annals of Applied Statistics, 7(1): 391–417, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527204.

Liang:2019:EBA

[Lia19] Kun Liang. Empirical Bayes analysis of RNA sequencing experiments with auxiliary information. Annals of Applied Statistics, 13(4):2452–2482, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910051.

Liebl:2013:MFE

[Lie13] Dominik Liebl. Modeling and forecasting electricity spot prices: A functional data perspective. Annals of Applied Statistics, 7(3):1562–1592, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804807.

Liebl:2019:NTD

[Lie19] Dominik Liebl. Nonparametric testing for differences in electricity prices: The case of the Fukushima nuclear accident. Annals of Applied Statistics, 13(2):1128–1146, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758440.

Lin:2013:ANR

[Lin13] Winston Lin. Agnostic notes on regression adjustments to experimental data: Reexamining Freedman’s critique.

Annals of Applied Statistics, 7(1):295–318, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1365527200. See [Fre08].

Langrock:2013:MLE

[LK13] Roland Langrock and Ruth King. Maximum likelihood estimation of mark–recapture–recovery models in the presence of continuous covariates. Annals of Applied Statistics, 7(3):1709–1732, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804813.

Li:2021:NIS

[LKB21] Shuoran Li, Young Myoung Ko, and Eunshin Byon. Nonparametric importance sampling for wind turbine reliability analysis with stochastic computer models. Annals of Applied Statistics, 15(4):1850–1871, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Nonparametric-
importance-sampling-for-wind-turbine-reliability-
analysis-with-stochastic/10.1214/21-AOAS1490.full.

Low-Kam:2015:BRT

[LKTJ^{+}15] Cecile Low-Kam, Donatello Telesca, Zhaoxia Ji, Haiyuan Zhang, Tian Xia, Jeffrey I. Zink, and Andre E. Nel. A Bayesian regression tree approach to identify the effect of nanoparticles’ properties on toxicity profiles. Annals of Applied Statistics, 9(1):383–401, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226097.

Luo:2009:MSI

[LL09] Ruiyan Luo and Bret Larget. Modeling substitution and indel processes for AFLP marker evolution and phylogenetic inference. Annals of Applied Statistics, 3(1):222–248, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888369.

Li:2010:VSR

[LL10] Caiyan Li and Hongzhe Li. Variable selection and regression analysis for graph-structured covariates with an application to genomics. Annals of Applied Statistics, 4(3):1498–1516, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409383.

Li:2011:ENN

[LL11] Mengxin Li and Wei-Liem Loh. Estimating the number of neurons in multi-neuronal spike trains. Annals of Applied Statistics, 5(1):176–200, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1300715187.

Lee:2016:QSI

[LL16] Duncan Lee and Andrew Lawson. Quantifying the spatial inequality and temporal trends in maternal smoking rates in Glasgow. Annals of Applied Statistics, 10(3):1427–1446, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069613.

Li:2019:PSW

[LL19] Fan Li and Fan Li. Propensity score weighting for causal inference with multiple treatments. Annals of Applied Statistics, 13(4):2389–2415, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1574910049.

Livsey:2018:MIV

[LLKP18] James Livsey, Robert Lund, Stefanos Kechagias, and Vladas Pipiras. Multivariate integer-valued time series with flexible autocovariances and their application to major hurricane counts. Annals of Applied Statistics, 12(1):408–431, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564478.

Lu:2010:MAC

[LLL10] QiQi Lu, Robert Lund, and Thomas C. M. Lee. An MDL approach to the climate segmentation problem. Annals of Applied Statistics, 4(1):299–319, March 2010. CODEN ????

ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584456.

Liu:2020:FSQ

[LLM20] Yusha Liu, Meng Li, and Jeffrey S. Morris. Function-on-scalar quantile regression with application to mass spectrometry proteomics data. Annals of Applied Statistics, 14(2): 521–541, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449314.

Levy-Leduc:2009:DLC

[LLR09] Céline Lévy-Leduc and François Roueff. Detection and localization of change-points in high-dimensional network traffic data. Annals of Applied Statistics, 3(2):637–662, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1245676189.

Lee:2010:SGA

[LLR10] Ann B. Lee, Diana Luca, and Kathryn Roeder. A spectral graph approach to discovering genetic ancestry. Annals of Applied Statistics, 4(1):179–202, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584452.

Liu:2015:NAA

[LLR15] Li Liu, Jing Lei, and Kathryn Roeder. Network assisted analysis to reveal the genetic basis of autism. Annals of Applied Statistics, 9(3):1571–1600, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1446488752.

Ling:2022:MEC

[LLS^{+}22] Yun Ling, Martin Lysy, Ian Seim, Jay Newby, David B. Hill, Jeremy Cribb, and M. Gregory Forest. Measurement error correction in particle tracking microrheology. Annals of Applied Statistics, 16(3):1747–1773, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Measurement-
error-correction-in-particle-tracking-microrheology/
10.1214/21-AOAS1565.full.

Li:2019:PMN

[LLZ19] Tianxi Li, Elizaveta Levina, and Ji Zhu. Prediction models for network-linked data. Annals of Applied Statistics, 13(1): 132–164, March 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1554861644.

Lau:2010:AMS

[LM10a] Ada Lau and Patrick McSharry. Approaches for multi-step density forecasts with application to aggregated wind power. Annals of Applied Statistics, 4(3):1311–1341, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1287409375. See correction [LM10b].

Lau:2010:CNA

[LM10b] Ada Lau and Patrick McSharry. Correction note: “Approaches for multi-step density forecasts with application to aggregated wind power”. Annals of Applied Statistics, 4(4): 2205, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167817. See [LM10a].

Lopez:2018:HOD

[LMB18] Michael J. Lopez, Gregory J. Matthews, and Benjamin S. Baumer. How often does the best team win? A unified approach to understanding randomness in North American sport. Annals of Applied Statistics, 12(4):2483–2516, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1542078053.

Lee:2015:BFA

[LMGJ15] Juhee Lee, Peter Müller, Kamalakar Gulukota, and Yuan Ji. A Bayesian feature allocation model for tumor heterogeneity. Annals of Applied Statistics, 9(2):621–639, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1437397104.

Liu:2012:SZI

[LMKC12] Hai Liu, Shuangge Ma, Richard Kronmal, and Kung-Sik Chan. Semiparametric zero-inflated modeling in multi-ethnic study of atherosclerosis (MESA). Annals of Applied Statistics, 6(3):1236–1255, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1346418581.

Li:2015:ECE

[LMM15] Fan Li, Alessandra Mattei, and Fabrizia Mealli. Evaluating the causal effect of university grants on student dropout: Evidence from a regression discontinuity design using principal stratification. Annals of Applied Statistics, 9(4):1906– 1931, December 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1453993098.

Li:2021:RDD

[LMMS21] Fan Li, Andrea Mercatanti, Taneli Mäkinen, and Andrea Silvestrini. A regression discontinuity design for ordinal running variables: Evaluating central bank purchases of corporate bonds. Annals of Applied Statistics, 15(1):304– 322, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/A-regression-discontinuity-design-for-
ordinal-running-variables--Evaluating/10.1214/20-
AOAS1396.full.

Li:2010:NMM

[LMS10] Qunhua Li, Michael J. MacCoss, and Matthew Stephens. A nested mixture model for protein identification using mass spectrometry. Annals of Applied Statistics, 4(2):962– 987, June 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1280842148.

Liang:2010:ECT

[LMW10] Hua Liang, Hongyu Miao, and Hulin Wu. Estimation of constant and time-varying dynamic parameters of HIV infection in a nonlinear differential equation model. Annals of Applied Statistics, 4(1):460–483, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584463.

Lund:2012:IDM

[LN12] Steven P. Lund and Dan Nettleton. The importance of distinct modeling strategies for gene and gene-specific treatment effects in hierarchical models for microarray data. Annals of Applied Statistics, 6(3):1118–1133, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418576.

Li:2019:BHM

[LNC^{+}19] Yang Li, Shaoyang Ning, Sarah E. Calvo, Vamsi K. Mootha, and Jun S. Liu. Bayesian hidden Markov tree models for clustering genes with shared evolutionary history. Annals of Applied Statistics, 13(1):606–637, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1554861662.

Li:2022:BBG

[LNR^{+}22] Yuliang Li, Yang Ni, Leah H. Rubin, Amanda B. Spence, and Yanxun Xu. BAGEL: a Bayesian graphical model for inferring drug effect longitudinally on depression in people with HIV. Annals of Applied Statistics, 16(1):21–39, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/BAGEL--A-Bayesian-graphical-model-for-
inferring-drug-effect/10.1214/21-AOAS1492.full.

Lee:2008:RTA

[LNW08a] Ann B. Lee, Boaz Nadler, and Larry Wasserman. Rejoinder of: “Treelets — An adaptive multi-scale basis for spare unordered data”. Annals of Applied Statistics, 2(2):494– 500, June 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1215118525. See [LNW08b].

Lee:2008:TAM

[LNW08b] Ann B. Lee, Boaz Nadler, and Larry Wasserman. Treelets — an adaptive multi-scale basis for sparse unordered data. Annals of Applied Statistics, 2(2):435–471, June 2008. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1215118518. See discussion [Mur08, BR08, MB08, Tib08, Qiu08, TvdL08] and rejoinder [LNW08a].

Loh:2009:IPC

[Loh09] Wei-Yin Loh. Improving the precision of classification trees. Annals of Applied Statistics, 3(4):1710–1737, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1267453961.

Lock:2022:BLM

[LPH22] Eric F. Lock, Jun Young Park, and Katherine A. Hoadley. Bidimensional linked matrix factorization for pan-omics pancancer analysis. Annals of Applied Statistics, 16(1):193– 215, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Bidimensional-linked-matrix-factorization-
for-pan-omics-pan-cancer-analysis/10.1214/21-AOAS1495.
full.

Loewinger:2022:HRB

[LPKP22] Gabriel Loewinger, Prasad Patil, Kenneth T. Kishida, and Giovanni Parmigiani. Hierarchical resampling for bagging in multistudy prediction with applications to human neurochemical sensing. Annals of Applied Statistics, 16(4):2145– 2165, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Hierarchical-resampling-for-bagging-in-
multistudy-prediction-with-applications-to/10.1214/
21-AOAS1574.full.

Long:2011:SSS

[LPT^{+}11] Christopher J. Long, Patrick L. Purdon, Simona Temereanca, Neil U. Desai, Matti S. Hämäläinen, and Emery N. Brown. State-space solutions to the dynamic magnetoencephalography inverse problem using high performance computing. Annals of Applied Statistics, 5(2B):1207–1228, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562719.

Liu:2019:IPS

[LQNM19] Lin Liu, Yuqi Qiu, Loki Natarajan, and Karen Messer. Imputation and post-selection inference in models with missing data: an application to colorectal cancer surveillance guidelines. Annals of Applied Statistics, 13(3):1370–1396, September 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1571277757.

Li:2020:EFS

[LR20a] Yicheng Li and Adrian E. Raftery. Estimating and forecasting the smoking-attributable mortality fraction for both genders jointly in over 60 countries. Annals of Applied Statistics, 14(1):381–408, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1587002679.

Liu:2020:AUA

[LR20b] Peiran Liu and Adrian E. Raftery. Accounting for uncertainty about past values in probabilistic projections of the total fertility rate for most countries. Annals of Applied Statistics, 14(2):685–705, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1593449321.

Li:2021:ASF

[LR21] Yicheng Li and Adrian E. Raftery. Accounting for smoking in forecasting mortality and life expectancy. Annals of Applied Statistics, 15(1):437–459, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Accounting-for-
smoking-in-forecasting-mortality-and-life-expectancy/
10.1214/20-AOAS1381.full.

Legramanti:2022:ESB

[LRDD22] Sirio Legramanti, Tommaso Rigon, Daniele Durante, and David B. Dunson. Extended stochastic block models with application to criminal networks. Annals of Applied Statistics, 16(4):2369–2395, December 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Extended-stochastic-
block-models-with-application-to-criminal-networks/
10.1214/21-AOAS1595.full.

Liu:2012:FFA

[LRHF12] Chong Liu, Surajit Ray, Giles Hooker, and Mark Friedl. Functional factor analysis for periodic remote sensing data. Annals of Applied Statistics, 6(2):601–624, June 2012. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1339419609.

Lim:2021:MBF

[LRI21] David K. Lim, Naim U. Rashid, and Joseph G. Ibrahim. Model-based feature selection and clustering of RNA-seq data for unsupervised subtype discovery. Annals of Applied Statistics, 15(1):481–508, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Model-based-
feature-selection-and-clustering-of-RNA-seq-data/
10.1214/20-AOAS1407.full.

Lofland:2017:ADL

[LRM17] Chelsea L. Lofland, Abel Rodríguez, and Scott Moser. Assessing differences in legislators’ revealed preferences: a case study on the 107th U.S. Senate. Annals of Applied Statistics, 11(1):456–479, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616888.

Letham:2015:ICU

[LRMM15] Benjamin Letham, Cynthia Rudin, Tyler H. McCormick, and David Madigan. Interpretable classifiers using rules and Bayesian analysis: Building a better stroke prediction model. Annals of Applied Statistics, 9(3):1350–1371, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488742.

Loredo:2009:IPA

[LRS09] Thomas J. Loredo, John Rice, and Michael L. Stein. Introduction to papers on astrostatistics. Annals of Applied Statistics, 3(1):1–5, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1239888360.

Luss:2012:ERI

[LRS12] Ronny Luss, Saharon Rosset, and Moni Shahar. Efficient regularized isotonic regression with application to gene–gene interaction search. Annals of Applied Statistics, 6(1):253– 283, March 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1331043396. See correction [LRS15].

Luss:2015:CER

[LRS15] Ronny Luss, Saharon Rosset, and Moni Shahar. Correction: Efficient regularized isotonic regression with application to gene-gene interaction search. Annals of Applied Statistics, 9(4):2266–2267, December 2015. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1453994200. See [LRS12].

Levina:2008:SEL

[LRZ08] Elizaveta Levina, Adam Rothman, and Ji Zhu. Sparse estimation of large covariance matrices via a nested Lasso penalty. Annals of Applied Statistics, 2(1):245–263, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1206367820.

Li:2018:SMG

[LS18] Meng Li and Armin Schwartzman. Standardization of multivariate Gaussian mixture models and background adjustment of PET images in brain oncology. Annals of Applied Statistics, 12(4):2197–2227, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078042.

Liang:2022:SML

[LS22] Jane W. Liang and ’Saunak Sen. Sparse matrix linear models for structured high-throughput data. Annals of Applied Statistics, 16(1):169–192, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Sparse-matrix-
linear-models-for-structured-high-throughput-data/
10.1214/21-AOAS1444.full.

Lahiri:2012:GBM

[LSAR12] S. N. Lahiri, C. Spiegelman, J. Appiah, and L. Rilett. Gap bootstrap methods for massive data sets with an application to transportation engineering. Annals of Applied Statistics, 6(4):1552–1587, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629051.

Lin:2015:MRF

[LSL^{+}15] Zhixiang Lin, Stephan J. Sanders, Mingfeng Li, Nenad Sestan, Matthew W. State, and Hongyu Zhao. A Markov random field-based approach to characterizing human brain development using spatial–temporal transcriptome data. Annals of Applied Statistics, 9(1):429–451, March 2015. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226099.

Lockwood:2015:ICE

[LSM15] J. R. Lockwood, Terrance D. Savitsky, and Daniel F. Mc-Caffrey. Inferring constructs of effective teaching from classroom observations: an application of Bayesian exploratory factor analysis without restrictions. Annals of Applied Statistics, 9(3):1484–1509, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1446488748.

Lopes:2012:MVU

[LSS^{+}12] Hedibert F. Lopes, Alexandra M. Schmidt, Esther Salazar, Mariana Gómez, and Marcel Achkar. Measuring the vulnerability of the Uruguayan population to vector-borne diseases via spatially hierarchical factor models. Annals of Applied Statistics, 6(1):284–303, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1331043397.

Li:2020:RRS

[LSS^{+}20] Qian Li, John Shamshoian, Damla Sentürk, Catherine Sugar, Shafali Jeste, Charlotte DiStefano, and Donatello Telesca. Region-referenced spectral power dynamics of EEG signals: a hierarchical modeling approach. Annals of Applied Statistics, 14(4):2053–2068, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Region-referenced-spectral-power-dynamics-
of-EEG-signals--A/10.1214/20-AOAS1374.full.

Lin:2022:BHR

[LSY^{+}22] Ruitao Lin, Haolun Shi, Guosheng Yin, Peter F. Thall, Ying Yuan, and Christopher R. Flowers. Bayesian hierarchical random-effects meta-analysis and design of phase I clinical trials. Annals of Applied Statistics, 16(4):2481–2504, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Bayesian-hierarchical-random-effects-meta-
analysis-and-design-of-phase/10.1214/22-AOAS1600.
full.

Liu:2022:BSN

[LSZL22] Yiwen Liu, Xiaoxiao Sun, Wenxuan Zhong, and Bing Li. Bscaling: a novel nonparametric data fusion method. Annals of Applied Statistics, 16(3):1292–1312, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/B-scaling-A-
novel-nonparametric-data-fusion-method/10.1214/21-
AOAS1537.full.

Li:2011:AWS

[LT11] Jia Li and George C. Tseng. An adaptively weighted statistic for detecting differential gene expression when combining multiple transcriptomic studies. Annals of Applied Statistics, 5(2A):994–1019, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1310562214.

Liu:2012:SRM

[LT12] Hai Liu and Wanzhu Tu. A semiparametric regression model for paired longitudinal outcomes with application in childhood blood pressure development. Annals of Applied Statistics, 6(4):1861–1882, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629063.

Lee:2019:BSJ

[LTL19] Juhee Lee, Peter F. Thall, and Steven H. Lin. Bayesian semiparametric joint regression analysis of recurrent adverse events and survival in esophageal cancer patients. Annals of Applied Statistics, 13(1):221–247, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1554861647.

Leday:2013:MAB

[LvdVvWvdW13] Gwenaël G. R. Leday, Aad W. van der Vaart, Wessel N. van Wieringen, and Mark A. van de Wiel. Modeling association between DNA copy number and gene expression with constrained piecewise linear regression splines. Annals of Applied Statistics, 7(2):823–845, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338469.

Liu:2017:GMD

[LW17] Shuyi Liu and Wei Wu. Generalized Mahalanobis depth in point process and its application in neural coding. Annals of Applied Statistics, 11(2):992–1010, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1500537732.

Luo:2018:NBL

[LW18] Xiangyu Luo and Yingying Wei. Nonparametric Bayesian learning of heterogeneous dynamic transcription factor networks. Annals of Applied Statistics, 12(3):1749–1772, September 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1536652973.

Liu:2016:VSP

[LWFW16] Ying Liu, Yuanjia Wang, Yang Feng, and Melanie M. Wall. Variable selection and prediction with incomplete highdimensional data. Annals of Applied Statistics, 10(1):418– 450, March 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1458909922.

Li:2015:BGL

[LWLW15] Jiahan Li, Zhong Wang, Runze Li, and Rongling Wu. Bayesian group Lasso for nonparametric varying-coefficient models with application to functional genome-wide association studies. Annals of Applied Statistics, 9(2):640–664, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1437397105.

Li:2019:BMI

[LWLX19] Qiwei Li, Xinlei Wang, Faming Liang, and Guanghua Xiao. A Bayesian mark interaction model for analysis of tumor pathology images. Annals of Applied Statistics, 13 (3):1708–1732, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277769.

Liu:2017:NEA

[LWSP17] Binghui Liu, Chong Wu, Xiaotong Shen, and Wei Pan. A novel and efficient algorithm for de novo discovery of mutated driver pathways in cancer. Annals of Applied Statistics, 11(3):1481–1512, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168837.

Liu:2019:HBM

[LWZ19] Yiyi Liu, Joshua L. Warren, and Hongyu Zhao. A hierarchical Bayesian model for single-cell clustering using RNA-sequencing data. Annals of Applied Statistics, 13 (3):1733–1752, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277771.

Liu:2018:PTB

[LX18] Yaowu Liu and Jun Xie. Powerful test based on conditional effects for genome-wide screening. Annals of Applied Statistics, 12(1):567–585, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564484.

Lai:2011:MVP

[LXC11] Tze Leung Lai, Haipeng Xing, and Zehao Chen. Mean– variance portfolio optimization when means and covariances are unknown. Annals of Applied Statistics, 5(2A):798– 823, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562206.

Lin:2016:BAC

[LY16] Ruitao Lin and Guosheng Yin. Bootstrap aggregating continual reassessment method for dose finding in drugcombination trials. Annals of Applied Statistics, 10(4):2349– 2376, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606863.

Liu:2022:PCW

[LYBA22] Bingjie Liu, Xubo Yue, Eunshin Byon, and Raed Al Kontar. Parameter calibration in wake effect simulation model with stochastic gradient descent and stratified sampling. Annals of Applied Statistics, 16(3):1795–1821, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Parameter-calibration-in-wake-effect-simulation-
model-with-stochastic-gradient/10.1214/21-AOAS1567.
full.

Liu:2016:SMA

[LYH^{+}16] Xiao Liu, Kyongmin Yeo, Youngdeok Hwang, Jitendra Singh, and Jayant Kalagnanam. A statistical modeling approach for air quality data based on physical dispersion processes and its application to ozone modeling. Annals of Applied Statistics, 10(2):756–785, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199892.

Lyons:2008:OSI

[Lyo08] Louis Lyons. Open statistical issues in particle physics. Annals of Applied Statistics, 2(3):887–915, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1223908045.

Larsen:2022:SCA

[LYRR22] Alexandra Larsen, Shu Yang, Brian J. Reich, and Ana G. Rappold. A spatial causal analysis of wildland firecontributed. Annals of Applied Statistics, 16(4):2714–2731, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

16/issue-4/A-spatial-causal-analysis-of-wildland-
fire-contributed-PM25-using/10.1214/22-AOAS1610.full.

Liu:2013:BDA

[LYY13] Suyu Liu, Guosheng Yin, and Ying Yuan. Bayesian data augmentation dose finding with continual reassessment method and delayed toxicity. Annals of Applied Statistics, 7(4):2138– 2156, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823313.

Li:2023:PEE

[LYY^{+}23] Yang Li, Haoyu Yang, Haochen Yu, Hanwen Huang, and Ye Shen. Penalized estimating equations for generalized linear models with multiple imputation. Annals of Applied Statistics, 17(3):2345–2363, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Penalized-estimating-
equations-for-generalized-linear-models-with-multiple-
imputation/10.1214/22-AOAS1721.full.

Loh:2007:ASC

[LZ07] Ji Meng Loh and Zhengyuan Zhu. Accounting for spatial correlation in the scan statistic. Annals of Applied Statistics, 1(2):560–584, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1196438031.

Luo:2011:BHM

[LZ11] Ruiyan Luo and Hongyu Zhao. Bayesian hierarchical modeling for signaling pathway inference from single cell interventional data. Annals of Applied Statistics, 5(2A):725– 745, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562203.

Loh:2013:RTL

[LZ13] Wei-Yin Loh and Wei Zheng. Regression trees for longitudinal and multiresponse data. Annals of Applied Statistics, 7(1):495–522, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1365527208.

Li:2021:TQR

[LZ21] Cai Li and Heping Zhang. Tensor quantile regression with application to association between neuroimages and human intelligence. Annals of Applied Statistics, 15(3):1455–1477, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Tensor-quantile-regression-with-application-
to-association-between-neuroimages-and/10.1214/21-
AOAS1475.full.

Ling:2021:ZIQ

[LZCW21] Wodan Ling, Wenfei Zhang, Bin Cheng, and Ying Wei. Zero-inflated quantile rank-score based test (ZIQRank) with application to scRNA-seq differential gene expression analysis. Annals of Applied Statistics, 15(4):1673–1696, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Zero-inflated-quantile-rank-score-based-
test-ZIQRank-with-application/10.1214/21-AOAS1442.
full.

Lee:2015:BBF

[LZK^{+}15] Eunjee Lee, Hongtu Zhu, Dehan Kong, Yalin Wang, Kelly Sullivan Giovanello, Joseph G. Ibrahim, and for the Alzheimer’s Disease Neuroimaging Initiative. BFLCRM: a Bayesian functional linear Cox regression model for predicting time to conversion to Alzheimer’s disease. Annals of Applied Statistics, 9(4):2153–2178, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453994196.

Li:2014:FAD

[LZLW14] Jiahan Li, Wei Zhong, Runze Li, and Rongling Wu. A fast algorithm for detecting gene–gene interactions in genome-wide association studies. Annals of Applied Statistics, 8(4):2292– 2318, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001744.

Liu:2016:MTU

[LZP16] Jie Liu, Chunming Zhang, and David Page. Multiple testing under dependence via graphical models. Annals of Applied Statistics, 10(3):1699–1724, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1475069624.

Liu:2016:BDF

[LZTB16] Yang Liu, James V. Zidek, Andrew W. Trites, and Brian C. Battaile. Bayesian data fusion approaches to predicting spatial tracks: Application to marine mammals. Annals of Applied Statistics, 10(3):1517–1546, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1475069617.

Li:2015:SBV

[LZW^{+}15] Fan Li, Tingting Zhang, Quanli Wang, Marlen Z. Gonzalez, Erin L. Maresh, and James A. Coan. Spatial Bayesian variable selection and grouping for high-dimensional scalaron-image regression. Annals of Applied Statistics, 9(2):687– 713, June 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1437397107.

Li:2018:MJM

[LZZL18] Wei Vivian Li, Anqi Zhao, Shihua Zhang, and Jingyi Jessica Li. MSIQ: Joint modeling of multiple RNA-seq samples for accurate isoform quantification. Annals of Applied Statistics, 12(1):510–539, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564482.

Milanzi:2014:PSS

[MAB^{+}14] Elasma Milanzi, Ariel Alonso, Christophe Buyck, Geert Molenberghs, and Luc Bijnens. A permutational-splitting sample procedure to quantify expert opinion on clusters of chemical compounds using high-dimensional data. Annals of Applied Statistics, 8(4):2319–2335, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1419001745.

Macnaughton:2009:LE

[Mac09] Donald B. Macnaughton. Letter to the Editor. Annals of Applied Statistics, 3(1):489, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1239888379.

MacDonald:2020:LEF

[Mac20a] Iain L. MacDonald. Letter to the Editor: Fitting a folded normal distribution without EM. Annals of Applied Statistics, 14(4):2096–2098, December 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-4/Letter-to-the-
Editor--Fitting-a-folded-normal-distribution/10.1214/

20-AOAS1410.full. See rejoinder [Mac20b].

MacDonald:2020:RFF

[Mac20b] Iain L. MacDonald. Rejoinder: Fitting a folded normal distribution without EM. Annals of Applied Statistics, 4 (2101):2099–2100, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-14/issue-4/Rejoinder-Fitting-a-folded-normal-
distribution-without-EM/10.1214/20-AOAS1412.full. See

[Mac20a].

Malmberg:2008:IFC

[MAE^{+}08] Anders Malmberg, Avelino Arellano, David P. Edwards, Natasha Flyer, Doug Nychka, and Christopher Wikle. Interpolating fields of carbon monoxide data using a hybrid statistical-physical model. Annals of Applied Statistics, 2 (4):1231–1248, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1231424208.

Miller:2022:MID

[MAL^{+}22] Andrew C. Miller, Lauren Anderson, Boris Leistedt, John P. Cunningham, David W. Hogg, and David M. Blei. Mapping interstellar dust with Gaussian processes. Annals of Applied Statistics, 16(4):2672–2692, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Mapping-interstellar-dust-with-Gaussian-
processes/10.1214/22-AOAS1608.full.

Monti:2017:LPS

[MAM17] Ricardo Pio Monti, Christoforos Anagnostopoulos, and Giovanni Montana. Learning population and subject-specific brain connectivity networks via mixed neighborhood selection. Annals of Applied Statistics, 11(4):2142–2164, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1514430280.

Mark:2008:GFS

[Mar08] Steven D. Mark. A general formulation for standardization of rates as a method to control confounding by measured and unmeasured disease risk factors. Annals of Applied Statistics, 2(3):1103–1122, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1223908054.

Mastrantonio:2022:MAM

[Mas22] Gianluca Mastrantonio. Modeling animal movement with directional persistence and attractive points. Annals of Applied Statistics, 16(3):2030–2053, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Modeling-animal-
movement-with-directional-persistence-and-attractive-
points/10.1214/21-AOAS1584.full.

Mandel:2013:VFE

[MAZM13] Micha Mandel, Manor Askenazi, Yi Zhang, and Jarrod A. Marto. Variance function estimation in quantitative mass spectrometry with application to iTRAQ labeling. Annals of Applied Statistics, 7(1):1–24, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527188.

Meinshausen:2008:DTA

[MB08] Nicolai Meinshausen and Peter Bühlmann. Discussion of: “Treelets — An adaptive multi-scale basis for sparse unordered data”. Annals of Applied Statistics, 2(2):478– 481, June 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1215118521. See [LNW08b].

Menchetti:2022:EEP

[MB22] Fiammetta Menchetti and Iavor Bojinov. Estimating the effectiveness of permanent price reductions for competing products using multivariate Bayesian structural time series models. Annals of Applied Statistics, 16(1):414–435, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Estimating-the-effectiveness-of-permanent-
price-reductions-for-competing-products/10.1214/21-
AOAS1498.full.

Mukhopadhyay:2011:BCC

[MBD11] Sabyasachi Mukhopadhyay, Sourabh Bhattacharya, and Kajal Dihidar. On Bayesian “central clustering”: Application to landscape classification of Western Ghats. Annals of Applied Statistics, 5(3):1948–1977, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1318514291.

McClintock:2014:PMC

[MBDL14] Brett T. McClintock, Larissa L. Bailey, Brian P. Dreher, and William A. Link. Probit models for capture–recapture data subject to imperfect detection, individual heterogeneity and misidentification. Annals of Applied Statistics, 8(4):2461– 2484, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001751.

Martinez-Beneito:2011:BJR

[MBGDS11] Miguel A. Martinez-Beneito, Gonzalo García-Donato, and Diego Salmerón. A Bayesian joinpoint regression model with an unknown number of break-points. Annals of Applied Statistics, 5(3):2150–2168, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1318514299.

Morris:2011:AAQ

[MBH^{+}11] Jeffrey S. Morris, Veerabhadran Baladandayuthapani, Richard C. Herrick, Pietro Sanna, and Howard Gutstein. Automated analysis of quantitative image data using isomorphic functional mixed models, with application to proteomics data. Annals of Applied Statistics, 5(2A):894–923, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562210.

Mohammed:2021:RRA

[MBK^{+}21] Shariq Mohammed, Karthik Bharath, Sebastian Kurtek, Arvind Rao, and Veerabhadran Baladandayuthapani. RA-DIOHEAD: Radiogenomic analysis incorporating tumor heterogeneity in imaging through densities. Annals of Applied Statistics, 15(4):1808–1830, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/RADIOHEAD--Radiogenomic-
analysis-incorporating-tumor-heterogeneity-in-imaging-
through/10.1214/21-AOAS1458.full.

Maruotti:2017:DMF

[MBL^{+}17] Antonello Maruotti, Jan Bulla, Francesco Lagona, Marco Picone, and Francesca Martella. Dynamic mixtures of factor analyzers to characterize multivariate air pollutant exposures. Annals of Applied Statistics, 11(3):1617–1648, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1507168842.

Meinshausen:2009:EBS

[MBR09] Nicolai Meinshausen, Peter Bickel, and John Rice. Efficient blind search: Optimal power of detection under computational cost constraints. Annals of Applied Statistics, 3(1): 38–60, March 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1239888362.

Marchetti-Bowick:2019:PRM

[MBYWX19] Micol Marchetti-Bowick, Yaoliang Yu, Wei Wu, and Eric P. Xing. A penalized regression model for the joint estimation of eQTL associations and gene network structure. Annals of Applied Statistics, 13(1):248–270, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1554861648.

Matechou:2017:MIM

[MC17] Eleni Matechou and François Caron. Modelling individual migration patterns using a Bayesian nonparametric approach for capture-recapture data. Annals of Applied Statistics, 11(1):21–40, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616870.

Merl:2009:IAC

[MCCW09] Daniel Merl, Julia Ling-Yu Chen, Jen-Tsan Chi, and Mike West. An integrative analysis of cancer gene expression studies using Bayesian latent factor modeling. Annals of Applied Statistics, 3(4):1675–1694, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453959.

McElroy:2009:ITM

[McE09] Tucker McElroy. Incompatibility of trends in multi-year estimates from the American community survey. Annals of Applied Statistics, 3(4):1493–1504, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453949.

Morvan:2021:PNT

[MDGM21] Marie Morvan, Emilie Devijver, Madison Giacofci, and Valérie Monbet. Prediction of the NASH through penalized mixture of logistic regression models. Annals of Applied Statistics, 15(2):952–970, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Prediction-of-
the-NASH-through-penalized-mixture-of-logistic-regression/
10.1214/20-AOAS1409.full.

Ma:2021:MPE

[MDP21] Chuoxin Ma, Hongsheng Dai, and Jianxin Pan. Modeling past event feedback through biomarker dynamics in the multistate event analysis for cardiovascular disease data. Annals of Applied Statistics, 15(3):1308–1328, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Modeling-past-event-feedback-through-biomarker-
dynamics-in-the-multistate/10.1214/21-AOAS1445.full.

Murphy:2010:VSU

[MDR10] Thomas Brendan Murphy, Nema Dean, and Adrian E. Raftery. Variable selection and updating in model-based discriminant analysis for high dimensional data with food authenticity applications. Annals of Applied Statistics, 4(1): 396–421, March 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1273584460.

Moran:2021:BJM

[MDWH21] Kelly R. Moran, David Dunson, Matthew W. Wheeler, and Amy H. Herring. Bayesian joint modeling of chemical structure and dose response curves. Annals of Applied Statistics, 15(3):1405–1430, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-3/Bayesian-joint-
modeling-of-chemical-structure-and-dose-response-
curves/10.1214/21-AOAS1461.full.

McLaughlin:2018:EAP

[ME18] Katherine R. McLaughlin and Joshua D. EmBree. Empirical assessment of programs to promote collaboration: a network model approach. Annals of Applied Statistics, 12(1): 654–682, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564488.

Meinshausen:2010:NH

[Mei10] Nicolai Meinshausen. Node harvest. Annals of Applied Statistics, 4(4):2049–2072, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167809.

Meng:2018:SPP

[Men18] Xiao-Li Meng. Statistical paradises and paradoxes in big data (I): Law of large populations, big data paradox, and the 2016 US presidential election. Annals of Applied Statistics, 12(2): 685–726, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743473.

Meyer:2008:IUS

[Mey08] Mary C. Meyer. Inference using shape-restricted regression splines. Annals of Applied Statistics, 2(3):1013–1033, September 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1223908050.

Mardia:2013:BAS

[MFB^{+}13] Kanti V. Mardia, Christopher J. Fallaize, Stuart Barber, Richard M. Jackson, and Douglas L. Theobald. Bayesian alignment of similarity shapes. Annals of Applied Statistics, 7(2):989–1009, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1372338476.

Moreira:2022:APO

[MG22] Guido A. Moreira and Dani Gamerman. Analysis of presenceonly data via exact Bayes, with model and effects identification. Annals of Applied Statistics, 16(3):1848–1867, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Analysis-of-presence-only-data-via-exact-
Bayes-with-model/10.1214/21-AOAS1569.full.

McParland:2014:CSA

[MGM^{+}14] Damien McParland, Isobel Claire Gormley, Tyler H. McCormick, Samuel J. Clark, Chodziwadziwa Whiteson Kabudula, and Mark A. Collinson. Clustering South African households based on their asset status using latent variable models. Annals of Applied Statistics, 8(2):747–776, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229513.

Mastrantonio:2019:NFL

[MGMB19] Gianluca Mastrantonio, Clara Grazian, Sara Mancinelli, and Enrico Bibbona. New formulation of the logistic-Gaussian process to analyze trajectory tracking data. Annals of Applied Statistics, 13(4):2483–2508, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1574910052.

Medous:2023:MOI

[MGRG^{+}23] Estelle Medous, Camelia Goga, Anne Ruiz-Gazen, Jean-François Beaumont, Alain Dessertaine, and Pauline Puech. Many-to-one indirect sampling with application to the French postal traffic estimation. Annals of Applied Statistics, 17(1): 838–859, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Many-to-One-indirect-sampling-with-application-
to-the-French/10.1214/22-AOAS1653.full.

McDavid:2019:GMZ

[MGSD19] Andrew McDavid, Raphael Gottardo, Noah Simon, and Mathias Drton. Graphical models for zero-inflated single cell gene expression. Annals of Applied Statistics, 13(2):848– 873, June 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1560758430.

McLain:2021:LBS

[MGTZ21] Alexander C. McLain, Siyuan Guo, Marie Thoma, and Jiajia Zhang. Length-biased semicompeting risks models for cross-sectional data: an application to current duration of pregnancy attempt data. Annals of Applied Statistics, 15(2): 1054–1067, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Length-biased-semicompeting-risks-models-
for-cross-sectional-data/10.1214/20-AOAS1428.full.

Meyer:2014:PLM

[MH14] Sebastian Meyer and Leonhard Held. Power-law models for infectious disease spread. Annals of Applied Statistics, 8(3):1612–1639, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1414091227.

Metelli:2019:BNE

[MH19] Silvia Metelli and Nicholas Heard. On Bayesian new edge prediction and anomaly detection in computer networks. Annals of Applied Statistics, 13(4):2586–2610, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1574910056.

Martinez:2009:UMS

[MHB^{+}09] Josue G. Martinez, Jianhua Z. Huang, Robert C. Burghardt, Rola Barhoumi, and Raymond J. Carroll. Use of multiple singular value decompositions to analyze complex intracellular calcium ion signals. Annals of Applied Statistics, 3(4):1467– 1492, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453948.

Murray:2015:CNF

[MHC15] Thomas A. Murray, Brian P. Hobbs, and Bradley P. Carlin. Combining nonexchangeable functional or survival data sources in oncology using generalized mixture commensurate priors. Annals of Applied Statistics, 9(3):1549–1570, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488751.

Mankad:2018:SSP

[MHG18] Shawn Mankad, Shengli Hu, and Anandasivam Gopal. Single stage prediction with embedded topic modeling of online reviews for mobile app management. Annals of Applied Statistics, 12(4):2279–2311, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078045.

Messick:2017:MSM

[MHH17] Rachel M. Messick, Matthew J. Heaton, and Neil Hansen. Multivariate spatial mapping of soil water holding capacity with spatially varying cross-correlations. Annals of Applied Statistics, 11(1):69–92, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616872.

Morris:2022:SNR

[MHK22] Emily L. Morris, Kevin He, and Jian Kang. Scalar on network regression via boosting. Annals of Applied Statistics, 16(4):2755–2773, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Scalar-on-network-
regression-via-boosting/10.1214/22-AOAS1612.full.

Masci:2022:SMM

[MIP22] Chiara Masci, Francesca Ieva, and Anna Maria Paganoni. Semiparametric multinomial mixed-effects models: a university students profiling tool. Annals of Applied Statistics, 16(3):1608–1632, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-16/issue-3/Semiparametric-multinomial-mixed-
effects-models--A-university-students-profiling/10.
1214/21-AOAS1559.full.

McCarthy:2016:PWD

[MJ16] Daniel McCarthy and Shane T. Jensen. Power-weighted densities for time series data. Annals of Applied Statistics, 10(1): 305–334, March 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1458909918.

Messier:2021:SPS

[MK21] Kyle P. Messier and Matthias Katzfuss. Scalable penalized spatiotemporal land-use regression for ground-level nitrogen dioxide. Annals of Applied Statistics, 15(2):688– 710, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Scalable-penalized-spatiotemporal-land-
use-regression-for-ground-level-nitrogen/10.1214/
20-AOAS1422.full.

Ma:2021:CMA

[MKKN21] Xiuyu Ma, Keegan Korthauer, Christina Kendziorski, and Michael A. Newton. A compositional model to assess expression changes from single-cell RNA-seq data. Annals of Applied Statistics, 15(2):880–901, June 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-2/A-
compositional-model-to-assess-expression-changes-
from-single-cell/10.1214/20-AOAS1423.full.

Mukhopadhyay:2023:EPE

[MKM23] Sabyasachi Mukhopadhyay, Wreetabrata Kar, and Gourab Mukherjee. Estimating promotion effects in email marketing using a large-scale cross-classified Bayesian joint model for nested imbalanced data. Annals of Applied Statistics, 17(1): 476–497, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Estimating-promotion-effects-in-email-
marketing-using-a-large-scale/10.1214/22-AOAS1638.
full.

Mitani:2022:ADU

[MKN22] Aya A. Mitani, Elizabeth K. Kaye, and Kerrie P. Nelson. Accounting for drop-out using inverse probability censoring weights in longitudinal clustered data with informative cluster size. Annals of Applied Statistics, 16(1):596– 611, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Accounting-for-drop-out-using-inverse-
probability-censoring-weights-in/10.1214/21-AOAS1518.
full.

Messer:2014:MFT

[MKS^{+}14] Michael Messer, Marietta Kirchner, Julia Schiemann, Jochen Roeper, Ralph Neininger, and Gaby Schneider. A multiple filter test for the detection of rate changes in renewal processes with varying variance. Annals of Applied Statistics, 8(4):2027–2067, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1419001734.

McCaffrey:2011:MDV

[ML11] Daniel F. McCaffrey and J. R. Lockwood. Missing data in value-added modeling of teacher effects. Annals of Applied Statistics, 5(2A):773–797, June 2011. CODEN ????

ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562205.

Mayrink:2013:SLF

[ML13] Vinicius Diniz Mayrink and Joseph Edward Lucas. Sparse latent factor models with interactions: Analysis of gene expression data. Annals of Applied Statistics, 7(2):799–822, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1372338468.

Mercatanti:2014:DDC

[ML14] Andrea Mercatanti and Fan Li. Do debit cards increase household spending? Evidence from a semiparametric causal analysis of a survey. Annals of Applied Statistics, 8(4):2485– 2508, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001752.

Ma:2023:TDM

[ML23] Siyuan Ma and Hongzhe Li. A tensor decomposition model for longitudinal microbiome studies. Annals of Applied Statistics, 17(2):1105–1126, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-2/A-
tensor-decomposition-model-for-longitudinal-microbiome-
studies/10.1214/22-AOAS1661.full.

Meister:2013:LND

[MLCW13] Arwen Meister, Ye Henry Li, Bokyung Choi, and Wing Hung Wong. Learning a nonlinear dynamical system model of gene regulation: a perturbed steady-state approach. Annals of Applied Statistics, 7(3):1311–1333, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1380804797.

Mews:2022:MCR

[MLKQ22] Sina Mews, Roland Langrock, Ruth King, and Nicola Quick. Multistate capture-recapture models for irregularly sampled data. Annals of Applied Statistics, 16(2):982– 998, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Multistate-capturerecapture-models-for-
irregularly-sampled-data/10.1214/21-AOAS1528.full.

Mattei:2013:EMO

[MLM13] Alessandra Mattei, Fan Li, and Fabrizia Mealli. Exploiting multiple outcomes in Bayesian principal stratification analysis with application to the evaluation of a job training program. Annals of Applied Statistics, 7(4):2336–2360, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823322.

Mastrantonio:2019:HMS

[MLP^{+}19] Gianluca Mastrantonio, Giovanna Jona Lasinio, Alessio Pollice, Giulia Capotorti, Lorenzo Teodonio, Giulio Genova, and Carlo Blasi. A hierarchical multivariate spatio-temporal model for clustered climate data with annual cycles. Annals of Applied Statistics, 13(2):797–823, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1560758428.

Moon:2023:UPH

[MLX23] Chul Moon, Qiwei Li, and Guanghua Xiao. Using persistent homology topological features to characterize medical images: Case studies on lung and brain cancers. Annals of Applied Statistics, 17(3):2192–2211, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Using-persistent-homology-topological-
features-to-characterize-medical-images/10.1214/22-
AOAS1714.full.

Manski:2008:SSD

[MM08] Charles F. Manski and Francesca Molinari. Skip sequencing: a decision problem in questionnaire design. Annals of Applied Statistics, 2(1):264–285, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1206367821.

McIntyre:2011:DSA

[MM11] Stephen McIntyre and Ross McKitrick. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):56–60, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715174. See [MW11b].

Mankad:2015:AML

[MM15] Shawn Mankad and George Michailidis. Analysis of multiview legislative networks with structured matrix factorization: Does Twitter influence translate to the real world? Annals of Applied Statistics, 9(4):1950–1972, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1453993100.

Mao:2022:DTM

[MM22] Jialiang Mao and Li Ma. Dirichlet-tree multinomial mixtures for clustering microbiome compositions. Annals of Applied Statistics, 16(3):1476–1499, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Dirichlet-tree-
multinomial-mixtures-for-clustering-microbiome-compositions/
10.1214/21-AOAS1552.full.

Mohler:2020:HBT

[MMBL20] George Mohler, Erin McGrath, Cody Buntain, and Gary LaFree. Hawkes binomial topic model with applications to coupled conflict — Twitter data. Annals of Applied Statistics, 14(4):1984–2002, December 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-4/Hawkes-binomial-
topic-model-with-applications-to-coupled-conflict-
Twitter/10.1214/20-AOAS1352.full.

Meyer:2022:OPF

[MMGC22] Mark J. Meyer, Jeffrey S. Morris, Regina Paxton Gazes, and Brent A. Coull. Ordinal probit functional outcome regression with application to computer-use behavior in rhesus monkeys. Annals of Applied Statistics, 16(1):537– 550, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Ordinal-probit-functional-outcome-regression-
with-application-to-computer-use/10.1214/21-AOAS1513.
full.

McDonald:2021:MSS

[MMGR21] Daniel J. McDonald, Michael McBride, Yupeng Gu, and Christopher Raphael. Markov-switching state space models for uncovering musical interpretation. Annals of Applied Statistics, 15(3):1147–1170, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-3/Markov-switching-
state-space-models-for-uncovering-musical-interpretation/
10.1214/21-AOAS1457.full.

Matechou:2016:OMR

[MMM^{+}16] Eleni Matechou, Rachel S. McCrea, Byron J. T. Morgan, Darryn J. Nash, and Richard A. Griffiths. Open models for removal data. Annals of Applied Statistics, 10(3):1572–1589, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069619.

Matteson:2011:FEM

[MMWH11] David S. Matteson, Mathew W. McLean, Dawn B. Woodard, and Shane G. Henderson. Forecasting emergency medical service call arrival rates. Annals of Applied Statistics, 5(2B): 1379–1406, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562725.

Marass:2016:PLF

[MMY^{+}16] Francesco Marass, Florent Mouliere, Ke Yuan, Nitzan Rosenfeld, and Florian Markowetz. A phylogenetic latent feature model for clonal deconvolution. Annals of Applied Statistics, 10(4):2377–2404, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606864.

Maroulas:2015:TRI

[MN15] Vasileios Maroulas and Andreas Nebenführ. Tracking rapid intracellular movements: a Bayesian random set approach. Annals of Applied Statistics, 9(2):926–949, June 2015. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1437397118.

Muralidharan:2012:DMM

[MNB^{+}12] Omkar Muralidharan, Georges Natsoulis, John Bell, Hanlee Ji, and Nancy R. Zhang. Detecting mutations in mixed sample sequencing data using empirical Bayes. Annals of Applied Statistics, 6(3):1047–1067, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418573.

Molina:2014:SAE

[MNR14] Isabel Molina, Balgobin Nandram, and J. N. K. Rao. Small area estimation of general parameters with application to poverty indicators: a hierarchical Bayes approach. Annals of Applied Statistics, 8(2):852–885, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229517.

Mohler:2013:DEH

[Moh13a] George Mohler. Discussion of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1866–1870, December 2013. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823296. See [CW13a].

Mohler:2013:MEM

[Moh13b] George Mohler. Modeling and estimation of multi-source clustering in crime and security data. Annals of Applied Statistics, 7(3):1525–1539, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804805.

McKennan:2020:EIM

[MON20] Chris McKennan, Carole Ober, and Dan Nicolae. Estimation and inference in metabolomics with nonrandom missing data and latent factors. Annals of Applied Statistics, 14(2):789– 808, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449326.

Mishchenko:2011:EMS

[MP11] Yuriy Mishchenko and Liam Paninski. Efficient methods for sampling spike trains in networks of coupled neurons. Annals of Applied Statistics, 5(3):1893–1919, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1318514289.

Mardia:2012:MMU

[MPT12] Kanti V. Mardia, Emma M. Petty, and Charles C. Taylor. Matching markers and unlabeled configurations in protein gels. Annals of Applied Statistics, 6(3):853–869, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1346418565.

Moghaddass:2015:LSH

[MR15] Ramin Moghaddass and Cynthia Rudin. The latent state hazard model, with application to wind turbine reliability. Annals of Applied Statistics, 9(4):1823–1863, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1453993095.

Moran:2021:SSL

[MRG21] Gemma E. Moran, Veronika Rocková, and Edward I. George. Spike-and-slab lasso biclustering. Annals of Applied Statistics, 15(1):148–173, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Spike-and-slab-
Lasso-biclustering/10.1214/20-AOAS1385.full.

McCormick:2012:BHR

[MRM12] Tyler H. McCormick, Cynthia Rudin, and David Madigan. Bayesian hierarchical rule modeling for predicting medical conditions. Annals of Applied Statistics, 6(2):652–668, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1339419611.

Maltiel:2015:EPS

[MRMB15] Rachael Maltiel, Adrian E. Raftery, Tyler H. McCormick, and Aaron J. Baraff. Estimating population size using the network scale up method. Annals of Applied Statistics, 9(3):1247–1277, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1446488738.

Marino:2019:SEB

[MRSA19] Maria Francesca Marino, Maria Giovanna Ranalli, Nicola Salvati, and Marco Alfò. Semiparametric empirical best prediction for small area estimation of unemployment indicators. Annals of Applied Statistics, 13(2):1166–1197, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758442.

Mariadassou:2010:ULS

[MRV10] Mahendra Mariadassou, Stéphane Robin, and Corinne Vacher. Uncovering latent structure in valued graphs: A variational approach. Annals of Applied Statistics, 4(2):715– 742, June 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1280842137.

Mandal:2009:SOS

[MRW09] Abhyuday Mandal, Pritam Ranjan, and C. F. Jeff Wu. G-SELC: Optimization by sequential elimination of level combinations using genetic algorithms and Gaussian processes. Annals of Applied Statistics, 3(1):398–421, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888376.

Mayer:2020:DRT

[MSG^{+}20] Imke Mayer, Erik Sverdrup, Tobias Gauss, Jean-Denis Moyer, Stefan Wager, and Julie Josse. Doubly robust treatment effect estimation with missing attributes. Annals of Applied Statistics, 14(3):1409–1431, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Doubly-robust-treatment-effect-estimation-
with-missing-attributes/10.1214/20-AOAS1356.full.

Molstad:2021:CEA

[MSH21] Aaron J. Molstad, Wei Sun, and Li Hsu. A covarianceenhanced approach to multitissue joint eQTL mapping with application to transcriptome-wide association studies. Annals of Applied Statistics, 15(2):998–1016, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-2/A-
covariance-enhanced-approach-to-multitissue-joint-
eQTL-mapping-with/10.1214/20-AOAS1432.full.

McGarrity:2014:NIS

[MSJ14] K. S. McGarrity, J. Sietsma, and G. Jongbloed. Nonparametric inference in a stereological model with oriented cylinders applied to dual phase steel. Annals of Applied Statistics, 8(4):2538–2566, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1419001754.

Mulrow:2009:AUA

[MSS09] Edward Mulrow, Hee-Choon Shin, and Fritz Scheuren. Assessing uncertainty in the American Indian trust fund. Annals of Applied Statistics, 3(4):1370–1381, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1267453944.

Mannshardt-Shamseldin:2010:DEC

[MSSS^{+}10] Elizabeth C. Mannshardt-Shamseldin, Richard L. Smith, Stephan R. Sain, Linda O. Mearns, and Daniel Cooley. Downscaling extremes: a comparison of extreme value distributions in point-source and gridded precipitation data. Annals of Applied Statistics, 4(1):484–502, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1273584464.

Mi:2021:DLS

[MTZZ21] Xinlei Mi, Patrick Tighe, Fei Zou, and Baiming Zou. A deep learning semiparametric regression for adjusting complex confounding structures. Annals of Applied Statistics, 15(3):1086–1100, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-3/A-
deep-learning-semiparametric-regression-for-adjusting-
complex-confounding-structures/10.1214/21-AOAS1481.
full.

Murtagh:2008:DTA

[Mur08] Fionn Murtagh. Discussion of: “Treelets — An adaptive multi-scale basis for sparse unordered data”. Annals of Applied Statistics, 2(2):472–473, June 2008. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1215118519. See [LNW08b].

Muralidharan:2010:EBM

[Mur10] Omkar Muralidharan. An empirical Bayes mixture method for effect size and false discovery rate estimation. Annals of Applied Statistics, 4(1):422–438, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584461.

Mortera:2008:DSA

[MV08] J. Mortera and P. Vicard. Discussion of: “Statistical analysis of an archeological find”. Annals of Applied Statistics, 2(1): 91–96, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367812. See [Feu08b].

Masarotto:2012:RLA

[MV12] Guido Masarotto and Cristiano Varin. The ranking lasso and its application to sport tournaments. Annals of Applied Statistics, 6(4):1949–1970, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629066.

Manrique-Vallier:2014:LMM

[MV14] Daniel Manrique-Vallier. Longitudinal Mixed Membership trajectory models for disability survey data. Annals of Applied Statistics, 8(4):2268–2291, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1419001743.

Mishchenko:2011:BAI

[MVP11] Yuriy Mishchenko, Joshua T. Vogelstein, and Liam Paninski. A Bayesian approach for inferring neuronal connectivity from calcium fluorescent imaging data. Annals of Applied Statistics, 5(2B):1229–1261, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562720.

Mortera:2013:OOB

[MVV13] Julia Mortera, Paola Vicard, and Cecilia Vergari. Objectoriented Bayesian networks for a decision support system for antitrust enforcement. Annals of Applied Statistics, 7(2):714– 738, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338465.

McNeely:2023:DDD

[MVW^{+}23] Trey McNeely, Galen Vincent, Kimberly M. Wood, Rafael Izbicki, and Ann B. Lee. Detecting distributional differences in labeled sequence data with application to tropical cyclone satellite imagery. Annals of Applied Statistics, 17(2):1260– 1284, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Detecting-distributional-differences-in-
labeled-sequence-data-with-application-to/10.1214/
22-AOAS1668.full.

McShane:2011:R

[MW11a] Blakeley B. McShane and Abraham J. Wyner. Rejoinder. Annals of Applied Statistics, 5(1):99–123, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1300715184.

McShane:2011:SAM

[MW11b] Blakeley B. McShane and Abraham J. Wyner. A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable? Annals of Applied Statistics, 5(1):5–44, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715170. See discussion [Ber11, Kap11, DL11a, MM11, HU11, SMR11, Hol11, Sme11, NL11, CR11, WA11, Rou11, Tin11].

Mercer:2015:STS

[MWP^{+}15] Laina D. Mercer, Jon Wakefield, Athena Pantazis, Angelina M. Lutambi, Honorati Masanja, and Samuel Clark. Space-time smoothing of complex survey data: Small area estimation for child mortality. Annals of Applied Statistics, 9(4):1889–1905, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453993097.

Martin:2020:MMA

[MWW20] Bryan D. Martin, Daniela Witten, and Amy D. Willis. Modeling microbial abundances and dysbiosis with beta-binomial regression. Annals of Applied Statistics, 14(1):94–115, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002666.

McCormick:2012:LDP

[MZ12] Tyler H. McCormick and Tian Zheng. Latent demographic profile estimation in hard-to-reach groups. Annals of Applied Statistics, 6(4):1795–1813, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629060.

Mou:2022:IIP

[MZA22] Xichen Mou, Hongmei Zhang, and S. Hasan Arshad. Identifying intergenerational patterns of correlated methylation sites. Annals of Applied Statistics, 16(1):521–536, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

16/issue-1/Identifying-intergenerational-patterns-
of-correlated-methylation-sites/10.1214/21-AOAS1511.
full.

Miranda:2018:TTP

[MZI18] Michelle F. Miranda, Hongtu Zhu, and Joseph G. Ibrahim. TPRM: Tensor partition regression models with applications in imaging biomarker detection. Annals of Applied Statistics, 12(3):1422–1450, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652960.

Nagaraja:2011:AAH

[NBZ11] Chaitra H. Nagaraja, Lawrence D. Brown, and Linda H. Zhao. An autoregressive approach to house price modeling. Annals of Applied Statistics, 5(1):124–149, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1300715185.

Newell:2013:ADN

[NCHJ13] Mark A. Newell, Dianne Cook, Heike Hofmann, and Jean-Luc Jannink. An algorithm for deciding the number of clusters and validation using simulated data with application to exploring crop population structure. Annals of Applied Statistics, 7(4):1898–1916, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1387823303.

Nicosia:2017:MSC

[NDRF17] Aurélien Nicosia, Thierry Duchesne, Louis-Paul Rivest, and Daniel Fortin. A multi-state conditional logistic regression model for the analysis of animal movement. Annals of Applied Statistics, 11(3):1537–1560, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1507168839.

Nguyen:2017:SAU

[NECS17] Trang Quynh Nguyen, Cyrus Ebnesajjad, Stephen R. Cole, and Elizabeth A. Stuart. Sensitivity analysis for an unobserved moderator in RCT-to-target-population generalization of treatment effects. Annals of Applied Statistics, 11(1): 225–247, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616879. See correction [NS20].

Newton:2009:IDP

[New09] Michael A. Newton. Introducing the discussion paper by Székely and Rizzo. Annals of Applied Statistics, 3(4):1233– 1235, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453932. See [SR09b].

Niu:2019:JMC

[NH19] Xiaoyue Niu and Peter D. Hoff. Joint mean and covariance modeling of multiple health outcome measures. Annals of Applied Statistics, 13(1):321–339, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1554861651.

Neto:2010:CGM

[NKAY10] Elias Chaibub Neto, Mark P. Keller, Alan D. Attie, and Brian S. Yandell. Causal graphical models in systems genetics: a unified framework for joint inference of causal network and genetic architecture for correlated phenotypes. Annals of Applied Statistics, 4(1):320–339, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584457.

Nychka:2011:DSA

[NL11] Doug Nychka and Bo Li. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):80–82, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715179. See [MW11b].

Nethery:2019:EPA

[NMD19] Rachel C. Nethery, Fabrizia Mealli, and Francesca Dominici. Estimating population average causal effects in the presence of non-overlap: The effect of natural gas compressor station exposure on cancer mortality. Annals of Applied Statistics, 13(2):1242–1267, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758445.

Ng:2021:MSM

[NMW^{+}21] Tin Lok James Ng, Thomas Brendan Murphy, Ted Westling, Tyler H. McCormick, and Bailey Fosdick. Modeling the social media relationships of Irish politicians using a generalized latent space stochastic blockmodel. Annals of Applied Statistics, 15(4):1923–1944, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Modeling-the-
social-media-relationships-of-Irish-politicians-using-
a/10.1214/21-AOAS1483.full.

Natarajan:2012:STI

[NPM12] Loki Natarajan, Minya Pu, and Karen Messer. Statistical tests for the intersection of independent lists of genes: Sensitivity, FDR, and type I error control. Annals of Applied Statistics, 6(2):521–541, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419606.

Newton:2007:RSM

[NQdB^{+}07] Michael A. Newton, Fernando A. Quintana, Johan A. den Boon, Srikumar Sengupta, and Paul Ahlquist. Random-set methods identify distinct aspects of the enrichment signal in gene-set analysis. Annals of Applied Statistics, 1(1):85– 106, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143730.

Niezink:2017:CES

[NS17] Nynke M. D. Niezink and Tom A. B. Snijders. Co-evolution of social networks and continuous actor attributes. Annals of Applied Statistics, 11(4):1948–1973, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1514430273.

Nguyen:2020:CSA

[NS20] Trang Quynh Nguyen and Elizabeth A. Stuart. Correction: Sensitivity analysis for an unobserved moderator in RCT-to-target-population generalization of treatment effects. Annals of Applied Statistics, 14(1):518–520, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1587002685. See [NECS17].

Naf:2023:IS

[NSMM23] Jeffrey Näf, Meta-Lina Spohn, Loris Michel, and Nicolai Meinshausen. Imputation scores. Annals of Applied Statistics, 17(3):2452–2472, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Imputation-scores/
10.1214/22-AOAS1727.full.

Nobles:2014:RSA

[NSS14a] Mallory Nobles, Nicoleta Serban, and Julie Swann. Rejoinder: “Spatial accessibility of pediatric primary healthcare: Measurement and inference”. Annals of Applied Statistics, 8(4):1961–1965, December 2014. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1419001731. See [NSS14b].

Nobles:2014:SAP

[NSS14b] Mallory Nobles, Nicoleta Serban, and Julie Swann. Spatial accessibility of pediatric primary healthcare: Measurement and inference. Annals of Applied Statistics, 8(4):1922–1946, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/

euclid.aoas/1419001727. See discussion [Hat14, Hav14, Wal14] and rejoinder [NSS14a].

Nalenz:2018:TER

[NV18] Malte Nalenz and Mattias Villani. Tree ensembles with rule structured horseshoe regularization. Annals of Applied Statistics, 12(4):2379–2408, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1542078049.

Nguyen:2023:SLZ

[NvdBCR23] Thi Kim Hue Nguyen, Koen van den Berge, Monica Chiogna, and Davide Risso. Structure learning for zero-inflated counts with an application to single-cell RNA sequencing data. Annals of Applied Statistics, 17(3):2555–2573, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Structure-learning-for-zero-inflated-counts-
with-an-application-to/10.1214/23-AOAS1732.full.

Nikooienejad:2020:BVS

[NWJ20] Amir Nikooienejad, Wenyi Wang, and Valen E. Johnson. Bayesian variable selection for survival data using inverse moment priors. Annals of Applied Statistics, 14(2):809–828, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1593449327.

Niu:2012:SRA

[NZ12] Yue S. Niu and Heping Zhang. The screening and ranking algorithm to detect DNA copy number variations. Annals of Applied Statistics, 6(3):1306–1326, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418584.

Nolde:2017:EBP

[NZ17a] Natalia Nolde and Johanna F. Ziegel. Elicitability and backtesting: Perspectives for banking regulation. Annals of Applied Statistics, 11(4):1833–1874, December 2017. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1514430265. See discussion [HK17a, Sch17, Dav17, Zho17a, Kra17] and rejoinder [NZ17b].

Nolde:2017:REB

[NZ17b] Natalia Nolde and Johanna F. Ziegel. Rejoinder: “Elicitability and backtesting: Perspectives for banking regulation”. Annals of Applied Statistics, 11(4):1901–1911, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1514430271. See [NZ17a].

Nie:2013:LRM

[NZRC13] Lei Nie, Zhiwei Zhang, Daniel Rubin, and Jianxiong Chu. Likelihood reweighting methods to reduce potential bias in noninferiority trials which rely on historical data to make inference. Annals of Applied Statistics, 7(3):1796–1813, September 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1380804817.

Opitz:2022:HRB

[OBHL22] Thomas Opitz, Haakon Bakka, Raphaël Huser, and Luigi Lombardo. High-resolution Bayesian mapping of landslide hazard with unobserved trigger event. Annals of Applied Statistics, 16(3):1653–1675, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/High-resolution-
Bayesian-mapping-of-landslide-hazard-with-unobserved-
trigger/10.1214/21-AOAS1561.full.

Owen:2012:BDA

[OE12] Art B. Owen and Dean Eckles. Bootstrapping data arrays of arbitrary order. Annals of Applied Statistics, 6(3):895– 927, September 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1346418567.

OBrien:2018:ENM

[OGP^{+}18] Jonathon J. O’Brien, Harsha P. Gunawardena, Joao A. Paulo, Xian Chen, Joseph G. Ibrahim, Steven P. Gygi, and Bahjat F. Qaqish. The effects of nonignorable missing data on label-free mass spectrometry proteomics experiments. Annals of Applied Statistics, 12(4):2075–2095, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078037.

Osthus:2017:FSI

[OHC^{+}17] Dave Osthus, Kyle S. Hickmann, Petruta C. Caragea, Dave Higdon, and Sara Y. Del Valle. Forecasting seasonal influenza with a state-space SIR model. Annals of Applied Statistics, 11(1):202–224, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616878.

Orlando:2009:BPM

[OIHH09] David A. Orlando, Edwin S. Iversen, Jr., Alexander J. Hartemink, and Steven B. Haase. A branching process model for flow cytometry and budding index measurements in cell synchrony experiments. Annals of Applied Statistics, 3(4):1521–1541, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1267453951.

Oates:2014:JEM

[OKGM14] Chris J. Oates, Jim Korkola, Joe W. Gray, and Sach Mukherjee. Joint estimation of multiple related biological networks. Annals of Applied Statistics, 8(3):1892–1919, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091238.

Olshen:2010:RLB

[Ols10] Richard A. Olshen. Remembering Leo Breiman. Annals of Applied Statistics, 4(4):1644–1648, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167789.

Oates:2012:NIB

[OM12] Chris. J. Oates and Sach Mukherjee. Network inference and biological dynamics. Annals of Applied Statistics, 6 (3):1209–1235, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1346418580.

OSullivan:2014:VLM

[OMM^{+}14] Finbarr O’Sullivan, Mark Muzi, David A. Mankoff, Janet F. Eary, Alexander M. Spence, and Kenneth A. Krohn. Voxellevel mapping of tracer kinetics in PET studies: a statistical approach emphasizing tissue life tables. Annals of Applied Statistics, 8(2):1065–1094, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229526.

Owen:2009:BCV

[OP09] Art B. Owen and Patrick O. Perry. Bi-cross-validation of the SVD and the nonnegative matrix factorization. Annals of Applied Statistics, 3(2):564–594, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676186.

Ostrovnaya:2015:USM

[OSB15] Irina Ostrovnaya, Venkatraman E. Seshan, and Colin B. Begg. Using somatic mutation data to test tumors for clonal relatedness. Annals of Applied Statistics, 9(3):1533–1548, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488750.

Olives:2014:RRS

[OSL^{+}14] Casey Olives, Lianne Sheppard, Johan Lindström, Paul D. Sampson, Joel D. Kaufman, and Adam A. Szpiro. Reducedrank spatio-temporal modeling of air pollution concentrations in the Multi-Ethnic Study of Atherosclerosis and Air Pollution. Annals of Applied Statistics, 8(4):2509–2537, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001753.

Ogburn:2017:VCS

[OV17] Elizabeth L. Ogburn and Tyler J. VanderWeele. Vaccines, contagion, and social networks. Annals of Applied Statistics, 11(2):919–948, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537729.

Olhede:2011:NTS

[OW11] Sofia C. Olhede and Brandon Whitcher. Nonparametric tests of structure for high angular resolution diffusion imaging in Q-space. Annals of Applied Statistics, 5(2B):1293–1327, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562722.

Owen:2007:PB

[Owe07] Art B. Owen. The pigeonhole bootstrap. Annals of Applied Statistics, 1(2):386–411, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1196438024.

Pal:2023:SPT

[PA23] Suvra Pal and Wisdom Aselisewine. A semiparametric promotion time cure model with support vector machine.

Annals of Applied Statistics, 17(3):2680–2699, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-3/A-
semiparametric-promotion-time-cure-model-with-support-
vector-machine/10.1214/23-AOAS1741.full.

Plumlee:2021:HFH

[PACB21] Matthew Plumlee, Taylor G. Asher, Won Chang, and Matthew V. Bilskie. High-fidelity hurricane surge forecasting using emulation and sequential experiments. Annals of Applied Statistics, 15(1):460–480, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/High-fidelity-
hurricane-surge-forecasting-using-emulation-and-sequential-
experiments/10.1214/20-AOAS1398.full.

Paddock:2013:E

[Pad13] Susan M. Paddock. Editorial. Annals of Applied Statistics, 7(4):1837, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1387823294.

Paddock:2014:ESA

[Pad14] Susan M. Paddock. Editorial: Spatial accessibility of pediatric primary healthcare: Measurement and inference. Annals of Applied Statistics, 8(4):1921, December 2014. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1419001726.

Pedone:2023:SSD

[PAS23] Matteo Pedone, Amedeo Amedei, and Francesco C. Stingo. Subject-specific Dirichlet-multinomial regression for multidistrict microbiota data analysis. Annals of Applied Statistics, 17(1):539–559, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-17/issue-1/Subject-specific-Dirichlet-multinomial-
regression-for-multi-district-microbiota-data/10.
1214/22-AOAS1641.full.

Pellin:2023:THS

[PBS^{+}23] Danilo Pellin, Luca Biasco, Serena Scala, Clelia Di Serio, and Ernst C. Wit. Tracking hematopoietic stem cell evolution in a Wiskott–Aldrich clinical trial. Annals of Applied Statistics, 17(3):1841–1860, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Tracking-hematopoietic-
stem-cell-evolution-in-a-WiskottAldrich-clinical-
trial/10.1214/22-AOAS1686.full.

Peterson:2023:HCH

[PBSVS23] Adam T. Peterson, Veronica J. Berrocal, Emma V. Sanchez-Vaznaugh, and Brisa N. Sánchez. How close and how much? linking health outcomes to built environment spatial distributions. Annals of Applied Statistics, 17(2):1641– 1662, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/How-close-and-how-much-Linking-health-
outcomes-to-built/10.1214/22-AOAS1687.full.

Park:2019:LAE

[PC19] Soyoung Park and Alicia Carriquiry. Learning algorithms to evaluate forensic glass evidence. Annals of Applied Statistics, 13(2):1068–1102, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758438.

Paul:2020:RES

[PC20] Subhadeep Paul and Yuguo Chen. A random effects stochastic block model for joint community detection in multiple networks with applications to neuroimaging. Annals of Applied Statistics, 14(2):993–1029, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1593449335.

Pircalabelu:2015:FIC

[PCJW15] Eugen Pircalabelu, Gerda Claeskens, Sara Jahfari, and Lourens J. Waldorp. A focused information criterion for graphical models in fMRI connectivity with high-dimensional data. Annals of Applied Statistics, 9(4):2179–2214, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1453994197.

Papadogeorgou:2020:CER

[PD20] Georgia Papadogeorgou and Francesca Dominici. A causal exposure response function with local adjustment for confounding: Estimating health effects of exposure to low levels of ambient fine particulate matter. Annals of Applied Statistics, 14(2):850–871, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1593449329.

Petersen:2019:FET

[PDM19] Alexander Petersen, Sean Deoni, and Hans-Georg Müller. Fréchet estimation of time-varying covariance matrices from sparse data, with application to the regional co-evolution of myelination in the developing brain. Annals of Applied Statistics, 13(1):393–419, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861654.

Pirinen:2013:ECL

[PDS13] Matti Pirinen, Peter Donnelly, and Chris C. A. Spencer. Efficient computation with a linear mixed model on large-scale data sets with applications to genetic studies. Annals of Applied Statistics, 7(1):369–390, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527203.

Pan:2013:ERM

[PG13] Qing Pan and Joseph L. Gastwirth. Estimating restricted mean job tenures in semi-competing risk data compensating victims of discrimination. Annals of Applied Statistics, 7(3):1474–1496, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804803.

Phillips:2014:TDH

[PG14] Daisy Phillips and Debashis Ghosh. Testing the disjunction hypothesis using Voronoi diagrams with applications to genetics. Annals of Applied Statistics, 8(2):801–823, June 2014.

CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229515.

Parzen:2011:GLM

[PGL^{+}11] Michael Parzen, Souparno Ghosh, Stuart Lipsitz, Debajyoti Sinha, Garrett M. Fitzmaurice, Bani K. Mallick, and Joseph G. Ibrahim. A generalized linear mixed model for longitudinal binary data with a marginal logit link function. Annals of Applied Statistics, 5(1):449–467, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1300715198.

Patel:2019:HMM

[PGL^{+}19] Lekha Patel, Nils Gustafsson, Yu Lin, Raimund Ober, Ricardo Henriques, and Edward Cohen. A hidden Markov model approach to characterizing the photo-switching behavior of fluorophores. Annals of Applied Statistics, 13 (3):1397–1429, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277758.

Pardy:2018:IEL

[PGW18] Christopher Pardy, Sally Galbraith, and Susan R. Wilson. Integrative exploration of large high-dimensional datasets. Annals of Applied Statistics, 12(1):178–199, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564469.

Park:2010:ADA

[PHCM^{+}10] Cheolwoo Park, Felix Hernández-Campos, J. S. Marron, Kevin Jeffay, and F. Donelson Smith. Analysis of dependence among size, rate and duration in Internet flows. Annals of Applied Statistics, 4(1):26–52, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1273584446.

Parker:2022:CEB

[PHJ22] Paul A. Parker, Scott H. Holan, and Ryan Janicki. Computationally efficient Bayesian unit-level models for non-Gaussian data under informative sampling with application to estimation of health insurance coverage. Annals of Applied Statistics, 16(2):887–904, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-2/Computationally-
efficient-Bayesian-unit-level-models-for-non-Gaussian-
data/10.1214/21-AOAS1524.full.

Potter:2011:EWH

[PHLH11] Gail E. Potter, Mark S. Handcock, Ira M. Longini, Jr., and M. Elizabeth Halloran. Estimating within-household contact networks from egocentric data. Annals of Applied Statistics, 5(3):1816–1838, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514286.

Potter:2012:EWS

[PHLH12] Gail E. Potter, Mark S. Handcock, Ira M. Longini, and M. Elizabeth Halloran. Estimating within-school contact networks to understand influenza transmission. Annals of Applied Statistics, 6(1):1–26, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1331043386.

Papakostas:2023:DFB

[PHM^{+}23] Demetrios Papakostas, P. Richard Hahn, Jared Murray, Frank Zhou, and Joseph Gerakos. Do forecasts of bankruptcy cause bankruptcy? A machine learning sensitivity analysis. Annals of Applied Statistics, 17(1):711–739, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Do-forecasts-
of-bankruptcy-cause-bankruptcy-A-machine-learning-
sensitivity/10.1214/22-AOAS1648.full.

Powers:2015:CTA

[PHT15] Scott Powers, Trevor Hastie, and Robert Tibshirani. Customized training with an application to mass spectrometric imaging of cancer tissue. Annals of Applied Statistics, 9(4):1709–1725, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453993091.

Paddock:2011:ARG

[PHWM11] Susan M. Paddock, Sarah B. Hunter, Katherine E. Watkins, and Daniel F. McCaffrey. Analysis of rolling group therapy data using conditionally autoregressive priors. Annals of Applied Statistics, 5(2A):605–627, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562197.

Panaretos:2011:SAP

[PK11] Victor M. Panaretos and Kjell Konis. Sparse approximations of protein structure from noisy random projections. Annals of Applied Statistics, 5(4):2572–2602, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1324399607.

Pan:2018:SMA

[PK18] Karen D. H. Pan and Karen Kafadar. Statistical modeling and analysis of trace element concentrations in forensic glass evidence. Annals of Applied Statistics, 12(2):788–814, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1532743477. See [PK19].

Pan:2019:CSM

[PK19] Karen D. H. Pan and Karen Kafadar. Correction to: Statistical modeling and analysis of trace element concentrations in forensic glass evidence. Annals of Applied Statistics, 13(2): 1319–1328, June 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1560758448. See [PK18].

Park:2023:SLI

[PKGG23] Beomjo Park, Mikael Kuusela, Donata Giglio, and Alison Gray. Spatiotemporal local interpolation of global ocean heat transport using Argo floats: a debiased latent Gaussian process approach. Annals of Applied Statistics, 17(2):1491– 1520, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Spatiotemporal-local-interpolation-of-
global-ocean-heat-transport-using-Argo/10.1214/22-
AOAS1679.full.

Park:2016:IAH

[PKP16] Seunghwan Park, Jae Kwang Kim, and Sangun Park. An imputation approach for handling mixed-mode surveys. Annals of Applied Statistics, 10(2):1063–1085, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1469199905.

Proietti:2008:RTE

[PL08] Tommaso Proietti and Alessandra Luati. Real time estimation in local polynomial regression, with application to trend-cycle analysis. Annals of Applied Statistics, 2(4):1523– 1553, December 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1231424221.

Pfeffermann:2011:PSB

[PL11] Danny Pfeffermann and Victoria Landsman. Are private schools better than public schools? Appraisal for Ireland by methods for observational studies. Annals of Applied Statistics, 5(3):1726–1751, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1318514283.

Pura:2023:TMT

[PLCX23] John A. Pura, Xuechan Li, Cliburn Chan, and Jichun Xie. TEAM: a multiple testing algorithm on the aggregation tree for flow cytometry analysis. Annals of Applied Statistics, 17(1):621–640, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/TEAM--A-multiple-
testing-algorithm-on-the-aggregation-tree/10.1214/
22-AOAS1645.full.

Phipson:2016:RHE

[PLM^{+}16] Belinda Phipson, Stanley Lee, Ian J. Majewski, Warren S. Alexander, and Gordon K. Smyth. Robust hyperparameter estimation protects against hypervariable genes and improves power to detect differential expression. Annals of Applied Statistics, 10(2):946–963, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1469199900.

Peng:2008:DBC

[PM08] Jie Peng and Hans-Georg Müller. Distance-based clustering of sparsely observed stochastic processes, with applications to online auctions. Annals of Applied Statistics, 2(3):1056–1077, September 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1223908052.

Poppick:2016:TTC

[PMMS16] Andrew Poppick, David J. McInerney, Elisabeth J. Moyer, and Michael L. Stein. Temperatures in transient climates: Improved methods for simulations with evolving temporal covariances. Annals of Applied Statistics, 10(1):477–505, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909924.

Palacios:2014:BMB

[PMQW14] A. Paula Palacios, J. Miguel Marín, Emiliano J. Quinto, and Michael P. Wiper. Bayesian modeling of bacterial growth for multiple populations. Annals of Applied Statistics, 8(3):1516–1537, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1414091223.

Parsons:2022:BHM

[PNB22] Jacob Parsons, Xiaoyue Niu, and Le Bao. A Bayesian hierarchical model for combining multiple data sources in population size estimation. Annals of Applied Statistics, 16(3):1550–1562, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-16/issue-3/A-Bayesian-hierarchical-model-for-
combining-multiple-data-sources-in/10.1214/21-AOAS1556.
full.

Paul:2011:SMA

[PPB11] Debashis Paul, Jie Peng, and Prabir Burman. Semiparametric modeling of autonomous nonlinear dynamical systems with application to plant growth. Annals of Applied Statistics, 5(3):2078–2108, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1318514296.

Presanis:2014:SEE

[PPB^{+}14] Anne M. Presanis, Richard G. Pebody, Paul J. Birrell, Brian D. M. Tom, Helen K. Green, Hayley Durnall, Douglas Fleming, and Daniela De Angelis. Synthesising evidence to estimate pandemic (2009) A/H1N1 influenza severity in 2009– 2011. Annals of Applied Statistics, 8(4):2378–2403, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1419001748.

Pimentel:2018:OMM

[PPLK18] Samuel D. Pimentel, Lindsay C. Page, Matthew Lenard, and Luke Keele. Optimal multilevel matching using network flows: An application to a summer reading intervention. Annals of Applied Statistics, 12(3):1479–1505, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1536652962.

Peng:2014:TWG

[PPM14] Jie Peng, Debashis Paul, and Hans-Georg Müller. Timewarped growth processes, with applications to the modeling of boom–bust cycles in house prices. Annals of Applied Statistics, 8(3):1561–1582, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1414091225.

Page:2021:DIU

[PQR21] Garritt L. Page, Fernando A. Quintana, and Gary L. Rosner. Discovering interactions using covariate informed random partition models. Annals of Applied Statistics, 15(1): 1–21, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Discovering-interactions-using-covariate-
informed-random-partition-models/10.1214/20-AOAS1372.
full.

Percival:2011:SSR

[PRRW11] Daniel Percival, Kathryn Roeder, Roni Rosenfeld, and Larry Wasserman. Structured, sparse regression with application to HIV drug resistance. Annals of Applied Statistics, 5(2A): 628–644, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562198.

Polson:2012:GGL

[PS12] Nicholas G. Polson and James G. Scott. Good, great, or lucky? Screening for firms with sustained superior performance using heavy-tailed priors. Annals of Applied Statistics, 6(1):161–185, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1331043392.

Polson:2015:BAT

[PS15] Nicholas Polson and Vadim Sokolov. Bayesian analysis of traffic flow on interstate I-55: The LWR model. Annals of Applied Statistics, 9(4):1864–1888, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453993096.

Pannekoek:2013:CIN

[PSD13] Jeroen Pannekoek, Natalie Shlomo, and Ton De Waal. Calibrated imputation of numerical data under linear edit restrictions. Annals of Applied Statistics, 7(4):1983–2006, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823307.

Pomann:2016:LFL

[PSL^{+}16] Gina-Maria Pomann, Ana-Maria Staicu, Edgar J. Lobaton, Amanda F. Mejia, Blake E. Dewey, Daniel S. Reich, Elizabeth M. Sweeney, and Russell T. Shinohara. A lag functional linear model for prediction of magnetization transfer ratio in multiple sclerosis lesions. Annals of Applied Statistics, 10(4):2325–2348, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1483606862.

Petersen:2018:SEN

[PSW18] Ashley Petersen, Noah Simon, and Daniela Witten. SCALPEL: Extracting neurons from calcium imaging data. Annals of Applied Statistics, 12(4):2430–2456, December 2018. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078051.

Phipps:2012:AEN

[PT12] Polly Phipps and Daniell Toth. Analyzing establishment nonresponse using an interpretable regression tree model with linked administrative data. Annals of Applied Statistics, 6(2): 772–794, June 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1339419616.

Patel:2012:TPE

[PTGN12] Trina Patel, Donatello Telesca, Saji George, and André E. Nel. Toxicity profiling of engineered nanomaterials via multivariate dose-response surface modeling. Annals of Applied Statistics, 6(4):1707–1729, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1356629057.

Passino:2022:GLP

[PTH22] Francesco Sanna Passino, Melissa J. M. Turcotte, and Nicholas A. Heard. Graph link prediction in computer networks using Poisson matrix factorisation. Annals of Applied Statistics, 16(3):1313–1332, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Graph-link-prediction-
in-computer-networks-using-Poisson-matrix-factorisation/
10.1214/21-AOAS1540.full.

Purdom:2011:ADM

[Pur11] Elizabeth Purdom. Analysis of a data matrix and a graph: Metagenomic data and the phylogenetic tree. Annals of Applied Statistics, 5(4):2326–2358, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1324399597.

Porter:2012:SEH

[PW12] Michael D. Porter and Gentry White. Self-exciting hurdle models for terrorist activity. Annals of Applied Statistics, 6(1):106–124, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1331043390.

Pfister:2021:SVS

[PWP^{+}21] Niklas Pfister, Evan G. Williams, Jonas Peters, Ruedi Aebersold, and Peter Bühlmann. Stabilizing variable selection and regression. Annals of Applied Statistics, 15(3):1220–1246, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Stabilizing-variable-selection-and-regression/
10.1214/21-AOAS1487.full.

Paciorek:2009:PLS

[PYP^{+}09] Christopher J. Paciorek, Jeff D. Yanosky, Robin C. Puett, Francine Laden, and Helen H. Suh. Practical large-scale spatio-temporal modeling of particulate matter concentrations. Annals of Applied Statistics, 3(1):370–397, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888375.

Park:2019:SPC

[PZ19] Seyoung Park and Hongyu Zhao. Sparse principal component analysis with missing observations. Annals of Applied Statistics, 13(2):1016–1042, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1560758436.

Peng:2010:RMR

[PZB^{+}10] Jie Peng, Ji Zhu, Anna Bergamaschi, Wonshik Han, Dong-Young Noh, Jonathan R. Pollack, and Pei Wang. Regularized multivariate regression for identifying master predictors with application to integrative genomics study of breast cancer. Annals of Applied Statistics, 4(1):53–77, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1273584447.

Peskoe:2023:EIE

[PZSW23] Sarah B. Peskoe, Ning Zhang, Donna Spiegelman, and Molin Wang. Estimation and inference for exposure effects with latency in the Cox proportional hazards model in the presence of exposure measurement error. Annals of Applied Statistics, 17(2):1574–1591, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Estimation-and-
inference-for-exposure-effects-with-latency-in-the/
10.1214/22-AOAS1682.full.

Quick:2013:MTG

[QBC13] Harrison Quick, Sudipto Banerjee, and Bradley P. Carlin. Modeling temporal gradients in regionally aggregated California asthma hospitalization data. Annals of Applied Statistics, 7(1):154–176, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1365527194.

Qin:2021:ERE

[QDN^{+}21] Jing Qin, Geng Deng, Jing Ning, Ao Yuan, and Yu Shen. Estrogen receptor expression on breast cancer patients’ survival under shape-restricted Cox regression model. Annals of Applied Statistics, 15(3):1291–1307, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Estrogen-receptor-expression-on-breast-
cancer-patients-survival-under-shape/10.1214/21-AOAS1446.
full.

Qin:2008:ASE

[QGFL08] Li Qin, Peter B. Gilbert, Dean Follmann, and Dongfeng Li. Assessing surrogate endpoints in vaccine trials with casecohort sampling and the Cox model. Annals of Applied Statistics, 2(1):386–407, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1206367826.

Qin:2014:CIR

[QGM^{+}14] Jing Qin, Tanya P. Garcia, Yanyuan Ma, Ming-Xin Tang, Karen Marder, and Yuanjia Wang. Combining isotonic regression and EM algorithm to predict genetic risk under monotonicity constraint. Annals of Applied Statistics, 8(2): 1182–1208, June 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1404229530.

Qian:2019:FDN

[QHPD19] Yanjun Qian, Jianhua Z. Huang, Chiwoo Park, and Yu Ding. Fast dynamic nonparametric distribution tracking in electron microscopic data. Annals of Applied Statistics, 13 (3):1537–1563, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277763.

Qiu:2008:DTA

[Qiu08] Xing Qiu. Discussion of: “Treelets — An adaptive multi-scale basis for sparse unordered data”. Annals of Applied Statistics, 2(2):484–488, June 2008. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1215118523. See [LNW08b].

Qiu:2023:EDR

[QM23] Yuqi Qiu and Karen Messer. An efficient doubly-robust imputation framework for longitudinal dropout, with an application to an Alzheimer’s clinical trial. Annals of Applied Statistics, 17(3):2473–2493, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-3/An-
efficient-doubly-robust-imputation-framework-for-
longitudinal-dropout-with/10.1214/23-AOAS1728.full.

Qian:2022:LSM

[QTL^{+}22] Junyang Qian, Yosuke Tanigawa, Ruilin Li, Robert Tibshirani, Manuel A. Rivas, and Trevor Hastie. Largescale multivariate sparse regression with applications to UK Biobank. Annals of Applied Statistics, 16(3):1891–1918, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Large-scale-multivariate-sparse-regression-
with-applications-to-UK-Biobank/10.1214/21-AOAS1575.
full.

Qin:2008:NSA

[QW08] Li Qin and Yuedong Wang. Nonparametric spectral analysis with applications to seizure characterization using EEG time series. Annals of Applied Statistics, 2(4):1432–1451, December 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1231424217.

Quick:2017:MSM

[QWC17] Harrison Quick, Lance A. Waller, and Michele Casper. Multivariate spatiotemporal modeling of age-specific stroke mortality. Annals of Applied Statistics, 11(4):2165–2177, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1514430281.

Qiu:2009:SMT

[QYP09] Peihua Qiu, Rong Yang, and Michael Potegal. Statistical modeling of the time course of tantrum anger. Annals of Applied Statistics, 3(3):1013–1034, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1254773276.

Ruiz:2020:SPM

[RAB20] Francisco J. R. Ruiz, Susan Athey, and David M. Blei. SHOPPER: a probabilistic model of consumer choice with substitutes and complements. Annals of Applied Statistics, 14 (1):1–27, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002662.

Rossell:2014:QAS

[RAKS14] David Rossell, Camille Stephan-Otto Attolini, Manuel Kroiss, and Almond Stöcker. Quantifying alternative splicing from paired-end RNA-sequencing data. Annals of Applied Statistics, 8(1):309–330, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1396966288. See cor-

rigendum [RAKS15].

Rossell:2015:CQA

[RAKS15] David Rossell, Camille Stephan-Otto Attolini, Manuel Kroiss, and Almond Stöcker. Corrigendum: Quantifying alternative splicing from paired-end RNA-seq data. Annals of Applied Statistics, 9(3):1706–1707, September 2015.

CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1446488757. See [RAKS14].

Raillard:2014:MEV

[RAY14] Nicolas Raillard, Pierre Ailliot, and Jianfeng Yao. Modeling extreme values of processes observed at irregular time steps: Application to significant wave height. Annals of Applied Statistics, 8(1):622–647, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1396966301.

Rambharat:2010:SMC

[RB10a] Bhojnarine R. Rambharat and Anthony E. Brockwell. Sequential Monte Carlo pricing of American-style options under stochastic volatility models. Annals of Applied Statistics, 4(1):222–265, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1273584454.

Reich:2010:LFM

[RB10b] Brian J. Reich and Dipankar Bandyopadhyay. A latent factor model for spatial data with informative missingness. Annals of Applied Statistics, 4(1):439–459, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1273584462.

Rambharat:2011:SMC

[RB11] Bhojnarine R. Rambharat and Anthony E. Brockwell. Sequential Monte Carlo pricing of American-style options under stochastic volatility models. Annals of Applied Statistics, 5(1):604, March 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1300715204.

Rocchetti:2011:PSE

[RBB11] Irene Rocchetti, John Bunge, and Dankmar Böhning. Population size estimation based upon ratios of recapture probabilities. Annals of Applied Statistics, 5(2B):1512–1533, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562731.

Ren:2020:BME

[RBF^{+}20] Boyu Ren, Sergio Bacallado, Stefano Favaro, Tommi Vatanen, Curtis Huttenhower, and Lorenzo Trippa. Bayesian mixed effects models for zero-inflated compositions in microbiome data analysis. Annals of Applied Statistics, 14(1): 494–517, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002684.

Ren:2023:KSI

[RC23a] Zhimei Ren and Emmanuel Candès. Knockoffs with side information. Annals of Applied Statistics, 17(2):1152– 1174, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Knockoffs-with-side-information/10.1214/
22-AOAS1663.full.

Rohrbeck:2023:SFE

[RC23b] Christian Rohrbeck and Daniel Cooley. Simulating flood event sets using extremal principal components. Annals of Applied Statistics, 17(2):1333–1352, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Simulating-flood-
event-sets-using-extremal-principal-components/10.
1214/22-AOAS1672.full.

Risser:2019:NSP

[RCBB19] Mark D. Risser, Catherine A. Calder, Veronica J. Berrocal, and Candace Berrett. Nonstationary spatial prediction of soil organic carbon: Implications for stock assessment decision making. Annals of Applied Statistics, 13(1):165–188, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1554861645.

Reich:2013:EVA

[RCF^{+}13] Brian Reich, Daniel Cooley, Kristen Foley, Sergey Napelenok, and Benjamin Shaby. Extreme value analysis for evaluating ozone control strategies. Annals of Applied Statistics, 7(2): 739–762, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338466.

Ramirez-Cobo:2010:BID

[RCLWW10] Pepa Ramirez-Cobo, Rosa E. Lillo, Simon Wilson, and Michael P. Wiper. Bayesian inference for double Pareto lognormal queues. Annals of Applied Statistics, 4(3):1533–1557, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409385.

Russell:2016:DMI

[RCP^{+}16] Brook T. Russell, Daniel S. Cooley, William C. Porter, Brian J. Reich, and Colette L. Heald. Data mining to investigate the meteorological drivers for extreme ground level ozone events. Annals of Applied Statistics, 10(3):1673–1698, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069623.

Rodriguez:2014:FCN

[RD14] Abel Rodriguez and David B. Dunson. Functional clustering in nested designs: Modeling variability in reproductive epidemiology studies. Annals of Applied Statistics, 8 (3):1416–1442, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1414091219.

Ruffieux:2020:GLA

[RDH^{+}20] Hélène Ruffieux, Anthony C. Davison, Jörg Hager, Jamie Inshaw, Benjamin P. Fairfax, Sylvia Richardson, and Leonardo Bottolo. A global-local approach for detecting hotspots in multiple-response regression. Annals of Applied Statistics, 14(2):905–928, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1593449331.

Rhodes:2023:DPR

[RDL23] Grace Rhodes, Marie Davidian, and Wenbin Lu. Dynamic prediction of residual life with longitudinal covariates using long short-term memory networks. Annals of Applied Statistics, 17(3):2039–2058, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Dynamic-prediction-
of-residual-life-with-longitudinal-covariates-using-
long/10.1214/22-AOAS1706.full.

Rohrbeck:2018:EVM

[REFT18] Christian Rohrbeck, Emma F. Eastoe, Arnoldo Frigessi, and Jonathan A. Tawn. Extreme value modelling of water-related insurance claims. Annals of Applied Statistics, 12(1):246– 282, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564472.

Reich:2011:CCD

[REG^{+}11] Brian J. Reich, Jo Eidsvik, Michele Guindani, Amy J. Nail, and Alexandra M. Schmidt. A class of covariate-dependent spatiotemporal covariance functions for the analysis of daily ozone concentration. Annals of Applied Statistics, 5(4):2425– 2447, December 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1324399601.

Remillard:2009:DBD

[Rém09] Bruno Rémillard. Discussion of: Brownian distance covariance. Annals of Applied Statistics, 3(4):1295–1298, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1267453939. See [SR09b].

Reich:2007:MSB

[RF07] Brian J. Reich and Montserrat Fuentes. A multivariate semiparametric Bayesian spatial modeling framework for hurricane surface wind fields. Annals of Applied Statistics, 1(1): 249–264, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143738.

Ren:2017:CCS

[RFB17] You Ren, Emily B. Fox, and Andrew Bruce. Clustering correlated, sparse data streams to estimate a localized housing price index. Annals of Applied Statistics, 11(2):808–839, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1500537725.

Rafei:2022:RBI

[RFWE22] Ali Rafei, Carol A. C. Flannagan, Brady T. West, and Michael R. Elliott. Robust Bayesian inference for Big Data: Combining sensor-based records with traditional survey data. Annals of Applied Statistics, 16(2):1038– 1070, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Robust-Bayesian-inference-for-Big-Data-
-Combining-sensor-based/10.1214/21-AOAS1531.full.

Risk:2021:SNG

[RG21] Benjamin B. Risk and Irina Gaynanova. Simultaneous non-Gaussian component analysis (SING) for data integration in neuroimaging. Annals of Applied Statistics, 15(3):1431–1454, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Simultaneous-non-Gaussian-component-analysis-
SING-for-data-integration-in/10.1214/21-AOAS1466.
full.

Ryan:2023:SSP

[RG23] Mary M. Ryan and Daniel L. Gillen. Sequential sampling in prospective observational studies. Annals of Applied Statistics, 17(1):153–177, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Sequential-sampling-
in-prospective-observational-studies/10.1214/22-AOAS1620.
full.

Reich:2020:ISM

[RGF^{+}20] Brian J. Reich, Yawen Guan, Denis Fourches, Joshua L. Warren, Stefanie E. Sarnat, and Howard H. Chang. Integrative statistical methods for exposure mixtures and health. Annals of Applied Statistics, 14(4):1945–1963, December 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-4/Integrative-
statistical-methods-for-exposure-mixtures-and-health/
10.1214/20-AOAS1364.full.

Roy:2019:BMS

[RGPC19] Arkaprava Roy, Subhashis Ghosal, Jeffrey Prescott, and Kingshuk Roy Choudhury. Bayesian modeling of the structural connectome for studying Alzheimer’s disease. Annals of Applied Statistics, 13(3):1791–1816, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1571277774.

Rodriguez-Girondo:2018:SDC

[RGSB^{+}18] Mar Rodríguez-Girondo, Perttu Salo, Tomasz Burzykowski, Markus Perola, Jeanine Houwing-Duistermaat, and Bart Mertens. Sequential double cross-validation for assessment of added predictive ability in high-dimensional omic applications. Annals of Applied Statistics, 12(3):1655–1678, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1536652969.

Raghavan:2013:HMM

[RGT13] Vasanthan Raghavan, Aram Galstyan, and Alexander G. Tartakovsky. Hidden Markov models for the activity profile of terrorist groups. Annals of Applied Statistics, 7(4):2402– 2430, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823325.

Rubinstein:2023:BWR

[RHC23] Max Rubinstein, Amelia Haviland, and David Choi. Balancing weights for region-level analysis: The effect of medicaid expansion on the uninsurance rate among states that did not expand medicaid. Annals of Applied Statistics, 17(2):1469– 1490, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Balancing-weights-for-region-level-analysis-
-The-effect-of/10.1214/22-AOAS1678.full.

Rusch:2013:MTT

[RHHH13] Thomas Rusch, Paul Hofmarcher, Reinhold Hatzinger, and Kurt Hornik. Model trees with topic model preprocessing: An approach for data journalism illustrated with the WikiLeaks Afghanistan war logs. Annals of Applied Statistics, 7(2):613–639, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1372338461.

Russell:2018:SVS

[RHHH18] James C. Russell, Ephraim M. Hanks, Murali Haran, and David Hughes. A spatially varying stochastic differential equation model for animal movement. Annals of Applied Statistics, 12(2):1312–1331, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1532743496.

Rahman:2022:SNB

[RHL^{+}22] Tanbin Rahman, Hsin-En Huang, Yujia Li, An-Shun Tai, Wen-Ping Hseih, Colleen A. McClung, and George Tseng. A sparse negative binomial classifier with covariate adjustment for RNA-seq data. Annals of Applied Statistics, 16(2):1071– 1089, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/A-sparse-negative-binomial-classifier-
with-covariate-adjustment-for-RNA/10.1214/21-AOAS1532.
full.

Riebler:2012:EET

[RHR12] Andrea Riebler, Leonhard Held, and Håvard Rue. Estimation and extrapolation of time trends in registry data — borrowing strength from related populations. Annals of Applied Statistics, 6(1):304–333, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1331043398.

Reiss:2015:WDR

[RHZ^{+}15] Philip T. Reiss, Lan Huo, Yihong Zhao, Clare Kelly, and R. Todd Ogden. Wavelet-domain regression and predictive inference in psychiatric neuroimaging. Annals of Applied Statistics, 9(2):1076–1101, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1437397124.

Radchenko:2011:IVS

[RJ11] Peter Radchenko and Gareth M. James. Improved variable selection with Forward-Lasso adaptive shrinkage. Annals of Applied Statistics, 5(1):427–448, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1300715197.

Ruddy:2016:SDP

[RJP16] Sean Ruddy, Marla Johnson, and Elizabeth Purdom. Shrinkage of dispersion parameters in the binomial family, with application to differential exon skipping. Annals of Applied Statistics, 10(2):690–725, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199890.

Ross:2022:SBF

[RK22] Gordon J. Ross and Aleksandar A. Kolev. Semiparametric Bayesian forecasting of SpatioTemporal earthquake occurrences. Annals of Applied Statistics, 16(4):2083–2100, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Semiparametric-Bayesian-forecasting-of-
SpatioTemporal-earthquake-occurrences/10.1214/21-
AOAS1554.full.

Relion:2019:NCA

[RKLT19] Jesús D. Arroyo Relión, Daniel Kessler, Elizaveta Levina, and Stephan F. Taylor. Network classification with applications to brain connectomics. Annals of Applied Statistics, 13(3):1648–1677, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277767.

Rahman:2023:EGD

[RKM^{+}23] Syed Rahman, Kshitij Khare, George Michailidis, Carlos Martínez, and Juan Carulla. Estimation of Gaussian directed acyclic graphs using partial ordering information with applications to DREAM3 networks and dairy cattle data. Annals of Applied Statistics, 17(2):929–960, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Estimation-of-Gaussian-directed-acyclic-
graphs-using-partial-ordering-information/10.1214/
22-AOAS1636.full.

Rusch:2013:IES

[RLH^{+}13] Thomas Rusch, Ilro Lee, Kurt Hornik, Wolfgang Jank, and Achim Zeileis. Influencing elections with statistics: Targeting voters with logistic regression trees. Annals of Applied Statistics, 7(3):1612–1639, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804809.

Ranta:2015:BAE

[RLH^{+}15] Jukka Ranta, Roland Lindqvist, Ingrid Hansson, Pirkko Tuominen, and Maarten Nauta. A Bayesian approach to the evaluation of risk-based microbiological criteria for campylobacter in broiler meat. Annals of Applied Statistics, 9 (3):1415–1432, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1446488745.

Roy:2021:PFA

[RLHD21] Arkaprava Roy, Isaac Lavine, Amy H. Herring, and David B. Dunson. Perturbed factor analysis: Accounting for group differences in exposure profiles. Annals of Applied Statistics, 15(3):1386–1404, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-3/Perturbed-factor-
analysis--Accounting-for-group-differences-in-exposure/
10.1214/20-AOAS1435.full.

Richards:2012:PSP

[RLSF12] Joseph W. Richards, Ann B. Lee, Chad M. Schafer, and Peter E. Freeman. Prototype selection for parameter estimation in complex models. Annals of Applied Statistics, 6(1):383– 408, March 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1331043401.

Rad:2017:RSB

[RMP17] Kamiar Rahnama Rad, Timothy A. Machado, and Liam Paninski. Robust and scalable Bayesian analysis of spatial neural tuning function data. Annals of Applied Statistics, 11(2):598–637, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537716.

Regier:2019:AIC

[RMS^{+}19] Jeffrey Regier, Andrew C. Miller, David Schlegel, Ryan P. Adams, Jon D. McAuliffe, and Prabhat. Approximate inference for constructing astronomical catalogs from images. Annals of Applied Statistics, 13(3):1884–1926, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1571277777.

Reimherr:2014:FDA

[RN14] Matthew Reimherr and Dan Nicolae. A functional data analysis approach for genetic association studies. Annals of Applied Statistics, 8(1):406–429, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1396966292.

Rutter:2019:MMC

[RODC19] Carolyn M. Rutter, Jonathan Ozik, Maria DeYoreo, and Nicholson Collier. Microsimulation model calibration using incremental mixture approximate Bayesian computation. Annals of Applied Statistics, 13(4):2189–2212, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1574910041.

Rossell:2009:GPF

[Ros09] David Rossell. GaGa: a parsimonious and flexible model for differential expression analysis. Annals of Applied Statistics, 3(3):1035–1051, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1254773277.

Rosenbaum:2012:EAT

[Ros12] Paul R. Rosenbaum. An exact adaptive test with superior design sensitivity in an observational study of treatments for ovarian cancer. Annals of Applied Statistics, 6(1):83– 105, March 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1331043389.

Rosenbaum:2016:USP

[Ros16] Paul R. Rosenbaum. Using Scheffé projections for multiple outcomes in an observational study of smoking and periodontal disease. Annals of Applied Statistics, 10(3):1447–1471, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069614.

Rosenbaum:2018:SAS

[Ros18] Paul R. Rosenbaum. Sensitivity analysis for stratified comparisons in an observational study of the effect of smoking on homocysteine levels. Annals of Applied Statistics, 12(4):2312–2334, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078046.

Rougier:2011:DSA

[Rou11] Jonathan Rougier. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):96–98, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715183. See [MW11b].

Reich:2013:DEH

[RP13] Brian J. Reich and Michael D. Porter. Discussion of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1871–1875, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823297. See [CW13a].

Regueiro:2016:DCC

[RRS16] Pedro Regueiro, Abel Rodríguez, and Juan Sosa. Discussion of “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1813–1819, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606837. See [JJ16a, JJ16b].

Reshef:2018:ESM

[RRSM18] David N. Reshef, Yakir A. Reshef, Pardis C. Sabeti, and Michael Mitzenmacher. An empirical study of the maximal and total information coefficients and leading measures of dependence. Annals of Applied Statistics, 12(1):123–155, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564467.

Rigat:2009:SPD

[RS09] Fabio Rigat and Jim Q. Smith. Semi-parametric dynamic time series modelling with applications to detecting neural dynamics. Annals of Applied Statistics, 3(4):1776–1804, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453964.

Rizzo:2010:DAN

[RS10] Maria L. Rizzo and Gábor J. Székely. DISCO analysis: a nonparametric extension of analysis of variance. Annals of Applied Statistics, 4(2):1034–1055, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1280842151.

Reich:2012:HMS

[RS12] Brian J. Reich and Benjamin A. Shaby. A hierarchical maxstable spatial model for extreme precipitation. Annals of Applied Statistics, 6(4):1430–1451, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629046.

Rodriguez:2014:BPS

[RS14] Abel Rodriguez and Scott C. Schmidler. Bayesian protein structure alignment. Annals of Applied Statistics, 8(4):2068– 2095, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001735.

Russo:2022:MMM

[RSD22] Massimiliano Russo, Burton H. Singer, and David B. Dunson. Multivariate mixed membership modeling: Inferring domainspecific risk profiles. Annals of Applied Statistics, 16(1):391– 413, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Multivariate-mixed-membership-modeling-
Inferring-domain-specific-risk-profiles/10.1214/21-
AOAS1496.full.

Reitan:2012:PES

[RSH12] Trond Reitan, Tore Schweder, and Jorijntje Henderiks. Phenotypic evolution studied by layered stochastic differential equations. Annals of Applied Statistics, 6(4):1531–1551, December 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1356629050.

Rashid:2016:SMA

[RSI16] Naim U. Rashid, Wei Sun, and Joseph G. Ibrahim. A statistical model to assess (allele-specific) associations between gene expression and epigenetic features using sequencing data. Annals of Applied Statistics, 10(4):2254–2273, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606859.

Retkute:2021:IGM

[RTB^{+}21] Renata Retkute, Panayiota Touloupou, María-Gloria Basáñez, T. Déirdre Hollingsworth, and Simon E. F. Spencer. Integrating geostatistical maps and infectious disease transmission models using adaptive multiple importance sampling. Annals of Applied Statistics, 15(4):1980–1998, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Integrating-
geostatistical-maps-and-infectious-disease-transmission-
models-using-adaptive/10.1214/21-AOAS1486.full.

Richards:2022:MES

[RTB22] Jordan Richards, Jonathan A. Tawn, and Simon Brown. Modelling extremes of spatial aggregates of precipitation using conditional methods. Annals of Applied Statistics, 16(4):2693–2713, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Modelling-extremes-

of-spatial-aggregates-of-precipitation-using-conditional-
methods/10.1214/22-AOAS1609.full.

Rubin:2008:OCI

[Rub08] Donald B. Rubin. For objective causal inference, design trumps analysis. Annals of Applied Statistics, 2(3):808–840, September 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1223908042.

Rubin:2018:SFD

[Rub18] Donald B. Rubin. On Stephen E. Fienberg as a discussant and a friend. Annals of Applied Statistics, 12(2):683– 684, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743472.

Ranciati:2020:IOT

[RVW20] Saverio Ranciati, Veronica Vinciotti, and Ernst C. Wit. Identifying overlapping terrorist cells from the Noordin Top actorevent network. Annals of Applied Statistics, 14(3):1516–1534, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Identifying-overlapping-terrorist-cells-
from-the-Noordin-Top-actorevent-network/10.1214/20-
AOAS1358.full.

Rodriguez:2017:ASR

[RWK17] Abel Rodríguez, Ziwei Wang, and Athanasios Kottas. Assessing systematic risk in the S&P500 index between 2000 and 2011: a Bayesian nonparametric approach. Annals of Applied Statistics, 11(2):527–552, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1500537714.

Rosenthal:2011:DMA

[RY11] Jeffrey S. Rosenthal and Albert H. Yoon. Detecting multiple authorship of United States Supreme Court legal decisions using function words. Annals of Applied Statistics, 5(1):283– 308, March 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1300715191.

Ryu:2022:LBB

[Ryu22] So Young Ryu. Likelihood-based bacterial identification approach for bimicrobial mass spectrometry data. Annals of Applied Statistics, 16(1):612–624, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Likelihood-based-
bacterial-identification-approach-for-bimicrobial-
mass-spectrometry-data/10.1214/21-AOAS1520.full.

Randolph:2018:KPR

[RZC^{+}18] Timothy W. Randolph, Sen Zhao, Wade Copeland, Meredith Hullar, and Ali Shojaie. Kernel-penalized regression for analysis of microbiome data. Annals of Applied Statistics, 12(1): 540–566, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564483.

Sadinle:2014:DDH

[Sad14] Mauricio Sadinle. Detecting duplicates in a homicide registry using a Bayesian partitioning approach. Annals of Applied Statistics, 8(4):2404–2434, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1419001749.

Sadinle:2018:BPR

[Sad18] Mauricio Sadinle. Bayesian propagation of record linkage uncertainty into population size estimation of human rights violations. Annals of Applied Statistics, 12(2):1013–1038, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743484.

Storlie:2014:SIM

[SAV^{+}14] Curtis Storlie, Blake Anderson, Scott Vander Wiel, Daniel Quist, Curtis Hash, and Nathan Brown. Stochastic identification of malware with dynamic traces. Annals of Applied Statistics, 8(1):1–18, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966276.

Savitsky:2016:BNM

[Sav16] Terrance D. Savitsky. Bayesian nonparametric multiresolution estimation for the American Community Survey. Annals of Applied Statistics, 10(4):2157–2181, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606855.

Sandholtz:2020:MDP

[SB20] Nathan Sandholtz and Luke Bornn. Markov decision processes with dynamic transition probabilities: an analysis of shooting strategies in basketball. Annals of Applied Statistics, 14(3):1122–1145, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Markov-decision-
processes-with-dynamic-transition-probabilities--
An-analysis/10.1214/20-AOAS1348.full.

Stival:2023:DOC

[SBD23] Mattia Stival, Mauro Bernardi, and Petros Dellaportas. Doubly-online changepoint detection for monitoring health status during sports activities. Annals of Applied Statistics, 17(3):2387–2409, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Doubly-online-
changepoint-detection-for-monitoring-health-status-
during-sports/10.1214/22-AOAS1724.full.

Sanchez:2009:EEA

[SBJR09] Brisa N. Sánchez, Esben Budtz-Jørgensen, and Louise M. Ryan. An estimating equations approach to fitting latent exposure models with longitudinal health outcomes. Annals of Applied Statistics, 3(2):830–856, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676197.

Steinrucken:2014:NSM

[SBS14] Matthias Steinrücken, Anand Bhaskar, and Yun S. Song. A novel spectral method for inferring general diploid selection from time series genetic data. Annals of Applied Statistics, 8(4):2203–2222, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1419001740.

Snoke:2018:PAM

[SBSH18] Joshua Snoke, Timothy R. Brick, Aleksandra Slavković, and Michael D. Hunter. Providing accurate models across private partitioned data: Secure maximum likelihood estimation. Annals of Applied Statistics, 12(2):877–914, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743480.

Sen:2009:SML

[SBW^{+}09] Bodhisattva Sen, Moulinath Banerjee, Michael Woodroofe, Mario Mateo, and Matthew Walker. Streaming motion in Leo I. Annals of Applied Statistics, 3(1):96–116, March 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1239888364.

Sidhu:2014:MEP

[SC14] Gagan Sidhu and Brian Caffo. MONEYBaRL: Exploiting pitcher decision-making using Reinforcement Learning. Annals of Applied Statistics, 8(2):926–955, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1404229520.

Spivack:2016:EOR

[SC16] John H. Spivack and Bin Cheng. Estimating odds ratios under a case-background design with an application to a study of Sorafenib accessibility. Annals of Applied Statistics, 10(4):2233–2253, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606858.

Songdechakraiwut:2023:TLB

[SC23] Tananun Songdechakraiwut and Moo K. Chung. Topological learning for brain networks. Annals of Applied Statistics, 17(1):403–433, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Topological-
learning-for-brain-networks/10.1214/22-AOAS1633.full.

Stein:2013:SAS

[SCA13] Michael L. Stein, Jie Chen, and Mihai Anitescu. Stochastic approximation of score functions for Gaussian processes. Annals of Applied Statistics, 7(2):1162–1191, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1372338483.

Sharma:2018:NMT

[SCDD18] Kshitij Sharma, Valérie Chavez-Demoulin, and Pierre Dillenbourg. Nonstationary modelling of tail dependence of two subjects’ concentration. Annals of Applied Statistics, 12(2): 1293–1311, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743495.

Simoiu:2017:PIM

[SCDG17] Camelia Simoiu, Sam Corbett-Davies, and Sharad Goel. The problem of infra-marginality in outcome tests for discrimination. Annals of Applied Statistics, 11(3):1193–1216, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1507168827.

Schwartzman:2008:ENF

[Sch08] Armin Schwartzman. Empirical null and false discovery rate inference for exponential families. Annals of Applied Statistics, 2(4):1332–1359, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1231424213.

Schoenberg:2013:DEH

[Sch13] Frederic Paik Schoenberg. Discussion of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1888–1890, December 2013. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823300. See [SR09b].

Schofield:2015:CME

[Sch15] Lynne Steuerle Schofield. Correcting for measurement error in latent variables used as predictors. Annals of Applied Statistics, 9(4):2133–2152, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453994195.

Schwartzman:2016:DFD

[Sch16] Armin Schwartzman. Discussion of “Fiber direction estimation in diffusion MRI”. Annals of Applied Statistics, 10(3):1157–1159, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1475069600.  See

[WLP^{+}16].

Schmidt:2017:DEB

[Sch17] Patrick Schmidt. Discussion of “Elicitability and backtesting: Perspectives for banking regulation”. Annals of Applied Statistics, 11(4):1883–1885, December 2017. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1514430267. See [NZ17a].

Schmid:2023:CMB

[Sch23] Matthias Schmid. Corrigendum: Modeling biomarker ratios with gamma distributed components. Annals of Applied Statistics, 3(2700):??, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Corrigendum-
Modeling-biomarker-ratios-with-gamma-distributed-
components/10.1214/23-AOAS1777.full. See [BWS19].

Steele:2019:MWH

[SCK19] Fiona Steele, Paul S. Clarke, and Jouni Kuha. Modeling within-household associations in household panel studies. Annals of Applied Statistics, 13(1):367–392, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1554861653.

Soiaporn:2013:MBF

[SCL^{+}13] Kunlaya Soiaporn, David Chernoff, Thomas Loredo, David Ruppert, and Ira Wasserman. Multilevel Bayesian framework for modeling the production, propagation and detection of ultra-high energy cosmic rays. Annals of Applied Statistics, 7(3):1249–1285, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1380804795.

Scott:2009:NBM

[Sco09] James G. Scott. Nonparametric Bayesian multiple testing for longitudinal performance stratification. Annals of Applied Statistics, 3(4):1655–1674, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453958.

Schliep:2020:DFM

[SCRS^{+}20] Erin M. Schliep, Sarah M. Collins, Shirley Rojas-Salazar, Noah R. Lottig, and Emily H. Stanley. Data fusion model for speciated nitrogen to identify environmental drivers and improve estimation of nitrogen in lakes. Annals of Applied Statistics, 14(4):1651–1675, December 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-4/Data-fusion-
model-for-speciated-nitrogen-to-identify-environmental-
drivers/10.1214/20-AOAS1371.full.

Stingo:2011:IBI

[SCTV11] Francesco C. Stingo, Yian A. Chen, Mahlet G. Tadesse, and Marina Vannucci. Incorporating biological information into linear models: a Bayesian approach to the selection of pathways and genes. Annals of Applied Statistics, 5(3):1978–2002, September 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1318514292.

Stingo:2010:BGM

[SCV^{+}10] Francesco C. Stingo, Yian A. Chen, Marina Vannucci, Marianne Barrier, and Philip E. Mirkes. A Bayesian graphical modeling approach to microRNA regulatory network inference. Annals of Applied Statistics, 4(4):2024–2048, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167808.

Sun:2023:DRP

[SCW^{+}23] Yifei Sun, Sy Han Chiou, Colin O. Wu, Meghan E. McGarry, and Chiung-Yu Huang. Dynamic risk prediction triggered by intermediate events using survival tree ensembles. Annals of Applied Statistics, 17(2):1375–1397, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Dynamic-risk-
prediction-triggered-by-intermediate-events-using-
survival-tree/10.1214/22-AOAS1674.full.

Suveges:2010:MMP

[SD10] Mária Süveges and Anthony C. Davison. Model misspecification in peaks over threshold analysis. Annals of Applied Statistics, 4(1):203–221, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584453.

Shin:2018:CMM

[SDH18] Yei Eun Shin, Yu Ding, and Jianhua Z. Huang. Covariate matching methods for testing and quantifying wind turbine upgrades. Annals of Applied Statistics, 12(2):1271– 1292, June 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1532743494.

Sabbaghi:2014:IDI

[SDHZ14] Arman Sabbaghi, Tirthankar Dasgupta, Qiang Huang, and Jizhe Zhang. Inference for deformation and interference in 3D printing. Annals of Applied Statistics, 8(3):1395–1415, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091218.

Saunders:2011:CEC

[SDL^{+}11] Christopher P. Saunders, Linda J. Davis, Andrea C. Lamas, John J. Miller, and Donald T. Gantz. Construction and evaluation of classifiers for forensic document analysis. Annals of Applied Statistics, 5(1):381–399, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1300715195.

Severn:2022:MVD

[SDP22] Katie E. Severn, Ian L. Dryden, and Simon P. Preston. Manifold valued data analysis of samples of networks, with applications in corpus linguistics. Annals of Applied Statistics, 16(1):368–390, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Manifold-valued-
data-analysis-of-samples-of-networks-with-applications/
10.1214/21-AOAS1480.full.

Schwartzman:2008:FDR

[SDT08] Armin Schwartzman, Robert F. Dougherty, and Jonathan E. Taylor. False discovery rate analysis of brain diffusion direction maps. Annals of Applied Statistics, 2(1):153–175, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1206367816.

Serban:2011:STV

[Ser11] Nicoleta Serban. A space–time varying coefficient model: The equity of service accessibility. Annals of Applied Statistics, 5(3):2024–2051, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514294.

Shang:2021:PMC

[SEX21] Zhuoran Shang, Elena A. Erosheva, and Gongjun Xu. Partial-mastery cognitive diagnosis models. Annals of Applied Statistics, 15(3):1529–1555, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-3/Partial-mastery-
cognitive-diagnosis-models/10.1214/21-AOAS1439.full.

Soneson:2011:MVI

[SF11] Charlotte Soneson and Magnus Fontes. A method for visual identification of small sample subgroups and potential biomarkers. Annals of Applied Statistics, 5(3):2131–2149, September 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1318514298.

Shi:2016:MMI

[SFB16] Peng Shi, Xiaoping Feng, and Jean-Philippe Boucher. Multilevel modeling of insurance claims using copulas. Annals of Applied Statistics, 10(2):834–863, June 2016. CODEN ????

ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1469199895.

Sain:2011:SAM

[SFC11] Stephan R. Sain, Reinhard Furrer, and Noel Cressie. A spatial analysis of multivariate output from regional climate models. Annals of Applied Statistics, 5(1):150–175, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1300715186.

Santos-Fernandez:2022:RID

[SFDMM22] Edgar Santos-Fernandez, Francesco Denti, Kerrie Mengersen, and Antonietta Mira. The role of intrinsic dimension in highresolution player tracking data-insights in basketball. Annals of Applied Statistics, 16(1):326–348, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/The-role-of-
intrinsic-dimension-in-high-resolution-player-tracking/
10.1214/21-AOAS1506.full.

Smith:2015:QRM

[SFGLR15] Luke B. Smith, Montserrat Fuentes, Penny Gordon-Larsen, and Brian J. Reich. Quantile regression for mixed models with an application to examine blood pressure trends in China. Annals of Applied Statistics, 9(3):1226–1246, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488737.

Schauer:2021:ESM

[SFPS^{+}21] Jacob M. Schauer, Kaitlyn G. Fitzgerald, Sarah Peko-Spicer, Mena C. R. Whalen, Rrita Zejnullahi, and Larry V. Hedges. An evaluation of statistical methods for aggregate patterns of replication failure. Annals of Applied Statistics, 15(1):208– 229, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/An-evaluation-of-statistical-methods-for-
aggregate-patterns-of-replication/10.1214/20-AOAS1387.
full.

Shi:2016:IDB

[SG16] Ran Shi and Ying Guo. Investigating differences in brain functional networks using hierarchical covariate-adjusted independent component analysis. Annals of Applied Statistics, 10(4):1930–1957, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606846.

Shirota:2017:SCT

[SG17] Shinichiro Shirota and Alan E. Gelfand. Space and circular time log Gaussian Cox processes with application to crime event data. Annals of Applied Statistics, 11(2):481–503, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1500537712.

Savitsky:2023:JPV

[SGC23] Terrance D. Savitsky, Julie Gershunskaya, and Mark Crankshaw. Joint point and variance estimation under a hierarchical Bayesian model for survey count data. Annals of Applied Statistics, 17(3):2002–2018, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Joint-point-
and-variance-estimation-under-a-hierarchical-Bayesian-
model/10.1214/22-AOAS1704.full.

Schliep:2017:BPU

[SGCT17] Erin M. Schliep, Alan E. Gelfand, James S. Clark, and Bradley J. Tomasek. Biomass prediction using a densitydependent diameter distribution model. Annals of Applied Statistics, 11(1):340–361, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1491616884.

Seo:2007:MMS

[SGCW07] David M. Seo, Pascal J. Goldschmidt-Clermont, and Mike West. Of mice and men: Sparse statistical modeling in cardiovascular genomics. Annals of Applied Statistics, 1(1):152– 178, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143733.

Sang:2008:ISO

[SGL^{+}08] Huiyan Sang, Alan E. Gelfand, Chris Lennard, Gabriele Hegerl, and Bruce Hewitson. Interpreting self-organizing maps through space–time data models. Annals of Applied Statistics, 2(4):1194–1216, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1231424206.

Sakov:2010:HTD

[SGLB10] Anat Sakov, Ilan Golani, Dina Lipkind, and Yoav Benjamini. High-throughput data analysis in behavior genetics. Annals of Applied Statistics, 4(2):743–763, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1280842138.

Spieker:2022:BLA

[SGNM22] Andrew J. Spieker, Robert A. Greevy, Lyndsay A. Nelson, and Lindsay S. Mayberry. Bounding the local average treatment effect in an instrumental variable analysis of engagement with a mobile intervention. Annals of Applied Statistics, 16(1):60–79, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Bounding-the-
local-average-treatment-effect-in-an-instrumental-
variable/10.1214/21-AOAS1476.full.

Shen:2008:FTS

[SH08] Haipeng Shen and Jianhua Z. Huang. Forecasting time series of inhomogeneous Poisson processes with application to call center workforce management. Annals of Applied Statistics, 2(2):601–623, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1215118530.

Su:2011:HDN

[SH11] Li Su and Joseph W. Hogan. HIV dynamics and natural history studies: Joint modeling with doubly interval-censored event time and infrequent longitudinal data. Annals of Applied Statistics, 5(1):400–426, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1300715196.

Stehlik:2015:LE

[SH15] Milan Stehlík and Philipp Hermann. Letter to the Editor. Annals of Applied Statistics, 9(4):2051, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1453994190.

Sabbaghi:2018:MTA

[SH18] Arman Sabbaghi and Qiang Huang. Model transfer across additive manufacturing processes via mean effect equivalence of lurking variables. Annals of Applied Statistics, 12(4):2409– 2429, December 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1542078050.

Shang:2016:MLE

[Sha16] Han Lin Shang. Mortality and life expectancy forecasting for a group of populations in developed countries: a multilevel functional data method. Annals of Applied Statistics, 10(3):1639–1672, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1475069622.

Saha:2022:BPM

[SHAB22] Abhisek Saha, Min Jin Ha, Satwik Acharyya, and Veerabhadran Baladandayuthapani. A Bayesian precision medicine framework for calibrating individualized therapeutic indices in cancer. Annals of Applied Statistics, 16(4):2055–2082, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/A-Bayesian-precision-medicine-framework-
for-calibrating-individualized-therapeutic-indices/
10.1214/21-AOAS1550.full.

Siddique:2012:AMD

[SHC12] Juned Siddique, Ofer Harel, and Catherine M. Crespi. Addressing missing data mechanism uncertainty using multiplemodel multiple imputation: Application to a longitudinal clinical trial. Annals of Applied Statistics, 6(4):1814–1837, December 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1356629061.

Scharf:2016:DSN

[SHF^{+}16] Henry R. Scharf, Mevin B. Hooten, Bailey K. Fosdick, Devin S. Johnson, Josh M. London, and John W. Durban. Dynamic social networks based on movement. Annals of Applied Statistics, 10(4):2182–2202, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1483606856.

Silva:2010:RRU

[SHGA10] Ricardo Silva, Katherine Heller, Zoubin Ghahramani, and Edoardo M. Airoldi. Ranking relations using analogies in biological and information networks. Annals of Applied Statistics, 4(2):615–644, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1280842133.

Sun:2022:OTD

[SHH22] Yifei Sun, Xuming He, and Jianhua Hu. An omnibus test for detection of subgroup treatment effects via data partitioning. Annals of Applied Statistics, 16(4):2266–2278, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-16/issue-4/An-
omnibus-test-for-detection-of-subgroup-treatment-
effects-via/10.1214/21-AOAS1589.full.

Spencer:2015:INS

[SHM15] Simon E. F. Spencer, Steven M. Hill, and Sach Mukherjee. Inferring network structure from interventional timecourse experiments. Annals of Applied Statistics, 9(1):507– 524, March 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1430226102.

Shan:2020:SWP

[SHM20] Qianqian Shan, Yili Hong, and William Q. Meeker. Seasonal warranty prediction based on recurrent event data. Annals of Applied Statistics, 14(2):929–955, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1593449332.

Sun:2022:MCM

[SHR^{+}22] Hu Sun, Zhijun Hua, Jiaen Ren, Shasha Zou, Yuekai Sun, and Yang Chen. Matrix completion methods for the total electron content video reconstruction. Annals of Applied Statistics, 16(3):1333–1358, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Matrix-completion-
methods-for-the-total-electron-content-video-reconstruction/
10.1214/21-AOAS1541.full.

Schlosser:2019:DRF

[SHSZ19] Lisa Schlosser, Torsten Hothorn, Reto Stauffer, and Achim Zeileis. Distributional regression forests for probabilistic precipitation forecasting in complex terrain. Annals of Applied Statistics, 13(3):1564–1589, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1571277764.

Sharma:2018:SDC

[SHW18] Amit Sharma, Jake M. Hofman, and Duncan J. Watts. Splitdoor criterion: Identification of causal effects through auxiliary outcomes. Annals of Applied Statistics, 12(4):2699– 2733, December 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1542078061.

Shi:2011:TSE

[SIL^{+}11] Xiaoyan Shi, Joseph G. Ibrahim, Jeffrey Lieberman, Martin Styner, Yimei Li, and Hongtu Zhu. Two-stage empirical likelihood for longitudinal neuroimaging data. Annals of Applied Statistics, 5(2B):1132–1158, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562716.

Silverman:2016:IDC

[Sil16] Bernard W. Silverman. Introduction to discussion of “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1777–1778, December 2016. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606835.

Singpurwalla:2009:URS

[Sin09] Nozer D. Singpurwalla. The utility of reliability and survival. Annals of Applied Statistics, 3(4):1581–1596, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1267453954.

Singpurwalla:2011:NRD

[Sin11] Nozer D. Singpurwalla. Network routing in a dynamic environment. Annals of Applied Statistics, 5(2B):1407– 1424, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562726. See [Sin11].

Stallrich:2020:OEP

[SIS^{+}20] Jonathan Stallrich, Md Nazmul Islam, Ana-Maria Staicu, Dustin Crouch, Lizhi Pan, and He Huang. Optimal EMG placement for a robotic prosthesis controller with sequential, adaptive functional estimation (SAFE). Annals of Applied Statistics, 14(3):1164–1181, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Optimal-EMG-
placement-for-a-robotic-prosthesis-controller-with-
sequential/10.1214/20-AOAS1324.full.

Sutton:2011:BIQ

[SJ11] Charles Sutton and Michael I. Jordan. Bayesian inference for queueing networks and modeling of Internet services. Annals of Applied Statistics, 5(1):254–282, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1300715190.

Sorrentino:2013:DFS

[SJA^{+}13] Alberto Sorrentino, Adam M. Johansen, John A. D. Aston, Thomas E. Nichols, and Wilfrid S. Kendall. Dynamic filtering of static dipoles in magnetoencephalography. Annals of Applied Statistics, 7(2):955–988, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338475.

Schwartzman:2013:MTL

[SJGM13] Armin Schwartzman, Andrew Jaffe, Yulia Gavrilov, and Clifford A. Meyer. Multiple testing of local maxima for detection of peaks in ChIP-seq data. Annals of Applied Statistics, 7(1): 471–494, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527207.

Sang:2011:CAL

[SJH11] Huiyan Sang, Mikyoung Jun, and Jianhua Z. Huang. Covariance approximation for large multivariate spatial data sets with an application to multiple climate model errors. Annals of Applied Statistics, 5(4):2519–2548, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399605.

Soltoft-Jensen:2020:MSP

[SJHJD20] Aleksander Søltoft-Jensen, Mads Peter Heide-Jørgensen, and Susanne Ditlevsen. Modelling the sound production of narwhals using a point process framework with memory effects. Annals of Applied Statistics, 14(4):2037–2052, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Modelling-the-sound-production-of-narwhals-
using-a-point-process/10.1214/20-AOAS1379.full.

Satopaa:2014:PAT

[SJM^{+}14] Ville A. Satopää, Shane T. Jensen, Barbara A. Mellers, Philip E. Tetlock, and Lyle H. Ungar. Probability aggregation in time-series: Dynamic hierarchical modeling of sparse expert beliefs. Annals of Applied Statistics, 8(2):1256– 1280, June 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1404229533.

Schneble:2022:IEG

[SK22] Marc Schneble and Göran Kauermann. Intensity estimation on geometric networks with penalized splines. Annals of Applied Statistics, 16(2):843–865, June 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-2/Intensity-estimation-
on-geometric-networks-with-penalized-splines/10.1214/
21-AOAS1522.full.

Shokoohi:2019:CHC

[SKAL19] Farhad Shokoohi, Abbas Khalili, Masoud Asgharian, and Shili Lin. Capturing heterogeneity of covariate effects in hidden subpopulations in the presence of censoring and large number of covariates. Annals of Applied Statistics, 13(1): 444–465, March 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1554861656.

Sun:2023:IRA

[SKBL23] Yuming Sun, Jian Kang, Chad Brummett, and Yi Li. Individualized risk assessment of preoperative opioid use by interpretable neural network regression. Annals of Applied Statistics, 17(1):434–453, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Individualized-
risk-assessment-of-preoperative-opioid-use-by-interpretable-
neural/10.1214/22-AOAS1634.full.

Su:2014:SAT

[SKKS14] Jingyong Su, Sebastian Kurtek, Eric Klassen, and Anuj Srivastava. Statistical analysis of trajectories on Riemannian manifolds: Bird migration, hurricane tracking and video surveillance. Annals of Applied Statistics, 8(1):530– 552, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966297.

Snijders:2010:MLE

[SKS10] Tom A. B. Snijders, Johan Koskinen, and Michael Schweinberger. Maximum likelihood estimation for social network dynamics. Annals of Applied Statistics, 4(2):567–588, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1280842131.

Sigrist:2012:DNS

[SKS12] Fabio Sigrist, Hans R. Künsch, and Werner A. Stahel. A dynamic nonstationary spatio-temporal model for short term prediction of precipitation. Annals of Applied Statistics, 6(4):1452–1477, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629047.

Sun:2014:SCQ

[SKZ14] Zhaonan Sun, Thomas Kuczek, and Yu Zhu. Statistical calibration of qRT-PCR, microarray and RNA-seq gene expression data with measurement error models. Annals of Applied Statistics, 8(2):1022–1044, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229524.

Sohn:2019:CMA

[SL19] Michael B. Sohn and Hongzhe Li. Compositional mediation analysis for microbiome studies. Annals of Applied Statistics, 13(1):661–681, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861664.

Sobel:2020:ECE

[SL20] Michael E. Sobel and Martin A. Lindquist. Estimating causal effects in studies of human brain function: New models, methods and estimands. Annals of Applied Statistics, 14(1): 452–472, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002682.

Sheng:2023:PHR

[SLBL23] Ben Sheng, Changcheng Li, Le Bao, and Runze Li. Probabilistic HIV recency classification — a logistic regression without labeled individual level training data. Annals of Applied Statistics, 17(1):108–129, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Probabilistic-
HIV-recency-classificationa-logistic-regression-without-
labeled-individual-level/10.1214/22-AOAS1618.full.

Shaddick:2008:EER

[SLZS08] Gavin Shaddick, Duncan Lee, James V. Zidek, and Ruth Salway. Estimating exposure response functions using ambient pollution concentrations. Annals of Applied Statistics, 2(4):1249–1270, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1231424209.

Singhal:2010:OED

[SM10] Harsh Singhal and George Michailidis. Optimal experiment design in a filtering context with application to sampled network data. Annals of Applied Statistics, 4(1):78–93, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1273584448.

Stadler:2013:PEH

[SM13] Nicolas Städler and Sach Mukherjee. Penalized estimation in high-dimensional hidden Markov models with state-specific graphical models. Annals of Applied Statistics, 7(4):2157– 2179, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823314.

Scheuerer:2015:PWS

[SM15] Michael Scheuerer and David Möller. Probabilistic wind speed forecasting on a grid based on ensemble model output statistics. Annals of Applied Statistics, 9(3):1328–1349, September 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1446488741.

Scharwachter:2020:DTT

[SM20a] Erik Scharwächter and Emmanuel Müller. Does terrorism trigger online hate speech? on the association of events and time series. Annals of Applied Statistics, 14(3):1285–1303, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/Does-terrorism-trigger-online-hate-speech-
On-the-association-of/10.1214/20-AOAS1338.full.

Spencer:2020:BHM

[SM20b] Neil A. Spencer and Jared S. Murray. A Bayesian hierarchical model for evaluating forensic footwear evidence. Annals of Applied Statistics, 14(3):1449–1470, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-14/issue-3/A-
Bayesian-hierarchical-model-for-evaluating-forensic-
footwear-evidence/10.1214/20-AOAS1334.full.

Starling:2020:BTS

[SMC^{+}20] Jennifer E. Starling, Jared S. Murray, Carlos M. Carvalho, Radek K. Bukowski, and James G. Scott. BART with targeted smoothing: an analysis of patient-specific stillbirth risk. Annals of Applied Statistics, 14(1):28–50, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1587002663.

Smerdon:2011:DSA

[Sme11] Jason E. Smerdon. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):76–79, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715178. See [MW11b].

Samarov:2011:LKC

[SML^{+}11] Daniel Samarov, J. S. Marron, Yufeng Liu, Christopher Grulke, and Alexander Tropsha. Local kernel canonical correlation analysis with application to virtual drug screening. Annals of Applied Statistics, 5(3):2169–2196, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1318514300.

Starling:2021:TSB

[SML^{+}21] Jennifer E. Starling, Jared S. Murray, Patricia A. Lohr, Abigail R. A. Aiken, Carlos M. Carvalho, and James G. Scott. Targeted smooth Bayesian causal forests: an analysis of heterogeneous treatment effects for simultaneous vs. interval medical abortion regimens over gestation. Annals of Applied Statistics, 15(3):1194–1219, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Targeted-Smooth-Bayesian-Causal-Forests-
-An-analysis-of-heterogeneous/10.1214/20-AOAS1438.
full.

Schmidt:2011:DSA

[SMR11] Gavin A. Schmidt, Michael E. Mann, and Scott D. Rutherford. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):65–70, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715176. See [MW11b].

Schumacher:2022:FBF

[SMW^{+}22] Austin E. Schumacher, Tyler H. McCormick, Jon Wakefield, Yue Chu, Jamie Perin, Francisco Villavicencio, Noah Simon, and Li Liu. A flexible Bayesian framework to estimate age- and cause-specific child mortality over time from sample registration data. Annals of Applied Statistics, 16(1):124–143, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-16/issue-1/A-
flexible-Bayesian-framework-to-estimate-age--and-
cause/10.1214/21-AOAS1489.full.

Song:2016:SRA

[SMZ16] Chi Song, Xiaoyi Min, and Heping Zhang. The screening and ranking algorithm for change-points detection in multiple samples. Annals of Applied Statistics, 10(4):2102–2129, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606853.

Sarkar:2021:TVF

[SMZ21] Shuchismita Sarkar, Volodymyr Melnykov, and Xuwen Zhu. Tensor-variate finite mixture modeling for the analysis of university professor remuneration. Annals of Applied Statistics, 15(2):1017–1036, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Tensor-variate-
finite-mixture-modeling-for-the-analysis-of-university/
10.1214/20-AOAS1420.full.

Savitsky:2013:BNH

[SP13] Terrance D. Savitsky and Susan M. Paddock. Bayesian nonparametric hierarchical modeling for multiple membership data in grouped attendance interventions. Annals of Applied Statistics, 7(2):1074–1094, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338479.

Sales:2019:RML

[SP19] Adam C. Sales and John F. Pane. The role of mastery learning in an intelligent tutoring system: Principal stratification on a latent variable. Annals of Applied Statistics, 13(1):420– 443, March 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1554861655.

Schnell:2020:MUS

[SP20] Patrick M. Schnell and Georgia Papadogeorgou. Mitigating unobserved spatial confounding when estimating the effect of supermarket access on cardiovascular disease deaths. Annals of Applied Statistics, 14(4):2069–2095, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Mitigating-unobserved-spatial-confounding-
when-estimating-the-effect-of-supermarket/10.1214/
20-AOAS1377.full.

Spitzer:2020:MCS

[SPF20] Kerstin Spitzer, Marta Pelizzola, and Andreas Futschik. Modifying the chi-square and the CMH test for population genetic inference: Adapting to overdispersion. Annals of Applied Statistics, 14(1):202–220, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1587002671.

Shi:2017:CHU

[SPH17] Xu Shi, Hristina Pashova, and Patrick J. Heagerty. Comparing healthcare utilization patterns via global differences in the endorsement of current procedural terminology codes. Annals of Applied Statistics, 11(3):1349–1374, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1507168832.

Sarti:2023:BAR

[SPI^{+}23] Danilo A. Sarti, Estevão B. Prado, Alan N. Inglis, Antônia A. L. dos Santos, Catherine B. Hurley, Rafael A. Moral, and Andrew C. Parnell. Bayesian additive regression trees for genotype by environment interaction models. Annals of Applied Statistics, 17(3):1936–1957, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Bayesian-additive-regression-trees-for-
genotype-by-environment-interaction-models/10.1214/
22-AOAS1698.full.

Scharpf:2008:HMM

[SPPR08] Robert B. Scharpf, Giovanni Parmigiani, Jonathan Pevsner, and Ingo Ruczinski. Hidden Markov models for the assessment of chromosomal alterations using high-throughput SNP arrays. Annals of Applied Statistics, 2(2):687–713, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1215118534.

Si:2020:BPM

[SPS20] Yajuan Si, Mari Palta, and Maureen Smith. Bayesian profiling multiple imputation for missing hemoglobin values in electronic health records. Annals of Applied Statistics, 14(4):1903–1924, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-4/Bayesian-profiling-
multiple-imputation-for-missing-hemoglobin-values-
in-electronic/10.1214/20-AOAS1378.full.

Sheu:2016:ATD

[SPsLC16] Ching-Fan Sheu, Émeline Perthame, Yuh shiow Lee, and David Causeur. Accounting for time dependence in largescale multiple testing of event-related potential data. Annals of Applied Statistics, 10(1):219–245, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1458909914.

Small:2009:EFM

[SR09a] Dylan S. Small and Paul R. Rosenbaum. Error-free milestones in error prone measurements. Annals of Applied Statistics, 3(3):881–901, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1254773270.

Szekely:2009:BDC

[SR09b] Gábor J. Székely and Maria L. Rizzo. Brownian distance covariance. Annals of Applied Statistics, 3(4):1236–1265, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://projecteuclid.

org/euclid.aoas/1267453933. See editor’s introduction

[New09], discussion [BX09, Kos09, Cop09, Feu09, GFS09, Rém09, Gen09, SR09c], and correction [Kos13].

Szekely:2009:RBD

[SR09c] Gábor J. Székely and Maria L. Rizzo. Rejoinder: Brownian distance covariance. Annals of Applied Statistics, 3(4):1303– 1308, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453941. See [SR09b].

Sottosanti:2023:CCS

[SR23]  Andrea Sottosanti and Davide Risso.  Co-clustering

of spatially resolved transcriptomic data. Annals of Applied Statistics, 17(2):1444–1468, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-2/Co-
clustering-of-spatially-resolved-transcriptomic-data/
10.1214/22-AOAS1677.full.

Scharfstein:2015:ATS

[SRA^{+}15] Daniel Scharfstein, Andrea Rotnitzky, Maria Abraham, Aidan McDermott, Richard Chaisson, and Lawrence Geiter. On the analysis of tuberculosis studies with intermittent missing sputum data. Annals of Applied Statistics, 9 (4):2215–2236, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453994198.

Schmidt:2015:PCA

[SRC15] Alexandra M. Schmidt, Marco A. Rodríguez, and Estelina S. Capistrano. Population counts along elliptical habitat contours: Hierarchical modeling using Poisson-lognormal mixtures with nonstationary spatial structure. Annals of Applied Statistics, 9(3):1372–1393, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1446488743.

Shaby:2016:MSM

[SRCK16] Benjamin A. Shaby, Brian J. Reich, Daniel Cooley, and Cari G. Kaufman. A Markov-switching model for heat waves. Annals of Applied Statistics, 10(1):74–93, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1458909908.

Si:2016:BLP

[SRH16] Yajuan Si, Jerome P. Reiter, and D. Sunshine Hillygus. Bayesian latent pattern mixture models for handling attrition in panel studies with refreshment samples. Annals of Applied Statistics, 10(1):118–143, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1458909910.

Shmueli:2007:BMB

[SRJ07] Galit Shmueli, Ralph P. Russo, and Wolfgang Jank. The BARISTA: a model for bid arrivals in online auctions. Annals of Applied Statistics, 1(2):412–441, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1196438025.

Szpiro:2010:MRR

[SRL10] Adam A. Szpiro, Kenneth M. Rice, and Thomas Lumley. Model-robust regression and a Bayesian “sandwich” estimator. Annals of Applied Statistics, 4(4):2099–2113, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167811.

Schildcrout:2015:BSD

[SRZ^{+}15] Jonathan S. Schildcrout, Paul J. Rathouz, Leila R. Zelnick, Shawn P. Garbett, and Patrick J. Heagerty. Biased sampling designs to improve research efficiency: Factors influencing pulmonary function over time in children with asthma. Annals of Applied Statistics, 9(2):731–753, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1437397109.

Sellers:2010:FRM

[SS10a] Kimberly F. Sellers and Galit Shmueli. A flexible regression model for count data. Annals of Applied Statistics, 4(2): 943–961, June 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1280842147.

Shlomo:2010:APP

[SS10b] Natalie Shlomo and Chris Skinner. Assessing the protection provided by misclassification-based disclosure limitation methods for survey microdata. Annals of Applied Statistics, 4(3):1291–1310, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409374.

Shim:2015:WBG

[SS15a] Heejung Shim and Matthew Stephens. Wavelet-based genetic association analysis of functional phenotypes arising from high-throughput sequencing assays. Annals of Applied Statistics, 9(2):665–686, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1437397106.

Sun:2015:SST

[SS15b] Ying Sun and Michael L. Stein. A stochastic space-time model for intermittent precipitation occurrences. Annals of Applied Statistics, 9(4):2110–2132, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453994194.

Skinner:2020:EDC

[SS20] Chris Skinner and Fiona Steele. Estimation of dyadic characteristics of family networks using sample survey data. Annals of Applied Statistics, 14(2):706–726, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1593449322.

Safo:2015:SSD

[SSD15] Sandra Safo, Xiao Song, and Kevin K. Dobbin. Sample size determination for training cancer classifiers from microarray and RNA-seq data. Annals of Applied Statistics, 9(2):1053– 1075, June 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1437397123.

Singh:2019:NST

[SSD^{+}19] Susheela P. Singh, Ana-Maria Staicu, Robert R. Dunn, Noah Fierer, and Brian J. Reich. A nonparametric spatial test to identify factors that shape a microbiome. Annals of Applied Statistics, 13(4):2341–2362, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910047.

Stenzel:2011:SMM

[SSH^{+}11] O. Stenzel, V. Schmidt, H. Hassfeld, R. Thiedmann, L. J. A. Koster, S. D. Oosterhout, S. S. van Bavel, M. M. Wienk, J. Loos, and R. A. J. Janssen. Spatial modeling of the 3d morphology of hybrid polymer-ZnO solar cells, based on electron tomography data. Annals of Applied Statistics, 5(3):1920–1947, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514290.

Shirley:2010:HMM

[SSL^{+}10] Kenneth E. Shirley, Dylan S. Small, Kevin G. Lynch, Stephen A. Maisto, and David W. Oslin. Hidden Markov models for alcoholism treatment trial data. Annals of Applied Statistics, 4(1):366–395, March 2010. CODEN ????

ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584459.

Spiegelman:2008:RLE

[SST^{+}08] Clifford Spiegelman, S. J. Sheather, W. A. Tobin, W. D. James, S. Wexler, and D. M. Roundhill. Response to the letter to the Editor. Annals of Applied Statistics, 2(1):432– 433, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367829.

Schwartzman:2019:SCE

[SSZT19] Armin Schwartzman, Andrew J. Schork, Rong Zablocki, and Wesley K. Thompson. A simple, consistent estimator of SNP heritability from genome-wide association studies. Annals of Applied Statistics, 13(4):2509–2538, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1574910053.

Speed:2011:SPN

[ST11] Doug Speed and Simon Tavaré. Sparse partitioning: Nonlinear regression with binary or tertiary predictors, with application to association studies. Annals of Applied Statistics, 5(2A):873–893, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1310562209.

Song:2014:HSO

[ST14] Chi Song and George C. Tseng. Hypothesis setting and order statistic for robust genomic meta-analysis. Annals of Applied Statistics, 8(2):777–800, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229514.

Stark:2008:CSP

[Sta08a] Philip B. Stark. Conservative statistical post-election audits. Annals of Applied Statistics, 2(2):550–581, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1215118528.

Stark:2008:SDM

[Sta08b] Philip B. Stark. A sharper discrepancy measure for postelection audits. Annals of Applied Statistics, 2(3):982–985, September 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1223908048.

Schmitt:2018:EBS

[STA18] Eric Schmitt, Christopher Tull, and Patrick Atwater. Extending Bayesian structural time-series estimates of causal impact to many-household conservation initiatives. Annals of Applied Statistics, 12(4):2517–2539, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1542078054.

Stark:2023:AAL

[Sta23] Philip B. Stark. ALPHA: Audit that learns from previously hand-audited ballots. Annals of Applied Statistics, 17(1): 641–679, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/ALPHA-Audit-that-learns-from-previously-
hand-audited-ballots/10.1214/22-AOAS1646.full.

Stephens:2013:FCA

[STD13] Alisa J. Stephens, Eric J. Tchetgen Tchetgen, and Victor De Gruttola. Flexible covariate-adjusted exact tests of randomized treatment effects with application to a trial of HIV education. Annals of Applied Statistics, 7(4):2106–2137, December 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1387823312.

Stein:2007:SVT

[Ste07] Michael L. Stein. Spatial variation of total column ozone on a global scale. Annals of Applied Statistics, 1(1):191– 210, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143735.

Stein:2009:SIH

[Ste09] Michael L. Stein. Spatial interpolation of high-frequency monitoring data. Annals of Applied Statistics, 3(1):272– 291, March 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1239888371.

Stein:2011:E

[Ste11] Michael L. Stein. Editorial. Annals of Applied Statistics, 5 (1):1–4, March 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1300715169.

Shan:2021:MIP

[STG21] Mingyang Shan, Kali S. Thomas, and Roee Gutman. A multiple imputation procedure for record linkage and causal inference to estimate the effects of home-delivered meals. Annals of Applied Statistics, 15(1):412–436, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-15/issue-1/A-
multiple-imputation-procedure-for-record-linkage-
and-causal-inference/10.1214/20-AOAS1397.full.

Stigler:2008:DSA

[Sti08] Stephen M. Stigler. Discussion of: “Statistical analysis of an archeological find”. Annals of Applied Statistics, 2(1): 55–56, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367806. See [Feu08b].

Spiegelman:2007:CFA

[STJ^{+}07] Cliff Spiegelman, William A. Tobin, William D. James, Simon J. Sheather, Stuart Wexler, and D. Max Roundhill. Chemical and forensic analysis of JFK assassination bullet lots: Is a second shooter possible? Annals of Applied Statistics, 1(2):287–301, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1196438019.

Salter-Townshend:2017:LSM

[STM17] Michael Salter-Townshend and Tyler H. McCormick. Latent space models for multiview network data. Annals of Applied Statistics, 11(3):1217–1244, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1507168828.

Schnell:2017:SIM

[STMC17] Patrick Schnell, Qi Tang, Peter Müller, and Bradley P. Carlin. Subgroup inference for multiple treatments and multiple endpoints in an Alzheimer’s disease treatment trial. Annals of Applied Statistics, 11(2):949–966, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1500537730.

Stone:2010:SRM

[Sto10] Charles J. Stone. Selected recollections of my relationship with Leo Breiman. Annals of Applied Statistics, 4(4):1652– 1655, December 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1294167791.

Sung:2022:EFP

[Sun22] Chih-Li Sung. Estimating functional parameters for understanding the impact of weather and government interventions on COVID-19 outbreak. Annals of Applied Statistics, 16(4):2505–2522, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Estimating-functional-
parameters-for-understanding-the-impact-of-weather-
and/10.1214/22-AOAS1601.full.

Schnitzer:2014:EBG

[SvdLMP14] Mireille E. Schnitzer, Mark J. van der Laan, Erica E. M. Moodie, and Robert W. Platt. Effect of breastfeeding on gastrointestinal infection in infants: a targeted maximum likelihood approach for clustered longitudinal data. Annals of Applied Statistics, 8(2):703–725, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229511.

Schisterman:2011:CED

[SVYP11] Enrique F. Schisterman, Albert Vexler, Aijun Ye, and Neil J. Perkins. A combined efficient design for biomarker data subject to a limit of detection due to measuring instrument sensitivity. Annals of Applied Statistics, 5(4):2651–2667, December 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1324399610.

Sun:2010:GIP

[SW10] Wei Sun and Fred A. Wright. A geometric interpretation of the permutation p-value and its application in eQTL studies. Annals of Applied Statistics, 4(2):1014–1033, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1280842150.

Song:2017:PTV

[SW17] Xiao Song and Li Wang. Partially time-varying coefficient proportional hazards models with error-prone timedependent covariates — an application to the AIDS Clinical Trial Group 175 data. Annals of Applied Statistics, 11(1): 274–296, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616881.

Schafer:2022:BIR

[SWH22] Toryn L. J. Schafer, Christopher K. Wikle, and Mevin B. Hooten. Bayesian inverse reinforcement learning for collective animal movement. Annals of Applied Statistics, 16(2): 999–1013, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Bayesian-inverse-reinforcement-learning-
for-collective-animal-movement/10.1214/21-AOAS1529.
full.

Shen:2011:EAC

[SWHO11] Hui Shen, William J. Welch, and Jacqueline M. Hughes-Oliver. Efficient, adaptive cross-validation for tuning and comparing models, with application to drug discovery. Annals of Applied Statistics, 5(4):2668–2687, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399611.

Sampson:2014:LLI

[SWLS14] Joshua N. Sampson, Bill Wheeler, Peng Li, and Jianxin Shi. Leveraging local identity-by-descent increases the power of case/control GWAS with related individuals. Annals of Applied Statistics, 8(2):974–998, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229522.

Shen:2013:SIC

[SWM13] Ronglai Shen, Sijian Wang, and Qianxing Mo. Sparse integrative clustering of multiple omics data sets. Annals of Applied Statistics, 7(1):269–294, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527199.

Shabalin:2009:FLA

[SWPN09] Andrey A. Shabalin, Victor J. Weigman, Charles M. Perou, and Andrew B. Nobel. Finding large average submatrices in high dimensional data. Annals of Applied Statistics, 3(3):985–1012, September 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1254773275.

Sohn:2009:HDP

[SX09] Kyung-Ah Sohn and Eric P. Xing. A hierarchical Dirichlet process mixture model for haplotype reconstruction from multi-population data. Annals of Applied Statistics, 3(2): 791–821, June 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1245676195.

Sun:2020:LCR

[SXC^{+}20] Zhe Sun, Wanli Xu, Xiaomei Cong, Gen Li, and Kun Chen. Log-contrast regression with functional compositional predictors: Linking preterm infants’ gut microbiome trajectories to neurobehavioral outcome. Annals of Applied Statistics, 14(3):1535–1556, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Log-contrast-
regression-with-functional-compositional-predictors-
-Linking-preterm/10.1214/20-AOAS1357.full.

Sun:2023:MFS

[SXZ23] Hong Sun, Maochao Xu, and Peng Zhao. A multivariate frequency-severity framework for healthcare data breaches. Annals of Applied Statistics, 17(1):240–268, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-1/A-
multivariate-frequency-severity-framework-for-healthcare-
data-breaches/10.1214/22-AOAS1625.full.

Siegmund:2011:DSV

[SYZ11] David Siegmund, Benjamin Yakir, and Nancy R. Zhang. Detecting simultaneous variant intervals in aligned sequences. Annals of Applied Statistics, 5(2A):645–668, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562199.

Shang:2015:TSA

[SYZ15] Hongwei Shang, Jun Yan, and Xuebin Zhang. A two-step approach to model precipitation extremes in California based on max-stable and marginal point processes. Annals of Applied Statistics, 9(1):452–473, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226100.

Shen:2012:CPM

[SZ12] Jeremy J. Shen and Nancy R. Zhang. Change-point model on nonhomogeneous Poisson processes with application in copy number profiling by next-generation DNA sequencing. Annals of Applied Statistics, 6(2):476–496, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1339419604.

Shi:2020:RCL

[SZ20] Peng Shi and Zifeng Zhao. Regression for copula-linked compound distributions with applications in modeling aggregate insurance claims. Annals of Applied Statistics, 14(1):357– 380, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002678.

Slawski:2010:FSG

[SzCT10] Martin Slawski, Wolfgang zu Castell, and Gerhard Tutz. Feature selection guided by structural information. Annals of Applied Statistics, 4(2):1056–1080, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1280842152.

Shi:2016:RAM

[SZL16] Pixu Shi, Anru Zhang, and Hongzhe Li. Regression analysis for microbiome compositional data. Annals of Applied Statistics, 10(2):1019–1040, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199903.

Sun:2012:MHT

[SZO12] Yunting Sun, Nancy R. Zhang, and Art B. Owen. Multiple hypothesis testing adjusted for latent variables, with an application to the AGEMAP gene expression data. Annals of Applied Statistics, 6(4):1664–1688, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629055.

Thibaud:2016:BIB

[TAC^{+}16] Emeric Thibaud, Juha Aalto, Daniel S. Cooley, Anthony C. Davison, and Juha Heikkinen. Bayesian inference for the Brown–Resnick process, with an application to extreme low temperatures. Annals of Applied Statistics, 10(4):2303–2324, December 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1483606861.

Thaweethai:2021:RIW

[TACH21] Tanayott Thaweethai, David E. Arterburn, Karen J. Coleman, and Sebastien Haneuse. Robust inference when combining inverse-probability weighting and multiple imputation to address missing data with application to an electronic health records-based study of bariatric surgery. Annals of Applied Statistics, 15(1):126–147, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Robust-inference-when-combining-inverse-

probability-weighting-and-multiple-imputation/10.
1214/20-AOAS1386.full.

Taddy:2015:DMR

[Tad15] Matt Taddy. Distributed multinomial regression. Annals of Applied Statistics, 9(3):1394–1414, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1446488744.

Talih:2013:RIH

[Tal13] Makram Talih. A reference-invariant health disparity index based on Rényi divergence. Annals of Applied Statistics, 7(2): 1217–1243, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338485.

Talih:2015:ESH

[Tal15] Makram Talih. Examining socioeconomic health disparities using a rank-dependent Rényi index. Annals of Applied Statistics, 9(2):992–1023, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1437397121.

Tullio:2022:CIT

[TB22] Federico Tullio and Francesco Bartolucci. Causal inference for time-varying treatments in latent Markov models: An application to the effects of remittances on poverty dynamics. Annals of Applied Statistics, 16(3):1962–1985, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Causal-inference-for-time-varying-treatments-
in-latent-Markov-models/10.1214/21-AOAS1578.full.

Tiberi:2023:DNA

[TCS^{+}23] Simone Tiberi, Helena L. Crowell, Pantelis Samartsidis, Lukas M. Weber, and Mark D. Robinson. distinct: a novel approach to differential distribution analyses. Annals of Applied Statistics, 17(2):1681–1700, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/distinct-A-novel-
approach-to-differential-distribution-analyses/10.
1214/22-AOAS1689.full.

Tian:2021:EBC

[TCW21] Tian Tian, Ruihua Cheng, and Zhi Wei. An empirical Bayes change-point model for transcriptome timecourse data. Annals of Applied Statistics, 15(1):509–526, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/An-empirical-Bayes-change-point-model-
for-transcriptome-time-course/10.1214/20-AOAS1403.
full.

Tan:2016:TAV

[TCZ16] Linda S. L. Tan, Aik Hui Chan, and Tian Zheng. Topicadjusted visibility metric for scientific articles. Annals of Applied Statistics, 10(1):1–31, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1458909905.

Tang:2023:FSE

[TDBM23] Mingwei Tang, Gytis Dudas, Trevor Bedford, and Vladimir N. Minin. Fitting stochastic epidemic models to gene genealogies using linear noise approximation. Annals of Applied Statistics, 17(1):1–22, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Fitting-stochastic-
epidemic-models-to-gene-genealogies-using-linear-
noise/10.1214/21-AOAS1583.full.

Tang:2014:ITP

[TDS^{+}14] Shaowu Tang, Ying Ding, Etienne Sibille, Jeffrey S. Mogil, William R. Lariviere, and George C. Tseng. Imputation of truncated p-values for meta-analysis methods and its genomic application. Annals of Applied Statistics, 8(4):2150– 2174, December 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1419001738.

Tveten:2022:SCP

[TEF22] Martin Tveten, Idris A. Eckley, and Paul Fearnhead. Scalable change-point and anomaly detection in crosscorrelated data with an application to condition monitoring. Annals of Applied Statistics, 16(2):721–743, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Scalable-change-point-and-anomaly-detection-
in-cross-correlated-data/10.1214/21-AOAS1508.full.

Towe:2017:SDF

[TETJ17] Ross Towe, Emma Eastoe, Jonathan Tawn, and Philip Jonathan. Statistical downscaling for future extreme wave heights in the North Sea. Annals of Applied Statistics, 11(4):2375–2403, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1514430290.

Thomas:2011:DNR

[TF11] Andrew C. Thomas and Stephen E. Fienberg. Discussion of “Network routing in a dynamic environment”. Annals of Applied Statistics, 5(2B):1425–1427, June 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1310562727. See [Sin11].

Tom:2014:MLP

[TFB14] Brian D. M. Tom, Vernon T. Farewell, and Sheila M. Bird. Maximum likelihood and pseudo score approaches for parametric time-to-event analysis with informative entry times. Annals of Applied Statistics, 8(2):726–746, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229512.

Touloupou:2020:BIM

[TFB^{+}20] Panayiota Touloupou, Bärbel Finkenstädt, Thomas E. Besser, Nigel P. French, and Simon E. F. Spencer. Bayesian inference for multistrain epidemics with application to Escherichia coli O157:H7 in feedlot cattle. Annals of Applied Statistics, 14(4):1925–1944, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Bayesian-inference-for-multistrain-epidemics-
with-application-to-ESCHERICHIA-COLI/10.1214/20-AOAS1366.
full.

Tilahun:2012:DME

[TFG12] Gelila Tilahun, Andrey Feuerverger, and Michael Gervers. Dating medieval English charters. Annals of Applied Statistics, 6(4):1615–1640, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1356629053.

Thioulouse:2011:SAS

[Thi11] Jean Thioulouse. Simultaneous analysis of a sequence of paired ecological tables: a comparison of several methods. Annals of Applied Statistics, 5(4):2300–2325, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399596.

Tipton:2019:PPC

[THN^{+}19] John R. Tipton, Mevin B. Hooten, Connor Nolan, Robert K. Booth, and Jason McLachlan. Predicting paleoclimate from compositional data using multivariate Gaussian process inverse prediction. Annals of Applied Statistics, 13(4):2363– 2388, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910048.

Tian:2012:TWR

[THSL12] Tian Siva Tian, Jianhua Z. Huang, Haipeng Shen, and Zhimin Li. A two-way regularization method for MEG source reconstruction. Annals of Applied Statistics, 6(3):1021–1046, September 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1346418572.

Tibshirani:2008:DTA

[Tib08] Robert Tibshirani. Discussion of: “Treelets — An adaptive multi-scale basis for sparse unordered data”. Annals of Applied Statistics, 2(2):482–483, June 2008. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1215118522. See [LNW08b].

Tingley:2011:SPR

[Tin11] Martin P. Tingley. Spurious predictions with random time series: The Lasso in the context of paleoclimatic reconstructions. Discussion of: A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable? Annals of Applied Statistics, 5(1):83–87, March 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715180. See [MW11b].

Tan:2017:BIM

[TJDE17] Linda S. L. Tan, Ajay Jasra, Maria De Iorio, and Timothy M. D. Ebbels. Bayesian inference for multiple Gaussian graphical models with application to metabolic association networks. Annals of Applied Statistics, 11(4):2222–2251, December 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1514430284.

Tian:2010:MAS

[TJW10] Tian Siva Tian, Gareth M. James, and Rand R. Wilcox. A multivariate adaptive stochastic search method for dimensionality reduction in classification. Annals of Applied Statistics, 4(1):340–365, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1273584458.

Tancredi:2011:HBA

[TL11] Andrea Tancredi and Brunero Liseo. A hierarchical Bayesian approach to record linkage and population size problems. Annals of Applied Statistics, 5(2B):1553–1585, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562733.

Tian:2019:CPE

[TLF^{+}19] Lu Tian, Yi Liu, Andrew Z. Fire, Scott D. Boyd, and Richard A. Olshen. Clonality: Point estimation. Annals of Applied Statistics, 13(1):113–131, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1554861643.

Tzeng:2014:GLP

[TLH14] Jung-Ying Tzeng, Wenbin Lu, and Fang-Chi Hsu. Gene-level pharmacogenetic analysis on survival outcomes using genetrait similarity regression. Annals of Applied Statistics, 8(2):

1232–1255, June 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1404229532.

Trimbur:2022:MAI

[TM22] Thomas M. Trimbur and Tucker S. McElroy. Modelled approximations to the ideal filter with application to GDP and its components. Annals of Applied Statistics, 16(2):627– 651, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Modelled-approximations-to-the-ideal-filter-
with-application-to-GDP/10.1214/21-AOAS1463.full.

Tang:2018:PST

[TMN18] Yunfan Tang, Li Ma, and Dan L. Nicolae. A phylogenetic scan test on a Dirichlet-tree multinomial model for microbiome data. Annals of Applied Statistics, 12(1):1–26, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1520564463.

Telesca:2012:MDG

[TMPF12] Donatello Telesca, Peter Müller, Giovanni Parmigiani, and Ralph S. Freedman. Modeling dependent gene expression. Annals of Applied Statistics, 6(2):542–560, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1339419607.

Tak:2017:BEA

[TMvD^{+}17] Hyungsuk Tak, Kaisey Mandel, David A. van Dyk, Vinay L. Kashyap, Xiao-Li Meng, and Aneta Siemiginowska. Bayesian estimates of astronomical time delays between gravitationally lensed stochastic light curves. Annals of Applied Statistics, 11(3):1309–1348, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168831.

Tang:2017:BCP

[TMY17] Xiaoying Tang, Michael I. Miller, and Laurent Younes. Biomarker change-point estimation with right censoring in longitudinal studies. Annals of Applied Statistics, 11 (3):1738–1762, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168846.

Terada:2020:COP

[TON20] Yoshikazu Terada, Issei Ogasawara, and Ken Nakata. Classification from only positive and unlabeled functional data. Annals of Applied Statistics, 14(4):1724–1742, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Classification-from-only-positive-and-
unlabeled-functional-data/10.1214/20-AOAS1404.full.

Trippa:2011:FDR

[TP11] Lorenzo Trippa and Giovanni Parmigiani. False discovery rates in somatic mutation studies of cancer. Annals of Applied Statistics, 5(2B):1360–1378, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562724.

Thibaudeau:2017:MLL

[TSG17] Yves Thibaudeau, Eric Slud, and Alfred Gottschalck. Modeling log-linear conditional probabilities for estimation in surveys. Annals of Applied Statistics, 11(2):680–697, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1500537719.

Tom:2010:RRR

[TSS10] Jennifer A. Tom, Janet S. Sinsheimer, and Marc A. Suchard. Reuse, recycle, reweigh: Combating influenza through efficient sequential Bayesian computation for massive data. Annals of Applied Statistics, 4(4):1722–1748, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167796.

Tang:2022:CFM

[TSY22] Chen Tang, Han Lin Shang, and Yanrong Yang. Clustering and forecasting multiple functional time series. Annals of Applied Statistics, 16(4):2523–2553, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Clustering-and-forecasting-multiple-functional-
time-series/10.1214/22-AOAS1602.full.

Tibshirani:2009:BCM

[TT09] Ryan J. Tibshirani and Robert Tibshirani. A bias correction for the minimum error rate in cross-validation. Annals of Applied Statistics, 3(2):822–829, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1245676196.

Tansey:2022:BMD

[TTB22] Wesley Tansey, Christopher Tosh, and David M. Blei. A Bayesian model of dose-response for cancer drug studies. Annals of Applied Statistics, 16(2):680–705, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-16/issue-2/A-
Bayesian-model-of-dose-response-for-cancer-drug-studies/
10.1214/21-AOAS1485.full.

Tai:2021:BBH

[TTH21] An-Shun Tai, George C. Tseng, and Wen-Ping Hsieh. BayICE: a Bayesian hierarchical model for semireferencebased deconvolution of bulk transcriptomic data. Annals of Applied Statistics, 15(1):391–411, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/BayICE--A-Bayesian-
hierarchical-model-for-semireference-based-deconvolution/
10.1214/20-AOAS1376.full.

Tuglus:2008:DTA

[TvdL08] Catherine Tuglus and Mark J. van der Laan. Discussion of: “Treelets — An adaptive multi-scale basis for sparse unordered data”. Annals of Applied Statistics, 2(2):489– 493, June 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1215118524. See [LNW08b].

Thomas:2013:CPH

[TVJM13] A. C. Thomas, Samuel L. Ventura, Shane T. Jensen, and Stephen Ma. Competing process hazard function models for player ratings in ice hockey. Annals of Applied Statistics, 7(3):1497–1524, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804804.

Tao:2018:TBR

[TWA18] Yebin Tao, Lu Wang, and Daniel Almirall. Tree-based reinforcement learning for estimating optimal dynamic treatment regimes. Annals of Applied Statistics, 12(3):1914–1938, September 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1536652980.

Tu:2013:EOR

[TWH13] I-Ping Tu, Shao-Hsuan Wang, and Yuan-Fu Huang. Estimating the occurrence rate of DNA palindromes. Annals of Applied Statistics, 7(2):1095–1110, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338480.

Trippa:2015:BNC

[TWHP15] Lorenzo Trippa, Levi Waldron, Curtis Huttenhower, and Giovanni Parmigiani. Bayesian nonparametric cross-study validation of prediction methods. Annals of Applied Statistics, 9(1):402–428, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1430226098.

Tomal:2015:ECM

[TWZ15] Jabed H. Tomal, William J. Welch, and Ruben H. Zamar. Ensembling classification models based on phalanxes of variables with applications in drug discovery. Annals of Applied Statistics, 9(1):69–93, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1430226085.

Uppala:2020:MWI

[UH20] Medha Uppala and Mark S. Handcock. Modeling wildfire ignition origins in southern California using linear network point processes. Annals of Applied Statistics, 14(1):339– 356, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002677.

Ursino:2021:REM

[URZF21] Moreno Ursino, Christian Röver, Sarah Zohar, and Tim Friede. Random-effects meta-analysis of Phase I dosefinding studies using stochastic process priors. Annals of Applied Statistics, 15(1):174–193, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Random-effects-
meta-analysis-of-Phase-I-dose-finding-studies/10.
1214/20-AOAS1390.full.

Vittert:2019:HCB

[VBK19] Liberty Vittert, Adrian W. Bowman, and Stanislav Katina. A hierarchical curve-based approach to the analysis of manifold data. Annals of Applied Statistics, 13(4):2539–2563, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910054.

Virkar:2014:PLD

[VC14] Yogesh Virkar and Aaron Clauset. Power-law distributions in binned empirical data. Annals of Applied Statistics, 8(1): 89–119, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966280.

Viaud:2022:FBI

[VCC22] Gautier Viaud, Yuting Chen, and Paul-Henry Cournède. Full Bayesian inference in hidden Markov models of plant growth. Annals of Applied Statistics, 16(4):2352–2368, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Full-Bayesian-inference-in-hidden-Markov-
models-of-plant-growth/10.1214/21-AOAS1594.full.

vandenBroek:2009:UEP

[vdBN09] Jan van den Broek and Hiroshi Nishiura. Using epidemic prevalence data to jointly estimate reproduction and removal. Annals of Applied Statistics, 3(4):1505–1520, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1267453950.

vandenBrakel:2010:IAS

[vdBR10] Jan van den Brakel and Joeri Roels. Intervention analysis with state-space models to estimate discontinuities due to a survey redesign. Annals of Applied Statistics, 4(2):1105– 1138, June 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1280842154.

vanDyk:2009:SAS

[vDDS^{+}09] David A. van Dyk, Steven DeGennaro, Nathan Stein, William H. Jefferys, and Ted von Hippel. Statistical analysis of stellar evolution. Annals of Applied Statistics, 3(1): 117–143, March 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1239888365.

vanderHeijden:2012:PBM

[vdHWC^{+}12] Peter G. M. van der Heijden, Joe Whittaker, Maarten Cruyff, Bart Bakker, and Rik van der Vliet. People born in the Middle East but residing in the Netherlands: Invariant population size estimates and the role of active and passive covariates. Annals of Applied Statistics, 6(3):831–852, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1346418564.

vandeKassteele:2017:EEA

[vdKvEW17] Jan van de Kassteele, Jan van Eijkeren, and Jacco Wallinga. Efficient estimation of age-specific social contact rates between men and women. Annals of Applied Statistics, 11(1): 320–339, March 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1491616883.

Venturini:2008:GSM

[VDP08] Sergio Venturini, Francesca Dominici, and Giovanni Parmigiani. Gamma shape mixtures for heavy-tailed distributions. Annals of Applied Statistics, 2(2):756–776, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1215118537.

Vandendijck:2016:PTE

[VFH16] Yannick Vandendijck, Christel Faes, and Niel Hens. Prevalence and trend estimation from observational data with highly variable post-stratification weights. Annals of Applied Statistics, 10(1):94–117, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1458909909.

Venturini:2017:REM

[VFMD17] Sergio Venturini, Jessica M. Franklin, Laura Morlock, and Francesca Dominici. Random effects models for identifying the most harmful medication errors in a large, voluntary reporting database. Annals of Applied Statistics, 11(2):504– 526, June 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1500537713.

Vsevolozhskaya:2014:PCT

[VGH14] Olga Vsevolozhskaya, Mark Greenwood, and Dmitri Holodov. Pairwise comparison of treatment levels in functional analysis of variance with application to erythrocyte hemolysis. Annals of Applied Statistics, 8(2):905–925, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1404229519.

Volfovsky:2014:HAP

[VH14] Alexander Volfovsky and Peter D. Hoff. Hierarchical array priors for ANOVA decompositions of cross-classified data. Annals of Applied Statistics, 8(1):19–47, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1396966277.

Vu:2013:MBC

[VHS13] Duy Q. Vu, David R. Hunter, and Michael Schweinberger. Model-based clustering of large networks. Annals of Applied Statistics, 7(2):1010–1039, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338477.

Valentini:2013:MUH

[VIF13] Pasquale Valentini, Luigi Ippoliti, and Lara Fontanella. Modeling US housing prices by spatial dynamic structural equation models. Annals of Applied Statistics, 7(2):763–798, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1372338467.

Voulgaraki:2012:SRT

[VKG12] Anastasia Voulgaraki, Benjamin Kedem, and Barry I. Graubard. Semiparametric regression in testicular germ cell data. Annals of Applied Statistics, 6(3):1185–1208, September 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1346418579.

Viles:2021:ICH

[VML^{+}21] Weston D. Viles, Juliette C. Madan, Hongzhe Li, Margaret R. Karagas, and Anne G. Hoen. Information content of high-order associations of the human gut microbiota network. Annals of Applied Statistics, 15(4):1788–1807, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Information-content-of-high-order-associations-
of-the-human-gut/10.1214/21-AOAS1449.full.

Vu:2011:EDV

[VRN^{+}11] Vincent Q. Vu, Pradeep Ravikumar, Thomas Naselaris, Kendrick N. Kay, Jack L. Gallant, and Bin Yu. Encoding and decoding V1 fMRI responses to natural images with sparse nonparametric models. Annals of Applied Statistics, 5(2B): 1159–1182, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562717.

Vinci:2018:ARL

[VVSK18] Giuseppe Vinci, Valérie Ventura, Matthew A. Smith, and Robert E. Kass. Adjusted regularization in latent graphical models: Application to multiple-neuron spike count data. Annals of Applied Statistics, 12(2):1068–1095, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743486.

Wahl:2011:DSA

[WA11] Eugene R. Wahl and Caspar M. Ammann. Discussion of: “A statistical analysis of multiple temperature proxies: Are reconstructions of surface temperatures over the last 1000 years reliable?”. Annals of Applied Statistics, 5(1):91–95, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1300715182. See [MW11b].

Wang:2020:ISA

[WACY20] Wenjie Wang, Robert Aseltine, Kun Chen, and Jun Yan. Integrative survival analysis with uncertain event times in application to a suicide risk study. Annals of Applied Statistics, 14(1):51–73, March 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1587002664.

Waller:2014:DSA

[Wal14] Lance A. Waller. Discussion of “Spatial accessibility of pediatric primary healthcare: Measurement and inference”. Annals of Applied Statistics, 8(4):1956–1960, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1419001730. See [NSS14b].

Wang:2011:QMC

[Wan11] Yazhen Wang. Quantum Monte Carlo simulation. Annals of Applied Statistics, 5(2A):669–683, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562200.

Wong:2014:AEF

[WBA^{+}14] Raymond K. W. Wong, Paul Baines, Alexander Aue, Thomas C. M. Lee, and Vinay L. Kashyap. Automatic estimation of flux distributions of astrophysical source populations. Annals of Applied Statistics, 8(3):1690–1712, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1414091230.

Wang:2013:BAD

[WBB13] Xiaojing Wang, James O. Berger, and Donald S. Burdick. Bayesian analysis of dynamic item response models in educational testing. Annals of Applied Statistics, 7(1):126– 153, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527193.

Wager:2015:WSC

[WBC15] Stefan Wager, Alexander Blocker, and Niall Cardin. Weakly supervised clustering: Learning fine-grained signals from coarse labels. Annals of Applied Statistics, 9(2):801–820, June 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL http://projecteuclid.org/euclid.
aoas/1437397112.

Wu:2019:CIC

[WBK^{+}19] Xiao Wu, Danielle Braun, Marianthi-Anna Kioumourtzoglou, Christine Choirat, Qian Di, and Francesca Dominici. Causal inference in the context of an error prone exposure: Air pollution and mortality. Annals of Applied Statistics, 13(1):520–547, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861659.

Whiteman:2022:BIB

[WBKJ22] Andrew S. Whiteman, Andreas J. Bartsch, Jian Kang, and Timothy D. Johnson. Bayesian inference for brain activity from functional magnetic resonance imaging collected at two spatial resolutions. Annals of Applied Statistics, 16(4):2626– 2647, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Bayesian-inference-for-brain-activity-
from-functional-magnetic-resonance-imaging/10.1214/
22-AOAS1606.full.

Wee:2023:EGP

[WCD23] Damien C. H. Wee, Feng Chen, and William T. M. Dunsmuir. Estimating GARCH(1) in the presence of missing data. Annals of Applied Statistics, 17(3):2596–2618, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Estimating-GARCH11-in-the-presence-of-
missing-data/10.1214/23-AOAS1734.full.

Wang:2023:BNH

[WCH^{+}23] Emily T. Wang, Sharon Chiang, Zulfi Haneef, Vikram R. Rao, Robert Moss, and Marina Vannucci. Bayesian nonhomogeneous hidden Markov model with variable selection for investigating drivers of seizure risk cycling. Annals of Applied Statistics, 17(1):333–356, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Bayesian-non-
homogeneous-hidden-Markov-model-with-variable-selection-
for/10.1214/22-AOAS1630.full.

Wang:2023:TRB

[WCL23] Wanjie Wang, Eric Chen, and Hongzhe Li. Truncated rank-based tests for two-part models with excessive zeros and applications to microbiome data. Annals of Applied Statistics, 17(2):1663–1680, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Truncated-rank-
based-tests-for-two-part-models-with-excessive/10.
1214/22-AOAS1688.full.

Wesolowski:2015:NFE

[WCW15] Sergiusz Wesolowski, Robert J. Contreras, and Wei Wu. A new framework for Euclidean summary statistics in the neural spike train space. Annals of Applied Statistics, 9 (3):1278–1297, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1446488739.

Warren:2022:CWV

[WCW^{+}22] Joshua L. Warren, Howard H. Chang, Lauren K. Warren, Matthew J. Strickland, Lyndsey A. Darrow, and James A. Mulholland. Critical window variable selection for mixtures: Estimating the impact of multiple air pollutants on stillbirth. Annals of Applied Statistics, 16(3):1633–1652, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Critical-window-variable-selection-for-
mixtures--Estimating-the-impact/10.1214/21-AOAS1560.
full.

Wang:2010:GEV

[WD10] Xia Wang and Dipak K. Dey. Generalized extreme value regression for binary response data: an application to B2B electronic payments system adoption. Annals of Applied Statistics, 4(4):2000–2023, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1294167807.

Wang:2022:BFR

[WDL22] Guoqing Wang, Abhirup Datta, and Martin A. Lindquist. Bayesian functional registration of fMRI activation maps. Annals of Applied Statistics, 16(3):1676–1699, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Bayesian-functional-registration-of-fMRI-
activation-maps/10.1214/21-AOAS1562.full.

Wang:2023:RIC

[WDSJ23] Bingkai Wang, Suzanne M. Dufault, Dylan S. Small, and Nicholas P. Jewell. Randomization inference for clusterrandomized test-negative designs with application to dengue studies: Unbiased estimation, partial compliance, and stepped-wedge design. Annals of Applied Statistics, 17(2): 1592–1614, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Randomization-inference-for-cluster-randomized-
test-negative-designs-with-application/10.1214/22-
AOAS1684.full.

Weir:2007:RDP

[Wei07] Bruce S. Weir. The rarity of DNA profiles. Annals of Applied Statistics, 1(2):358–370, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1196438022.

Wen:2016:MQD

[Wen16] Xiaoquan Wen. Molecular QTL discovery incorporating genomic annotations using Bayesian false discovery rate control. Annals of Applied Statistics, 10(3):1619–1638, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL http://projecteuclid.org/euclid.
aoas/1475069621.

Won:2023:IMB

[WESVS23] Jung Yeon Won, Michael R. Elliott, Emma V. Sanchez-Vaznaugh, and Brisa N. Sánchez. Integrating multiple built environment data sources. Annals of Applied Statistics, 17(2):1722–1739, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Integrating-
multiple-built-environment-data-sources/10.1214/22-
AOAS1692.full.

Wang:2021:OSB

[WEWX21] Lin Wang, Jake Elmstedt, Weng Kee Wong, and Hongquan Xu. Orthogonal subsampling for big data linear regression. Annals of Applied Statistics, 15(3):1273–1290, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Orthogonal-subsampling-for-big-data-linear-
regression/10.1214/21-AOAS1462.full.

White:2022:SFD

[WFC^{+}22] Philip A. White, Henry Frye, Michael F. Christensen, Alan E. Gelfand, and John A. Silander. Spatial functional data modeling of plant reflectances. Annals of Applied Statistics, 16(3):1919–1936, September 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Spatial-functional-
data-modeling-of-plant-reflectances/10.1214/21-AOAS1576.
full.

Wang:2022:ESR

[WFH^{+}22] Zhengfan Wang, Miranda J. Fix, Lucia Hug, Anu Mishra, Danzhen You, Hannah Blencowe, Jon Wakefield, and Leontine Alkema. Estimating the stillbirth rate for 195 countries using a Bayesian sparse regression model with temporal smoothing. Annals of Applied Statistics, 16(4):2101–2121, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

16/issue-4/Estimating-the-stillbirth-rate-for-195-
countries-using-a-Bayesian/10.1214/21-AOAS1571.full.

Walsh:2023:BHM

[WFHZ23] Stephen Walsh, Marco A. R. Ferreira, David Higdon, and Stephanie Zick. A Bayesian hierarchical model framework to quantify uncertainty of tropical cyclone precipitation forecasts. Annals of Applied Statistics, 17(3):1984–2001, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-3/A-
Bayesian-hierarchical-model-framework-to-quantify-
uncertainty-of-tropical/10.1214/22-AOAS1703.full.

Wang:2019:TWC

[WFS19] Miaoyan Wang, Jonathan Fischer, and Yun S. Song. Threeway clustering of multi-tissue multi-individual gene expression data using semi-nonnegative tensor decomposition. Annals of Applied Statistics, 13(2):1103–1127, June 2019. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758439.

Wang:2023:LRB

[WG23] Yikai Wang and Ying Guo. LOCUS: a regularized blind source separation method with low-rank structure for investigating brain connectivity. Annals of Applied Statistics, 17(2): 1307–1332, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/LOCUS--A-regularized-blind-source-separation-
method-with-low/10.1214/22-AOAS1670.full.

Weber:2018:BAA

[WGL^{+}18a] Sebastian Weber, Andrew Gelman, Daniel Lee, Michael Betancourt, Aki Vehtari, and Amy Racine-Poon. Bayesian aggregation of average data: an application in drug development. Annals of Applied Statistics, 12(3):1583–1604, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1536652966.

Wu:2018:AWB

[WGL^{+}18b] Xiaowei Wu, Ting Guan, Dajiang J. Liu, Luis G. León Novelo, and Dipankar Bandyopadhyay. Adaptive-weight burden test for associations between quantitative traits and genotype data with complex correlations. Annals of Applied Statistics, 12(3):1558–1582, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652965.

Westveld:2011:MEM

[WH11] Anton H. Westveld and Peter D. Hoff. A mixed effects model for longitudinal relational and network data, with applications to international trade and conflict. Annals of Applied Statistics, 5(2A):843–872, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562208.

Wu:2021:BSJ

[WHAW21] Guohui Wu, Scott H. Holan, Alexis Avril, and Jonas Waldenström. A Bayesian semiparametric Jolly–Seber model with individual heterogeneity: an application to migratory mallards at stopover. Annals of Applied Statistics, 15(2):813– 830, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/A-Bayesian-semiparametric-JollySeber-model-
with-individual-heterogeneity--An/10.1214/20-AOAS1421.
full.

Wilson:2022:KMD

[WHC^{+}22] Ander Wilson, Hsiao-Hsien Leon Hsu, Yueh-Hsiu Mathilda Chiu, Robert O. Wright, Rosalind J. Wright, and Brent A. Coull. Kernel machine and distributed lag models for assessing windows of susceptibility to environmental mixtures in children’s health studies. Annals of Applied Statistics, 16(2): 1090–1110, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Kernel-machine-and-distributed-lag-models-
for-assessing-windows-of/10.1214/21-AOAS1533.full.

White:2013:DEH

[Whi13] Gentry White. Discussion of “Estimating the historical and future probabilities of large terrorist events” by Aaron Clauset and Ryan Woodard. Annals of Applied Statistics, 7(4):1876–1880, December 2013. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1387823298. See [CW13a].

Wang:2015:MFA

[WHLN15] Zhishi Wang, Qiuling He, Bret Larget, and Michael A. Newton. A multi-functional analyzer uses parameter constraints to improve the efficiency of model-based gene-set analysis. Annals of Applied Statistics, 9(1):225–246, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226091.

Wu:2015:BBM

[WHNW15] Guohui Wu, Scott H. Holan, Charles H. Nilon, and Christopher K. Wikle. Bayesian binomial mixture models for estimating abundance in ecological monitoring studies. Annals of Applied Statistics, 9(1):1–26, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1430226082.

Wu:2007:SFA

[WI07] Zhijin Wu and Rafael A. Irizarry. A statistical framework for the analysis of microarray probe-level data. Annals of Applied Statistics, 1(2):333–357, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1196438021.

Wilson:2010:BMS

[WIC^{+}10] Melanie A. Wilson, Edwin S. Iversen, Merlise A. Clyde, Scott C. Schmidler, and Joellen M. Schildkraut. Bayesian model search and multilevel inference for SNP association studies. Annals of Applied Statistics, 4(3):1342–1364, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409376.

Wang:2022:IQR

[WILW22] Tianying Wang, Iuliana Ionita-Laza, and Ying Wei. Integrated Quantile RAnk Test (iQRAT) for gene-level associations. Annals of Applied Statistics, 16(3):1423–1444, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-3/Integrated-Quantile-RAnk-Test-iQRAT-for-
gene-level-associations/10.1214/21-AOAS1548.full.

Witten:2011:CCS

[Wit11] Daniela M. Witten. Classification and clustering of sequencing data using a Poisson model. Annals of Applied Statistics, 5(4):2493–2518, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1324399604.

Wang:2015:IGG

[WJF^{+}15] Y. X. Rachel Wang, Keni Jiang, Lewis J. Feldman, Peter J. Bickel, and Haiyan Huang. Inferring gene–gene interactions and functional modules using sparse canonical correlation analysis. Annals of Applied Statistics, 9(1):300–323, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226094.

Wulsin:2016:NML

[WJL16] Drausin F. Wulsin, Shane T. Jensen, and Brian Litt. Nonparametric multi-level clustering of human epilepsy seizures. Annals of Applied Statistics, 10(2):667–689, June 2016. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1469199889.

Watson:2021:EAU

[WJT^{+}21] Joe Watson, Ruth Joy, Dominic Tollit, Sheila J. Thornton, and Marie Auger-Méthé. Estimating animal utilization distributions from multiple data types: a joint spatiotemporal point process framework. Annals of Applied Statistics, 15(4):1872–1896, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Estimating-animal-

utilization-distributions-from-multiple-data-types-
-A/10.1214/21-AOAS1472.full.

Woodworth:2010:ATV

[WK10] George Woodworth and Joseph Kadane. Age- and timevarying proportional hazards models for employment discrimination. Annals of Applied Statistics, 4(3):1139–1157, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409367.

Winderbaum:2015:FEP

[WKG^{+}15] Lyron J. Winderbaum, Inge Koch, Ove J. R. Gustafsson, Stephan Meding, and Peter Hoffmann. Feature extraction for proteomics imaging mass spectrometry data. Annals of Applied Statistics, 9(4):1973–1996, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453994187.

Wong:2016:DAC

[WKLvD16] Raymond K. W. Wong, Vinay L. Kashyap, Thomas C. M. Lee, and David A. van Dyk. Detecting abrupt changes in the spectra of high-energy astrophysical sources. Annals of Applied Statistics, 10(2):1107–1134, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1469199907.

White:2021:HIS

[WKR21] Philip A. White, Durban G. Keeler, and Summer Rupper. Hierarchical integrated spatial process modeling of monotone West Antarctic snow density curves. Annals of Applied Statistics, 15(2):556–571, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Hierarchical-
integrated-spatial-process-modeling-of-monotone-West-
Antarctic-snow/10.1214/21-AOAS1443.full.

Wei:2008:HST

[WL08a] Zhi Wei and Hongzhe Li. A hidden spatial-temporal Markov random field model for network-based analysis of time course gene expression data. Annals of Applied Statistics, 2(1):408– 429, March 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1206367827.

Wu:2008:CDA

[WL08b] Tong Tong Wu and Kenneth Lange. Coordinate descent algorithms for lasso penalized regression. Annals of Applied Statistics, 2(1):224–244, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1206367819.

Wu:2010:MVD

[WL10] Tong Tong Wu and Kenneth Lange. Multicategory vertex discriminant analysis for high-dimensional data. Annals of Applied Statistics, 4(4):1698–1721, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1294167795.

Wu:2022:EHG

[WL22] Qiuyu Wu and Xiangyu Luo. Estimating heterogeneous gene regulatory networks from zero-inflated single-cell expression data. Annals of Applied Statistics, 16(4):2183–2200, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Estimating-heterogeneous-gene-regulatory-
networks-from-zero-inflated-single-cell/10.1214/21-
AOAS1582.full.

West:2021:ASB

[WLA^{+}21] Brady T. West, Roderick J. Little, Rebecca R. Andridge, Philip S. Boonstra, Erin B. Ware, Anita Pandit, and Fernanda Alvarado-Leiton. Assessing selection bias in regression coefficients estimated from nonprobability samples with applications to genetics and demographic surveys. Annals of Applied Statistics, 15(3):1556–1581, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Assessing-selection-bias-in-regression-
coefficients-estimated-from-nonprobability-samples/
10.1214/21-AOAS1453.full.

Wang:2017:PCP

[WLG17] Chenkun Wang, Hai Liu, and Sujuan Gao. A penalized Cox proportional hazards model with multiple time-varying exposures. Annals of Applied Statistics, 11(1):185–201, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1491616877.

Wong:2018:ECS

[WLK18] Samuel W. K. Wong, Jun S. Liu, and S. C. Kou. Exploring the conformational space for protein folding with sequential Monte Carlo. Annals of Applied Statistics, 12(3):1628–1654, September 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1536652968.

Wang:2017:DPM

[WLL17] Jue Wang, Sheng Luo, and Liang Li. Dynamic prediction for multiple repeated measures and event time data: an application to Parkinson’s disease. Annals of Applied Statistics, 11(3):1787–1809, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168848.

Wang:2021:BRC

[WLM^{+}21] Zhenxun Wang, Lifeng Lin, Thomas Murray, James S. Hodges, and Haitao Chu. Bridging randomized controlled trials and single-arm trials using commensurate priors in arm-based network meta-analysis. Annals of Applied Statistics, 15(4):1767–1787, December 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Bridging-randomized-
controlled-trials-and-single-arm-trials-using-commensurate/
10.1214/21-AOAS1469.full.

Wang:2023:BAI

[WLML23] Ru Wang, Ye Liang, Zhuqi Miao, and Tieming Liu. Bayesian analysis for imbalanced positive-unlabelled diagnosis codes in electronic health records. Annals of Applied Statistics, 17(2): 1220–1238, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

17/issue-2/Bayesian-analysis-for-imbalanced-positive-
unlabelled-diagnosis-codes-in-electronic/10.1214/
22-AOAS1666.full.

Wong:2016:FDE

[WLP^{+}16] Raymond K. W. Wong, Thomas C. M. Lee, Debashis Paul, Jie Peng, and Alzheimer’s Disease Neuroimaging Initiative. Fiber direction estimation, smoothing and tracking in diffusion MRI. Annals of Applied Statistics, 10(3):1137–1156, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/

euclid.aoas/1475069599. See discussion [Sch16, Laz16, KL16] and rejoinder [WLPP16].

Wong:2016:RFD

[WLPP16] Raymond K. W. Wong, Thomas C. M. Lee, Debashis Paul, and Jie Peng. Rejoinder: “Fiber direction estimation, smoothing and tracking in diffusion MRI”. Annals of Applied Statistics, 10(3):1166–1169, September 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1475069603. See [WLP^{+}16].

Wraith:2014:UIP

[WMA^{+}14] Darren Wraith, Kerrie Mengersen, Clair Alston, Judith Rousseau, and Tareq Hussein. Using informative priors in the estimation of mixtures over time with application to aerosol particle size distributions. Annals of Applied Statistics, 8(1): 232–258, March 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1396966285.

Wang:2017:VEM

[WME17] Y. Samuel Wang, Ross L. Matsueda, and Elena A. Erosheva. A variational EM method for mixed membership models with multivariate rank data: an analysis of public policy preferences. Annals of Applied Statistics, 11(3):1452–1480, September 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1507168836.

Wang:2023:SBD

[WMGB23] Nanwei Wang, Hélène Massam, Xin Gao, and Laurent Briollais. The scalable birth-death MCMC algorithm for mixed graphical model learning with application to genomic data integration. Annals of Applied Statistics, 17(3):1958–1983, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/The-scalable-birthdeath-MCMC-algorithm-
for-mixed-graphical-model-learning/10.1214/22-AOAS1701.
full.

Worthington:2019:EAM

[WMKG19] Hannah Worthington, Rachel McCrea, Ruth King, and Richard Griffiths. Estimating abundance from multiple sampling capture-recapture data via a multi-state multi-period stopover model. Annals of Applied Statistics, 13(4):2043– 2064, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910035.

Warren:2021:SDL

[WMT^{+}21] Joshua L. Warren, Marie Lynn Miranda, Joshua L. Tootoo, Claire E. Osgood, and Michelle L. Bell. Spatial distributed lag data fusion for estimating ambient air pollution. Annals of Applied Statistics, 15(1):323–342, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Spatial-distributed-lag-data-fusion-for-
estimating-ambient-air-pollution/10.1214/20-AOAS1399.
full.

Wang:2011:RL

[WNRZ11] Sijian Wang, Bin Nan, Saharon Rosset, and Ji Zhu. Random lasso. Annals of Applied Statistics, 5(1):468–485, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1300715199.

Wang:2023:EEE

[WNX^{+}23] Shikun Wang, Jing Ning, Ying Xu, Ya-Chen Tina Shih, Yu Shen, and Liang Li. An extension of estimating equations to model longitudinal medical cost trajectory with Medicare claims data linked to SEER cancer registry. Annals of Applied Statistics, 17(1):881–899, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/An-extension-
of-estimating-equations-to-model-longitudinal-medical-
cost/10.1214/22-AOAS1659.full.

Wang:2014:RFR

[WNZK14] Xuejing Wang, Bin Nan, Ji Zhu, and Robert Koeppe. Regularized 3D functional regression for brain image data via Haar wavelets. Annals of Applied Statistics, 8(2):1045–1064, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229525.

Wang:2018:TDA

[WOC18] Yuan Wang, Hernando Ombao, and Moo K. Chung. Topological data analysis of single-trial electroencephalographic signals. Annals of Applied Statistics, 12(3):1506–1534, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1536652963.

Williams:2023:GFF

[WOH23] Jonathan P. Williams, Danica M. Ommen, and Jan Hannig. Generalized fiducial factor: An alternative to the Bayes factor for forensic identification of source problems. Annals of Applied Statistics, 17(1):378–402, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Generalized-
fiducial-factor--An-alternative-to-the-Bayes-factor/
10.1214/22-AOAS1632.full.

Worby:2016:RTT

[WOK^{+}16] Colin J. Worby, Philip D. O’Neill, Theodore Kypraios, Julie V. Robotham, Daniela De Angelis, Edward J. P. Cartwright, Sharon J. Peacock, and Ben S. Cooper. Reconstructing transmission trees for communicable diseases using densely sampled genetic data. Annals of Applied Statistics, 10(1):395–417, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1458909921.

Wei:2012:BJM

[WP12] Peng Wei and Wei Pan. Bayesian joint modeling of multiple gene networks and diverse genomic data to identify target genes of a transcription factor. Annals of Applied Statistics, 6(1):334–355, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1331043399.

Wang:2012:MIS

[WR12] Hao Wang and Jerome P. Reiter. Multiple imputation for sharing precise geographies in public use data. Annals of Applied Statistics, 6(1):229–252, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1331043395.

Wang:2016:DCC

[WR16] Song Wang and Karl Rohe. Discussion of “Coauthorship and citation networks for statisticians”. Annals of Applied Statistics, 10(4):1820–1826, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1483606838. See [JJ16a,

JJ16b].

Wilson:2014:MET

[WRNR14] Ander Wilson, Ana G. Rappold, Lucas M. Neas, and Brian J. Reich. Modeling the effect of temperature on ozone-related mortality. Annals of Applied Statistics, 8(3):1728–1749, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091232.

Wehbe:2015:RBR

[WRSS15] Leila Wehbe, Aaditya Ramdas, Rebecca C. Steorts, and Cosma Rohilla Shalizi. Regularized brain reading with shrinkage and smoothing. Annals of Applied Statistics, 9 (4):1997–2022, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1453994188.

Warton:2010:CNP

[WS10a] David I. Warton and Leah C. Shepherd. Correction note: Poisson point process models solve the “pseudo-absence problem” for presence-only data in ecology. Annals of Applied Statistics, 4(4):2203–2204, December 2010. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/

1294167816. See [WS10b].

Warton:2010:PPP

[WS10b] David I. Warton and Leah C. Shepherd. Poisson point process models solve the “pseudo-absence problem” for presence-only data in ecology. Annals of Applied Statistics, 4(3):1383–1402, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1287409378. See correc-

tion [WS10a].

Wen:2010:ULP

[WS10c] Xiaoquan Wen and Matthew Stephens. Using linear predictors to impute allele frequencies from summary or pooled genotype data. Annals of Applied Statistics, 4(3):1158–1182, September 2010. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1287409368.

Wen:2014:BMG

[WS14] Xiaoquan Wen and Matthew Stephens. Bayesian methods for genetic association analysis with heterogeneous subgroups: From meta-analyses to gene–environment interactions. Annals of Applied Statistics, 8(1):176–203, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1396966283.

Wang:2012:TML

[WSGH12] Hao Wang, Saul Shiffman, Sandra D. Griffith, and Daniel F. Heitjan. Truth and memory: Linking instantaneous and retrospective self-reported cigarette consumption. Annals of Applied Statistics, 6(4):1689–1706, December 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1356629056.

Wu:2014:DBH

[WSH^{+}14] Zheyang Wu, Yiming Sun, Shiquan He, Judy Cho, Hongyu Zhao, and Jiashun Jin. Detection boundary and Higher Criticism approach for rare and weak genetic effects. Annals of Applied Statistics, 8(2):824–851, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1404229516.

Wang:2021:SPI

[WSK^{+}21] Bruce Wang, Timothy Sudijono, Henry Kirveslahti, Tingran Gao, Douglas M. Boyer, Sayan Mukherjee, and Lorin Crawford. A statistical pipeline for identifying physical features that differentiate classes of 3D shapes. Annals of Applied Statistics, 15(2):638–661, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/A-statistical-pipeline-for-identifying-
physical-features-that-differentiate-classes/10.1214/
20-AOAS1430.full.

Wang:2016:ERP

[WSM^{+}16] Le Wang, Pamela A. Shaw, Hansie M. Mathelier, Stephen E. Kimmel, and Benjamin French. Evaluating risk-prediction models using data from electronic health records. Annals of Applied Statistics, 10(1):286–304, March 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1458909917.

Wang:2019:NMT

[WSU^{+}19] Y. X. Rachel Wang, Purnamrita Sarkar, Oana Ursu, Anshul Kundaje, and Peter J. Bickel. Network modelling of topological domains using Hi-C data. Annals of Applied Statistics, 13(3):1511–1536, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277762.

Witten:2008:TSF

[WT08] Daniela M. Witten and Robert Tibshirani. Testing significance of features by lassoed principal components. Annals of Applied Statistics, 2(3):986–1012, September 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1223908049.

Winter:2016:MEN

[WTB16] Hugo C. Winter, Jonathan A. Tawn, and Simon J. Brown. Modelling the effect of the El Niño–Southern Oscillation on extreme spatial temperature events over Australia. Annals of Applied Statistics, 10(4):2075–2101, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606852.

Wang:2010:NIP

[WTCW10] Rui Wang, Lu Tian, Tianxi Cai, and L. J. Wei. Nonparametric inference procedure for percentiles of the random effects distribution in meta-analysis. Annals of Applied Statistics, 4(1):520–532, March 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1273584466.

Wadsworth:2010:ACM

[WTJ10] J. L. Wadsworth, J. A. Tawn, and P. Jonathan. Accounting for choice of measurement scale in extreme value modeling. Annals of Applied Statistics, 4(3):1558–1578, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1287409386.

Wu:2022:MMH

[WWCZ22] Jing Wu, Owen G. Ward, James Curley, and Tian Zheng. Markov-modulated Hawkes processes for modeling sporadic and bursty event occurrences in social interactions. Annals of Applied Statistics, 16(2):1171–1190, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Markov-modulated-Hawkes-processes-for-
modeling-sporadic-and-bursty-event/10.1214/21-AOAS1539.
full.

Wang:2022:SBS

[WWL22] Jianqiao Wang, Wanjie Wang, and Hongzhe Li. Sparse block signal detection and identification for shared cross-trait association analysis. Annals of Applied Statistics, 16(2):866– 886, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Sparse-block-signal-detection-and-identification-
for-shared-cross-trait/10.1214/21-AOAS1523.full.

Wilson:2014:TBE

[WWM^{+}14] James D. Wilson, Simi Wang, Peter J. Mucha, Shankar Bhamidi, and Andrew B. Nobel. A testing based extraction algorithm for identifying significant communities in networks. Annals of Applied Statistics, 8(3):1853–1891, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091237.

Westgate:2013:TTE

[WWMH13] Bradford S. Westgate, Dawn B. Woodard, David S. Matteson, and Shane G. Henderson. Travel time estimation for ambulances using Bayesian data augmentation. Annals of Applied Statistics, 7(2):1139–1161, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338482.

Wang:2014:RES

[WYH^{+}14] Xin Wang, Ke Yuan, Christoph Hellmayr, Wei Liu, and Florian Markowetz. Reconstructing evolving signalling networks by hidden Markov nested effects models. Annals of Applied Statistics, 8(1):448–480, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1396966294.

Wu:2007:SAM

[WYKH07] Haiyan Wu, Ming Yuan, Susan M. Kaech, and M. Elizabeth Halloran. A statistical analysis of memory CD8 t cell differentiation: an application of a hierarchical state space model to a short time course microarray experiment. Annals of Applied Statistics, 1(2):442–458, December 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1196438026.

Wang:2022:CFC

[WYL^{+}22] Tianhao Wang, Lei Yu, Sue E. Leurgans, Robert S. Wilson, David A. Bennett, and Patricia A. Boyle. Conditional functional clustering for longitudinal data with heterogeneous nonlinear patterns. Annals of Applied Statistics, 16(2):1191– 1214, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

16/issue-2/Conditional-functional-clustering-for-
longitudinal-data-with-heterogeneous-nonlinear-patterns/
10.1214/21-AOAS1542.full.

Wang:2020:EMC

[WYT^{+}20] Yan Wang, Xiaowei Yue, Rui Tuo, Jeffrey H. Hunt, and Jianjun Shi. Effective model calibration via sensible variable identification and adjustment with application to composite fuselage simulation. Annals of Applied Statistics, 14(4):1759– 1776, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Effective-model-calibration-via-sensible-
variable-identification-and-adjustment-with/10.1214/
20-AOAS1353.full.

Wang:2023:VBA

[WYW^{+}23] Yaotian Wang, Guofen Yan, Xiaofeng Wang, Shuoran Li, Lingyi Peng, Dana L. Tudorascu, and Tingting Zhang. A variational Bayesian approach to identifying whole-brain directed networks with fMRI data. Annals of Applied Statistics, 17(1):518–538, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-17/issue-1/A-
variational-Bayesian-approach-to-identifying-whole-
brain-directed-networks/10.1214/22-AOAS1640.full.

Wu:2016:DBP

[WZ16] Han Wu and Yu Zhu. Deconvolution of base pair level RNA-Seq read counts for quantification of transcript expression levels. Annals of Applied Statistics, 10(3):1195–1216, September 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1475069605.

Wang:2017:SSS

[WZ17] Tao Wang and Hongyu Zhao. Structured subcomposition selection in regression and its application to microbiome data analysis. Annals of Applied Statistics, 11(2):771–791, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1500537723.

Wang:2018:BAC

[WZ18] Zheyu Wang and Xiao-Hua Zhou. Biomarker assessment and combination with differential covariate effects and an unknown gold standard, with an application to Alzheimer’s disease. Annals of Applied Statistics, 12(2):1204–1227, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743491.

Wang:2019:CIS

[WZD19] Lu Wang, Zhengwu Zhang, and David Dunson. Common and individual structure of brain networks. Annals of Applied Statistics, 13(1):85–112, March 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1554861642.

Wang:2013:MAS

[WZF^{+}13] Jiaping Wang, Hongtu Zhu, Jianqing Fan, Kelly Giovanello, and Weili Lin. Multiscale adaptive smoothing models for the hemodynamic response function in fMRI. Annals of Applied Statistics, 7(2):904–935, June 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1372338473.

Wu:2018:NBF

[WZF18] Mengyun Wu, Liping Zhu, and Xingdong Feng. Networkbased feature screening with applications to genome data. Annals of Applied Statistics, 12(2):1250–1270, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743493.

Wang:2012:TMD

[WZHC12] Yong Wang, Ilze Ziedins, Mark Holmes, and Neil Challands. Tree models for difference and change detection in a complex environment. Annals of Applied Statistics, 6(3):1162–1184, September 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1346418578.

Wei:2012:PCC

[WZL12] Ying Wei, Zhibiao Zhao, and Dennis K. J. Lin. Profile control charts based on nonparametric L − 1 regression methods. Annals of Applied Statistics, 6(1):409–427, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1331043402.

Wei:2020:GAR

[WZLP20] Bo Wei, Zhumin Zhang, HuiChuan J. Lai, and Limin Peng. Generalized accelerated recurrence time model in the presence of a dependent terminal event. Annals of Applied Statistics, 14(2):956–976, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1593449333.

Watson:2019:GTP

[WZS19] Joe Watson, James V. Zidek, and Gavin Shaddick. A general theory for preferential sampling in environmental networks. Annals of Applied Statistics, 13(4):2662–2700, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1574910059.

Xu:2023:LML

[XBS23] Zhixing Xu, Jonathan R. Bradley, and Debajyoti Sinha. Latent multivariate log-gamma models for high-dimensional MultiType responses with application to daily fine particulate matter and mortality counts. Annals of Applied Statistics, 17(2):1175–1198, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Latent-multivariate-
log-gamma-models-for-high-dimensional-MultiType-responses/
10.1214/22-AOAS1664.full.

Xu:2020:ALL

[XCCL20] Xiaofei Xu, Ying Chen, Cathy W. S. Chen, and Xiancheng Lin. Adaptive log-linear zero-inflated generalized Poisson autoregressive model with applications to crime counts. Annals of Applied Statistics, 14(3):1493–1515, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Adaptive-log-
linear-zero-inflated-generalized-Poisson-autoregressive-
model-with/10.1214/20-AOAS1360.full.

Xu:2011:BMO

[XCS11] Lizhen Xu, Radu V. Craiu, and Lei Sun. Bayesian methods to overcome the winner’s curse in genetic studies. Annals of Applied Statistics, 5(1):201–231, March 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1300715188.

Xiong:2015:VHU

[XDM15] Jie Xiong, D. P. Dittmer, and J. S. Marron. “Virus hunting” using radial distance weighted discrimination. Annals of Applied Statistics, 9(4):2090–2109, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453994193.

Xu:2010:ESC

[XDO10] Ya Xu, Justin S. Dyer, and Art B. Owen. Empirical stationary correlations for semi-supervised learning on graphs. Annals of Applied Statistics, 4(2):589–614, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1280842132.

Xing:2010:SSM

[XFS10] Eric P. Xing, Wenjie Fu, and Le Song. A state-space mixed membership blockmodel for dynamic network tomography. Annals of Applied Statistics, 4(2):535–566, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1280842130.

Xu:2019:SIP

[XKG^{+}19] Jason Xu, Samson Koelle, Peter Guttorp, Chuanfeng Wu, Cynthia Dunbar, Janis L. Abkowitz, and Vladimir N. Minin. Statistical inference for partially observed branching processes with application to cell lineage tracking of in vivo hematopoiesis. Annals of Applied Statistics, 13(4):2091–2119, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910037.

Xiao:2015:MSM

[XKS15] Sai Xiao, Athanasios Kottas, and Bruno Sansó. Modeling for seasonal marked point processes: An analysis of evolving hurricane occurrences. Annals of Applied Statistics, 9(1):353– 382, March 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1430226096.

Xie:2013:IEI

[XLDO13] Minge Xie, Regina Y. Liu, C. V. Damaraju, and William H. Olson. Incorporating external information in analyses of clinical trials with binary outcomes. Annals of Applied Statistics, 7(1):342–368, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1365527202.

Xu:2019:ICD

[XLS^{+}19] Huiping Xu, Xiaochun Li, Changyu Shen, Siu L. Hui, and Shaun Grannis. Incorporating conditional dependence in latent class models for probabilistic record linkage: Does it matter? Annals of Applied Statistics, 13(3):1753–1790, September 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1571277772.

Xie:2023:CCD

[XQ23] Xiulin Xie and Peihua Qiu. Control charts for dynamic process monitoring with an application to air pollution surveillance. Annals of Applied Statistics, 17(1):47–66, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Control-charts-for-dynamic-process-monitoring-
with-an-application-to/10.1214/22-AOAS1615.full.

Xu:2011:PPM

[XS11] Haiyong Xu and Frederic Paik Schoenberg. Point process modeling of wildfire hazard in Los Angeles County, California. Annals of Applied Statistics, 5(2A):684–704, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562201.

Xin:2017:CTS

[XZC17] Lu Xin, Mu Zhu, and Hugh Chipman. A continuous-time stochastic block model for basketball networks. Annals of Applied Statistics, 11(2):553–597, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1500537715.

Xie:2021:INL

[XZW21] Shanghong Xie, Donglin Zeng, and Yuanjia Wang. Integrative network learning for multimodality biomarker data. Annals of Applied Statistics, 15(1):64–87, March 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Integrative-
network-learning-for-multimodality-biomarker-data/
10.1214/20-AOAS1382.full.

Xie:2018:BBD

[XZX18] Fangzheng Xie, Mingyuan Zhou, and Yanxun Xu. BayCount: a Bayesian decomposition method for inferring tumor heterogeneity using RNA-Seq counts. Annals of Applied Statistics, 12(3):1605–1627, September 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1536652967.

Yuan:2017:PPM

[YBL^{+}17] Yuan Yuan, Fabian E. Bachl, Finn Lindgren, David L. Borchers, Janine B. Illian, Stephen T. Buckland, Håvard Rue, and Tim Gerrodette. Point process models for spatiotemporal distance sampling data from a large-scale survey of blue whales. Annals of Applied Statistics, 11(4):2270– 2297, December 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1514430286.

Yang:2023:GTD

[YD23] Yang Yang and Ke Deng. Generalized theme dictionary models for association pattern discovery. Annals of Applied Statistics, 17(1):269–293, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Generalized-
theme-dictionary-models-for-association-pattern-discovery/
10.1214/22-AOAS1626.full.

Yao:2014:SAI

[YE14] Zhigang Yao and William F. Eddy. A statistical approach to the inverse problem in magnetoencephalography. Annals of Applied Statistics, 8(2):1119–1144, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1404229528.

Yao:2020:QTV

[YFHE20] Zhigang Yao, Zengyan Fan, Masahito Hayashi, and William F. Eddy. Quantifying time-varying sources in magnetoencephalography — a discrete approach. Annals of Applied Statistics, 14(3):1379–1408, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Quantifying-
time-varying-sources-in-magnetoencephalographyA-discrete-
approach/10.1214/19-AOAS1321.full.

Yuan:2019:IEP

[YFM19] Lo-Hua Yuan, Avi Feller, and Luke W. Miratrix. Identifying and estimating principal causal effects in a multi-site trial of Early College High Schools. Annals of Applied Statistics, 13(3):1348–1369, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277756.

Yan:2020:SBM

[YGAT20] Dongyan Yan, Subharup Guha, Chul Ahn, and Ram Tiwari. Semiparametric Bayesian Markov analysis of personalized benefit-risk assessment. Annals of Applied Statistics, 14(2): 768–788, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449325.

Yang:2008:BFE

[YGLH08] Yang Yang, Peter Gilbert, Ira M. Longini, Jr., and M. Elizabeth Halloran. A Bayesian framework for estimating vaccine efficacy per infectious contact. Annals of Applied Statistics, 2(4):1409–1431, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1231424216.

Yau:2013:DTA

[YH13] Christopher Yau and Christopher C. Holmes. A decisiontheoretic approach for segmental classification. Annals of Applied Statistics, 7(3):1814–1835, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1380804818.

Yanchenko:2020:HMS

[YH20] Anna K. Yanchenko and Peter D. Hoff. Hierarchical multidimensional scaling for the comparison of musical performance styles. Annals of Applied Statistics, 14(4):1581–1603, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/Hierarchical-multidimensional-scaling-
for-the-comparison-of-musical-performance-styles/
10.1214/20-AOAS1391.full.

Ye:2020:JEF

[YHE20] Zi Ye, Giles Hooker, and Stephen P. Ellner. The Jensen effect and functional single index models: Estimating the ecological implications of nonlinear reaction norms. Annals of Applied Statistics, 14(3):1326–1341, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-3/The-Jensen-effect-and-functional-single-
index-models--Estimating/10.1214/20-AOAS1349.full.

Ye:2013:HDH

[YHX13] Zhi-Sheng Ye, Yili Hong, and Yimeng Xie. How do heterogeneities in operating environments affect field failure predictions and test planning? Annals of Applied Statistics, 7(4):2249–2271, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1387823318.

Yang:2021:RDM

[YJD21] Yang Yang, Chunlin Ji, and Ke Deng. Rapid design of metamaterials via multitarget Bayesian optimization. Annals of Applied Statistics, 15(2):768–796, June 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Rapid-design-
of-metamaterials-via-multitarget-Bayesian-optimization/
10.1214/20-AOAS1426.full.

Yuan:2009:SVS

[YJZ09] Ming Yuan, V. Roshan Joseph, and Hui Zou. Structured variable selection and estimation. Annals of Applied Statistics, 3(4):1738–1757, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1267453962.

Yu:2021:DGS

[YKHS21] Menggang Yu, Chensheng Kuang, Jared D. Huling, and Maureen Smith. Diagnosis-group-specific transitional care program recommendations for 30-day rehospitalization reduction. Annals of Applied Statistics, 15(3):1478–1498, September 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-3/Diagnosis-group-specific-transitional-
care-program-recommendations-for-30-day/10.1214/21-
AOAS1473.full.

Yu:2023:RMC

[YKLK23] Ruoqi Yu, Rachel Kelz, Scott Lorch, and Luke J. Keele. The risk of maternal complications after cesarean delivery: Near-far matching for instrumental variables study designs with large observational datasets. Annals of Applied Statistics, 17(2):1701–1721, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/The-risk-of-
maternal-complications-after-cesarean-delivery--Near/
10.1214/22-AOAS1691.full.

Yin:2011:SCG

[YL11] Jianxin Yin and Hongzhe Li. A sparse conditional Gaussian graphical model for analysis of genetical genomics data. Annals of Applied Statistics, 5(4):2630–2650, December 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1324399609.

Yang:2013:RPL

[YL13] Jingyuan Yang and Shili Lin. Robust partial likelihood approach for detecting imprinting and maternal effects using case-control families. Annals of Applied Statistics, 7(1):249– 268, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527198.

Yen:2017:ELN

[YLC^{+}17] Tso-Jung Yen, Zong-Rong Lee, Yi-Hau Chen, Yu-Min Yen, and Jing-Shiang Hwang. Estimating links of a network from time to event data. Annals of Applied Statistics, 11(3):1429–1451, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168835.

Ye:2015:JML

[YLG15] Jun Ye, Yehua Li, and Yongtao Guan. Joint modeling of longitudinal drug using pattern and time to first relapse in cocaine dependence treatment data. Annals of Applied Statistics, 9(3):1621–1642, September 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1446488754.

Yang:2007:RBT

[YLH07] Yang Yang, Ira M. Longini, and M. Elizabeth Halloran. A resampling-based test to detect person-to-person transmission of infectious disease. Annals of Applied Statistics, 1(1): 211–228, June 2007. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1183143736.

Yan:2017:DPD

[YLH17] Fangrong Yan, Xiao Lin, and Xuelin Huang. Dynamic prediction of disease progression for leukemia patients by functional principal component analysis of longitudinal expression levels of an oncogene. Annals of Applied Statistics, 11(3):1649–1670, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168843.

Yue:2012:MAF

[YLL12] Yu Ryan Yue, Martin A. Lindquist, and Ji Meng Loh. Metaanalysis of functional neuroimaging data using Bayesian nonparametric binary regression. Annals of Applied Statistics, 6(2):697–718, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419613.

Yang:2021:MFE

[YLLS21] Jenny Yang, Yang Liu, Yufeng Liu, and Wei Sun. Model free estimation of graphical model using gene expression data. Annals of Applied Statistics, 15(1):194–207, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/Model-free-estimation-of-graphical-model-
using-gene-expression-data/10.1214/20-AOAS1380.full.

Yang:2014:ECE

[YLS14] Fan Yang, Scott A. Lorch, and Dylan S. Small. Estimation of causal effects using instrumental variables with nonignorable missing covariates: Application to effect of type of delivery NICU on premature infants. Annals of Applied Statistics, 8(1):48–73, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966278.

Yu:2011:BSC

[YMP11] Qingzhao Yu, Steven N. MacEachern, and Mario Peruggia. Bayesian synthesis: Combining subjective analyses, with an application to ozone data. Annals of Applied Statistics, 5(2B):1678–1698, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1310562738.

Ye:2014:AIF

[YN14] Zhisheng Ye and Hon Keung Tony Ng. On analysis of incomplete field failure data. Annals of Applied Statistics, 8(3):1713–1727, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1414091231.

Yao:2023:BHB

[YOZC23] Yujing Yao, R. Todd Ogden, Chubing Zeng, and Qixuan Chen. Bivariate hierarchical Bayesian model for combining summary measures and their uncertainties from multiple sources. Annals of Applied Statistics, 17(2):1782– 1800, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Bivariate-hierarchical-Bayesian-model-
for-combining-summary-measures-and-their/10.1214/
22-AOAS1699.full.

Yu:2021:SVM

[YR21] Xingchen Yu and Abel Rodríguez. Spatial voting models in circular spaces: a case study of the U.S. House of Representatives. Annals of Applied Statistics, 15(4):1897–1922, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Spatial-voting-models-in-circular-spaces-
-A-case-study/10.1214/21-AOAS1454.full.

Young:2017:MBC

[YRY17] William Chad Young, Adrian E. Raftery, and Ka Yee Yeung. Model-based clustering with data correction for removing artifacts in gene expression data. Annals of Applied Statistics, 11(4):1998–2026, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1514430275.

Ylitalo:2016:WWL

[YSG16] Anna-Kaisa Ylitalo, Aila Särkkä, and Peter Guttorp. What we look at in paintings: a comparison between experienced and inexperienced art viewers. Annals of Applied Statistics, 10(2):549–574, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1469199884.

Yarger:2022:FDA

[YSH22]  Drew Yarger, Stilian Stoev, and Tailen Hsing.  A

functional-data approach to the Argo data. Annals of Applied Statistics, 16(1):216–246, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/journals/
annals-of-applied-statistics/volume-16/issue-1/A-
functional-data-approach-to-the-Argo-data/10.1214/
21-AOAS1477.full.

Yu:2008:BMA

[YSL08] Qingzhao Yu, Elizabeth A. Stasny, and Bin Li. Bayesian models to adjust for response bias in survey data for estimating rape and domestic violence rates from the NCVS. Annals of Applied Statistics, 2(2):665–686, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1215118533.

Yu:2021:ICI

[YSR21] Ruoqi Yu, Dylan S. Small, and Paul R. Rosenbaum. The information in covariate imbalance in studies of hormone replacement therapy. Annals of Applied Statistics, 15(4):2023– 2042, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/The-information-in-covariate-imbalance-
in-studies-of-hormone-replacement/10.1214/21-AOAS1448.
full.

Ye:2022:DPF

[YSR22] Ting Ye, Dylan S. Small, and Paul R. Rosenbaum. Dimensions, power and factors in an observational study of behavioral problems after physical abuse of children. Annals of Applied Statistics, 16(4):2732–2754, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Dimensions-power-and-factors-in-an-observational-
study-of-behavioral/10.1214/22-AOAS1611.full.

Yuan:2018:DAU

[YTHY18] Miao Yuan, Cheng Yong Tang, Yili Hong, and Jian Yang. Disentangling and assessing uncertainties in multiperiod corporate default risk predictions. Annals of Applied Statistics, 12(4):2587–2617, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078057.

Yu:2010:RL

[Yu10] Bin Yu. Remembering Leo. Annals of Applied Statistics, 4(4):1657–1659, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167793.

Yuan:2009:SPD

[Yua09] Ming Yuan. State price density estimation via nonparametric mixtures. Annals of Applied Statistics, 3(3):963–984, September 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1254773274.

Yao:2023:PLT

[YWB^{+}23] Tsung-Hung Yao, Zhenke Wu, Karthik Bharath, Jinju Li, and Veerabhadran Baladandayuthapani. Probabilistic learning of treatment trees in cancer. Annals of Applied Statistics, 17(3):1884–1908, September 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Probabilistic-
learning-of-treatment-trees-in-cancer/10.1214/22-
AOAS1696.full.

Yan:2012:SMT

[YWL^{+}12] Donghui Yan, Pei Wang, Michael Linden, Beatrice Knudsen, and Timothy Randolph. Statistical methods for tissue array images — algorithmic scoring and co-training. Annals of Applied Statistics, 6(3):1280–1305, September 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1346418583.

Yan:2022:BBC

[YWLL22] Han Yan, Jiexing Wu, Yang Li, and Jun S. Liu. Bayesian bi-clustering methods with applications in computational biology. Annals of Applied Statistics, 16(4):2804–2831, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Bayesian-bi-clustering-methods-with-applications-
in-computational-biology/10.1214/22-AOAS1622.full.

Yu:2023:RJM

[YWQG23] Tingting Yu, Lang Wu, Jin Qiu, and Peter B. Gilbert. Robust joint modelling of left-censored longitudinal data and survival data with application to HIV vaccine studies. Annals of Applied Statistics, 17(2):1017–1037, June 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-2/Robust-joint-
modelling-of-left-censored-longitudinal-data-and-
survival/10.1214/22-AOAS1656.full.

Yuan:2011:BPI

[YY11] Ying Yuan and Guosheng Yin. Bayesian phase I/II adaptively randomized oncology trials with combined drugs. Annals of Applied Statistics, 5(2A):924–942, June 2011. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1310562211.

Yu:2013:LTI

[YZAD13] Tao Yu, Chunming Zhang, Andrew L. Alexander, and Richard J. Davidson. Local tests for identifying anisotropic diffusion areas in human brain with DTI. Annals of Applied Statistics, 7(1):201–225, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527196.

Yuan:2013:VCM

[YZS^{+}13] Ying Yuan, Hongtu Zhu, Martin Styner, John H. Gilmore, and J. S. Marron. Varying coefficient model for modeling diffusion tensors along white matter tracts. Annals of Applied Statistics, 7(1):102–125, March 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1365527192.

Zanella:2015:RPM

[Zan15] Giacomo Zanella. Random partition models and complementary clustering of Anglo-Saxon place-names. Annals of Applied Statistics, 9(4):1792–1822, December 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1453993094.

Zhang:2012:MAL

[ZASM12] Zhiwei Zhang, Paul S. Albert, and Bruce Simons-Morton. Marginal analysis of longitudinal count data in long sequences: Methods and applications to a driving study. Annals of Applied Statistics, 6(1):27–54, March 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1331043387.

Zigler:2011:PBP

[ZB11] Corwin M. Zigler and Thomas R. Belin. The potential for bias in principal causal effect estimation when treatment received depends on a key covariate. Annals of Applied Statistics, 5(3):1876–1892, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514288.

Zhai:2016:IRP

[ZBC16] Yongliang Zhai and Alexandre Bouchard-Côté. Inferring rooted population trees using asymmetric neighbor joining. Annals of Applied Statistics, 10(4):2047–2074, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1483606851.

Zhu:2014:IGW

[ZBG14] Xinxin Zhu, Kenneth P. Bowman, and Marc G. Genton. Incorporating geostrophic wind information for improved space–time short-term wind speed forecasting. Annals of Applied Statistics, 8(3):1782–1799, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1414091234.

Zheng:2017:IEB

[ZBLC17] Yingye Zheng, Marshall Brown, Anna Lok, and Tianxi Cai. Improving efficiency in biomarker incremental value evaluation under two-phase designs. Annals of Applied Statistics, 11(2):638–654, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1500537717.

Zhou:2020:EII

[ZBT^{+}20] Shuo Zhou, Howard Bondell, Antoinette Tordesillas, Benjamin I. P. Rubinstein, and James Bailey. Early identification of an impending rockslide location via a spatially-aided Gaussian mixture model. Annals of Applied Statistics, 14(2): 977–992, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449334.

Zheng:2020:APC

[ZCD^{+}20] Yingye Zheng, Douglas A. Corley, Chyke Doubeni, Ethan Halm, Susan M. Shortreed, William E. Barlow, Ann Zauber, Tor Devin Tosteson, and Jessica Chubak. Analyses of preventive care measures with incomplete historical data in electronic medical records: an example from colorectal cancer screening. Annals of Applied Statistics, 14(2):1030– 1044, June 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1593449336.

Zhou:2009:SMA

[ZCG^{+}09] Xi Kathy Zhou, Merlise A. Clyde, James Garrett, Viridiana Lourdes, Michael O’Connell, Giovanni Parmigiani, David J. Turner, and Tim Wiles. Statistical methods for automated drug susceptibility testing: Bayesian minimum inhibitory concentration prediction from growth curves. Annals of Applied Statistics, 3(2):710–730, June 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1245676192.

Zhang:2021:ARW

[ZCGC21] Jiachen Zhang, Paola Crippa, Marc G. Genton, and Stefano Castruccio. Assessing the reliability of wind power operations under a changing climate with a non-Gaussian bias correction. Annals of Applied Statistics, 15(4):1831–1849, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-4/Assessing-the-reliability-of-wind-power-
operations-under-a-changing/10.1214/21-AOAS1460.full.

Zuo:2016:HFS

[ZCH^{+}16] Chandler Zuo, Kailei Chen, Kyle J. Hewitt, Emery H. Bresnick, and Sündüz Keles. A hierarchical framework for statespace matrix inference and clustering. Annals of Applied Statistics, 10(3):1348–1372, September 2016. CODEN ????

ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1475069610.

Zhang:2011:NMM

[ZCM^{+}11] Saijuan Zhang, Raymond J. Carroll, Douglas Midthune, Patricia M. Guenther, Susan M. Krebs-Smith, Victor Kipnis, Kevin W. Dodd, Dennis W. Buckman, Janet A. Tooze, and Laurence Freedman. A new multivariate measurement error model with zero-inflated dietary data, and its application to dietary assessment. Annals of Applied Statistics, 5(2B): 1456–1487, June 2011. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1310562729.

Zhu:2014:PRM

[ZCP14] Fukang Zhu, Zongwu Cai, and Liang Peng. Predictive regressions for macroeconomic data. Annals of Applied Statistics, 8(1):577–594, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966299.

Zhang:2018:ANC

[ZCRC18] Hong Zhang, Nilanjan Chatterjee, Daniel Rader, and Jinbo Chen. Adjustment of nonconfounding covariates in casecontrol genetic association studies. Annals of Applied Statistics, 12(1):200–221, March 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1520564470.

Zhang:2013:BLM

[ZCS13] Ran Zhang, Claudia Czado, and Karin Sigloch. A Bayesian linear model for the high-dimensional inverse problem of seismic tomography. Annals of Applied Statistics, 7(2):1111– 1138, June 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1372338481.

Zhang:2013:PTA

[ZD13] Yuping Zhang and Ronald Davis. Principal trend analysis for time-course data with applications in genomic medicine. Annals of Applied Statistics, 7(4):2205–2228, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1387823316.

Zhou:2010:SAM

[ZDL10] Yijie Zhou, Francesca Dominici, and Thomas A. Louis. A smoothing approach for masking spatial data. Annals of Applied Statistics, 4(3):1451–1475, September 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1287409381.

Zaman:2014:BAP

[ZFB14] Tauhid Zaman, Emily B. Fox, and Eric T. Bradlow. A Bayesian approach for predicting the popularity of tweets. Annals of Applied Statistics, 8(3):1583–1611, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1414091226.

Zhang:2022:BSD

[ZGJ^{+}22] Boya Zhang, Robert B. Gramacy, Leah R. Johnson, Kenneth A. Rose, and Eric Smith. Batch-sequential design and heteroskedastic surrogate modeling for delta smelt conservation. Annals of Applied Statistics, 16(2):816–842, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-2/Batch-sequential-design-and-heteroskedastic-
surrogate-modeling-for-delta-smelt/10.1214/21-AOAS1521.
full.

Zhang:2013:EDN

[ZGLH13] Lixun Zhang, Yongtao Guan, Brian P. Leaderer, and Theodore R. Holford. Estimating daily nitrogen dioxide level: Exploring traffic effects. Annals of Applied Statistics, 7(3):1763–1777, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804815.

Zhang:2023:MNB

[ZGM23] Wenyu Zhang, Maryclare Griffin, and David S. Matteson. Modeling a nonlinear biophysical trend followed by longmemory equilibrium with unknown change point. Annals of Applied Statistics, 17(1):860–880, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Modeling-a-nonlinear-
biophysical-trend-followed-by-long-memory-equilibrium/
10.1214/22-AOAS1655.full.

Zipunnikov:2014:LHD

[ZGS^{+}14] Vadim Zipunnikov, Sonja Greven, Haochang Shou, Brian S. Caffo, Daniel S. Reich, and Ciprian M. Crainiceanu. Longitudinal high-dimensional principal components analysis with application to diffusion tensor imaging of multiple sclerosis. Annals of Applied Statistics, 8(4):2175–2202, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1419001739.

Zhang:2016:SNB

[ZGV^{+}16] Linlin Zhang, Michele Guindani, Francesco Versace, Jeffrey M. Engelmann, and Marina Vannucci. A spatiotemporal nonparametric Bayesian model of multi-subject fMRI data. Annals of Applied Statistics, 10(2):638–666, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1469199888.

Zhang:2009:SAS

[ZHB09] Yufen Zhang, James S. Hodges, and Sudipto Banerjee. Smoothed ANOVA with spatial effects as a competitor to MCAR in multivariate spatial smoothing. Annals of Applied Statistics, 3(4):1805–1830, December 2009. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1267453965.

Zhang:2023:EAT

[ZHFN23] Zhiwei Zhang, Zonghui Hu, Dean Follmann, and Lei Nie. Estimating the average treatment effect in randomized clinical trials with all-or-none compliance. Annals of Applied Statistics, 17(1):294–312, March 2023. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Estimating-the-
average-treatment-effect-in-randomized-clinical-trials-
with/10.1214/22-AOAS1627.full.

Zhou:2015:MCL

[ZHJZ15] Haiming Zhou, Timothy Hanson, Alejandro Jara, and Jiajia Zhang. Modeling county level breast cancer survival data using a covariate-adjusted frailty proportional hazards model. Annals of Applied Statistics, 9(1):43–68, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226084.

Zhu:2019:BIV

[ZHM^{+}19] Li Zhu, Zhiguang Huo, Tianzhou Ma, Steffi Oesterreich, and George C. Tseng. Bayesian indicator variable selection to incorporate hierarchical overlapping group structure in multiomics applications. Annals of Applied Statistics, 13(4):2611– 2636, December 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1574910057.

Zhou:2017:DEB

[Zho17a] Chen Zhou. Discussion on “Elicitability and backtesting: Perspectives for banking regulation”. Annals of Applied Statistics, 11(4):1888–1893, December 2017. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/

1514430269. See [NZ17a].

Zhou:2017:UFV

[Zho17b] Xiang Zhou. A unified framework for variance component estimation with summary statistics in genome-wide association studies. Annals of Applied Statistics, 11(4):2027–2051, December 2017. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1514430276.

Zhong:2022:MNT

[ZHO22] Peng Zhong, Raphaël Huser, and Thomas Opitz. Modeling nonstationary temperature maxima based on extremal dependence changing with event magnitude. Annals of Applied Statistics, 16(1):272–299, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Modeling-nonstationary-

temperature-maxima-based-on-extremal-dependence-changing-
with/10.1214/21-AOAS1504.full.

Zhang:2023:SDC

[ZHYS23] Bo Zhang, Siyu Heng, Ting Ye, and Dylan S. Small. Social distancing and COVID-19: Randomization inference for a structured dose-response relationship. Annals of Applied Statistics, 17(1):23–46, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-1/Social-distancing-
and-COVID-19--Randomization-inference-for-a/10.1214/
22-AOAS1613.full.

Zhao:2021:BDS

[ZJBS21] Qingyuan Zhao, Nianqiao Ju, Sergio Bacallado, and Rajen D. Shah. BETS: The dangers of selection bias in early analyses of the coronavirus disease (COVID-19) pandemic. Annals of Applied Statistics, 15(1):363–390, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-1/BETS--The-dangers-of-selection-bias-in-
early-analyses/10.1214/20-AOAS1401.full.

Zhang:2008:QMR

[ZJLC08] Xiaoxi Zhang, Timothy D. Johnson, Roderick J. A. Little, and Yue Cao. Quantitative magnetic resonance image analysis via the EM algorithm with stochastic variation. Annals of Applied Statistics, 2(2):736–755, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1215118536.

Zhang:2010:NID

[ZK10] Tingting Zhang and S. C. Kou. Nonparametric inference of doubly stochastic Poisson process data via the kernel method. Annals of Applied Statistics, 4(4):1913–1941, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1294167804.

Zhang:2015:END

[ZKS15] Yaonan Zhang, Eric D. Kolaczyk, and Bruce D. Spencer. Estimating network degree distributions under sampling: an inverse problem, with applications to monitoring social media networks. Annals of Applied Statistics, 9(1):166–199, March 2015. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1430226089.

Zhao:2014:BNM

[ZKY14] Yize Zhao, Jian Kang, and Tianwei Yu. A Bayesian nonparametric mixture model for selecting genes and gene subnetworks. Annals of Applied Statistics, 8(2):999–1021, June 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1404229523.

Zhou:2012:BMA

[ZLD12] Xi Kathy Zhou, Fei Liu, and Andrew J. Dannenberg. A Bayesian model averaging approach for observational gene expression studies. Annals of Applied Statistics, 6(2):497– 520, June 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1339419605.

Zhu:2017:THD

[ZLDR17] Lingxue Zhu, Jing Lei, Bernie Devlin, and Kathryn Roeder. Testing high-dimensional covariance matrices, with application to detecting schizophrenia risk genes. Annals of Applied Statistics, 11(3):1810–1831, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1507168849.

Zhu:2018:USF

[ZLDR18] Lingxue Zhu, Jing Lei, Bernie Devlin, and Kathryn Roeder. A unified statistical framework for single cell and bulk RNA sequencing data. Annals of Applied Statistics, 12(1):609– 632, March 2018. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1520564486.

Zhang:2023:SNR

[ZLJW23] Hong Zhang, Ming Liu, Jiashun Jin, and Zheyang Wu. Signal-noise ratio of genetic associations and statistical power of SNP-set tests. Annals of Applied Statistics, 17 (3):2410–2431, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-statistics/
volume-17/issue-3/Signal-noise-ratio-of-genetic-associations-
and-statistical-power-of/10.1214/22-AOAS1725.full.

Zhang:2010:RDC

[ZLOS10] Zhongyang Zhang, Kenneth Lange, Roel Ophoff, and Chiara Sabatti. Reconstructing DNA copy number by penalized estimation and imputation. Annals of Applied Statistics, 4(4):1749–1773, December 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1294167797.

Zeldow:2019:SMA

[ZLR19] Bret Zeldow, Vincent Lo Re III, and Jason Roy. A semiparametric modeling approach using Bayesian additive regression trees with an application to evaluate heterogeneous treatment effects. Annals of Applied Statistics, 13(3):1989–2010, September 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1571277780.

Zou:2020:ELT

[ZLR20] Jingjing Zou, David J. Lederer, and Daniel Rabinowitz. Efficiency in lung transplant allocation strategies. Annals of Applied Statistics, 14(3):1088–1121, September 2020. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-14/issue-3/Efficiency-in-
lung-transplant-allocation-strategies/10.1214/20-
AOAS1350.full.

Zablocki:2017:SCM

[ZLS^{+}17] Rong W. Zablocki, Richard A. Levine, Andrew J. Schork, Shujing Xu, Yunpeng Wang, Chun C. Fan, and Wesley K. Thompson. Semiparametric covariate-modulated local false discovery rate for genome-wide association studies. Annals of Applied Statistics, 11(4):2252–2269, December 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1514430285.

Zhang:2018:MHT

[ZLZB18] Heping Zhang, Dungang Liu, Jiwei Zhao, and Xuan Bi. Modeling hybrid traits for comorbidity and genetic studies of alcohol and nicotine co-dependence. Annals of Applied Statistics, 12(4):2359–2378, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078048.

Zhou:2016:PMA

[ZM16] Zhengyi Zhou and David S. Matteson. Predicting Melbourne ambulance demand using kernel warping. Annals of Applied Statistics, 10(4):1977–1996, December 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1483606848.

Zhang:2019:RIR

[ZMA^{+}19] Youyi Zhang, Jeffrey S. Morris, Shivali Narang Aerry, Arvind U. K. Rao, and Veerabhadran Baladandayuthapani. RadioiBAG: Radiomics-based integrative Bayesian analysis of multiplatform genomic data. Annals of Applied Statistics, 13 (3):1957–1988, September 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1571277779.

Zhang:2020:ECR

[ZMA^{+}20] Hong Zhang, Bhramar Mukherjee, Victoria Arthur, Gang Hu, Hagit Hochner, and Jinbo Chen. An efficient and computationally robust statistical method for analyzing casecontrol mother-offspring pair genetic association studies. Annals of Applied Statistics, 14(2):560–584, June 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1593449316.

Zhang:2023:SMS

[ZMB23] Bo Zhang, Emily J. Mackay, and Mike Baiocchi. Statistical matching and subclassification with a continuous dose: Characterization, algorithm, and application to a health outcomes study. Annals of Applied Statistics, 17(1):454– 475, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Statistical-matching-and-subclassification-

with-a-continuous-dose--Characterization/10.1214/
22-AOAS1635.full.

Zemplenyi:2021:FFR

[ZMC^{+}21] Michele Zemplenyi, Mark J. Meyer, Andres Cardenas, Marie-France Hivert, Sheryl L. Rifas-Shiman, Heike Gibson, Itai Kloog, Joel Schwartz, Emily Oken, Dawn L. DeMeo, Diane R. Gold, and Brent A. Coull. Function-on-function regression for the identification of epigenetic regions exhibiting windows of susceptibility to environmental exposures. Annals of Applied Statistics, 15(3):1366–1385, September 2021. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-3/Function-on-
function-regression-for-the-identification-of-epigenetic-
regions/10.1214/20-AOAS1425.full.

Zhang:2022:TST

[ZMLS22] Jingru Zhang, Kathleen R. Merikangas, Hongzhe Li, and Haochang Shou. Two-sample tests for multivariate repeated measurements of histogram objects with applications to wearable device data. Annals of Applied Statistics, 16(4):2396–2416, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-4/Two-sample-tests-
for-multivariate-repeated-measurements-of-histogram-
objects/10.1214/21-AOAS1596.full.

Zhang:2022:MBD

[ZMO22] Yuping Zhang, Disheng Mao, and Zhengqing Ouyang. Model-based distance embedding with applications to chromosomal conformation biology. Annals of Applied Statistics, 16(3):1253–1267, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-3/Model-based-
distance-embedding-with-applications-to-chromosomal-
conformation-biology/10.1214/21-AOAS1479.full.

Zhang:2021:LSI

[ZNB^{+}21] Zhenyu Zhang, Akihiko Nishimura, Paul Bastide, Xiang Ji, Rebecca P. Payne, Philip Goulder, Philippe Lemey, and Marc A. Suchard. Large-scale inference of correlation among mixed-type biological traits with phylogenetic multivariate probit models. Annals of Applied Statistics, 15(1):230–251, March 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-1/Large-scale-
inference-of-correlation-among-mixed-type-biological-
traits/10.1214/20-AOAS1394.full.

Zhang:2014:UCR

[ZNSL14] Zhiwei Zhang, Lei Nie, Guoxing Soon, and Aiyi Liu. The use of covariates and random effects in evaluating predictive biomarkers under a potential outcome framework. Annals of Applied Statistics, 8(4):2336–2355, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1419001746.

Zhang:2017:SFD

[ZOZ17] Yuping Zhang, Zhengqing Ouyang, and Hongyu Zhao. A statistical framework for data integration through graphical models with application to cancer genomics. Annals of Applied Statistics, 11(1):161–184, March 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1491616876.

Zhang:2018:DPT

[ZPBW^{+}18] Yilin Zhang, Marie Poux-Berthe, Chris Wells, Karolina Koc-Michalska, and Karl Rohe. Discovering political topics in Facebook discussion threads with graph contextualization. Annals of Applied Statistics, 12(2):1096–1123, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743487.

Zamberletti:2021:MRF

[ZPGO21] Patrizia Zamberletti, Julien Papaïx, Edith Gabriel, and Thomas Opitz. Markov random field models for vectorbased representations of landscapes. Annals of Applied Statistics, 15(4):1628–1651, December 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-4/Markov-random-

field-models-for-vector-based-representations-of-
landscapes/10.1214/21-AOAS1447.full.

Zanghi:2010:SOI

[ZPMA10] Hugo Zanghi, Franck Picard, Vincent Miele, and Christophe Ambroise. Strategies for online inference of model-based clustering in large and growing networks. Annals of Applied Statistics, 4(2):687–714, June 2010. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1280842136.

Zubizarreta:2014:MBP

[ZPR14] José R. Zubizarreta, Ricardo D. Paredes, and Paul R. Rosenbaum. Matching for balance, pairing for heterogeneity in an observational study of the effectiveness of for-profit and notfor-profit high schools in Chile. Annals of Applied Statistics, 8(1):204–231, March 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1396966284.

Zeng:2021:CMA

[ZRA^{+}21] Shuxi Zeng, Stacy Rosenbaum, Susan C. Alberts, Elizabeth A. Archie, and Fan Li. Causal mediation analysis for sparse and irregular longitudinal data. Annals of Applied Statistics, 15(2):747–767, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-15/issue-2/Causal-mediation-
analysis-for-sparse-and-irregular-longitudinal-data/
10.1214/20-AOAS1427.full.

Zhang:2021:ILM

[ZRCC21] Rong Zhang, Zhao Ren, Juan C. Celedón, and Wei Chen. Inference of large modified Poisson-type graphical models: Application to RNA-seq data in childhood atopic asthma studies. Annals of Applied Statistics, 15(2):831– 855, June 2021. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
15/issue-2/Inference-of-large-modified-Poisson-type-
graphical-models--Application/10.1214/20-AOAS1413.
full.

Zhou:2009:FDA

[ZS09] Yingchun Zhou and Nell Sedransk. Functional data analytic approach of modeling ECG t-wave shape to measure cardiovascular behavior. Annals of Applied Statistics, 3(4):1382– 1402, December 2009. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1267453945.

Zhu:2017:BLS

[ZS17] Xiang Zhu and Matthew Stephens. Bayesian large-scale multiple regression with summary statistics from genomewide association studies. Annals of Applied Statistics, 11 (3):1561–1592, September 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1507168840.

Zheng:2018:CPP

[ZS18] Yuchen Zheng and Nicoleta Serban. Clustering the prevalence of pediatric chronic conditions in the United States using distributed computing. Annals of Applied Statistics, 12(2):915–939, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1532743481.

Zhao:2022:NDC

[ZS22] Sen Zhao and Ali Shojaie. Network differential connectivity analysis. Annals of Applied Statistics, 16(4):2166–2182, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Network-differential-connectivity-analysis/
10.1214/21-AOAS1581.full.

Zhang:2023:BMI

[ZSCL23] Shuangjie Zhang, Yuning Shen, Irene A. Chen, and Juhee Lee. Bayesian modeling of interaction between features in sparse multivariate count data with application to microbiome study. Annals of Applied Statistics, 17(3):1861–1883, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-3/Bayesian-modeling-of-interaction-between-

features-in-sparse-multivariate-count/10.1214/22-
AOAS1690.full.

Zhou:2022:SEM

[ZSFS22] Ling Zhou, Shiquan Sun, Haoda Fu, and Peter X.-K. Song. Subgroup-effects models for the analysis of personal treatment effects. Annals of Applied Statistics, 16(1):80–103, March 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-1/Subgroup-effects-models-for-the-analysis-
of-personal-treatment-effects/10.1214/21-AOAS1503.
full.

Zhou:2011:DMA

[ZSG11] Rensheng R. Zhou, Nicoleta Serban, and Nagi Gebraeel. Degradation modeling applied to residual lifetime prediction using functional data analysis. Annals of Applied Statistics, 5(2B):1586–1610, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1310562734.

Zubizarreta:2013:SII

[ZSG^{+}13] José R. Zubizarreta, Dylan S. Small, Neera K. Goyal, Scott Lorch, and Paul R. Rosenbaum. Stronger instruments via integer programming in an observational study of late preterm birth outcomes. Annals of Applied Statistics, 7(1):25–50, March 2013. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1365527189.

Zhou:2014:DBR

[ZSG14] Rensheng Zhou, Nicoleta Serban, and Nagi Gebraeel. Degradation-based residual life prediction under different environments. Annals of Applied Statistics, 8(3):1671–1689, September 2014. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1414091229.

Zhang:2013:RRS

[ZSH13] Lingsong Zhang, Haipeng Shen, and Jianhua Z. Huang. Robust regularized singular value decomposition with application to mortality data. Annals of Applied Statistics, 7(3):1540–1561, September 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1380804806.

Zhang:2023:BCS

[ZSLH23] Bohai Zhang, Huiyan Sang, Zhao Tang Luo, and Hui Huang. Bayesian clustering of spatial functional data with application to a human mobility study during COVID-19. Annals of Applied Statistics, 17(1):583–605, March 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-1/Bayesian-clustering-of-spatial-functional-
data-with-application-to-a/10.1214/22-AOAS1643.full.

Zhou:2019:TRT

[ZSMJ19] Tianjian Zhou, Subhajit Sengupta, Peter Müller, and Yuan Ji. TreeClone: Reconstruction of tumor subclone phylogeny based on mutation pairs using next generation sequencing data. Annals of Applied Statistics, 13(2):874–899, June 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1560758431.

Zhou:2020:RTS

[ZSMJ20] Tianjian Zhou, Subhajit Sengupta, Peter Müller, and Yuan Ji. RNDClone: Tumor subclone reconstruction based on integrating DNA and RNA sequence data. Annals of Applied Statistics, 14(4):1856–1877, December 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
14/issue-4/RNDClone--Tumor-subclone-reconstruction-
based-on-integrating-DNA-and/10.1214/20-AOAS1368.
full.

Zhang:2019:FDN

[ZSP19] Ningshan Zhang, Kyle Schmaus, and Patrick O. Perry. Fitting a deeply nested hierarchical model to a large book review dataset using a moment-based estimator. Annals of Applied Statistics, 13(4):2260–2288, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910044.

Zubizarreta:2014:ICN

[ZSR14] José R. Zubizarreta, Dylan S. Small, and Paul R. Rosenbaum. Isolation in the construction of natural experiments. Annals of Applied Statistics, 8(4):2096–2121, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1419001736.

Zhang:2023:LHW

[ZSS23] Lin Zhang, Lisa J. Strug, and Lei Sun. Leveraging Hardy– Weinberg disequilibrium for association testing in casecontrol studies. Annals of Applied Statistics, 17(2):1764– 1781, June 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
17/issue-2/Leveraging-HardyWeinberg-disequilibrium-
for-association-testing-in-case-control-studies/10.
1214/22-AOAS1695.full.

Zidek:2014:REB

[ZST14] James V. Zidek, Gavin Shaddick, and Carolyn G. Taylor. Reducing estimation bias in adaptively changing monitoring networks with preferential site selection. Annals of Applied Statistics, 8(3):1640–1670, September 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1414091228.

Zanini:2016:URM

[ZST16] Paolo Zanini, Haipeng Shen, and Young Truong. Understanding resident mobility in Milan through independent component analysis of Telecom Italia mobile usage data. Annals of Applied Statistics, 10(2):812–833, June 2016. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

http://projecteuclid.org/euclid.aoas/1469199894.

Zhang:2020:AMF

[ZTCS20] Chelsea Zhang, Sean J. Taylor, Curtiss Cobb, and Jasjeet Sekhon. Active matrix factorization for surveys. Annals of Applied Statistics, 14(3):1182–1206, September 2020. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-

14/issue-3/Active-matrix-factorization-for-surveys/
10.1214/20-AOAS1322.full.

Zimmerman:2019:OAC

[ZTH19] Dale L. Zimmerman, Jun Tang, and Rui Huang. Outline analyses of the called strike zone in Major League Baseball. Annals of Applied Statistics, 13(4):2416–2451, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330

(electronic). URL https://projecteuclid.org/euclid.
aoas/1574910050.

Zhang:2020:TPT

[ZTLW20] Hong Zhang, Tiejun Tong, John Landers, and Zheyang Wu. TFisher: a powerful truncation and weighting procedure for combining p-values. Annals of Applied Statistics, 14(1):178– 201, March 2020. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1587002670.

Zhou:2007:CHM

[ZW07] Qing Zhou and Wing Hung Wong. Coupling hidden Markov models for the discovery of cis-regulatory modules in multiple species. Annals of Applied Statistics, 1(1):36–65, June 2007. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1183143728.

Zhou:2008:REL

[ZW08] Qing Zhou and Wing Hung Wong. Reconstructing the energy landscape of a distribution from Monte Carlo samples. Annals of Applied Statistics, 2(4):1307–1331, December 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL http://projecteuclid.org/euclid.aoas/
1231424212.

Zhou:2012:ISB

[ZW12] Baiyu Zhou and Alice S. Whittemore. Improving sequencebased genotype calls with linkage disequilibrium and pedigree information. Annals of Applied Statistics, 6(2):457– 475, June 2012. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1339419603.

Zhang:2015:RBP

[ZW15] Wenfei Zhang and Ying Wei. Regression based principal component analysis for sparse functional data with applications to screening growth paths. Annals of Applied Statistics, 9(2): 597–620, June 2015. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1437397103.

Zhu:2018:RDM

[ZW18] Zhe Zhu and Roy E. Welsch. Robust dependence modeling for high-dimensional covariance matrices with financial applications. Annals of Applied Statistics, 12(2):1228–1249, June 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (elec-

tronic). URL https://projecteuclid.org/euclid.aoas/
1532743492.

Zhang:2019:JMA

[ZW19] Hongbin Zhang and Lang Wu. Joint model of accelerated failure time and mechanistic nonlinear model for censored covariates, with application in HIV/AIDS. Annals of Applied Statistics, 13(4):2140–2157, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https:

//projecteuclid.org/euclid.aoas/1574910039.

Zhang:2008:TFB

[ZWS08] Nancy R. Zhang, Mary C. Wildermuth, and Terence P. Speed. Transcription factor binding site prediction with multivariate gene expression data. Annals of Applied Statistics, 2(1):332–365, March 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1206367824.

Zhao:2013:EBT

[ZWW13] Zhigen Zhao, Wei Wang, and Zhi Wei. An empirical Bayes testing procedure for detecting variants in analysis of next generation sequencing data. Annals of Applied Statistics, 7(4):2229–2248, December 2013. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1387823317.

Zeng:2019:PBT

[ZWZ19] Li Zeng, Joshua L. Warren, and Hongyu Zhao. Phylogenybased tumor subclone identification using a Bayesian feature allocation model. Annals of Applied Statistics, 13(2):1212– 1241, June 2019. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL https://projecteuclid.org/
euclid.aoas/1560758444.

Zhu:2022:STP

[ZX22] Shixiang Zhu and Yao Xie. Spatiotemporal-textual point processes for crime linkage detection. Annals of Applied Statistics, 16(2):1151–1170, June 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-2/Spatiotemporal-
textual-point-processes-for-crime-linkage-detection/
10.1214/21-AOAS1538.full.

Zhang:2012:BML

[ZY12] Guangyu Zhang and Ying Yuan. Bayesian modeling longitudinal dyadic data with nonignorable dropout, with application to a breast cancer study. Annals of Applied Statistics, 6(2):753–771, June 2012. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1339419615.

Zhang:2017:BIH

[ZYC^{+}17] Tingting Zhang, Qiannan Yin, Brian Caffo, Yinge Sun, and Dana Boatman-Reich. Bayesian inference of highdimensional, cluster-structured ordinary differential equation models with applications to brain connectivity studies. Annals of Applied Statistics, 11(2):868–897, June 2017. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1500537727.

Zhang:2019:ERC

[ZYFF19] Ye Zhang, Zhigang Yao, Patrik Forssén, and Torgny Fornstedt. Estimating the rate constant from biosensor data via an adaptive variational Bayesian approach. Annals of Applied Statistics, 13(4):2011–2042, December 2019. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL

https://projecteuclid.org/euclid.aoas/1574910034.

Zhang:2016:SSP

[ZYXS16] Nancy R. Zhang, Benjamin Yakir, Li C. Xia, and David Siegmund. Scan statistics on Poisson random fields with applications in genomics. Annals of Applied Statistics, 10(2): 726–755, June 2016. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1469199891.

Zheng:2008:MCC

[ZZ08] Lu Zheng and Marvin Zelen. Multi-center clinical trials: Randomization and ancillary statistics. Annals of Applied Statistics, 2(2):582–600, June 2008. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1215118529.

Zhang:2018:VSE

[ZZ18] Baqun Zhang and Min Zhang. Variable selection for estimating the optimal treatment regimes in the presence of a large number of covariates. Annals of Applied Statistics, 12(4):2335–2358, December 2018. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/euclid.aoas/1542078047.

Zhang:2022:SIV

[ZZ22] Baqun Zhang and Min Zhang. Subgroup identification and variable selection for treatment decision making. Annals of Applied Statistics, 16(1):40–59, March 2022. CO-DEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-16/issue-1/Subgroup-identification-
and-variable-selection-for-treatment-decision-making/
10.1214/21-AOAS1468.full.

Zeisel:2011:FCA

[ZZD11] Amit Zeisel, Or Zuk, and Eytan Domany. FDR control with adaptive procedures and FDR monotonicity. Annals of Applied Statistics, 5(2A):943–968, June 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http:

//projecteuclid.org/euclid.aoas/1310562212.

Zhu:2022:SSD

[ZZD22] Weicheng Zhu, Zhengyuan Zhu, and Xiongtao Dai. Spatiotemporal satellite data imputation using sparse functional data analysis. Annals of Applied Statistics, 16(4):2291–2313, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Spatiotemporal-satellite-data-imputation-
using-sparse-functional-data-analysis/10.1214/21-
AOAS1591.full.

Zou:2008:NMB

[ZZH08] Hui Zou, Ji Zhu, and Trevor Hastie. New multicategory boosting algorithms based on multicategory Fisherconsistent losses. Annals of Applied Statistics, 2(4):1290– 1306, December 2008. CODEN ???? ISSN 1932-6157 (print),

1941-7330 (electronic). URL http://projecteuclid.org/
euclid.aoas/1231424211.

Zhang:2011:BBB

[ZZL11] Yu Zhang, Jing Zhang, and Jun S. Liu. Block-based Bayesian epistasis association mapping with application to WTCCC type 1 diabetes data. Annals of Applied Statistics, 5(3):2052–2077, September 2011. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL http://

projecteuclid.org/euclid.aoas/1318514295.

Zhang:2022:DAE

[ZZTL22] Wei Zhang, Zhiwei Zhang, James F. Troendle, and Aiyi Liu. Data-adaptive efficient estimation strategies for biomarker studies embedded in randomized trials. Annals of Applied Statistics, 16(4):2250–2265, December 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-statistics/volume-
16/issue-4/Data-adaptive-efficient-estimation-strategies-
for-biomarker-studies-embedded-in/10.1214/21-AOAS1588.
full.

Zou:2023:BID

[ZZXL23] Haotian Zou, Donglin Zeng, Luo Xiao, and Sheng Luo. Bayesian inference and dynamic prediction for multivariate longitudinal and survival data. Annals of Applied Statistics, 17(3):2574–2595, September 2023. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL https://projecteuclid.org/journals/annals-of-
applied-statistics/volume-17/issue-3/Bayesian-inference-
and-dynamic-prediction-for-multivariate-longitudinal-
and-survival/10.1214/23-AOAS1733.full.